A Complete Bibliography of IEEE Transactions on Computers (1990–1999)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

10 November 2025 Version 1.20

Title word cross-reference

(12, 8) [BBT94, Joh93, SW90]. (16, 8) [GB93, KGB95]. (2^{n} + 1) [Ma98]. (2 log N − 1) [HSY97]. (d ≤ t) [SAR91]. (m, k) [HR95b, HR97]. (x/d) [ALB98]. {0, 1} [SV94]. 1 [LT90, PEH90, PV94a, THL92, VRK93a, VCB96, WL94]. 16 [CR90a]. 2 [ABZ96, CZSM99, CSA94, KR93, LM92c, MC93, MC95, SB90, SLH97, SPM97, Tse97, TNC95, VCB96]. 2^{N} − 1 [SR92]. 3 [COK95, LT90, PEH90, PV94a, VRK93a, WL94]. 4 [ABZ96, EL90b, EL90d, MC92, Tak92a]. ^{2} [ABK90, IO98]. _{DDQ} [ZWL98]. _{K} [MLK98]. m [YL98]. A + B = K [CL92b, Par94]. b [XSFR96]. cos^{−1} [MMM93]. d [Nik91, NK92, SAR91]. d > k > t [NK92]. d > t [Nik91]. GD(2^{n}) [Ara93]. GF (2^{4n}) [PFR98]. GF (2^{m}) [BCH93b, Dro98, FBT96, GW98, HB92, HWB92, Has97, Has98, WP90, WL93, Wan94a, Wei94, WH97a]. GF (q) [CM96a]. I [CC96a]. I^{2}L [DS94]. K [GP97, She95b, BG90, BBKA95, CSF90, CLC92, CRS98, Dal90, Dal91, DH90, FBM93, FR97b, GY95a, GS93, HSL96, JHK97, KFYA93, LH91, NK92, OK99, RHT95, YC94]. log N [YA90]. m [AM91, CW99a, DSPN95, SB92b, WG98b]. m_{j} [WS95]. N [CA91b, ENS91, THL92, BG90, BBKA95, CZSM99, CSF90, CW99a, Dal90, Dal91, GY95a, Lat93, Lin93b, LH91, OK99, PR99, RBG98, WG98b]. O(log_{2} N) [DFM95]. O(n) [HTYP97]. O(n^{2.5}) [RS94]. O(N · log N) [Szy97]. O(N lg N) [She95a]. P [YM92]. p^{t} [HTCT91]. sin^{−1} [MMM93]. t [LB90, Nik91, NK92, SAR91, XH95, YL98]. T.H = O(2^{n/2}) [Fer91]. t/s [LB90]. Z [Cor96]. Z_{2}(m) [FRK94].

-ary [BBKA95, Dal90, Dal91, GY95a, GS93, LH91, OK99, SB92b, WG98b]. -Bit [XSFR96]. -by- [CZSM99]. -color [CRS98]. -coloring [CRS98]. -Coterie [YC94, KFYA93]. -cube [Dal90, Dal91, Lat93, Lin93b, CA91b, ENS91]. -cubes [BBKA95, GY95a, LH91, OK99]. -D [CSA94, KR93]. -depth [HTYP97]. -detections [PR99]. -diagnosability [LB90]. -diagnosable [XH95]. -dimensional [RBG98, WG98b]. -EC [YL98]. -EC/AUED [YL98]. -error [Nik91, NK92]. -extra-stage [HSL96, She95b]. -fault-tolerant [DH90]. -firm [HR95b, HR97]. -functions [YM92]. -II [ABK90]. -matrices [SV94]. -nearest [CLC92]. -out-of- [CSF90, CW99a, LT90, PEH90, WL94, THL92]. -out-of-n [DSPN95]. -pairwise [GP97]. -projection [WS95]. -resilient [FR97b, RHT95]. -shuffle-exchange [HTCT91]. -sorter [BG90]. -spare-approach [VRK93a]. -stage [HSY97, YA90]. -track- [VRK93a]. -trees [FBM93]. -UEC [SAR91]. -UED [SAR91]. -unidirectional [NK92, SAR91].

13th [Ano98b]. 17th [Tri98].

2-bit [Nai95]. 2-level [vdBT93].

32-bit [MFH91].

754 [ABCW92].

802.6 [DL97]. 86^{T M} [MLK98].

90/150 [NC96].

ABFT [NAB96]. Absolute [HM91]. Abstract [Wel94]. Abstraction [MAH98]. Abstractions [Oik92]. Accelerated [Kan93a]. accelerator [WSEZ90]. accelerators [SJ95]. acceptable [SC90b]. access [ACZD94, CF94, COK95, Dow92, HL91b, HC93, KMD97, KT91a, LS91b, MM91a, MMM96, Maz93, SMV91, TPDB98, VLPA95, WK97]. accesses [Har92, RH93]. accessing [AMS96]. accumulation [DM97a, Has97]. Accumulator [RT93a]. Accumulator-based [RT93a]. Accuracy [AB98, DM94, KC93b, CK99, Egg91, KE97]. Accurate [KTM^{+}94, Lew94, WF92]. achievable [PM98]. achieve [LF96]. Achieving [TYJ99, KWF95, RYT95, Wu95]. acknowledgment [Bal98]. Active [TRS^{+}99]. acyclic [BHS92, CL90]. adaptable [FNS95]. Adaptation [CV92]. Adaptive [BK90, CS90b, CK90b, FBL96, LYS95, SS96, BFS97, DN98, HB95, KM97b, LK95, LH91, OK99, UVM97, ZC98]. Add [JP98, HNR90]. Addendum [DBAP94]. Adder [Lo98, CK94, GM98a, Kan93c, Lo97b, LS92, MJ92, TH99, WT90]. adders [CS90a, CSTO92, DP96, DO96, Kan93a, Kan93b, PLM91, Pie94, RB99, Tya93]. addition [ABC92, CM96b, EL90a, Fag92, KOIH92, Lew90, QF92, Tho97, VCB96, Wal97, ZC90]. additions [Jun95, Ung94]. additive [DC93, GRT96]. address [AC98, CK99, CT91a, Har92, HLP93, Kat97, WSY97, WHL95]. Addressing [CSK90]. adjacency [CT94b]. adjacent [BBT94, GB93, KGB95, SW90, TB98b]. adjustment [CK99, KM97b]. admissibility [SXW95, She95a]. adopted [Ara93]. advanced [Leu97a]. advances [JBA^{+}97]. affine [SHC96]. aggregation [BT90]. Agrawal [CLL91]. agreement [AK96, VP95]. aided [BM94]. air [KBMP99]. algebra [CFV91, Jun95, TCI98, TLROL93, Ung94]. algebraic [CT94a, LK91b]. Algorithm [Ano96b, BKM94, BRS^{+}90, CK90b, CV92, DM96, DM93, HHB91, Kos91, OY91, RB90, RCBB96, RCB96, SR96, WJ94, ABLZ97, AVBZ97, AHT92, Ara90, BDK98, BB90, Bar98, BHS92, BB91, BM96, BW95, CCGS98, CRC96, CLC92, CPK90, CB96, Cor96, DS94, DDGS97, DN98, DA96, EW93, FRK94, FS94a, GDY90, GMP94, GPS93, GRR94, GW98, HTYP97, Han91, HB98, HL90, HN93, HC96, HTCT91, HBAB95, Hwa97, Kam92, KOIH92, KSL93, KMM91, KK92, KYM97, KM90, KM92b, Kum91, Leu97a, LHL99, LC92, LS94b, MM97a, MSS99, NS96, NAB96, OYR95, OK99, PK93, Pha98b, PLJ94, RS94, RDZ95, Red98, RR93, Saa95, SS90a, SJ90, SS93a, SBP90, SBP91, SXW95, She95a, SC95b, SJP96, SP95, SLS95, SL90b, SW97, Tak92a, TCI98, TK97, Tem99, TY92b, VRK93b, VT94]. algorithm [VJ93, WL90, WZXX93, WeBM97, YL95b, Zar95, ZB97, dM90]. Algorithm-based [BRS^{+}90, RB90, RCBB96, RCB96, WJ94, BB90, CRC96, CB96, FRK94, GRR94, NAB96, OYR95, Red98, RR93, TK97, VJ93]. Algorithmic [AEBVZ91, LPP91, JBA^{+}97, LMT95, OVL96]. Algorithms [AC98, CSS90, DL94, Fow93, Has90, Kor94b, Lo94b, OF97b, SL95, THH92, WG94, AC95, BB92a, BBB^{+}94, BC95, BP95, CRMM98, CT98, CG96b, CRS98, CCL92, CG94, DB91, Esh91, GGD96, HL92a, HJ92a, HD95, HRP99, JK92, Joh95, KM97a, KMD97, KT99a, KY92, KS93b, LM92b, LS90b, LCT97, LL96b, NO94, NO91, OM91, PFSR99, PI93, PGS95, PK91, PB99a, Qui90, RT91b, Rao93b, RF91, RBK90, SM92a, SF91, SF92, SHC96, SA92, SBM^{+}98, SW94b, SA98, TY92a, TB91, TC95, VJ98, WPS97, YLC90, YMG99]. Aliasing [IN96, PR95a, PGK90, Rob91, CMH98, ESS98, FFKI95, LG95, MI93, PG91, SFM92, SM97]. alignment [PNDN99]. all-optical [YMG99]. All-to-all [PB97a]. Allasing [NAH98]. allocated [TF92]. Allocation [HS97, MYD96, AK92, CS90c, CU93a, CT92, DH91, FM91, FBM93, HHS97, HKL98, KM97a, KLL97, KF92, SWG92, SM96b, TSS95, TNC95, Vog90]. allocator [CG96a]. allow [CLL90]. Almost [FHA94, Weg94b]. alpha [CRP^{+}96]. alpha-particle-induced [CRP^{+}96]. alphabet [PG92]. Alternative [MB93, Lin91]. alternatives [DP99]. ALU [PV94a, VPB93]. always [BNP91]. ambiguities [YM92]. AMD5 [MLK98]. among [LYW94, RH93]. AMULET1 [WDF^{+}97]. analyses [SISU90]. Analysis [CM96a, CS90a, CL94b, DOR91, DW90, DHL96, GPTT98, HM93, KKKM99, Kip94, LD90, LHS^{+}98, LZM90, MMSS96, MMJ91, NC96, NAH98, Qui90, SM90a, SSS97, SD96, SL90a, Smi95, TI92, UP93, YB91, YLL90, ZB98, ABLZ97, BR93, BJ91, BC99, BAA97, BM90, Bok96, Bos90, BD97, BCLA91, CPV91, CPRV95, CB95a, CS91, CU93b, CL94a, CSSP90, CW99b, CI92, CS96c, CB96, CMST90, Cor94b, DDS98, DKL90, Dal90, DC93, Dav90, DZLH93, DB94, DRS91, DM92, GOH92, GS95, GY95b, Gos97, GW91, GL92, HSHG99, HH93, HMC90, HJ92b, HP95, IA91, KPI93, Kan90, KACV94, KC93a, KLV94, KS96a, KS96c, KKS93, KM94, LNS94, LY96, Lat93, LP98, LK91a, LKK^{+}94, LG92, LP91, LBF92, Mah94b, MS93, MDyF94, MYD96, MAH98, MTG93, MY94a, NS96, NAB96, NB93, NS90, OI90a]. analysis [PMF96, PGS95, PLM91, QSA91, Rao93b, RW98, RF91, RS95, SFM92, SM97, SB97, Ser99, SB94, SIKS95, SGP92, SB92b, SRHI98, SY99, VC92, VT94, Wil92a, WSB90, WL92, WI95, vdLS93]. Analytic [MY95, MY96, IOP99]. Analytical [HC93, HKN99, JCSM96, PMHY97, SVSB97, ESS98, JCMS97, MB93, ACH^{+}97a]. analyzer [MI93, PGK90]. analyzers [Hla92, MMJ91]. Analyzing [dG90, PG91]. AND/OR [BOH90]. angle [ABLZ97, HN93, HC96, LL92a]. annealing [ACH^{+}97b]. ANNs [MBR96, MBR95]. anonymous [LW94]. any [She95b]. aperiodic [BS99, SC95b, SLS95]. Application [KWF95, Lew94, NC97, PR94a, SF96b, Ara93, Bry91, CK90a, CZSM99, CCC96, CT93a, CW99b, CKY^{+}99, Cor94b, DBD93, DC93, EL90c, ENS91, GYY98, Jha93, KB93, Kle90, MS98b, PR99, RS95, SG90, TY92a, WP90, WG98a]. Applications [DJM99, BR94, BMP^{+}97b, CDFH93, CRS98, DW90, DPO95, FA96, HLH96, JTY^{+}99, Kal95, Kor94a, Lim92, LP99, LM99b, MS98a, MJ91, NKC94, PDBH97, PR95b, RB90, Sch96, Sch98, SSS97, ST95b, TCI98, Vij91, WPS96, WHK97, WH97b, XH95]. applied [Hei91]. Applying [ABCW92, SIKS95]. approach [ACH^{+}97b, BB92b, BJK93, BBRL98, CK90a, CPV91, CL90, CT98, CRC96, CF90, CL99a, CFV91, CT91b, FP98, FL93, Fri94, HS98a, HJ92b, KS91, KA94, KB93, KT99a, KF97, KR93, KK91c, Lin90, LMT95, NSUR98, OVL96, RYT95, RBG98, SY91a, SSL96, SWWW94, SB97, SJB91, SD92, SRL90, SK94, SA98, VRK93a, WM96, Wil93, WMD96, vdLS93]. approaches [PHK94]. approximants [GY95b]. approximate [AK96, HP95, SR98, SGP92]. Approximating [SS99, KT91a]. Approximation [HV91, SWT93, SF96b, CCGS98, CRS98, ITY97, KM92b, ZPP91]. Approximations [KZ90, WF92, ZVDF96]. approximators [Kos94]. ARB [FS96]. arbiter [VBH^{+}95]. arbitrarily [RF92]. arbitrary [AMS96, BDN96, BNG92, LA98, RDZ95, TS99a]. arbitration [Kip94, Wil92a]. architectural [CWM^{+}95, DP99, IS90]. Architecture [CS96b, Dek98, HB95, Lee92, ABC92, AN90, BGJ^{+}96, BS90a, BP92b, CRMM98, FP98, GB90a, GE92, HLB^{+}97, Har94, IO98, IO91, IJN94, JS92, KT99b, LVB^{+}93, LAY96, LD94, Lew90, LKL98, LS91b, LZ94, Ma98, Mat90, MF92, NO94, Paa96, Pou97, PV94b, RK96, RHS98, SJ90, SR98, SJB91, SY91b, SB92b, SP95, TRM95, TB99b, THA^{+}99, Tze93, UP93, WW91, WL93, Wei96]. Architectures [Ara93, AB98, AVBZ97, BOI94, BBBM95, CJ90, CJ91, CU93a, CT96b, DBAP93, DBAP94, DH90, HC93, KKL90, KHC94, McA92, MNL99, NLV91, PFR98, SES^{+}92]. arcs [SV95]. Area [BOI94, CK94, Sar90, WT90, KIN^{+}94, PKI^{+}96, SC90c, Tya93, WSY97]. area-efficient [KIN^{+}94, PKI^{+}96]. Area-time [CK94, WT90, SC90c]. Argument [Smi95, HTHR94]. Arithmetic [ABCW98, GRT96, Hei91, KL97, KC93b, Lew94, Lo94b, PG92, SJ95, Tan92, ABCC90, ABCW92, CR90b, DOP93, Dro98, EL92, Kat96b, KM90, LW91, LO97a, MM97b, Mul94, PFSR99, Par93, PB97b, Pie90, RF99, SKA92, Sri90, VCB96, Vui90, WJ90, Ano98b]. ARM [WDF^{+}97]. arrangement [DT93]. Array [Kor94b, LTR96, BOI94, BP93a, CPRV95, CCBC98, CU93b, CUC97, FL98, GW98, IJW92, IDI95, JLLR94, KRM95, Lee92, LLK92, LK91b, LC90, LK90, LHL99, LTR93, Meg92, Mie93, OFW92, Pek99, ST95b, Sto93, SCW91, TK97, Wan94a, Zar95]. arrays [BB92b, CT94a, CT98, CA91b, CF90, CT91b, FR97a, Fri94, Hug93, JBS90, JK92, KT99a, KR93, KT91c, KL92, LaF99, LK93, LR97, LH90, LBF92, LL96b, MRB90, Mel92, MY95, NA90, NM94, QM93, RM94, RBK90, SS93b, SM96b, SW94a, TB91, TC95, VJ98, VRK93a, WD92, WC90, ZWL98]. arrival [TH99]. arrivals [HS94d]. arsenide [MFH91]. artificial [RF91]. ary [BBKA95, Dal90, Dal91, GY95a, GS93, LH91, OK99, SB92b, WG98b]. ASCA [NO94]. aspects [PS90, TY92b, TC94]. Assertions [BGM^{+}95, Fit97]. Assessing [CDHK93]. assessment [LS98, PS91]. assigning [BLOS95]. Assignment [TS94, AC98, BM91, Bon90, BNG92, BS99, GS97, HL91a, HR95b, HB98, KA94, LLK92, Leu97b, NST96]. assignments [FW93, KWF95, Ros90, SC90b, TK91]. assisted [ACFH95, BB90, RKS90]. assists [SM90b]. Associative [Mer97, CFV91, CGC95b, HKN99, RS96, SKA92, SST92, WZXX93, Wan96, WHL95]. associativity [BS97]. Asymmetric [ABB94a, Rag93, Rag96, TB98b]. Asymmetric/unidirectional [ABB94a]. Asymptotic [KT92]. Asymptotically [Fra92, DFM95, JO93]. Asynchronous [AVR98, KT91a, MT98, AAL94, AS92a, AK96, BBD97, DB91, FW93, KA94, Kip94, LB99, LKL97, McA92, OYY97, SI94, TLE97, TKS94, VBH^{+}95, VT97, WM92, WDF^{+}97]. asynchronous/internally [VT97]. at-speed [HG96]. ATM [Hou99, RKJZ99]. atomic [DLW91]. ATPG [Dae91, WG98a]. attack [Kuh98]. attacks [GP96a]. AUED [BB92d, KB96, Kat96a, LY96, YL98]. Augmented [LAY96, MB95]. Augmenting [PNDN99]. Authenticating [Mit92]. authentication [RS98]. Author [Ano98a, Ano99a, KIN^{+}96]. Automata [CBGC94, NC97, BMP^{+}97b, BK95, CK90a, CM96a, CZSM99, CCC96, CCBC98, CGC95b, DC93, HMC90, KMM91, NKC94, NC96, NK97, OdSC96, SCC96]. automata-based [HMC90]. Automated [Cor94a]. Automatic [BDL^{+}99, IYI90, LM99b, MBGT91, OK96, LL91a]. automation [PR95b]. automaton [IO98, TW95a]. automotive [HLB^{+}97]. autonomous [LS96b]. Availability [KFYA93, KBMP99, Ser99, YC94, HY99, RS95]. available [SKK^{+}90]. Avalanche [HT95]. Average [BHSB97, BS93b, DM97b, Man90, Sas91, WH99]. avoidance [Bel90, BBB^{+}94].

B [LP98, LY90, PK90]. B-network [LY90]. B-networks [LP98]. B-trees [PK90]. back [EZ92]. back-recovery [EZ92]. backpropagation [RF91]. backup [HS98a]. Backward [BGM^{+}95, Fit97, HC96, LY90]. Balanced [TB99a, ABB93, ABB94b, TB98a, UVM97, WH97b]. balancing [BK90, BHS92, CL90, Dae91, HS93, KK92, LK98, NO94]. Ballast [GB90b]. Band [Tat97]. Bandwidth [KH90, Kam92, Soh93]. banks [RR95]. banyan [YA90]. BAR [HC96]. Base [DJM99, WSB90]. Based [CBGC94, KZ90, WG94, ASS94, AVR98, AVBZ97, AS92b, Bae97, BB90, BRS^{+}90, BDL^{+}99, Ber96, BMP97a, BK95, CJ90, CC92, CA91b, CRC96, CCBC98, CU93a, CU93b, CU94, CB95c, CL99a, CAK90, CYH91, CGC95a, CB96, DKL90, DL97, DTB96, DA96, FM95, FRK94, GPTT98, Gos97, GRR94, HM95, HHMC91, HRT^{+}95, HS94b, HMC90, HKC97, Hou99, JW91, Jun95, KE97, KT96, KKKM99, KKA95, KR93, KK91c, LS96b, LCFH95, LL91b, LS94a, LCL95, LO97a, LKL97, LZ94, LM99b, MY96, Mer97, MDyF94, MYD96, MTG93, NAB96, OYR95, ORPP93, Paa96, PGS95, Pek99, PG97, PW95, Pin99, PR97a, PR98, PR99, RT93a, RCS92, Rao93a, RFKL93, RFS92, RB90, Red98, RR93, RCBB96, RCB96, SCC96, SLD95, SC95b, SE98, SP96, SR91, SR96, ST96, TK97, TB99b, TS99b]. based [TLROL93, THZ^{+}99, Tse97, TYJ99, Ung94, VJ93, Vog90, WJ94, WS95, Wan96, WY95, WH97b, WG98b, YLC90]. baseline [LFH92, NL90]. BASEMENT [HLB^{+}97]. bases [PB97b, WHB98]. basic [SV95]. Basis [Kun92, FBT96, Has98, KS98, Lu97]. batch [YL95a]. BDDs [JBA^{+}97]. be [GM94, Kop95]. behavior [BJ91, BP96, GK92, HKN99, KKKM99, SV95]. behavioral [Cor94a]. Being [OI90b]. benchmark [Ber92, SS95]. benchmarking [IOP99]. bends [LYW94]. benefits [DP99]. Benes [RB91, Sch92]. Benes-like [Sch92]. Berger [LTR96, LTR93, RFKL93]. Berlekamp [Zar95]. better [Kop95]. between [CL92a, CL99b, FC97, JH95, OFW92, OZL96, SLH97, YA90]. Bhuyan [CLL91]. Bi [Mel92]. Bi-level [Mel92]. biased [SS91a]. biased-random [SS91a]. bidirectional [Mit92, Wan96]. Binary [KTM^{+}94, Kos91, Lo98, Ara92, BDN96, BAA95, Cha93, CU93a, DE92, FC97, FS90, GH91, HM95, JM96, JH95, Kip94, KM90, KB95, LV99, LPV96, LAY96, Lin93b, Lo97b, OCVSV98, PK93, Pou97, Pur91, RB99, SG90, SY91b, Tho97, Weg94a]. binary-reflected [JH95]. binary-tree [CU93a]. binomial [KKS93]. Bipartite [SS99, KWF95]. birthday [YF92a]. bisection [Saa95]. bisectional [Ho92]. BIST [CG96b, ITZ96, Nic96, PG91, RT99, Sav98b, WI95, ZI92]. Bit [HB92, IV94, SKA92, Wan94a, XSFR96, Coh93, KH93, KS98, KM90, Man90, MFH91, Nai95, PHK94, Szy97, Vai95, WH98, WHB98]. bit-complexity [Szy97]. Bit-level [Wan94a]. Bit-parallel [SKA92, KS98, WH98, WHB98]. bit-pipelined [KM90]. Bit-serial [HB92, IV94]. bit/byte [Vai95]. bitonic [LC98a]. BKM [BKM94]. block [CSTO92, DW91, Mon91, WdB92]. blocking [DR96]. board [CCBC98]. boards [KC93a]. Boltzmann [ZPP91]. Boolean [BC99, BBP91, BJK93, Bry91, DGY92a, DLN95, FC97, Fal99, GM94, JBA^{+}97, LL92b, LP99, LM92c, Sch96, Sch98, TMS97, WM96, WM98, Weg94b]. Borden [HPN95, Pie96]. Borel [TA95]. Borrow [Sto93]. both [SC95b]. Bottleneck [SW92]. bound [AM93, BOH90, KS90, LH98, PRS94, Qui90, WNS96]. boundary [NC96, RTZ98]. Bounded [PK94, ACH^{+}97b, Aln93, SISU90]. bounded-degree [Aln93]. bounded-difference [SISU90]. Bounding [HAM^{+}99, KM92a, LMT95]. Bounds [Ber93b, NAH98, Sas91, AR94, HM94, HBAB95, Krä98, Li97, MB95, RR93, SFM92, SM97, Sem95]. BPC [She95b]. Branch [DF91, PS93, SAMC99, CK99, EAK98, FM95, Fag97, GL93, KE97, MBGT91, Nai95, Qui90, RW98]. branch-and-bound [Qui90]. branch-and-combine [EAK98]. branching [Pha98a, Pha98b, Weg94b]. breakdowns [RS95]. bridged [DMS94]. bridging [CT96a, CL98b, KM94, PR98, ZWL98]. Broadcast [CA91a, YM95, BH92, BB99b, Fuj99, JL91, MMM96, OYY97, SSA94, Tse97, TS97, YLC90, YM91, YY97]. Broadcasting [THY93, BNL93, Bal98, CSK90, Fra92, GS93, LCT97, PB97a, Wu95]. broadcasts [TS95]. broadcasts/multicasts [TS95]. Bruijn [DHL96, PMF96, SP91, Ann97, DZLH93, LS96c, LLJ97, RB93, RB97, SR91]. BSP [GLR^{+}99, SN99]. buffer [ACFH95, DYD94, DB94, FNS95, Har94, PS93, Vog90, Wei96]. buffer-based [Vog90]. Buffered [RR95, MY94a]. buffering [BD97, MF92]. bufferless [BBP^{+}96]. buffers [AS95, CR93, FM95, Fag97, Kus90, TF92]. Bug [BW96a]. Built [HRT^{+}95, RFM90, BK95, GPZ99, GP96b, HMC90, IN96]. Built-in [HRT^{+}95, RFM90, BK95, GPZ99, GP96b, HMC90, IN96]. bus [BHR94, CS91, CL99a, CS92a, FY96, HHS97, Ish95, JS92, JL96, JW91, Kuh98, KEA96, LD94, Mah94b, OI90a, Wil92b, YB91, vdLS93]. bus-based [CL99a]. bus-control [JW91]. bus-encryption [Kuh98]. bused [KKL90, KEA95]. buses [BB95, KH97, Raj96]. busses [Kip94]. butterfly [GH91, Tze94]. bypassing [JCMH99]. Byte [BAB92, CGC95a, DDV90, HF97, SI93, SCC96, Vai95, XSFR96]. Byzantine [VP95].

C [LH90]. C-testable [LH90]. C2SC [SSA94]. CA [CGC95a]. CA-based [CGC95a]. CAA [SCC96]. Cache [GW92, MV99, Tak92b, CL98a, CB99, CCCH93, CS99, CHH98, DDS98, DW91, Egg91, FL93, HKN99, HAM^{+}99, HSHG99, HS98c, IS94, JLC98, JCMH99, KKKM99, KRC99, KLH^{+}99, LHS^{+}98, Mah94a, MTP93, NB93, NV99, PNDN99, PFP99, PHS99, PG97, PBND98, PH93, RBS99, SS95, SS93a, SAMC99, SVSB97, STW92, SY99, TSW92, TWS92, TWS93, TG99, TXD98, TS98, WSY97, WHL93, Yan93, ZC98]. cache-coherent [PBND98]. cache-related [LHS^{+}98]. caches [KHW94, OMB97, Sez97, SST92, SV97, TRS^{+}99, TLH94, VJ98, WHL95]. caching [SV97, WSY97]. CAECC [CBGC94]. calculating [CY95]. calculation [ABLZ97, LL92a]. calculations [ABK90]. call [DL97]. CAM [IO98]. Cambridge [ZVN92]. Can [Kop95, GM94]. canonical [CRC96, KS98, LM92c]. capabilities [LW91]. capability [CC97b, KS92, MMJ91]. capable [YM92]. capacity [CF94, CGC95b, TSS95]. Carry [CM96b, Lo98, PK94, CS90a, CSTO92, CL92b, DO96, EL90a, Kan93a, Kan93b, Kan93c, Lo97b, LS92, Par94, Pie94, Tya93]. carry-lookahead [CSTO92, Kan93c]. carry-lookahead/carry-select [Kan93c]. carry-propagate [EL90a]. Carry-save [CM96b, Pie94]. carry-select [Kan93c, Tya93]. carry-skip [CS90a, CSTO92, Kan93a, Kan93b]. case [ACH^{+}97a, ACH^{+}97b, Ber92, BFS97, BHSB97, CK91, KPB92, Krä98, Vai98]. cases [ZPP91]. Causal [AV97, Bal98]. caused [JLC98]. CAUTRA [KBMP99]. Cayley [Als92, MPS94, TA95]. CCD [AEBV90, AEBVZ91]. CD [KM94]. cell [CA91b, HD98]. cells [CA90a, LCS93, RT99, SMV91]. Cellular [CCBC98, CBGC94, HMC90, NC97, NK97, BMP^{+}97b, BK95, CM96a, CZSM99, CCC96, CGC95b, DC93, GS97, IO98, NKC94, NC96, PV90, SCC96, SLN97]. Cellular-automata-array-based [CCBC98]. central [BK90]. centralized [BBB^{+}94]. CEQRNS [ST95b]. Certification [SWM95]. certifier [BSM97]. Chains [PK94, BT90, NGB93, Sem95, Wal98]. Channel [FCL98, CF94, Dow92, GS97, HS98a, MM91a, WL90]. channels [CA91a, LH98]. Chaos [KS94]. chaotic [BFS97, PBB97]. characteristic [LKL97]. characteristics [HT95]. characterization [CF95b, Cha93, DAF94, LB90, PG97, Weg94a, XH95]. characterizations [Mul94]. Characterizing [Joh95]. Check [LTR96, GRR94, LTR93]. checked [LS90a, MLK98]. checker [BJ94, CW99a, PEH90, RFKL93, THL92, Tao97, WL94]. checkers [DSPN95, GP96b, HPN95, Jha93, JW94, LT90, Nik98, Pie90, Pie96]. checking [BB90, BJ94, CW99a, HPN95, Jha93, KPB92, KT99a, KSGT96, LT90, Lo96, RFKL93, SM90b, THL92, UWZ97]. checkpoint [Vai97, ZB97]. checkpointing [Gai90, PV94b, Vai97, ZB98]. checkpoints [BP92b, Wan97]. Checks [SM90a, CB96, NHA92, SJ93a]. Checksums [SM90a, SM90b]. Chinese [HP95]. Chinese-Remainder-Theorem [HP95]. chip [FTG92, KT99b, LS91b, Maz93, MYK^{+}99, NV99, SL91c, TW95b, WSY97]. chip-multiprocessor [KT99b]. chips [SL92, Sta93]. choice [DBJ98]. Choices [BW95, CS92a]. chordal [HW95]. Cipher [Kuh98, BMP97a, LCL95]. circle [Hua92]. Circuit [CVMS99, CCGS98, DOP93, DR96, HTYP97, HRP99, KA94, KSL93, Lim92, Maz93, MY94b, PR94a, PR94b, PR98, RFM90, SAF90, Wei94]. circuit-switched [DR96]. circuitry [LH96, Tho97]. Circuits [Vui94, AR94, Bar92, BH96, BS92, BOH90, BS90c, Cha96, CT96a, CMS99, CR97, CA90b, CAK90, CL98b, Dam92, Dav90, DGY92a, DGY92b, Dro98, FW93, Fuj90, HM93, HRT^{+}95, Hou99, HG96, KPB92, KIN^{+}94, KM95, KKS93, LL96a, LIO94, LB99, LKL97, Lo96, NSUR98, PKI^{+}96, Pie93, PK91, PR93a, PR95c, PR96b, PR96a, PR97a, PR99, RD96, RT94, SISU90, SB92a, SB97, SSDK95, SMOR98, TSH94, TSH95, WM92, XK92]. circular [Ara92, MS93]. circulating [AD97]. circumvent [MS91]. Class [Kor94b, PDBH97, TA95, Aln93, AHT92, CU94, Cor94b, DKT96, FS94a, HF97, HPN95, HWB92, HWB93, Joh93, KS98, LS96a, LL96b, MM97a, MM95a, OSST96, SCG94, WH98, Yan98, YF92b]. Class-congruence [TA95]. classes [DBD94, KT92]. Classification [PR93a, DBD94, LS94a, TMS97]. classifiers [CFV91]. client [DM92, WNPM95]. client-server [DM92]. client-server-like [WNPM95]. Clock [Fis90, Chi94, CPK90, CL94b, EAK98, GP96a, KS90, LZM90, OS94, Pal94, RKS90]. clock-controlled [GP96a]. clocked [VT97]. Clos [MY94b, YM99, YW99]. Clos-type [YM99]. Closed [BHR94, WH99, BR90, KT92, LP91]. Closed-form [WH99]. closure [SM92a]. cluster [DYD94, GP97, MDyF94, TNC95]. cluster-based [MDyF94]. clustered [AHT92, BP93a, Hog99, MY96]. clustering [SK96]. CMOS [BSS94, CS90a, CL98b, HM93, KA94, QF92]. co [ABCW98, DMS95, DJ99, RT91a]. co-NP [RT91a]. co-processor [DMS95]. co-synthesis [DJ99]. co-transformations [ABCW98]. Coda [SKK^{+}90]. code [BBT94, CLM^{+}95, CW99a, CCCH93, CGC95a, DRW92, GB93, JH95, KGB95, LT90, PEH90, RFKL93, RCSSS94, SCC96, SW90, THL92, WL94, CBGC94]. Coded [Lo94b]. codes [AVR98, ABB93, ABB94a, ABB94b, Bos91, BAB92, BB92d, Che92, CT94b, CC96a, DSPN95, DDV90, DM97b, FM91, FRK94, FRK98, HF97, HPN95, HBA98, Joh93, KB96, Kat96a, KY92, KS93b, KR90b, LY96, Lin93a, MK90, Mon91, NA90, Nik91, NK92, Pie90, Pie96, SI93, SAR91, SW95, TB98a, TB98b, TB99a, THL91, WdB92, XSFR96, YL98, ZC90]. coding [BBRL98, BCH95, BJ94]. coefficient [WMD96]. coefficients [ORPP93]. COFTA [DJ99]. Cohen [KH93]. Coherence [MNL99, CB99, DW91, HSHG99, SVSB97]. coherency [Egg91]. coherent [NB93, PBND98]. Cohorts [JHK97]. coincident [KS96b]. collaborative [SBM^{+}98]. collapsing [PV94a, VPB93]. color [CRS98]. colored [CDFH93]. coloring [CRS98, FS94b]. column [SM96b, WJM95]. combination [KRC99]. combinational [AR94, BSS94, CT96a, Fuj90, Hei91, HG96, KM95, LT90, PR96b, PR99, RT94]. Combinatorial [Lat93, CB95a, DPO95, TY92b, VT94]. combine [EAK98]. combined [ALB98, HSY97, LM99a]. Combining [CHH98, HLHC91, Tze92]. Comment [MM97a, BMP97a, JCMS97]. Comments [ABC92, BMP^{+}97b, CK94, CL94b, Dav90, HNG94, Jun95, Kat94, KGB95, KY92, Mah94a, NC97, PMF96, Par94, PKI^{+}96, PT92, Pha98a, RS94, SB94, SW94b, Weg94a, Wei96, WH94, Wil92a, YC94, JBS90]. commit [ALPS90, EZ92]. commitment [DLW91]. common [DMS95, GM98b]. communicating [CK90a, KS96c]. Communication [CB97, LD90, SS90b, BB95, DW97, DL97, Fuj99, GM91, GM98b, HR95a, HS98a, HHMC91, Ish95, KD94, LH92, LK96, MS99, PBB97, PS91, RHS98, TF92, TS94]. communications [IIK^{+}91, LL97, LT93, LLJ97, QM93, YLC90, ZR97]. commutative [RS96]. compacted [LCS93]. compaction [CMH98, HRP99, PW95, RT93a, RP99]. compactors [ITZ96, Kop95]. Comparative [SC94, SY99]. comparator [KSGT96]. compare [PLJ94]. compare-and-swap [PLJ94]. Comparing [HW97]. Comparison [JP98, NT92, SSA94, Vai96, AVR98, Ber96, BP92a, BB99b, DIP93, KHW94, KK91c, Met91, Pel92, SD92, Wan99]. comparison-based [AVR98, Ber96, KK91c]. Comparisons [FTL93, DYD94, MY95]. compensation [LD94]. Competitive [BH98, CRMM98]. Compile [HL91a, HL97]. Compile-time [HL91a, HL97]. Compiler [ACFH95, BB90, LCFH95, CWM^{+}95, LM99b, WDS^{+}95]. Compiler-assisted [ACFH95, BB90]. Compiler-based [LCFH95]. compiler-controlled [CWM^{+}95]. compiler-inserted [LM99b]. complement [SG90]. complementary [PHK94]. complementation [WM96]. Complete [LW95a, LW95b, AAL94, BBP91, Bok96, BW96b, CT92, GH91, RT91a, TNS99, ZPP91]. completely [HM91]. completion [CT93b]. Complex [BKM94, CV92, Kat94, ABCW98, DHK93, FSS99, LW91, SS92, WJ90]. Complexity [BP92a, SB97, AM93, Bar98, Bry91, CCL92, Dro98, DLN95, EAK98, Fuj90, HWB92, HB95, Hua92, KS98, KEA95, Lim92, LH90, Paa96, PK91, Rao93a, SB90, Szy97, TW95c, WH98, WHB98]. component [BB92a, CSS90, TS94]. components [LT95]. Composed [Wel94]. composite [Paa96, PFSR99]. Comprehensive [MOM91, XT99, CUC97]. compress [LP90]. Compressed [Zhe94]. Compression [PG97, KN90, PG91, Rob91, Tat97, WJM95]. Compression-based [PG97]. computable [Mul94]. Computation [ALB98, CL97, KL97, Kos91, TAY91, BM94, CK90a, CJ90, CB96, GPS93, Gre94, MT98, PR95a, SRK91, UP93, WD92, WMD96, WPZZ96]. Computational [Fuj90, RM94, Rao93a, CMD93, PS90, SWM95, dM90]. Computations [WG94, BH98, DA96, Fit97, Krä98, LO94a, MvdS90, MPKRS93, NR90, NO91, NLV91, QSA91, RSS97, RCB96, SL95, TRM95, TH91, WJ90, WJR94]. computed [DLN95]. Computer [Ano98b, Ber92, BR90, BNG92, CLL91, Che92, DG91, DDV90, FRK98, GL92, Han91, KK92, LK98, PS90, PLB93, PT92, Ser99, SWG92, SP96, Vui90, WW91, WHL93, dG90]. computers [BHBA91, CT92, CSS90, Dek98, DH91, FRU96, FY96, Hog99, HS93, IS94, KS96a, KS96b, LH92, OM91, PB99b, Raj96, RFB97, RF92, SIKS95, SSK^{+}96, Soh90, TS99a, WH94, Yan93, Ano96a, Ano98a]. Computing [Ano97b, BT90, CL97, EL90e, MMM93, MS94, RHT95, Tri98, Ber93a, BCH93b, CH90, CL94a, EL92, FCL98, GLR^{+}99, HL92a, HS96, JTK97, Kan96, KM97a, LW94, LS90c, LT95, Mer97, Pur91, SC90b, WP90, WL93, ZVDF96, ZR97]. concentrators [GO98]. concepts [Kin91]. Conceptual [Cyr94]. Concurrency [SBM^{+}98, Uht92]. Concurrent [CR97, DA90, Hug93, KCCD99, KL92, OYY97, UR94, WK97, BM90, BL99, GP96b, HBAB95, JTY^{+}99, LS91b, OYR95, OPLP93]. concurrently [LS90a]. condition [XK92]. Conditional [LHNP94, Lo98, KM92a, Lo97b]. conditions [CL92b, Par94, RN95, WdB92, YM99]. cone [PR99]. cone-based [PR99]. Conference [Tri98]. configurable [HS98b, SSH^{+}99, TB99b]. Configuration [LaF99, CZSM99]. configurations [WL92]. configuring [NGB93]. Conflict [ASS94, HL91b, VLPA95, COK95, DBD93, MM91a, MMM96, VB96]. Conflict-free [HL91b, VLPA95, COK95]. conflicts [CL99b, KT99a, MS94, TG99]. congestion [GC97]. congruence [TA95]. congruential [OSST96]. conjugate [CL96]. connected [ASS94, AMB96, BB92a, BCS94, CS96a, CSS90, LT95, Mat90, MD95, OM91, Raj96, RF92, SC95a, Tze93, WH94, Zar95]. connecting [Hua92]. connection [MS98b, EL90e, MM95a]. connections [Ish95]. Connectivity [DHL96, DZLH93, LHNP94, PMF96, TS99a, Wil92b]. conquer [LZ94]. consensus [Kum91]. conservation [LS97]. consideration [CA92, HS94d]. consistency [CL98a, ZC98]. Consistent [Wan97, MRJ97]. Constant [CC94, DM96, LL92a, Lin93a, LO94a, Mie93, TAY91, NL90, SW94a, TB98a, TW95c]. Constant-factor [LL92a]. constant-time [TW95c]. constrained [Cha99, HLH96, HHS97, ZSR90]. Constraint [Sut91]. constraints [ASWW92, CL94a, DTAL93, HS97, Li97, SL95, SS94c, TW95a, WNS96]. Constructing [CL92a, Lin91, Yan90, ZS99, BB92d, LT93, YL98]. Construction [GRR94, SW95, HWB92, TYJ99, Wil93]. constructions [KB96]. constructs [HL97]. consumption [ZR97]. contain [Wan97]. containment [CS99]. Contention [TNC95, AMS96, DE92, DW91, Fri95, JW91]. contention-based [JW91]. Contention-free [TNC95]. context [HC94b]. Continued [Smi95, Vui90]. Continued-Fraction [Smi95]. continuous [KK91b, RBG98]. Control [SM90b, ACZD94, FRK98, FTL93, GS95, HF97, JW91, KBMP99, KMD97, LK95, LL91b, McC95, MMSS96, PR93b, Sav98a, SS94a, SW95, SBM^{+}98, Vai95, ZR97]. Control-flow [SM90b, SS94a]. controllability [Fuj90]. controllability/observability [Fuj90]. controlled [Ber96, CWM^{+}95, GP96a]. controller [KS96a, KS96b, MNL99, TKS94]. controllers [FSS99, LS90a]. Convergence [GGD96, HHB91, HTHR94]. convergent [BH96]. conversion [AM93, JH95]. conversions [Kor94a]. converter [CN99, RM91]. Cooperative [TH91]. coordinated [SC95a]. copies [HW97]. Coping [LRG91]. Coprocessor [KZ90]. CORBA [GS98]. CORDIC [ABZ96, ABLZ97, AVBZ97, DM96, DM93, EL90c, HTHR94, HC94a, HD95, HHB91, HN93, HC96, KC93b, LA98, LL92a, LS90b, MMM93, PGS95, Pha98a, Pha98b, TAY91, THH92, WPS97]. CORDIC-based [PGS95]. correct [Lam97]. Correcting [CBGC94, DM96, ABB93, ABB94a, BBT94, BAB92, Che92, CGC95a, DM97b, FM91, FRK94, GB93, Joh93, KGB95, KR90b, Lin93a, Maz93, MK90, Nik91, NK92, RCSSS94, RNA93, SI93, SAR91, SCC96, SW90, WdB92, XSFR96, YM92]. correcting/ [Nik91]. correcting/detecting [WdB92]. Correction [Ano96b, Bou90, LTR96, Lo98, MBR96, NAH98, SP91, SBP91, VS93, BL99, DOP93, Kat96b, Kat97, KS93b, KL92, Red98, SL90b, WS95]. correctly [Lam97, LMT98]. correctness [KYM97]. correlated [TI92]. Correlation [GP96a, FM95]. correlation-based [FM95]. Corrigendum [DHL96, TWS93]. Cosine [TAY91, CJ90, Pha98b]. Cost [AEBV90, CSA94, AZ98, BS90c, CW99a, CGC95b, DDS98, GS93, JO93, JS92, LKK^{+}94, PZ99, Pie90, QSA91, SJ95, San93, SB92a, SB97, SIKS95, Tze92, WSY97, dM90]. cost-effective [Pie90, Tze92]. cost-effectiveness [JS92, LKK^{+}94]. Cost-performance [CSA94]. cost-table [BS90c, SB97]. cost-tables [SB92a]. costs [PZ99]. Coterie [YC94, KFYA93]. Counter [PW95, FM95, KLPS93]. counter- [FM95]. Counter-based [PW95]. counters [CA90a, CW94, KTM96, STE98, Tok94]. Counting [SM96a]. coupled [SMV91]. Coverage [CPA99, MV95, MAH98, PMW92, PMAC95]. covering [DM97b, HBA98, ZPP91]. covers [SM96b]. CPU [NV99, PHS99, TS98]. Crash [DW97, TS95]. Creating [TCW93]. criterion [Sch96]. critical [JHK97, Ung95]. crossbar [GO98, HM95, Wil92b]. crossbar-based [HM95]. crossed [KB95]. Crossing [MNKF90]. crosspoint [VC92]. CRT [CN99]. Cryptanalysis [LG97, O’C95]. Cryptography [NC97, BMP^{+}97b, NKC94]. cryptosystem [LG97, TY92a]. CSMA [KM94]. CSMA/CD [KM94]. cube [ASS94, AMB96, BCS94, CA91b, Dal90, Dal91, ENS91, HSL96, Lat93, Lin93b, OM91, She95a, Tze93, WH97b]. cube-based [ASS94, WH97b]. cube-connected [BCS94, OM91, Tze93]. cube-type [HSL96, She95a]. cubes [BBKA95, CL95, Dal91, GY95a, HBA98, JH95, KB95, LH91, OK99]. cumulative [BT90, DG91]. current [KIN^{+}94, PKI^{+}96, SISU90]. current-mode [KIN^{+}94, PKI^{+}96]. cut [DRS91, HV91, KS92, KM94, SWK90, Tra96, Vij91]. cut-through [DRS91, KS92, KM94]. CVS [Jha93]. Cycle [LD94, BCLR97, GC97]. cycle-stealing [BCLR97]. cycles [AMB96, BCS94, DT93, Tze93]. Cyclic [HKL98, SM90a, GK92].

d [RCSSS94, CSA94, COK95, KR93, SLH97, Tse97, TNC95]. D-mesh [TNC95]. d-UED [THL91]. d-unidirectional [RCSSS94]. D. [ABC92, CLL91]. Damage [LS98]. Data [BH92, JK92, Wel94, WS91, AMS96, AHT92, BHS92, CL90, CB95c, HL91a, HBA98, IYI90, JLC98, KRC99, KR90a, Kum91, MS93, MRT97, MF92, NR90, PNDN99, PM91, PB99b, RTZ98, RNA93, Rob91, SJ93b, Tak92b, TRS^{+}99, TH91, WPS96, WK97, YY97]. data-access [WK97]. data-dependent [HL91a]. data-driven [SJ93b]. data-flow [HL91a, PM91]. data-parallel [TH91]. Database [SL90a, FNS95, GS95, HS93, LLS91]. databases [LL90, Leu97a]. dataflow [AN90, GB90a, HL97]. David [Dav90]. DB [LLS91]. DBEC [KY92]. DC [YL98]. DCC [OSST96]. DCVS [KPB92]. de-skewing [COK95]. dead [MTP93]. deadline [BS99, HS97]. deadlines [ACZD94, HR95b, HR97, RKJZ99]. Deadlock [KS93a, Bel90, BBB^{+}94, CG94, GDYS96, LH98, Pin99, SS96]. Deadlock-free [KS93a, CG94, GDYS96, LH98, SS96]. decision [BAA95, BOH90, BHSB97, CK90a, Cha93, FC97, FS90, LV99, LPV96, LL91b, MM91b, OCVSV98, Pur91, Weg94a, ZV98]. decision-making [CK90a]. decoded [IS94]. decoder [IDI95, SCC96, SLN97]. decoders [AVR98]. decoding [HP95, KY92, TB99a]. decomposable [LNS94]. Decomposition [Kat94, SS92, SV94, CR90b, CJ90, CL90, CYH91, DBJ98, EL90e, Hwa97, LIO94, LW95a, LW95b, LP91, NS90]. decompression [RTZ98]. Decoupled [Sez97, KHC94]. deeply [McC91, McC95]. defect [KKS93, LKK^{+}94, SK96, TSH94, TSH95]. defect-tolerant [KKS93, TSH94, TSH95]. deferrable [SLS95]. Deferring [HR95a]. defined [LL91a]. Degradable [VP95, LS95]. degrading [CSSP90, SN99]. degree [Aln93, BHBA91]. Delaunay [SBP91, SBP90]. Delay [BB95, CSTO92, PW95, AVR98, BAA95, BAA98, BE92, Cha96, CS96b, CKC96, Coh93, FFS92, GL93, Gre94, HG96, KH93, KM95, KD94, LHS^{+}98, LH96, LMT95, McC91, McC95, PMW92, PR96b, THL92]. Delay-insensitive [BB95, AVR98, LH96]. delay-sensitivity [BE92]. delay-verifiable [KM95]. delays [KT91a, PLM91]. Deleting [PRS94]. Delimiting [Yok92]. delivery [HKC97]. delta [CC96a]. delta- [CC96a]. DEMUS [McC95]. denotational [Bou90]. denumerable [RS95]. DEPEND [Gos97]. Dependability [AKL90, SRHI98, TI93, ACC^{+}93, Gos97, MD95, Rai95, SJP96, TS99b]. dependable [GSH^{+}92, HS98a, LKL98, NNHG93]. dependence [KT99a, SHC96]. dependencies [RT96, SF91, SF92, Uht91]. dependent [HL91a]. Depth [SRK91, DFM95, HTYP97]. Depth-size [SRK91]. derivation [Pal94]. derivations [Sch98]. Deriving [BAA98, CKLY98, SA98]. description [DA90, KSNM91, Mil91, SS91b]. Design [ABB93, ABB94b, BP94, BAA97, CPRV95, CSSP90, CS96c, CBGC94, FTG92, GW91, KPB92, KC93a, KS96a, Maz93, MRJ97, MC92, MRT97, NB93, NAH98, OF97a, Pie90, Pie94, Pie96, PBND98, SM92a, SHHJ93, SB92b, Szy97, TB98a, TC95, WS98, Wil92a, AE94, Ber92, BZ92, BS90c, CCGS98, CS90a, CW99a, Cd93, CCCH93, CK94, CG96b, CL99a, CS92a, CGC95b, DOR91, DH92, FL92, GB90b, HK90, Hla92, Hon90, IDI95, KS91, KA94, KIN^{+}94, KH97, LY96, LLS91, LHF93, LC90, LK90, LO97a, LH96, LT90, LH90, LS94b, MM95b, MFH91, MC94, MJ92, Nik91, NK92, OFW92, PKI^{+}96, PFP99, PS93, PR95b, PZ93, RCSSS94, SJ95, SM97, SB92a, SB97, SJB91, SS91b, SJ93a, TLE97, TB98b, TK96, TS99b, TY92b, VP93, WC96, WP90, WJM95, WT90, WDS^{+}95]. design [WC99, XK92, Yan93, ZV98]. Designing [HLP93, JCSM96, Kan93b, MM91b, CT98, DH90, EAK98, JCMS97, PG91, THL92]. Designs [SS94b, GP98, LB90, RTZ98, RFKL93, SY99]. Destage [VJ98]. Destination [RFS92]. Detailed [LBF92]. detect [PR96b, TMS96]. detectable [Oik92]. Detecting [LB99, ABB94a, BBT94, BAB92, DDV90, GB93, Joh93, KGB95, KR90b, Lin93a, MK90, Nik91, NK92, RCSSS94, SS90a, SI93, SAR91, SW90, WdB92, XSFR96]. Detection [JBS90, LFH92, SF96a, AR94, BBRL98, BM90, BS92, CH90, CR97, Dam92, DW90, Fri94, Hei91, Hug93, Jha93, JW94, KRM95, KL92, LL96a, LH90, MMJ91, OYR95, RB90, RCBB96, Ser99, SJ93a, TSS95, WeBM97, WK97]. detections [PR99]. Determinacy [SK92b]. Determination [HS96, CT97b, HNY93, TL96]. Determining [WRS95]. deterministic [BK95, KTM96, KMM91, Kus90, Sch96]. development [IIK^{+}91]. device [TPDB98]. devices [GM98a]. DFT [Ara90, SA98]. Diagnosabilities [KK91a]. Diagnosability [VJ93, Wan94b, Wan99, LB90, NL90, RT91b, RT91a, SP96, SR93]. diagnosable [RS94, SD92, XH95]. Diagnosing [RF92]. Diagnosis [BP93b, KLV94, RT99, TS99a, BB92c, BP92a, BSM92a, BSM92b, BB99b, CSF90, CCBC98, DTAL93, DF90, DP97, DN98, FBL96, FHA94, KK91a, KF97, KK91c, KF92, LS93, LYS95, Pel91, Pel92, Pel98, Rag96, RDZ95, Rao93a, Rao93b, SD92, SLD95, Som90, SA92, SCW91, VP94, VJ93, Wan99]. diagnostic [KC93a]. diagnostics [Sav98b]. Diagonal [TP94, LS94b, Pea96, Tse97]. diagram [Pur91]. diagrams [BAA95, BHSB97, Cha93, FC97, FS90, LV99, LPV96, OCVSV98, Weg94a, ZV98]. diameter [BS93b, CCD^{+}92, Efe91, KTS90, Lat93]. dictionary [LP90, PR97a]. dictionary-based [PR97a]. difference [PBB97, SISU90]. different [San93]. Differential [CK90b, DM96, GM98a, O’C95, RCBB96]. Digit [Kor94a, Par90, PK94, AW93, CR90a, CT91a, GM98a, MC93, MC94, MC95, Par93, Pou97, WH99]. digit-by-digit [MC94]. Digit-set [Kor94a]. Digital [Ano96b, AB98, BP98, BS92, Bou90, Cd93, CR97, Kip94, MM95b, PK91, PR97a, RD96, SAF90, YL95b, YL95a]. Digraph [DHL96, DZLH93, FL92, PMF96]. digraphs [Kno96, PS99]. Dijkstra [Bar98]. dilated [Tse97]. dilated-diagonal-based [Tse97]. dilation [MH90]. dimension [SWT93]. dimensional [CM96a, FL98, IO98, IO91, KT99a, Maz93, RBG98, SF96a, ST96, TW95b, UVM97, Ver92, WM91, WG98b]. dimensions [BCH95, RYT95, SBP90, SBP91, YTR94]. Diogenes [BB92b]. Diophantine [BMP97a, LG97, LCL95]. direct [BB91, GDYS96, SV97]. direct-form [BB91]. direct-mapped [SV97]. directed [BS93b, PB99a]. disagreeing [HW97]. discovered [KOIH92]. Discrete [KK91b, CJ90, CW94]. disjoint [CT97a, FA96, TCW93, WL94]. Disk [FM91, BBBM95, CPRV95, DKL90, FR97a, Gai90, KT91a, LK93, MY95, Ng91, NM94, TSW92, VJ98]. disks [BBRL98]. dispatching [GPS95]. display [MS94]. dissemination [BH94, KT91b]. dissimilarity [OZL96]. dissipation [WG98a]. Distance [HLH96, BHBA91, BBKA95, BS93b, KTS90]. Distance-constrained [HLH96]. distances [LYW94]. Distortion [LVB^{+}93]. Distributed [AG94, GDYS96, LW94, LK99, RB97, SS90b, Sav99, SLD95, SL90a, SA92, SW94b, Tri98, WW91, WDS^{+}95, YMG99, AV97, ALPS90, AHT92, BBD97, BHS93, BH94, BB92c, BNG92, BES^{+}96, CK90a, CB99, CL94a, DJ99, DN98, FP98, GHH93, GYY98, Gun97, HL92a, HLB^{+}97, HSHG99, HJ92a, HS94c, HS94d, HS97, IO91, IA91, IJN94, KM97a, KK92, KS96c, KA96, LS93, LYS95, LK98, MvdS90, MS95, OS94, RKS90, RDZ95, RPW98, SKK^{+}90, ST95a, SC94, SB94, SWG92, SC90b, SES^{+}92, SI94, Tze94, TS95, WNPM95, WF90]. distributed-memory [Tze94]. distributed/parallel [LK98]. distributing [CS96c]. distribution [Chi94, DG91, KLL97, KKS93, NM94, YY97]. divide [LZ94]. Divider [NL99, HB92]. Dividers [LTR96, LTR93, PV90]. divisible [BHR94]. Division [BW96a, CL94c, ELM94, Has97, Man95, OF97b, OF97a, WF92, ALB98, BW95, CR90a, Dow92, EL90d, FBT96, GW98, HTHR94, HK95, ITY97, LM99a, LC92, Man90, Man93, MPG98, MC92, MC93, MC94, MC95, MLK98, QM93, SF93a, SP95, SPM97, Wal98]. Division-and-accumulation [Has97]. division-free [HTHR94]. division/square [ALB98, LM99a]. division/square-root [ALB98]. divisions [WL93]. DOACROSS [HKL98]. Doing [Pea96]. domain [HTHR94]. dot [DM97a]. Double [DJM99, Has98, Pha98b, BBBM95, CC92, CT97a, GB93, HW95, Joh93, KGB95, PT92, XSFR96]. Double-Base [DJM99].

Double-basis [Has98]. double-integer [CC92]. double-loop [HW95, PT92]. down [STE98]. DQDB [MB93]. drawing [PT98]. drawings [SY91a]. Driven [Tan92, GL92, Mah94a, SJ93b, TWS92, TWS93]. dropping [KACV94, PR95a]. DS5002FP [Kuh98]. DSM [DP99]. DSP [CKY^{+}99, DPO95, ST95b]. DSPs [MYK^{+}99]. Dual [JP98, FBT96, HHS97, SSS97, WHB98]. dual-bus [HHS97]. dual-link [SSS97]. Dual-Pass [JP98]. Duato [OK99]. duplex [Fuj99, Vai96]. duplication [ZB98]. Duprat [Pha98a]. during [GP96b, Har92, WG98a]. Dynamic [FTYZ90, HS93, JTY^{+}99, MMM96, CSTO92, CK99, CCCC96, DRW92, DW97, FS96, HL97, HR95b, HR97, HL91b, LVB^{+}93, Maz93, NR90, RW98, RP99, SSH^{+}99, SS93a, WJR94]. Dynamically [TF92, CMS99]. Dynamically-allocated [TF92].

Easily [Sas97, WC90, CCC96]. easy

[CL99a, CW94]. easy-to-use [CL99a]. EC [BB92d]. EC/AUED [BB92d, KB96, YL98]. EC/d [THL91]. ED [HF97]. Edge [FA96, KT91b, LNS94, LPV96]. Edge-disjoint [FA96]. edge-valued [LPV96]. editing [IJW92]. Editor [Tri98]. Editors [BC97, GL99b, KP98, MV99]. Effect [HJ92b, MS98b, BP96, CCCH93, SS95, SK96]. Effective [CB95c, GPZ99, Pie90, RYT95, Tze92, ZI92]. effectiveness [JS92, KA96, LKK^{+}94]. Effects [KLH^{+}99, ATF98, CS92b, EANPH93, KS96b, KHC94]. efficiency [CPV91, DB91, TS94]. Efficient [Ber96, BSM92a, CL96, CT94b, FE97, Fow93, GM94, HPN95, HJ92a, HC92, ITY97, LS95, MBR95, MBR96, Met91, Mon91, NAB96, OM95, PFR98, RFB97, RBK90, RP99, SBP90, SZN92, TL96, WHL95, WH97a, ABB93, ABB94b, Ann97, AMB96, Bel90, BBBM95, CF95a, CT94a, CB99, CW94, CHH98, CG94, DGY92a, DDGS97, DL97, Fri94, GPS93, GLR^{+}99, HTCT91, JLC98, Kam92, KIN^{+}94, KSL93, LS93, LK95, LL96b, MM91b, OYR95, PKI^{+}96, PA96, PEH90, Pin99, Pur91, RD96, WL90, Wu95]. eigenvalue [BC99, GPS93]. Election [AAL94]. elections [MS98a]. electronic [Mit92]. element [FHA94, ST95b, TRM95]. Elementary [BKM94, KZ90, SS94b, SS99, WG94, WG95]. elements [Dro98, Mat90, SJ93b]. Eliminating [McC91, TG99]. ELM [KOIH92]. embedded [CKY^{+}99, DJ99, LKL98, Nik98]. Embedding [ARS95, BDN96, BCGS96, BCS94, DT93, Ell91, GH91, KB95, MH90, YTR94, CDP98, GC97, LK99, RB93, RB97, SLH97, TC94]. embeddings [MNKF90, MPS94]. empirical [ST96]. Emulation [Mat90]. encoded [VSS91, VS93]. encoder [HB95]. Encoding [Kat94, Lo96, SS92, TB99a]. encryption [HT95, Kuh98]. end [RKJZ99]. end-to-end [RKJZ99]. energy [ZR97]. enforced [McC91, McC95]. engineering [Kin91]. Enhanced [TW91, PI93, SLS95, Wan94b, Wan99]. enhancement [GOH92, KF92]. enhances [BS97]. Enhancing [MV99]. Enriched [BBD97]. Ensure [CK90b]. entries [JHK97]. enumeration [LC90]. Environment [SS90b, CCG^{+}91, CRP^{+}96, CS92a, CW99b, CI92, CCCC96, Dae91, DYD94, Gos97, IOP99, KSNM91, RT99, SKK^{+}90, SZTN91]. environments [BS99, PDBH97, SLS95]. equation [LG97, WM91]. equations [Ber93a, BMP97a, Jun95, LCL95, PBB97, RCBB96, SISU90, TLROL93, Ung94, WM98]. equipartitioning [GDY90]. Equivalence [TSH95, YA90, BS90c, Cor94a, Lio93]. erroneous [LRG91]. Error [ABLZ97, BGM^{+}95, CBGC94, FRK94, HP95, Lo94b, PGS95, ZR97, ABB93, ABB94a, BM90, BAB92, CR97, Che92, CGC95a, CB96, DOP93, DDV90, DM97b, ESS98, FM91, Fit97, FRK98, HF97, HH93, Hug93, IN96, KKA95, Kat96b, Kat97, Krä98, KS93b, KR90b, KL92, Maz93, MK90, Mon91, Nik91, NK92, OYR95, RCSSS94, Red98, RCBB96, SAR91, SCC96, SL95, SJ93a, SW95, SL90b, TB98b, Vai95, WS95, XSFR96]. Error-Coded [Lo94b]. error-correcting [ABB93, Che92, CGC95a, DM97b, Maz93, RCSSS94]. error-detecting [RCSSS94]. error-detection [RCBB96]. errors [BBT94, Fit97, GB93, Joh93, KGB95, Lin93a, SI93, SW90, WdB92, YS96]. Estelle [CA91a, LL91a]. Estimation [GOH92, Cor96, CPA99, HC93, MV95, Red98]. Estimators [PMAC95]. Ethernet [CK91]. Euclid [GW98]. Euler [TYJ99]. evaluate [GP98, TSH94]. Evaluating [CS92a, HR97, KZ90, Rai95, TK97, DG91, Maj96]. Evaluation [CL92b, Par94, PFP99, AKL90, ACC^{+}93, BB97, BP94, BD95, Ber92, CPV91, CT96b, DDS98, DP99, FTG92, GB90a, HHMC91, HS94b, IS94, KA96, LLS91, MP90, MRJ97, MD95, MY94b, NO94, NHA92, PLB93, PMAC95, RBS99, SN99, SJP96, TN93, TS98, WG95, WHL93, ZVN92]. even [Tho97]. EVENODD [BBBM95]. event [CW94, KCCD99]. events [HBAB95]. every [FHA94]. Exact [LG92, OCVSV98, Vui90, HBAB95]. Exactly [SS94b]. Excel [ZCT99]. Excel-NUMA [ZCT99]. Exception [DL94]. exchange [Bok96, Hua90, HTCT91, MB95, RB91, Sch92, TNS99]. exclusion [HY99]. executable [PD99, RM91]. execute [Lam97]. Executing [AN90, Uht92]. Execution [KS96c, CWM^{+}95, CL94a, Hog99, IOP99, Leu97b, Li97, SSK^{+}96, TD93]. exhaustive [GRT96]. exitory [LB99]. EXOR [BS93a]. Expanding [HHB91, CCCH93]. expansion [WPZZ96]. expansions [KT92, LM92c, Pur91, RF99, VR94]. Expected [Rao93b, CK91, KS90, LMT95]. Expected-value [Rao93b]. experiment [LKK^{+}94]. Experimental [NP90, CI92, IYI90, Pal94]. experiments [BCSS90, Ber92, DF90]. exploit [XT99]. exploitation [CL96]. Exploiting [DP99, MV99, SS94a, PRASA99, TRS^{+}99, Tem99, TS99b, TSS95]. Expm1 [Tan92]. Exponent [Yok92]. exponential [Kan96, Man95]. Exponentiation [Wal98, Ara93, HTYP97, Tak92a, Wan94a, WH97a]. exponentiations [WP90]. exponents [PFSR99]. Express [Dal91]. expression [ESS98, WH99]. expressions [BS93a, DTB96, Dre99, MP90, Sas91, Sas97, SV95, Sri90]. Extended [SM90a, AMB96, GM94, LS94b, SM90b]. Extended-Precision [SM90a, SM90b]. Extending [Fit97]. extensible [CT98]. Extension [KL97, AS92b]. extensions [DDS98]. external [BL92, ESS98]. external- [ESS98]. externally [VT97]. extra [HSL96, She95a, She95b]. extra-stage [She95a]. extraction [Uht92]. Extreme [SC90c].

F. [Wil92a]. Factor [DM96, TAY91, CJ90, LL92a, LS94b, WS91]. Faddeev [Meg92]. Fail [Nic98, BES^{+}96, LC98b]. Fail-safe [Nic98, LC98b]. fail-silent [BES^{+}96]. failure [BP93a, CT93b, FR97b, HR97, NNHG93, WHK97]. failures [AAL94, AA92, BGJ^{+}96, BBBM95, CS96a, IYI90, NM94, PT92, SK94, SCW91, TI92, TS94, TS95, YM92]. fair [LD94]. fairness [MB93]. False [TLH94]. family [CT94a, CG97, DMS94, KT91c, SES^{+}92]. fan [LL96a]. fan-out [LL96a]. fanout [Lio93]. fanout-point [Lio93]. Fast [CRS98, CY95, CN99, DS94, DPO95, DTB96, DM93, EL90a, Fag92, FY96, HRP99, JO93, Kus90, LM92b, Lo98, NNHG93, PFSR99, RF99, WF92, WG94, WG95, ZVN92, AZ98, AHT92, ANR96, CF95a, CT91a, CT92, Dal92, DIP93, EZ92, Fio99, Kan93a, KOIH92, KS93b, Lo97b, Meg92, Mie93, OVL96, Pha98b, Saa95, SP95, STE98, TSH94, Wan94a, WPZZ96]. Faster [DL94]. fat [Gre94]. fat-pyramid [Gre94]. Fault [ACC^{+}93, BCSS90, BGM^{+}95, BC95, BCH93a, CDHT99, CS99, DF90, Jha93, KRM95, KF92, LCT97, LH90, LKW97, MRB90, MV95, MS95, RB93, SCD^{+}95, SR91, SCG94, TK96, VP93, VC92, WG98b, ASS94, Als92, AKL90, AK96, BR90, BR93, BR94, BRS^{+}90, Bar92, Ber96, BP98, BP92a, BSM92b, BB99b, BP95, BP96, BCH95, Cha96, CB95b, Cd93, CF90, CS90b, CU94, CUC97, CW99b, CW96, CC97b, CPK90, CI92, CS96c, CPA99, CS92b, Dam92, DTAL93, DJ99, DG91, DH90, DH92, DA96, DM97b, FP98, FRK94, FFS92, FA96, Fri94, Fuj99, GB91, GRR94, GP97, Hei91, HJ92b, HNG94, JHK97, JW94, KKA95, Kan90, KACV94, KS93a, KLL97, KK91c, KH97, LaF99, LNS94, LL96a, LMJ93, Lat93, LS96a, LH92, LK96, LK99, LS94a, LH91, LBF92, Maz93, Mel92, NA90, NHA92]. fault [NAB96, NG90, OPT92, PM98, PMW92, PS90, PLB93, Pel91, Pel92, PR93b, PR95a, PR95c, PR97a, PMAC95, PV94b, RS94, RYT95, Rai95, Rao93b, RB90, RNA93, Red98, RR93, RCB96, ST95a, SC94, SJB91, SLD95, SK94, SS91a, SY91b, SJ93a, ST95b, SJP96, Som90, SP96, SL92, Sta93, SR93, SRHI98, Sto93, SS96, SW97, SR96, TT90, TK97, TSS95, THZ^{+}99, TYJ99, TS95, VJ93, WJ94, WL94, WRS95, WH97b, YF92a, MM95a]. Fault-containment [CS99]. fault-diameter [Lat93]. fault-free [ASS94]. fault-injection [CPA99]. fault-state [SK94]. Fault-tolerance [VC92, SL92, Sto93]. Fault-tolerant [BC95, BCH93a, LKW97, MS95, RB93, SCD^{+}95, SR91, TK96, VP93, AKL90, ACC^{+}93, BR90, BR93, BR94, Bar92, BP98, BCH95, Cd93, CF90, CS90b, CU94, CUC97, CW99b, CW96, CC97b, CPK90, CS96c, DG91, DH92, FP98, Fuj99, GB91, JHK97, Kan90, KS93a, KLL97, KH97, LS96a, LH92, LK96, LK99, LBF92, Maz93, NA90, NHA92, PM98, PS90, PLB93, PR93b, PV94b, RNA93, ST95a, SY91b, SJ93a, ST95b, Sta93, SS96, SW97, TT90, TYJ99, TS95, VJ93, WH97b]. faults [AR94, BP94, BAA95, BCS92, CRP^{+}96, CH90, CT96a, CB97, CCBC98, CKC96, CL98b, DP97, HD98, HS94a, HG96, JBS90, KS96b, LM92b, LG95, LB99, Lio93, LFH92, Oik92, Pel98, PW95, Pol95, PR93a, PR96b, PR97b, PR98, PH93, SF96a, SW94a, TL96, Ver92, WK97, ZWL98]. faulty [BH94, BNL93, BJ91, BCS94, CA90a, CT97b, Fra92, GDYS96, GMP94, LS90c, MS98a, PB97a, RS96, RB97, RN95, WY95, Wu97, WG98b, YTR94]. FBDD [GM94]. FDDI [HR95a]. features [ABCW92, SCD^{+}95, SES^{+}92]. feedback [CA90b, GK92, HRT^{+}95, Rob91, SM92b]. feedbacks [Hla92]. Feng [LHL99]. FERRARI [KKA95]. fetch [HNR90, PFP99]. fetch-and-add [HNR90]. FFT [OYR95, WJ94]. FIAT [BCSS90]. Field [Yok92, IYI90, Paa96, PK93, WHB98, ZWL98]. fields [Dro98, HB92, HWB92, HWB93, KS98, Paa96, PFR98, PFSR99, WL93, WH98]. FIFO [vdBT93]. file [CL96, HW97, Met91, SKK^{+}90]. filters [BB91, LHF93, MM95b]. final [CM96b]. Finding [FS90]. fine [RSS97]. fine-grain [RSS97]. Finite [DB94, HH93, CK90a, CCC96, CKLY98, DOR91, Dro98, HB92, HWB92, HWB93, HB98, IJW92, KMM91, KS98, LY94, MY94a, Oik92, Paa96, PR97b, Ros90, TRM95, WL93, WJR94, WMD96, WH98, WHB98]. finite-state [IJW92, LY94, Oik92, PR97b]. firm [HR95b, HR97]. first [NT92]. first-in-first-out [NT92]. Fixed [KSS95, CRC96, CR93, DTB96, LHS^{+}98, MRT97, Raj96, TMS97, WMD96]. fixed-coefficient [WMD96]. fixed-polarity [CRC96]. fixed-priority [LHS^{+}98]. fixed-size [CR93]. fixed-width [MRT97]. flattened [PB99a]. Flexible [FNS95, Pin99, ACH^{+}97a, ACH^{+}97b, KKA95, RM94, RPW98]. Floating [BGM^{+}95, CV92, Has90, JP98, Lo94b, OY91, OF97a, SLW98, Tan92, Yok92, Fit97, Krä98, MLK98]. Floating-Point [CV92, Has90, JP98, Lo94b, OY91, OF97a, Tan92, Yok92, Krä98, MLK98]. flow [BHS92, CL90, CT91b, HL91a, JK92, LK95, PM91, PB99a, PB99b, SM90b, SS94a, Tak92b]. fly [EL92]. FOCUS [CI92]. folding [HLHC91]. forbidden [WG98b]. force [PB99a]. Foresighted [ASWW92]. fork [LP91]. fork/join [LP91]. form [BHR94, BB91, Fal99, Man95, Sch98, WH99]. Formal [LPV96, Win95, CW99b, KSNM91, Mil91]. Formalization [Kin91]. formally [BSM97, HHMC91]. formed [CDFH93]. forms [TMS96, TMS97, WM98]. formulas [HNY93]. Forward [FC97, BBB^{+}94, PV94b, PV97]. foundation [SS91b]. foundations [Nic98]. Four [McA92, DMS95]. Fourier [CT91a, Mie93]. FPD [WC99]. FPGAs [CMS99, HD98, RN95, ZV98]. Fraction [Smi95]. fractions [Vui90]. frame [Har94, Wei96]. Frames [CF95b]. Framework [Par90, PK94, GSH^{+}92, PLB93, PG91]. Free [RYT95, SC90a, Yok92, ASS94, COK95, CG94, FW93, GDYS96, HTHR94, HL91b, Joh95, KWF95, KS93a, LL96a, LH98, LKL97, SS96, TNC95, VLPA95, YTR94]. freeze [CR97]. French [KBMP99]. frequently [CF95b]. friendly [SZTN91]. FRIENDS [FP98]. FRR [SJ90]. FTRE [Rai95]. full [CF90]. full-use-of-suitable-spares [CF90]. fully [OK99, SST92]. Function [Lew94, OY91, Tan92, WG94, CS92b, JM96, LLS91, LP99, VR94, WM96]. Functional [PHS99, SS91a, Fri94, FSS99, Gun97, Soh90]. functionality [YA90]. Functions [BKM94, Fow93, KZ90, SS94b, SS99, AEBV90, AEBVZ91, BC99, Bry91, BS90c, BHSB97, DMS95, DGY92a, Dre99, DLN95, FC97, Fal99, HTHR94, HNY93, HM91, JBA^{+}97, Kan96, KEA96, Kun92, LL92b, Lin91, Lu97, MBNSV93, MMM93, Mul94, PI93, Par93, RFB97, San93, Sas91, Sch96, Sch98, TMS97, Weg94b, WG95, YM92, ZV95]. fundamental [HNY93]. Further [JBS90]. Fused [JP98]. FUSS [CF90]. future [HS94d]. Fuzzy [Kos94, SISU90, LL91b].

GaAs [PLM91]. gallium [MFH91]. Galois [PFR98, PFSR99]. Gamma [TCW93]. gate [BE92, CRP^{+}96, HM93, THL92, ZWL98]. gate-level [CRP^{+}96]. gates [BSS94]. gathering [BPRR93]. Gaussians [SWT93]. General [CC92, FTYZ90, LL92b, Lin93b, NNHG93, RCS92, THL92, vdLS93]. Generalization [Vij91, XSFR96]. generalizations [Kor94a]. Generalized [CLL91, Dam92, OPLP93, Par90, Red98, TMS96, WPZZ96, BL98, BL92, BM90, CA90a, Cor94b, LW95a, LW95b, Par93, Pur91, RFKL93, SG90, Sas97, SB92b, SK92b]. generate [WP90]. generated [LIO94, RT96]. Generating [Ann97, CAK90]. Generation [CKC96, CT91a, Lo98, OY91, SM96b, VR94, BBP91, BAA95, BAA98, CT96a, CA91b, CCCC96, DC93, KSL93, LL91a, Lo97b, MBGT91, MAH98, NK97, OVL96, ORPP93, Pha98b, PR94b, PR97b, PR99, RD96, RT93b, Sav99, SS91a, SS91b, SJP96, TS91]. generator [BK95, LS96b, WS98]. generators [CG96b, GP96a, Gol96, GRT96, Pie94, ZVDF96]. Genetic [BM96, CRC96, MYK^{+}99, PR99]. genus [HB94]. geometric [BB92a, LW91, RBG98].

Geometrical [PK93]. geometry [Mie93]. given [Wan97]. Global [RS96, DYD94, SWWW94, Wan97, WY95]. Globally [DP97, Ber96]. goal [LF96]. good [FHA94]. Gordon [Hwa97]. Gossiping [BHS93, Fra92, FY96]. gracefully [CSSP90, LS95, SN99]. grain [BOI94, RSS97]. Graph [OdSC96, BDN96, BB99a, BC99, Bry91, BM96, CYH91, Fuj99, KF97, KTS90, LKL97, LCT97, Mer97, PT98, Pel91, Rao93a, Sch92, SWK90, SR91, Tra96, TS97, Wil93, YLC90, YK98]. graph-based [Mer97, Rao93a]. graph-construction [Wil93]. Graphs [DHL96, Als92, ARS95, BNL93, Bar98, BHS92, BCS94, CL90, DBD93, DT93, DZLH93, GB91, GE92, GS93, HL91a, HL97, HB94, KWF95, MLD93, MNKF90, MPS94, OSST96, PMF96, PB99a, PB99b, TA95]. GRASP [MSS99]. Gray [JH95]. GRD [MYK^{+}99]. Greedy [BB99a]. grids [Ell91, MH90, RBG98]. Group [Bos90, CG97, LY96, MPS94, NM94]. groups [CCD^{+}92]. Growing [SM93]. Guaranteeing [ACZD94]. guard [TLE97]. Guest [BC97, GL99b, KP98, MV99, Tri98]. Guibaly [Wil92a]. guide [SS91a].

H. [PT92]. Haar [FC97]. HaL [SCD^{+}95]. Half [Ros90, Fuj99]. half-duplex [Fuj99]. Half-hot [Ros90]. Hamming [AW93]. Handling [DL94, KCCD99, TD93]. handoff [FCL98]. Hard [VSS91, VS93, AS93, BS99, GPS95, Pol95, SLS95]. Hard-wired [VSS91, VS93]. Hardware [BKM94, EW93, Fow93, KC93b, OY91, RKS90, RPW98, SS94b, SF96b, WG94, BP94, CM92, CF90, CB95c, CPK90, DJ99, DA90, Fer91, FS96, FTG92, HH93, Kan96, Mil91, PR93b, PR94b, RFB97, TY92a, Tak92a, TS98, Uht92, WZXX93, WK97, WSEZ90, XT99, YS96]. Hardware-assisted [RKS90]. Hardware-Based [WG94, CB95c]. Hardware-Oriented [OY91]. hardware-software [DJ99]. Hartley [CJ90]. HARTS [DRS91]. Harvest [SCF96]. hashing [RFB97]. hazard [LKL97]. hazard-free [LKL97]. Hazards [Ung95]. heat [WG98a]. Hensel [ZC90]. heterogeneous [BNG92, DJ99, GHH93, HS94d, IOP99, LL97, Pel98]. Heuristic [BM91]. heuristics [BB99a, Ber93b]. hexagonal [CSK90, DRS91]. HGDM [HKC97]. Hierarchical [DE90, DBD94, DBAP93, DBAP94, HHMC91, Kum91, SSK^{+}96, ZC98, CS91, DA99, DN98, HS94b, Mah94b, PA96, PMHY97, PB99a]. hierarchically [AHT92]. Hierarchies [JCSM96, JCMS97, XT99]. High [AVBZ97, BP98, CL94c, ELM94, KIN^{+}94, PKI^{+}96, PV94a, QF92, Soh93, TT90, TKS94, VT97, ALB98, BAA97, CG96a, CB95c, CKC96, CGC95b, CCCC96, DBAP93, DBAP94, DM97b, GS98, HK90, HLP93, LM99a, LKL98, Man90, MM95b, Pie90, RFB97, RF92, SJ90, SK92a, SF93a, SS91b, Soh90, SRHI98, TH99, Tze93, VP93, WS98, WHL93, XSFR96, ZCT99]. High-bandwidth [Soh93]. high-capacity [CGC95b]. high-level [SS91b]. High-performance [PV94a, TT90, CB95c, DBAP93, DBAP94, HLP93, Soh90, WHL93]. High-Radix [CL94c, SF93a]. High-speed [KIN^{+}94, PKI^{+}96, QF92, TKS94, GS98, Pie90, SRHI98, TH99, WS98, XSFR96]. high-yield [HK90]. Higher [CM90, LM92a, Wal97]. highly [CRC96, FSS99, GSH^{+}92, IO98, NO94, NNHG93, Pol95, SKK^{+}90, XK92]. highly-parallel [IO98]. HIPPI [CGK93]. history [KE97, RW98]. history-based [KE97]. hit [TSW92]. hit-ratios [TSW92]. homing [PR94a]. homogeneous [FE97, RKS90]. Hopfield [OPLP93]. horizontal [MS91]. Hot [DE92, Ros90]. Hot-spot [DE92]. Householder [HD95]. Hybrid [PK94, WPS97, AK96, BP95, CM96a, CZSM99, Dae91, Hla92, Kan93c, KK91c, LW91, LL90, NT92, PA96, SD96, SF93b]. Hyeti [LKK^{+}94]. hyper [Lo96]. hyperbolic [HTHR94]. hyperbus [CLL91]. Hypercube [Ish95, KD94, MBR96, AC95, BRS^{+}90, BB92a, CB95b, CS96a, CS90b, CS90c, CLL91, CW96, CC97b, CT92, DE92, Dow92, DH91, Efe91, FBL96, Fra92, HS94a, HS93, IPS91, LS91a, LAY96, LH92, LK96, MBR95, Mat90, Qui90, SC95a, TS94, VD92, WH97b]. hypercube-connected [CS96a, SC95a]. hypercubes [BP94, BCGS96, BCS92, BCH93a, BCS94, BCH95, CB95a, CT94b, CT97a, CT97b, DMS94, GS93, GP97, HM95, KK91a, KB93, KS93a, Kus90, LK99, MPS94, PB97a, RYT95, RS96, SS96, THY93, TNC95, TW91, TC94, TL96, VB96, Wan94b, Wan99, WY95, Wu95, Wu97, WG98b, YTR94]. Hypergraph [FS94b, GC97]. Hyperneural [RD96]. HYTREM [LL90].

I/O [GP98, NLV91]. IADM [RFS92]. IC [SK96]. Ichimori [PT92]. IDDQ [CT96a]. identification [LY94, RS94]. Identifying [Tok94]. IEEE [Ano98b, Tri98, Ano96a, Ano98a, ABCW92, DL97, Tan92]. II [ACH^{+}97b, ABK90]. ILAs [SF96a, Ver92]. image [AS92b, CSS90, Esh91]. images [Tat97]. immediately [Oik92]. immediately-detectable [Oik92]. Impact [Vai97, PRASA99]. Implementation [BB97, BB92c, DMS95, DM96, DM93, RW98, Tan92, AEBVZ91, Ann97, BB99b, CT93a, CPK90, DGY92a, EW93, GW98, HN93, HC96, KM94, LS95, LHL99, MS99, Nic98, Par93, PHS99, Qui90, SJ90, SD96, WZXX93, Zar95]. Implementations [OF97b, Bry91, HH93, JLLR94, RF91, SG90, SK92a, TS98]. implemented [Pie93, SRHI98]. Implementing [BES^{+}96, DGY92b, Vin95, Lin90]. implications [PH93]. implicit [OCVSV98]. importance [CLM^{+}95]. imprecise [BH98, CLL90, SL95]. improve [SV97]. Improved [Ano96b, HL92a, NO91, RT91b, RR93, Sta93, YL95b, YL95a, CPK90, HR95a, JS92, KKKM99, LT93, PNDN99, SL91b, SW97, Tze93]. improvement [KB96]. Improvements [GDY90, TK97]. improves [BS97]. Improving [BW96b, DRT97, HY99, KE97, KRC99, Ng91, TSW92, TS99b, CK99, Dal91, Ish95, MB93, PG97]. including [HTHR94]. incomplete [HY99, THY93, TC94, TL96]. incompletely [ZV95]. Incorporation [HS94c]. Increased [HDS91, Har92]. Increasing [DKT96, PK91]. INCREDYBLE [PR95b]. Incremental [KTS90, KM97b, IIK^{+}91]. independence [RS98]. Independent [SF92, Gre94, RF99, vdBT93]. Index [Ano96a, Ano98a, Ano98d, Ano99a, Ano99b, SA98]. Indexed [JBA^{+}97]. Indirect [LY91, BB97]. induced [CRP^{+}96]. inference [NO94]. InfoPad [TPDB98]. Information [BH94, KT91b, LRG91, TRS^{+}99, TPDB98, WY95]. inherent [Tho97]. inheritance [SRL90]. inhibitor [SV95]. initial [EL90b, ITY97]. Initializability [CA92]. initialization [SSDK95]. injection [ACC^{+}93, BCSS90, CPA99, KKA95, SRHI98, THZ^{+}99, YS96]. injured [KS93a, TYJ99]. inline [HC92]. inner [LO94a, WJ90, WMD96]. Innovative [BSS94]. input [ESS98, FFKI95, KPI93, LIO94, MI93, PGK90, SISU90, Sas91, TH99, YM92]. inputs [PR93b]. insecure [Mit92]. insensitive [AVR98, BB95, LH96, VBH^{+}95]. inserted [LM99b]. insertion [FL98, HC92, PK91]. inside-out [KYM97]. instantaneous [SI94]. Instruction [DAF94, Soh90, ACFH95, ASWW92, CL96, CCCH93, HAM^{+}99, HS98c, HC92, IS94, KKKM99, KS96a, Kuh98, LCFH95, PRASA99, SS94a, SAMC99, TXD98, Uht92, WW91]. instruction-level [CL96, PRASA99, SS94a]. instruction-retry [KS96a]. instruction-window [SAMC99]. instructions [CKY^{+}99, LP90]. Integer [Has90, HK95, Bry91, CC92]. Integers [Zur94, Fag92]. integrate [SS94c]. integrated [CL98b, CMD93, KS91, RFM90, SS94a, TSH94, TSH95]. integration [IJN94, TH91]. intelligent [Leu97a]. Interactive [PT98]. interchangers [JLLR94]. interconnect [HD98, SSL96]. interconnected [BS93b, SA92]. interconnection [AMB96, Ber96, BAA97, BS90a, BCLA91, CS91, CGK93, CDHK93, Dal90, Dal91, DE90, DBD93, DBD94, DR96, DB94, Dou93, Dow92, DBAP93, DBAP94, FR94, FL92, FTL93, GDYS96, GB91, Ho92, HSY97, KEA96, LY90, LYW94, LK95, LFH92, MvdS90, MJ91, MYD96, MOM91, MY94a, NB93, NL90, OSST96, SXW95, TCW93, Wil92b, WL92, Yan98, YMG99]. interconnections [KKL90, KEA95]. interface [RNA93]. interfaces [Nic98]. interference [RH93]. Interleaved [Lew94, CSSP90, CL99b, LKW97, RH93, Soh93]. interleaving [KT91a]. Interlock [VPB93, PV94a, SSK^{+}96]. interlocks [McC91]. intermediate [NC96, WI95]. Intermittent [BSM92b, AAL94, IYI90]. internal [AR94, ESS98]. internal-XOR [ESS98]. internally [VT97]. International [Tri98]. Internet [Mit92]. interpolation [PG92, RBG98]. Interpolators [Lew94]. interprocessor [MS99]. Interrupt [TD93]. interruptible [Soh90]. interrupts [SE98]. Interval [RS95, TL99]. intervals [RHT95]. intractability [HV91]. intrainverted [GK92]. Introducing [Yan93]. Introduction [BC97, KP98, MV99, Tri98, GL99b, KR90a]. invalidation [GW92, Rag93, Rag96]. invariant [LVB^{+}93]. inverse [FC97, HTHR94, Kal95]. inverses [BCH93b, WL93]. inversion [GW98, Has98]. inverted [LS96b]. investigations [RCS92]. IRIS [DKT96]. ISO [Mei91]. isomorphic [Pea96]. Isomorphism [DBD93]. Issue [Ano97b, BSMC98, Soh90]. Issues [OF97a, CU93a, DH92, IS90, Rao93a, SJ95]. iteration [FL93, HL91a, MC93, MC95]. Iterations [DM96, DHL96, WF92, DZLH93, PMF96]. iterative [CA91b, Fri94, JK92, PM91, RCBB96, SW94a, WC90].

J. [JBS90]. Jack [CV92]. Jacobi [GPS93, PGS95]. Jacobi-like [GPS93]. jitter [Kle90]. job [BHR94, BK90, Bon90, KT92]. jobs [CLL90, LK98]. jog [Tua92]. join [LP91]. joining [LCS93]. jukebox [DKL90]. jump [MS91].

K-term [WS95]. Kalman [Red98]. Karnaugh [OM91]. Kautz [DHL96, PMF96, DZLH93, PS99]. key [BMP97a, LG97, LC90, LCL95, Lu97, PFSR99]. keystream [GP96a, Gol96]. Kleenean [HNY93]. knapsack [Fer91]. knee [BS90b, SWWW94]. knock [BS90b, SWWW94]. knock-knee [BS90b, SWWW94]. knowledge [WSB90]. knowledge-base [WSB90]. Kronecker [Dre99]. KYKLOS [MJ91].

L. [CLL91]. L.0 [CCG^{+}91]. labeling [BB92a, CSS90, SY91a]. LAN [MTG93]. language [CCG^{+}91, CMD93, ST95a]. Large [SS90b, WJR94, Zur94, ABK90, BDL^{+}99, CL90, CPK90, Fag92, FL92, FW93, Fit97, HS93, KT92, KF92, LHNP94, LL90, LY91, MBNSV93, OS94, PK90, Sem95, SP96, TRM95, Tze92]. large-scale [ABK90, CL90, FW93, Tze92]. Latch [CBRS90]. Latch-to-latch [CBRS90]. Latency [THH92, GS98, KHC94, Ng91, TK98, Vai97]. lattice [NLV91]. laxity [NT92]. laxity/first [NT92]. laxity/first-in-first-out [NT92]. layer [LYW94]. layout [BS90b, Tua92]. layouts [FE97, SWWW94]. Lazy [MM97b, DLN95]. LC [HSL96]. Leading [BL99]. Leading-one [BL99]. Learning [CGTN92, CRMM98, OdSC96, PK93, TCI98, WZXX93]. Least [HM94]. Lee [BBKA95]. left [TH99]. left-to-right [TH99]. Length [Yok92, ABC92, AC98, BBP^{+}96, IS94, Lew90, MV95, PRS94]. lengths [UWZ97]. level [BB92c, BOH90, CRP^{+}96, CL96, CCBC98, CU93a, CU93b, CU94, DN98, FBL96, FRK98, GP98, Gos97, Kan93a, Kan93b, LKW97, MBNSV93, Mel92, PRASA99, Pel91, RDZ95, SS94a, SS91b, Som90, VP94, Vai98, Wan94a, vdBT93]. leveled [BBP^{+}96]. levels [KSS95, Tem99, WRS95, Wu95, Wu97, YF92a]. Lewis [ABC92]. LFSR [ESS98, RT96]. LFSR-generated [RT96]. library [PDBH97]. lightwave [PS99]. like [GPS93, Sch92, WNPM95]. likelihoods [SK94]. limited [CDP98, EZ92, KSS95, WY95]. limited-global-information-based [WY95]. Line [DZLH93, DHL96, HC94a, PMF96, BH98, BB99b, CT97a, DKT96, EL90c, FL92, GP96b, HL92b, Kip94, LS90b, LT95, Mul94, SY91a, SSL96, ST96, ZB97]. Linear [CFV91, Gol96, Kor94b, Ber93a, CCD^{+}92, CM96a, CT98, Cd93, DOR91, DD95, GK92, Har92, HRT^{+}95, Hla92, HTCT91, IJW92, Kop95, KT91c, LLK92, LK90, MRB90, MNKF90, OSST96, RT96, RBG98, SWT93, SM92b, SF91, ST95b, SR96, TC95, VD92, Wan96]. Linear-Array [Kor94b]. linearization [CR97]. Linearizer [dM90]. linearly [RF99, Zar95]. lines [AR94, MTP93]. link [AAL94, BM91, KT99a, LVB^{+}93, SSS97]. linked [Han91]. links [DRT97, LY90, SY91b]. List [Ano97a, Ano97b, Ano98c, Han91, SJP96]. live [MTP93]. LNS [DMS95, Lew94]. Load [CS96a, Dae91, HS94d, LK98, BHR94, BK90, CS96c, HS94c, HS93, KK92, KLL97, KM97b, NO94, SC95a]. load-distributing [CS96c]. local [DTAL93, FL93, Han91, KMM91, Wan97]. Locality [GAFN96, MV99, CR93, KRC99, Tem99, TLH94]. locally [LaF99]. Location [LS90c, PR98, Ber96, LFH92, PR97a, RCB96, SC95a, SJ93a, TSS95]. lock [Joh95]. lock-free [Joh95]. Locking [SL90a, SRSC91]. log [SE98]. log-based [SE98]. logarithm [ABCW92]. logarithmic [ABCC90, ABC92, ABCW98, Lew90, Man95, MST98]. logarithmic-exponential [Man95]. Logarithms [Kos91, CW94]. Logic [CL98b, Bar92, BAA98, CA91b, DA90, FW93, FHA94, Fri94, HNY93, HLHC91, IO91, JBS90, KPB92, LL91b, LH96, PR96b, PR97a, RB99, Sav98a, SB97, Soh90, SW94a, TCI98, TB91, VCB96, WC90, XK92, YM92, ZWL98]. Logically [SI94]. Long [STE98, ABC92, Lew90]. look [GL99a, Tak98]. look-up [GL99a, Tak98]. lookahead [CSTO92, DO96, LS92]. lookahead/carry [Kan93c]. Loop [SC90a, HKL98, HW95, KRC99, MS98a, MS95, PNDN99, PT92]. Loop-Free [SC90a]. loops [FTYZ90, HKL98, SMC91]. lossless [Tat97]. LOTOS [SCV91, SV95, TS91]. Low [CW99a, KS98, NL99, THH92, TSS95, WH98, AZ98, BHBA91, CGC95b, DJ99, EZ92, HWB92, HB95, Paa96, PB97b, Pie90, SHHJ93, TW95c, WHB98]. Low-complexity [KS98, TW95c, WHB98]. Low-cost [CW99a, CGC95b, Pie90]. Low-Power [NL99]. lower [AR94, AM93, BOH90, Efe91, KT99a, LH98, WNS96]. LRU [vdBT93].

M [ABC92]. MAC [Mit92]. Machine [Wel94, CA92, DRW92, LL90, NO94, ZPP91, EL90e]. Machines [MBR96, CCC96, DOR91, DGY92b, HB98, IJW92, LY94, LS96b, LP90, LPP91, MBR95, Oik92, PR97b, Ros90, TW95a, Tak92b, Zhe94].

Mactaggart [CV92]. mail [Mit92]. mainframe [SIKS95, SSK^{+}96]. Maintaining [CL98a]. maintenance [GPTT98, dG90]. majorization [NST96]. make [BZ92, Lam97]. making [CK90a]. Management [AS95, BP95, DYD94, FNS95, LAY96, SCD^{+}95, TRS^{+}99]. Managing [CR93, Kum91]. Manchester [CS90a]. Manetho [EZ92]. Manhattan [CSA94]. manifestation [CS92b]. Manipulating [Lin93b]. manipulation [GM94]. Mantissa [DA96]. Mantissa-preserving [DA96]. mapped [SV97]. Mapping [MBR96, SV95, KB93, KT99a, LPP91, MBR95, MY96, MYD96, NO91, SA98, WS91]. maps [OM91]. Markov [BT90, CMST90, JG99, RS95, Sem95]. Markovian [GSH^{+}92, OI90a]. Martin [TLE97]. mask [IN96]. masking [TB98b]. mass [PMHY97]. Massey [HWB93, Zar95]. Massively [MBR96, AMB96, MBR95, SN99, SKA92]. massively-parallel [SKA92]. master [GW91, Sah96]. master-slave [Sah96]. master/slave [GW91]. Mastrovito [SK99]. Masuyama [PT92]. match [SL91c]. matching [IS90, SR98]. mathematical [Cor96]. matrices [Cor94b, SV94, WC99]. matrix [ABK90, DA96, EL90c, KT91c, LLS92, LO94a, LM92c, NA90, RCB96, TY92b, WM91]. maxcut [CRS98]. Maximal [CW94]. Maximizing [FR97b, KM97a, SWG92]. Maximum [EAT95, FFKI95, HV91, RBG98, SISU90, SWK90, SL95, TL96]. maximum-speed [RBG98]. maze [WSEZ90]. mean [CS96b, FR97b, LMT95]. mean-time [FR97b]. means [CL94b, LZM90]. measure [HNG94, NG90]. measurement [KM92a, TI93]. measurements [DRW92]. measures [BT90, KTS90, LHNP94, OZL96, Rai95, WG98b]. Measuring [DM94, GS98, SS95]. mechanically [MLK98]. mechanism [AA92, FS96, MMSS96, PFP99, Wu95]. mechanisms [TK91]. medium [ACZD94]. meet [RKJZ99]. Meets [CL97]. megabyte [KHW94]. membership [HS98b]. memorie [TWS93]. memories [CS99, CFV91, Mah94a, Mon91, PHS99, RH93, SMV91, Soh93, Tak92b, TWS92, VLPA95, WZXX93, WSY97]. Memory [JCSM96, KHC94, Lew94, MV99, AMS96, BGJ^{+}96, BP93a, BP95, CG96a, CB99, Che92, CS96b, CSSP90, CGC95b, CL99b, DDS98, DA99, Dek98, DDV90, FL93, FS96, Fri95, GW92, HF97, Han91, Har92, HC93, HSHG99, HS94b, JCMS97, Kat97, LL96b, LKW97, Maz93, MTP93, MNL99, PRASA99, PD99, PG97, PBND98, RPW98, SS93a, Sav98a, SCD^{+}95, SL92, Sta93, STW92, Tem99, Tze92, Tze94, Vai96, Vin95, Wan96, WF90, WDS^{+}95, XT99, XSFR96]. merging [GL91]. Mesh [Raj96, BC95, CSK90, CSS90, DDGS97, DRS91, FY96, Mat90, MD95, TP94, TNC95, WH94]. mesh-bus [FY96]. mesh-connected [CSS90, Mat90, MD95, WH94]. meshes [BP94, BNP91, BCH93a, BCH95, CC94, DDS99, MPKRS93, Pea96, SLH97, SS96, SW97, ST96, TYJ99, UVM97, YTR94, MM95a]. Message [TS95, ACZD94, Ber93b, SI94, WHK97, WSB90]. Message-optimal [TS95]. message-oriented [WSB90]. message-passing [WHK97]. messages [BPRR93, BBP^{+}96, HHS97, Mit92, ZSR90]. MetaNet [OYY97]. metaobject [FP98]. metastability [Kle90, Ung95, VBH^{+}95]. Method [KTM^{+}94, SF96b, CA90b, CAK90, CB96, Dal92, DD95, HTHR94, HN93, LK90, LT93, Man90, OVL96, Pur91, SIKS95, SGP92, SL91b, SV97, ST96, TLE97, TSH94, Vog90, ZPP91]. Methodologies [HD98, OFW92]. methodology [GB90b, HLB^{+}97, MM95b, TLE97, YS96]. Methods [TAY91, CHH98, CPA99, FM91, HS98c, TSH95, Uht92]. Micro [QSA91, BOI94, TT90]. micro-grain [BOI94]. microarchitecture [RBS99]. microcode [PHK94]. microcontroller [Kuh98]. microprocessor [AS92b, LKK^{+}94, MFH91, WDF^{+}97]. microprocessor-based [AS92b]. microprocessors [Cor94a, Win95]. microprogrammable [TKS94]. microsequencing [PI93]. migration [CS90c, CT97a, DL97, PI93]. Min [Tra96, Vij91, SSA94]. Min-cut [Tra96, Vij91]. mini [KMD97]. mini-slot [KMD97]. Minimal [SC90a, AW93, BCH93a, DD95, LV99, LL96a, Lu97, Pie93, SM96b, Tok94, Uht91]. Minimization [AMS96, BS93a, TB91, Bar92, DTB96, HM91, HLHC91, LM92c, MBNSV93, MNKF90, OCVSV98, Sar90, SL91a, WG98a]. minimize [KS96b, SL95]. Minimizing [BL92, CL99b, WSY97, UWZ97]. Minimum [FL91, GC97, PM98, BDL^{+}99, JO93, LMT95, NT92, SISU90, Sas91, SWK90, THL92]. Minimum-congestion [GC97]. MISR [Sav96]. miss [BSMC98, SST92, vdBT93]. miss-path [BSMC98]. miss-rate [SST92]. mixed [ABZ96, WPZZ96]. Mobile [Ano97b, CL97, JTK97, AV97, FCL98, LL97, LS97, MS98b]. Mobility [CL97]. Möbius [CL95]. mod [SB90]. mod- [SB90]. mode [KIN^{+}94, PKI^{+}96, SISU90, SSA94]. Model [JCSM96, AD97, BR93, BP93a, BB99b, CA91b, CS91, CR93, CMD93, Cor96, Egg91, Fuj99, GY95a, GLR^{+}99, GYY98, JCMS97, JL96, Kle90, LYW94, Mer97, MB93, OK99, PMW92, PZ99, RD96, RCS92, RFS92, SN99, SST92, SR93, SF93b, TW95a, TS99b, Wan99, WNPM95, vdBT93]. model-based [TS99b]. modeled [LG95]. Modeling [CU94, DF91, MTP93, AKL90, BR90, CK90a, CDFH93, EANPH93, HKN99, LBF92, MY95, MY96, MY94b, PMHY97, PLM91, TI92, TI93, TSH95, Win95, ZVN92]. modelling [DYD94]. models [AK96, BR94, BP92a, BSM92a, CWM^{+}95, FTL93, GP98, GS95, Gol96, GSH^{+}92, GL92, KK91b, LPP91, OI90a, PS90, Pel91, Pel92, Sta93, VRK93b, VRK93a, VT94]. modification [Hwa97, ITY97, Tak98]. modified [BB92b, HWB93, RK96]. Modular [HWB92, PB97b, SJ90, TY92a, VBH^{+}95, BDK98, CT98, CW99a, EW93, HTYP97, Pie94, SY91b, Tak92a, Wal93, Wal97]. Module [LVAG98, AS93, FBM93, HS96, LS90c, PK91]. modules [CYH91, HS97, Sar90, TKS94, TS94]. moduli [CN99]. modulo [AM91, LM92c, Ma98, SR92]. modulo- [LM92c]. modulus [WJ90]. moment [CT93a]. moment-preserving [CT93a]. monitoring [SS94a, UR94, Wil93]. monotonicity [Sch96]. Montgomery [BDK98, EW93, Kal95]. Moore [Kno96]. most [HW97]. moving [Ara90, CY95, FL91]. MP [XT99]. MPS [BSMC98]. Muller [DTB96, ORPP93, Pha98a, Pur91, Sas97, TMS96, TMS97, VR94, WPZZ96, ZV95]. Multi [NO94, DDGS97, KHW94, TF92]. Multi-ASCA [NO94]. multi-megabyte [KHW94]. multi-mesh [DDGS97]. multi-queue [TF92]. multiaccess [Har94, Wei96]. Multibit [RNA93, SG90]. Multibuffered [YLL90]. multibutterflies [LM92b]. multibutterfly [GMP94]. multicast [ASS94, Mit92, TYJ99, YM99]. multicast-connected [ASS94]. multicasting [WY95, Yan98]. multicasts [TS95]. multiclass [LK98, MTG93]. multicomputer [ALPS90, BB99b, DE90, DRS91, KS92, KK91c, MJ91, Qui90, RHS98, TI92, TI93]. multicomputers [BM91, CB97, CS96a, CS90b, CW96, CC97b, Fra92, KLL97, SC95a, SZN92, TS94]. multidimensional [CJ91, CSTO92, MS99, WPS96]. multihop [HS98a]. multilayer [BS90b, PK93]. Multilevel [OMB97, Mah94b]. multimedia [GYY98, TPDB98]. multimode [PT92]. multimodule [VLPA95]. multioperand [Pie94]. multioutput [Pie93]. Multiphase [Bok96, LT93]. Multiple [GM98a, JW94, KLPS93, LL96a, San93, AA92, ACFH95, AE94, BSMC98, BHSB97, CA90a, CS91, CU93a, CU93b, CT94b, CU94, CF94, DP99, Dam92, DDS99, Dav90, Dow92, HNY93, HRT^{+}95, HJ92a, HS96, JS92, JL96, KPI93, KACV94, KC93a, KIN^{+}94, KMD97, KT92, KEA96, LaF99, LS96b, LCFH95, LFH92, MM91a, MMM96, MI93, NGB93, NM94, PKI^{+}96, PR92, PR97b, PGK90, RTZ98, SISU90, Sas91, SM92b, SF96a, SB92a, SB97, Soh90, TCI98, TB91, TS97, VRK93b, VRK93a, Ver92, Wil92b, WI95, YB91, ZV95]. multiple-adjacency [CT94b]. multiple-bus [JL96, YB91]. multiple-input [SISU90]. multiple-issue [BSMC98]. multiple-level [CU93a, CU93b, CU94]. multiple-module [HS96]. multiple-path [DP99]. multiple-polynomial [HRT^{+}95]. multiple-sequence [LS96b]. multiple-track [VRK93b, VRK93a]. multiple-value [Sas91]. Multiple-valued [GM98a, KLPS93, BHSB97, HNY93, KIN^{+}94, PKI^{+}96, SB92a, SB97, TCI98, TB91, ZV95]. Multiple-way [San93]. multiplexed [YMG99]. Multiplexer [Pek99]. Multiplexer-based [Pek99]. Multiplication [CV92, BDK98, Bry91, CM96b, DS94, EW93, EL90a, FBT96, Fio99, ITY97, KT91c, Lin90, Ma98, TY92a, Tak92a, Tak98, TY92b, Vin95, Wal93, Wal97, ZC90]. multiplicative [BCH93b, Has98, ITY97]. Multiplier [Kor94b, AM91, HB92, HWB93, Hon90, KIN^{+}94, Lim92, MJ92, Paa96, PFR98, PKI^{+}96, SG90, SK99]. Multipliers [Kat94, LTR96, WG94, YGI91, ATF98, GPZ99, HWB92, IV94, KS98, LTR93, Lu97, MRT97, OVL96, OI90b, Pek99, SS92, SR92, SMOR98, VSS91, VS93, WJM95, WH98, WHB98]. Multiply [JP98]. Multiply-Add [JP98]. Multiplying [Zur94]. multiprocess [Lam97]. multiprocessing [ENS91]. Multiprocessor [DTAL93, YLL90, ABK90, BRS^{+}90, BD95, BB92a, BSM92a, BSM92b, BP93b, BLOS95, CPK90, DA99, FR94, JS92, JL96, JW91, KRM95, KM92b, KT99b, Lam97, LHNP94, LY91, LS95, NHA92, OI90a, QM93, RCS92, RR95, RSD^{+}98, RCB96, RS95, Sah96, SP91, SD92, SJ93a, TLH94, TSS95, YB91, vdLS93]. multiprocessors [BB90, BGJ^{+}96, CB95b, CB99, CSK90, CS90c, CS92a, DDS98, DH92, DM97b, FBL96, Fri95, GW91, GW92, HS94b, Ish95, KH97, Mah94b, MNL99, MDyF94, NT92, PRASA99, SLD95, Tze92, Tze94]. Multirate [LR97]. multirendezvous [SCV91]. multiserver [BK90]. multispectral [Tat97]. multistage [Ber96, BAA97, BS90a, CDHK93, DBD93, DBD94, DB94, FTL93, KK91b, LY90, MYD96, MOM91, MY94a, NB93, SXW95, SLD95, ST96, WL92]. multistate [VT94]. multistream [McC95]. multistreaming [McC91]. Multithreading [Gun97, KCCD99, KT99b, KLH^{+}99, TK98]. multivalued [DS94, MBNSV93]. multiversion [Leu97b]. mutation [GOH92]. mutual [HY99]. MVL [AEBVZ91]. MVMT [AEBV90].

N [CLL91, DSPN95]. natural [CMD93]. Near [MPS94, BDL^{+}99, CW94, KS96b]. near-coincident [KS96b]. near-maximal [CW94]. near-minimum [BDL^{+}99]. nearest [CLC92, Mat90]. nearest-neighbor [Mat90]. nearly [MY96]. Necessary [WdB92, Sch96, YM99]. Negation [WM96]. negative [GM98a, KKS93]. negative-binomial [KKS93]. neighbor [Mat90]. neighbors [CLC92]. Nest [LMJ93]. nested [FTYZ90, LMJ93]. nested-predicate [LMJ93]. net [GYY98]. nets [CDFH93, CL94b, CGTN92, CKLY98, LG92, LW95a, LW95b, LZM90, Mer97, SV95]. Network [CK90b, HNG94, NG90, AZ98, AMS96, BHS93, BR94, BJK93, BS90a, CF95a, CLL91, CT93a, CC96b, CS96c, CT91b, DP99, DDS99, DDGS97, DR96, FA96, FTL93, GMP94, HH93, JO93, KTS90, LY90, LPP91, LL91b, LO94a, LCT97, MvdS90, MJ91, MB93, MYK^{+}99, OI90a, OPLP93, PS99, PWG94, PZ93, RD96, RDZ95, RFS92, Saa95, SY91a, SP91, San93, SBP90, SBP91, SJB91, SLD95, SRHI98, SF93b, SW94b, TCI98, TW95c, WH94, WNPM95, WL92, CSA94, MM95a]. network-based [SLD95]. Networks [YLL90, ASS94, AAL94, Aln93, AHT92, ANR96, Bae97, BH92, BHR94, BHBA91, BJ91, Ber96, BCGS96, BP98, BPRR93, BBP^{+}96, BCLR97, BAA97, BC95, BS93b, BE92, BCLA91, CF95b, CCD^{+}92, CDP98, CDHT99, CRC96, CS91, CC97a, Coh93, CDHK93, CG94, Dal90, Dal91, DE90, DE92, DBD93, DBD94, DB94, DW97, DL97, Dou93, EAK98, FCL98, FS94a, FE97, FR94, FL92, Fio99, FTL93, GDYS96, GS97, GB91, GS98, GE92, GM98b, GH91, HR95a, HHS97, HS98a, HY99, HS94a, Hei91, HT95, Ho92, Hou99, HSL96, HSY97, HTCT91, HW95, JL91, KH93, KS92, KT92, KM94, LNS94, LW94, LS96a, LP98, LY91, LLK92, LL94, LM92b, Leu97a, LW95a, LW95b, LK95, LL97, LO97a, LS96c, LLJ97, LFH92, Mah94b, MS98a, MM97a, MMM96, MB95, MS98b, MYD96, MOM91, MS95, MY94a]. networks [MY94b, NB93, NL90, O’C95, OYR95, OSST96, PK93, PA96, PT92, Pie93, RB91, RKJZ99, RHS98, RF91, RB93, RB97, SM93, SC94, SSS97, SXW95, She95a, She95b, SR91, SCG94, Sut91, Szy97, TCP91, TP94, TK96, TCW93, VC92, WM91, WJ94, Wil92b, WS91, YM91, YM95, Yan98, YM99, YW99, YF92b, YA90, YMG99, dM90]. networks-architecture [GE92]. Neural [LL91b, BJ91, BP98, BJK93, CT93a, CFV91, FTL93, HH93, LPP91, MYK^{+}99, OPLP93, PWG94, PK93, PZ93, RF91, SM93, SY91a, SJ95, SRK91, SF93b, TCP91, TH91, TK96, TW95c, ZVDF96]. Neural-network-based [LL91b]. neurocomputer [CRMM98]. neuron [WZXX93]. neuron-weighted [WZXX93]. Newton [KTM^{+}94]. NFS [DRT97]. nibble [Joh93]. no [UR94]. Node [DM97b, CS96a, PK90, SES^{+}92, TK98]. Node-covering [DM97b]. nodes [BES^{+}96, BHSB97, LW94]. Noise [EANPH93, CC96a, CFV91]. nomadic [ZR97]. non [HR95a]. non-real-time [HR95a]. nonbinary [PZ93]. Nonblocking [YM91, DFM95, PLJ94, Szy97, YM99, YW99]. noncoherent [UP93]. nonconvex [CB97, SWT93]. noncrossing [Yan97]. nongroup [CCC96]. nonhomogeneous [BK90]. nonincreasing [Tao97]. nonlinear [CR97, Kop95, Lee92, LS96b]. nonminimal [LK95]. Nonprime [Kat97]. nonrestoring [SF93a]. nonrobust [FFS92]. nonuniform [KT91b]. nonvolatile [VJ98]. norm [DD95]. normal [KS98, Lu97]. note [Bar98, BBT94, CLL91, Fal99, Kat96a, Kno96, LK90, MLD93, MI93, SAR91, THL91, dM90]. Notes [Hua90, KPI93]. nothing [HS93]. notification [SBM^{+}98]. Novel [RFKL93, HC96, LC92, NSUR98, PV94b]. NP [AS93, BW96b, RT91a, ZPP91]. NP-complete [BW96b, ZPP91]. NP-hard [AS93]. NUMA [ZCT99]. Number [DJM99, Par90, PK94, WF92, Yok92, AM93, AK91, ABC92, ABCW98, BHSB97, DIP93, HK95, KS93b, LW91, Lew90, LH98, LC92, MST98, NA90, OPT92, Par93, PR96b, Sas91, SC90b, SL90b, SW94a, TSH94, WP90]. Numbers [Has90, Vui94, ABC92, BCH93a, DHK93, Lew90, SG90]. Numerical [DL94, KZ90, KC93b, OZL96].

O [GP98, NLV91]. OBDD [GM94, HM94, LL92b, WeBM97, Weg94b]. OBDD-representation [LL92b]. OBDDs [BW96b]. object [CG96a, GHH93, LVB^{+}93, Vog90]. object-oriented [CG96a, Vog90]. objects [ACH^{+}97a, ACH^{+}97b, BG90, Joh95, SBM^{+}98]. observability [BP96, CCCC96, Fuj90]. observable [XK92]. observation [Ho92, PR92]. Observations [CS92b]. obstacles [LYW94]. occupancy [FCL98]. occurrence [Som90]. odd [GB91]. OFDD [DTB96]. OFDD-based [DTB96]. off [PV97, ST96]. off-line [ST96]. office [AS92b, DKL90]. offs [BOI94, CS96b, CSA94, DAF94, SAMC99, Wal97]. offsets [MBNSV93]. OKFDDs [DBJ98]. OLAP [HBA98]. OMC [Esh91]. Omega [She95b]. Omura [HWB93]. on-chip [Maz93, NV99, WSY97]. On-Line [HC94a, CT97a, DKT96, HL92b, LS90b, BH98, BB99b, EL90c, GP96b, Mul94, ZB97]. On-the-fly [EL92]. once [BBRL98, Gai90, Weg94b]. One [Coh93, KH93, BL99, CM96a, FL98, IJW92, Kan93b, KK91a, KSGT96, VC92, Ver92]. One-bit [Coh93, KH93]. one-dimensional [CM96a, FL98, Ver92]. one-level [Kan93b]. one-sided [VC92]. one-step [KK91a]. one-way [IJW92]. ones [Kop95]. online [BB92c, NHA92]. only [Vin95]. onto [ARS95, LPP91, SV95, YTR94]. OODBMS [ZC98]. open [PW95, TS99b]. operand [BW95, ITY97, RB99, Tak98]. Operands [Lo94b, EL90d]. operating [SSA94, TXD98, XT99]. Operation [SF96b, Kan90]. Operations [OF97a, DA96, LLS91, NA90, RS96, SR96, TB99b, VCB96, WL92]. operative [Rao93a]. operator [TLE97]. Operators [CV92]. OPSILA [BA93]. optical [BBRL98, CDP98, DKL90, Gai90, GM98b, JLLR94, LLJ97, QM93, YMG99]. Optimal [ACH^{+}97a, ACH^{+}97b, AK91, AC95, BHBA91, BBB^{+}94, BS99, CMH98, CSF90, CC97a, Cor94b, FRK98, HKC97, KN90, KEA96, LK91a, LLK92, LS93, LP90, LCS93, LL97, MS98a, MM91a, Nai95, NGB93, NR90, Nik98, Pel92, Pel98, She95b, SJ93a, SMOR98, STW92, WPS96, Aln93, Als92, BNP91, BCLR97, BK90, CK94, CF94, DBD93, DP97, EAK98, Fra92, FS90, GL91, Han91, HS94c, HS96, KK92, KS96a, KEA95, KT91c, LK91b, LS94a, LS98, Lo96, PM91, SS90a, SM92a, SF91, SXW95, She95a, TC95, TS97, TNS99, Tua92, TS95, WL90, Wan96, WNS96, Weg94b, WT90, WH94, Wil93, Yan90, ZS99]. Optimally [HSL96, SJB91, Tem99]. Optimistic [BZ92]. Optimization [TCP91, WD92, AE94, CSTO92, CVMS99, DF91, Fis90, GGD96, OMB97, PHK94, PS93, PR99, SWT93, SK96, SF93b, TK96, YK98]. optimizations [CCCH93]. Optimized [LS90a, OVL96]. Optimizing [TXD98, BS93b, GS98, Maj96]. optimum [Kan93b, PM91, TSH94]. options [PFP99]. optoelectronic [LO97a]. OR-parallel [NO94]. oracles [DLN95]. Order [SC90a, PD99, PB97b, Tao97, TD93]. Ordered [LV99, FC97]. ordering [AV97, BW96b, CRS98, DBJ98, FS90, Tat97]. organization [KKL90]. organized [HF97]. orientation [CYH91]. Oriented [OY91, CG96a, CS92a, LG97, Vog90, WSB90]. orthogonal [MLD93, PT98, SBP90, SBP91]. Other [OF97a]. out-of-order [TD93]. output [Dav90, EZ92, KSGT96, Oik92, Sas91]. outputs [Tho97]. Over-redundant [MC94, SPM97]. Overflow [Yok92]. Overflow/Underflow [Yok92]. Overflow/Underflow-Free [Yok92]. overhead [DJ99, Egg91, EZ92, NLV91, SHHJ93, TSS95, Vai97]. overlapping [Wil92b]. overlaying [Ber93a]. overload [BH97]. Overturned [MJ92]. Overturned-stair [MJ92]. oxide [HM93].

P [CLL91]. Packet [YLL90, CG94, Raj96, ST96, YB91]. packet-switched [YB91]. Packet-Switching [YLL90]. packing [DLC93, YW99, ZPP91]. page [Gai90, ZC98]. pages [HW97]. paging [Dal92]. pair [Yan97]. pairwise [GP97]. paper [Pha98a]. Papers [Ano98b]. paradigm [BBD97, LH96]. Parallel [BB92a, DD95, Esh91, Fio99, Hla92, IPS91, MBR96, MS99, MPKRS93, NAH98, PB99a, PK90, SS90b, SBP91, SM97, SF93a, Sri90, Wel94, YLC90, Zar95, AD97, AC95, AMB96, BHBA91, BB95, BA93, Bon90, BB91, BJ94, CL92a, CC97a, CMD93, Cor94b, Cor96, DS94, Dek98, DB91, FD90, FTYZ90, FL93, FRU96, Fer91, FFS92, FL98, GPZ99, GMP94, GE92, GPS93, GLR^{+}99, Gre94, GL91, HL92a, HC93, HWB92, HWB93, HB98, HL90, Hog99, Hon90, IO98, KS98, Li97, LK98, LPP91, MBR95, MT98, Mer97, MMJ91, NO94, NR90, NO91, NLV91, OVL96, OM91, Paa96, PWG94, QSA91, RF92, SP91, SN99, SBP90, SKA92, SC94, SMOR98, TB99a, TS99a, TH91, TY92b, WNS96, WH98, WHB98, HNR90]. parallelism [CL96, DAF94, PRASA99, SS94a]. parallelization [SMC91]. Parameter [BM94]. Parameters [KA96]. Parametric [FBM93, GPS95, LG92]. parity [FR97a, Jha93, JW94, LK93, NM94, Nik98, Tho97]. parsing [IPS91, Sri90]. Partial [Cha99, Fag97, HBA98, PL99, CA90b, FBM93, FL92, GB90b, JS92, KT96, OVL96, PR93b, Rag93, RCBB96, VSS91, VS93]. partial-multiple-bus [JS92]. Partial-sum [HBA98]. Partially [WL94, AS95, WPZZ96]. partially-mixed-polarity [WPZZ96]. particle [CRP^{+}96]. partition [FL93, KM92b]. partitionable [BBD97, KM92b]. partitioned [GM98b]. Partitioning [CMS99, BB99a, BCS92, BM96, HL92a, KF97, LL97, OM95, OdSC96, PR98, PB99b, RFKL93, San93, SS93a, SF92, STW92, TSW92, Tra96, Vij91]. Pass [JP98, CHH98, WM92]. passing [SI94, WHK97]. passive [CDP98, GM98b]. path [ASS94, AHT92, BSMC98, BAA95, BAA98, CT94a, CKC96, DP99, DL97, FFS92, Hou99, HG96, LK91b, PRS94, PR96b, RW98, RS98, THZ^{+}99, TYJ99, WH94, YMG99]. path-based [Hou99, THZ^{+}99]. paths [CL92a, CT97a, LYW94, MRT97, TCW93, Yan97]. pattern [CG96b, DC93, MV95, Maj96, MBGT91, NSUR98, NK97, ORPP93, RT93b, Sav98a, SS91a]. patterns [AMS96, FFS92, GM98b, GW92, GRT96, Sav99, SW94a]. PCS [FCL98]. Pentium [BW96a]. Per-node [TK98]. Perfect [FR94]. Performability [CMST90, IA91, NS96, SB94, PS90, PLB93]. Performance [AS92a, BD95, BM90, BD97, BCLA91, CS91, DDS98, Dal90, DYD94, DA99, DRS91, DM92, GB90a, GL92, HS94b, IS94, IS90, JL96, Kan90, LD90, Mah94b, MDyF94, MY94a, MY94b, PH93, SL90a, SJ93b, TN93, WSB90, WL92, YLL90, AVBZ97, BB97, Ber93b, BP98, BAA97, BS97, BP92b, CPRV95, CG96a, CB95c, CT96b, CFV91, CS96c, CSA94, DP99, Dal91, DG91, DBAP93, DBAP94, DRT97, EAT95, FD90, FM95, GY95a, GS95, GY95b, HM95, Har92, HC93, HAM^{+}99, HHMC91, HSHG99, HS98c, HLP93, Ish95, Joh95, KKKM99, KK91b, KTS90, KLH^{+}99, KM94, LLS91, LK93, LKL98, MM95b, Mei91, MS98b, NO94, Ng91, OI90a, OK99, PNDN99, PMHY97, PG97, PV94a, PV97, RFB97, RCS92, RF91, SS95, SJ95, SJ90, SK92a, SIKS95, SGP92, SAMC99, Soh90, SVSB97, SV97]. performance [TT90, TXD98, Tze93, VB96, VT94, WNPM95, WHL93, ZVN92, ZCT99, vdLS93]. performance-reliability [PV97]. periodic [CLL90, HS97, KSGT96, NC96, PM98, SC95b]. permutation [AZ98, CF95a, CC96b, DBD94, HT95, Hua90, JO93, KKL90, LL94, LO94a, LO97a, MY96, O’C95, SXW95, She95a, She95b, ST96]. permutations [CF95b, HSL96, HTCT91]. personal [DL97]. pessimistic [KK91a]. Petri [CL94b, CKLY98, GYY98, LG92, LW95a, LW95b, LZM90, SV95]. Pfair [BL98]. phase [TA95]. Phased [LH96]. photonic [DBAP93, DBAP94]. piecewise [ZVDF96]. pinwheel [BL98, CC92]. Pipeline [MFH91, DF91, HAM^{+}99, LS94b, SSK^{+}96, WS91]. Pipelined [DP96, BB95, Cor94b, GY95a, GDYS96, GL93, HL92a, KM90, LHF93, McC91, McC95, NO91, OMB97, Soh90, VD92]. pipelines [EAT95, SCF96]. PLA [AE94, EL90b, HK90, SB90, SHHJ93]. Placement [LS91a, Bae97, CT94b, TG99, ZB97]. placements [ACH^{+}97a, ACH^{+}97b, LK93]. placing [Wil93]. planar [SWK90]. plane [FL91, SW92]. PLAs [HLHC91, KF92]. PLAYTHROUGH [SGP92]. PM [SA98]. PMC [SR93]. Point [BGM^{+}95, CV92, Has90, JP98, Lo94b, OY91, OF97a, Tan92, Yok92, Fit97, Krä98, Lio93, MLK98]. pointer [LM99b]. pointer-based [LM99b]. points [CCCC96, FL91]. polarity [CRC96, DTB96, TMS97, WPZZ96]. policies [DYD94, DKT96, NT92, RSD^{+}98, dG90]. policy [GS97, HKL98, KS96a, LS94a, LMT95, SC95a, TN93, WH94]. polling [GL92]. polygon [AC95, SS90a, Yan97]. Polynomial [Kat94, LIO94, PK91, BHS92, Fal99, HRT^{+}95, KMM91, MP90, RT94, Sch98, SS92, VRK93b]. polynomial-time [RT94]. POPS [GM98b]. port [Fuj99]. Portable [GLR^{+}99, TPDB98]. position [BL99]. positive [Bal98]. Power [NL99, Wei94]. power-sum [Wei94]. Powering [Tak98]. PPMB [JS92]. Practical [McC95, PV90, CL99a, DH92, LKK^{+}94, Szy97]. precedence [Cha99, HS97, SS94c]. Precision [SM90a, AB98, HH93, Lim92, SM90b, WG95]. predicate [LMJ93]. predictability [BS97]. Prediction [LTR96, BL99, CK99, HLP93, IOP99, KE97, LTR93, RW98, SST92, SAMC99, SVSB97]. predictors [JG99, Nai95]. preemption [LHS^{+}98]. preemptive [LHS^{+}98]. Prefetching [JG99, CB95c, HS98c, LM99b, MS91, TS98]. prefix [Han91, MT98, WNS96]. prepass [CLM^{+}95]. Prescaling [ELM94, LM92a, LM99a]. preselected [PL99]. presence [CS96a, HS94a, LaF99]. preserving [CT93a, DA96]. pressure [LVAG98]. preventive [GPTT98]. primary [HS98a, HJ92b]. primary-backup [HS98a]. prime [CJ90, LS94b, WS91]. primitive [WH97a]. Primitives [LD90]. Principal [SES^{+}92]. principles [Szy97]. priori [CPV91, Krä98]. prioritized [GYY98, JL96, MM91a, MMM96]. Priority [SRL90, GS97, HR95b, KSS95, LHS^{+}98, MTG93, RCS92]. priority-based [MTG93]. private [BR94]. Probabilistic [NHA92, OS94, BSM92a, CGTN92, LS93]. probabilities [AR94, FFKI95, KM92a, Rob91]. Probability [LF96, CH90, DW90, ESS98, HR97, MI93, PGK90, RF92]. probe [CCCC96]. problem [Bar92, BC99, BBP91, BJK93, BP93a, BNG92, CC92, CT94a, CLC92, Fer91, GGD96, HV91, IJN94, JLC98, KMM91, LK91b, MS91, SM92a]. Problems [MV99, AS93, AC95, CCGS98, DLC93, Fri95, Fuj90, FTL93, GY95b, HL92a, Hua92, LK98, LS96c, PR95b, SF93b, SW94b, TK96, WDS^{+}95, YK98, ZPP91]. procedural [Uht91]. procedure [BM90, LP91, PR99]. procedures [CSF90, Yan90]. process [AA92, BNG92, KLL97, UR94]. processes [CMST90, NNHG93, RS95, SCV91]. processing [AB98, AS92b, Cor94b, DF91, Esh91, FL93, FFS92, Lim92, LKL98, Mat90, MYK^{+}99, PM98, PK90, RB90, SP91, SC94, ST95b, SY99]. Processor [Leu97b, MF92, RSD^{+}98, ABZ96, AD97, BGJ^{+}96, BB92b, BA93, BM91, Bon90, CF90, CS99, CT92, DA99, DMS95, Dow92, DBAP93, DBAP94, FTYZ90, HJ92a, KT99a, KR93, LW91, LAY96, Lee92, Li97, LO97a, LBF92, MMSS96, NA90, SKA92, THA^{+}99, VRK93a, Zar95]. Processors [KC93b, AK91, BSMC98, BHR94, BPRR93, CLM^{+}95, CL96, CU93b, CB95c, CUC97, CT93b, Cor96, DPO95, GL93, HLP93, KRM95, LS91a, LO94a, McC91, McC95, Mie93, MD95, NV99, RSS97, Soh93, Sto93, SCW91, TK97, TD93]. Producing [SSDK95]. product [FE97, KS93b, LO94a, OVL96, WMD96]. products [DM97a, Sas91, VSS91, VS93, WJ90]. program [AN90, BS97, BP96, CL94a, DAF94, HL91a, HL97, KKKM99, KOIH92, Lam97, Li97, MLK98, PG97, Wil93]. programmability [ZCT99]. Programmable [ITZ96, Hug93, JBS90, TB99b, TB91, ZWL98]. Programming [ST95a, BBD97, BBP91, CSTO92, DA90, SS93a]. Programs [Wel94, CT93b, FD90, HC94b, PM91, Weg94b]. Progressive [WHK97, Pin99]. prohibition [BB99a]. projection [WS95]. Prolog [FD90]. Proof [NAH98, MLK98, SG90]. proofs [TSH95]. propagate [EL90a]. Propagation [PK94, CL92b, Par94, PLM91]. Properties [vdBT93, BBKA95, CDHT99, Dou93, GGD96, HNY93, LC98a, MJ91]. property [CN99, TA95]. propositional [MSS99]. Protected [GHH93]. protection [FR97a]. protocol [ACZD94, ALPS90, Bal98, CCG^{+}91, CB99, CF94, CK91, DDS98, DL97, DW91, KS91, Kin91, LD94, Mei91, RM91, SRSC91, SZTN91, SCV91, VB96, ZSR90]. protocols [CG97, GM91, HHMC91, HSHG99, Lin91, LT93, PS91, RHT95, SRL90, SVSB97, TS95, Wil92a]. prototyping [CCG^{+}91]. pruning [SM93]. Pseudo [Dre99, BBP91, GRT96]. pseudo-Boolean [BBP91]. pseudo-exhaustive [GRT96]. Pseudo-Kronecker [Dre99]. pseudoexhaustive [DC93, RT93b, SHHJ93]. Pseudorandom [YGI91, SSDK95, WP90]. public [BMP97a, LG97, LCL95, PFSR99]. public-key [BMP97a, LCL95, PFSR99]. Purpose [KC93b, FTG92]. pyramid [CDHT99, Gre94].

QoS [Hou99]. quality [CKC96]. quantitative [HSHG99, WHL93]. quantizers [PWG94]. Quasi [WC99]. Quasi-universal [WC99]. queries [HBA98]. queue [TF92]. Queueing [DKL90, GL92, KT92, LP91, OI90a, dM90]. queues [Bon90, PLJ94]. Quicksort [HNR90]. quorum [Kum91]. Quotient [WF92, SPM97].

R. [Dav90, KH93]. Race [FW93, KWF95]. Race-free [FW93, KWF95]. races [Ung95]. radially [Kno96]. Radix [CR90a, CL94c, EL90b, ELM94, SPM97, ABZ96, AVBZ97, ALB98, CM90, EL90d, LM92a, LM99a, MC92, MC93, MC95, NK99, SF93a, SP95, Tak92a, Wal97]. Radix- [CR90a, EL90b, EL90d, MC95, Tak92a]. radix-4 [AVBZ97, SP95]. RAID [BBBM95, CT96b, MY96]. rail [Jha93]. RAM [CGTN92, FS94b, LS91b]. RAMs [DF90, HDS91, Nic96]. Random [KMD97, Sav98a, DF90, DP97, MV95, Maj96, Maz93, MY96, MK90, NSUR98, NK97, Pel98, SMV91, Sav99, SFM92, SS91a]. random-access [Maz93, SMV91]. Randomized [TG99]. randomly [LM92b]. randomly-wired [LM92b]. Range [HHB91, CDP98, WJR94]. Rao [LHL99]. Raphson [KTM^{+}94]. rapid [YS96]. rate [FCL98, HB95, PM91, SCF96, SST92]. rate-adaptive [HB95]. rate-optimal [PM91]. ratio [Vai97, vdBT93]. Rational [GY95b, KZ90, KM90]. ratios [TSW92]. rays [COK95]. reactive [BB99a]. read [ACFH95, Vin95, Weg94b]. read-once [Weg94b]. read-only [Vin95]. readers [TN93]. Real [NA90, ABCW98, BH97, Ber93a, BLOS95, BS99, DKT96, GOH92, GPS95, HR95a, HLH96, HS98a, HLB^{+}97, HL92b, HS94c, HS94d, HS97, HLP93, IA91, JL91, Kan90, KL92, KM97b, NT92, Pol95, RHS98, SS93a, SRL90, SRSC91, SSS97, SB94, SC95b, SS94c, SLS95, Vui90, WRS95, WHL95]. Real-number [NA90]. real-time [BH97, Ber93a, BLOS95, BS99, DKT96, GOH92, GPS95, HR95a, HLH96, HS98a, HLB^{+}97, HL92b, HS94c, HS94d, HS97, IA91, JL91, Kan90, KL92, NT92, Pol95, RHS98, SS93a, SRL90, SRSC91, SSS97, SC95b, SS94c, SLS95, WRS95]. real-time-systems [SB94]. realization [AEBV90, KEA96, LS94b, She95b, WS91]. realizations [Sas97]. realized [CF95b]. Realizing [GM98b, HTCT91, SB92a]. Rearrangeable [Dou93, FS94a, MM97a, YF92b]. reasoning [Cyr94]. reception [CF94]. Reciprocal [DM94, Man95]. recoding [HC96, SG90]. recognition [CT92, IPS91, IYI90, LVB^{+}93, SW94b]. recognition-complete [CT92]. Reconfigurability [SS93b]. reconfigurable [BS90a, CMS99, CU93b, CC94, HK90, KK91b, KF92, MPKRS93, PB99b, Raj96, RSS97, SCF96, SB92b, ZWL98]. Reconfiguration [BB92b, BP93a, CF90, CU93a, CUC97, CT91b, JTY^{+}99, KR93, KM97b, LK99, LL96b, MYK^{+}99, RBK90, SSA94, Som90]. reconfigurations [Mel92]. reconfigured [FS94b]. Reconfiguring [VRK93a, BP94, DH90, GMP94, VRK93b]. recording [HN93]. Recoverable [WF90].

Recovering [AA92, OK96]. recovery [ACFH95, EZ92, GS95, LS98, Pin99, PV97, RCB96, SE98, Vai98, WHK97]. Rectangular [WG94, Ell91, MH90, RBG98]. rectilinear [Yan97]. Recursive [Ara90, RT93b, BB91, Kan93c, LHF93, MvdS90, MM95b, OM91]. recursively [LNS94]. Reduce [WF92]. reduced [AK91, CM92, Lim92, LVAG98, TH99, Tya93, Uht91, Weg94b]. reduced-area [Tya93]. reduced-size [TH99]. Reducing [GL93, MC93, MC95, RH93, Sav96]. Reduction [Smi95, AEBV90, ABLZ97, CC92, CC96a, HTHR94, Ng91, OVL96, RS96, SWT93, SL91b, TLE97]. redundancies [PR96a]. Redundancy [SM90a, CU93a, CU93b, CU94, FSS99, HDS91, LKW97, MRB90, RM94, SK94, WRS95]. Redundant [ABCC90, DM96, EL90c, HC94a, KTM^{+}94, NK99, Par90, PK94, TAY91, ABZ96, CS99, DOP93, DHK93, EANPH93, Kle90, KM90, LL92a, MC93, MC94, MC95, PB97b, SPM97, TY92a]. Reed [DTB96, HB95, IDI95, KY92, ORPP93, Pur91, Sas97, SLN97, TMS96, TMS97, VR94, WPZZ96, ZV95]. Referee [Ano98c]. reference [UR94, WI95]. references [FS96, GHH93, vdBT93]. Refinable [Sem95]. reflected [JH95]. Reflections [BW96a]. register [AK92, CL96, CW94, LVAG98, SM92b]. registers [FFKI95, GP96a, GK92, HRT^{+}95]. Regular [GO98, CZSM99, CT94a, CT98, LNS94, LW94, SA92, SP96, TC95, WH94]. regulation [RKJZ99]. Related [MV99, Fal99, LHS^{+}98, VCB96]. Relating [DRW92, GK92]. relationship [OFW92]. relaxed [PD99]. release [HKC97]. Reliability [BR94, CU93a, BR90, BR93, CB95a, CL94a, CU94, CDHK93, GOH92, HW95, KM97a, LBF92, MB95, MYD96, PV97, RHT95, SS93b, SWG92, TK97, Tze93, UP93, VP93, Vai96, VT94, YF92a]. Reliable [KL97, Lo94b, Tze94, Wu97, LC98b, SES^{+}92, SSDK95, Wu95]. Remainder [HP95]. remainders [MC93, MC95]. remapping [NR90]. remark [MC95]. remote [Met91, TK98]. removal [FSS99]. removing [PR96a]. rendering [COK95]. rendezvous [WNPM95]. reordering [FS96, NP90]. repair [BP93b, CT93b, FR97b, HDS91, NNHG93, RS95]. repairable [BR93, Ser99, SM96b, dG90]. repeated [Wal97]. replacement [vdBT93]. Replicated [KLL97, Kum91, Leu97a, MY95, Met91, SBM^{+}98]. replication [MRJ97, WJ90]. Reply [KIN^{+}96, NC97, BMP^{+}97b, CG97, HNG94, JBS90, KGB95]. represent [JBA^{+}97]. Representation [KTM^{+}94, AM93, Cyr94, Dro98, JK92, JM96, LL92b, TY92a]. Representational [Bou90]. Representations [Par90, PK94, Yok92, AW93, Bry91, DHK93, NK99, WH99]. representing [LW95a, LW95b]. reproducing [KTM96]. Request [DR96, CG97]. requesting [CS91]. required [LH98]. requirement [TL99]. requirements [Hou99]. rerouting [MS98b]. reseeding [HRT^{+}95]. reservation [EAT95, SC95b, YMG99]. reservation-based [SC95b]. reset [AG94, PR94b]. Residue [KL97, AM93, DOP93, DIP93, HK95, Kat96b, KS93b, LC92, OPT92, Pie94, Pou97, SL90b, WS95, WJ90]. resilience [HNG94, NG90]. Resilient [RS98, DW97, FR97b, RHT95]. resistance [BS92, GM98a]. resolution [ASS94, Fag97, MM91a, MMM96]. Resource [Bae97, LS97, CT94b, WNS96]. resources [SS94c, ZWL98]. Response [KH93, HJ92b, LMT95]. responses [CMH98, KN90, KLV94, RT93a]. responsive [Pol95]. responsiveness [SLS95]. resubmission [DR96]. result [MC93, MC95]. Results [DM93, ACH^{+}97a, CLL90, HV91, Man93, NP90, SWM95].

Retiming [KT96]. Retiming-based [KT96]. retrieval [LL90]. retry [HS96, KS96a, LCFH95, LS94a, WHK97]. return [YF92a]. reuse [TRS^{+}99]. reversed [CT91a]. Reviewers [Ano97a, Ano97b]. revisited [SL91a]. Reward [DKT96, CMST90]. rewiring [CVMS99]. rewrite [BS93a]. RIDDLE [SS91b]. Right [Kor94b, TH99]. Right-Shift [Kor94b]. Ring [CK90b, KH90, BD95, Coh93, HS94b, Kam92, KH93, LK99, MTG93, RSS97, RB93, RB97, YM95, ZVN92]. ring-based [HS94b]. rings [HW95, NK99, SGP92, SJ93b, WJR94, WMD96, YTR94]. risk [PZ99]. river [Tua92]. RNS [BDK98, CN99, DPO95, GL99a]. Robust [FFS92, DA96, HG96, KR90a, KN90, LS96a, Saa95]. robustly [CKC96]. role [PR94b]. Roll [PV94b, PV97, EZ92]. Roll-forward [PV94b, PV97]. rollback [ACFH95, EZ92, FTG92, LS98, PV97, SE98, TT90]. rollback-recovery [SE98]. ROM [DM94, IN96, ZI92]. Root [CL94c, SF96b, ALB98, EL90b, ITY97, LM92a, LM99a, MC93, MC95, WH97a]. Rooting [Has90, CM90]. ROSOP [WM98]. rotation [AS93, AVBZ97, LL92a]. Rounded [SS94b, LMT98]. Rounding [ELM94, KTM^{+}94, YGI91, ALB98, EL92, LM99a]. roundings [DM97a]. routability [RN95]. routed [Tse97, TYJ99, TNS99]. router [KS94, RHS98]. routers [BD97]. Routing [ANR96, Hou99, LS96c, SC90a, TW95b, AZ98, BNL93, Ber93b, BBP^{+}96, BL92, BFS97, BC95, CF95a, CS90b, CSK90, CC97a, CC96b, CW96, CC97b, CG97, CG94, DBD93, Dou93, FRU96, FS94a, GDYS96, GP97, HSL96, HTCT91, JO93, KS92, KS93a, KD94, KYM97, Kus90, LL94, LM92b, LK95, LH98, LH91, LMT95, MM97a, MS95, OK99, PA96, Pin99, RB91, Raj96, RFS92, SS96, SW97, ST96, Szy97, TA95, TL99, Tua92, UVM97, VB96, WL90, WH94, WSEZ90, WS91].

Rover [JTK97]. row [BL92, SM96b].

row/column [SM96b]. RSA

[BM94, TY92a]. rule [KR93]. rule-based [KR93]. rules [BS93a, CBRS90]. Run [JCMH99, SMC91, DB91]. Run-time [JCMH99, SMC91, DB91]. runtimes [SS95].

S [HF97]. S3.mp [PBND98]. Safe [VP94, AS95, LC98b, Nic98]. Safety [Wu95, VP93, Wu97]. salesman [BJK93]. Salphasic [Chi94]. Salvaging [Sav98b]. same [SLH97]. sampling [CHH98, KHW94]. sandwich [YM95]. SAT [GGD96]. satisfiability [MSS99]. save [CM96b, Pie94]. saving [RSD^{+}98]. Sb [HF97]. SbEC [HF97]. scalability [Dek98, GS98, HSHG99]. scalable [DBAP93, DBAP94, MNL99, PA96, TB99b]. Scale [DM96, TAY91, ABK90, CL90, FW93, LKL98, Tze92]. scaling [ATF98, ALB98, EL90d, GL99a, SL90b]. scan [CA90b, GB90b, HRT^{+}95, KT96, NGB93, PL99, RTZ98, RT99, TB99b]. scan-based [TB99b]. scatter [NS90]. Scattering [BPRR93]. scheduled [HKC97, dG90]. scheduler [BK90]. schedulers [CC92]. schedules [SF91, TC95, WNS96]. Scheduling [BH97, CLL90, Sah96, ASWW92, BSMC98, BH98, BL98, BS99, Cha99, CLM^{+}95, DA99, DKT96, FTYZ90, HL91a, HL97, HLH96, HL92b, IOP99, KSS95, KM92b, LK91a, LHS^{+}98, LVAG98, NT92, PM91, PB99a, PB99b, RSD^{+}98, SMC91, SK92a, SL95, SC95b, SS94c, TW95a, TN93, WPS96]. schema [SK92b]. Scheme [KTM^{+}94, BNL93, BBBM95, BJ94, CL96, CUC97, COK95, Fuj99, GL99a, GPZ99, HL91b, HKL98, JW91, Kat96b, LMJ93, Lee92, LH92, LYS95, LK96, Lo96, Man93, PA96, PV94b, Sto93, Tse97, Tya93, UVM97, Vai97, WY95, ZI92]. schemes [AS92a, BD97, BM94, CM96b, FM95, LS93, NAB96, RW98, RCBB96, SSA94, Vai98, ZB98].

Schur [Zar95]. Scientific [KL97]. scope [FR97a]. SDE [IIK^{+}91]. search [Ara92, CCL92, Kuh98, Lu97, MSS99, PG92, PR95b, Tok94, ZS99]. Searching [HMM^{+}90]. SEC [Kat96a]. SEC/AUED [Kat96a]. Section [Ano98c, JHK97]. sectored [Sez97]. Secure [PBB97, WL94]. security [Kuh98]. seed [SM92b]. segmentation [Dal92]. segments [LT95]. select [Kan93c, Tya93]. Selection [ELM94, ASS94, ALB98, BM94, MC92, SPM97]. Selective [SV97, YY97]. Self [KSGT96, SCW91, WM92, Yok92, AZ98, BK95, BJ94, CF95a, CW99a, CC96b, CL94b, DGY92a, DGY92b, FTYZ90, GPZ99, GP96b, HPN95, HMC90, HJ92a, IN96, JO93, Jha93, KPB92, KC93a, LL94, LYS95, LT90, Lo96, LZM90, Nik98, OK96, Pie90, Pie96, RB91, RFKL93, SD92, SM96b, Szy97, THL92, Tao97]. Self-checking [KSGT96, BJ94, CW99a, HPN95, Jha93, KPB92, LT90, Lo96, RFKL93, THL92]. Self-Delimiting [Yok92]. self-diagnosable [SD92]. Self-diagnosis [SCW91, LYS95]. self-diagnostic [KC93a]. self-recovering [OK96]. self-repairable [SM96b]. self-routing [AZ98, CF95a, CC96b, JO93, LL94, RB91, Szy97]. self-scheduling [FTYZ90]. self-stabilizing [CL94b, LZM90]. self-test [BK95, GPZ99, GP96b, HMC90, IN96]. self-testing [Nik98, Pie90, Pie96, Tao97]. self-timed [DGY92a, DGY92b]. semantics [Bou90]. Semi [MST98, CMST90]. Semi-logarithmic [MST98]. semi-Markov [CMST90]. semimodular [LB99]. sense [VB96, YW99]. sensitivity [BE92, CI92]. sensor [IJN94, Pol95]. separation [HBAB95]. sequence [LS96b, Leu97b, RP99]. sequences [Ann97, CT91a, CW94, KTS90, OM95, PR94a, PR96a, RT96, SM96a, UWZ97, WP90]. Sequencing [KS96b, HC92]. Sequential [PWG94, RT91a, Som90, CMS99, CA90b, CAK90, CA92, DGY92b, FW93, HRP99, KSL93, KF97, LB90, NSUR98, PR93a, PR94a, PR94b, PR95c, PR96a, SSDK95, SW94a, Ver92, WM92]. Sequentially [XH95]. Serial [JLLR94, HB92, IV94, SFM92]. series [Yan90]. server [BM94, DM92, SLS95, WNPM95, ZC98]. server-aided [BM94]. Service [CK90b, DKT96, HS98b, LW95a, LW95b, MRJ97, vdLS93]. set [BBP91, CR90b, HKN99, Kor94a, LL96a, Pie93, SR93, WW91, Wan97, WHL95]. set-associative [HKN99, WHL95]. sets [CR93, GRR94, Hua90, KTM96, KEA96, Kun92, MC94, SP96, WG98b]. Setup [OM91, DL97]. seven [FTL93]. SFU [LLS91]. Shared [DW91, BGJ^{+}96, CB99, CS92a, DDS98, DYD94, DA99, Dek98, GW92, HSHG99, HS94b, HS93, KH97, MNL99, PRASA99, PBND98, PLJ94, RPW98, SS94c, Tze92, WF90]. shared-memory [BGJ^{+}96, DDS98, DA99, GW92, MNL99, PRASA99, PBND98, Tze92]. shared-nothing [HS93]. Sharing [KH90, BHR94, Bon90, CS96a, HS94c, HS94d, JLC98, SC95a, TLH94]. Shift [Kor94b, CW94, GP96a, GK92, HRT^{+}95, SM92b, Tok94]. shifters [FR94]. short [Man95, SY91b]. shortest [AHT92, WH94]. shortest-path [WH94]. shorts [HM93]. shuffle [Hua90, HTCT91, MB95, RB91, Sch92]. shuffle-exchange [MB95, RB91, Sch92]. shuffle/exchange [Hua90]. ShuffleNets [LLJ97]. sided [Sar90, VC92]. Sigmoid [ZVDF96]. sign [ABCW92]. sign/logarithm [ABCW92]. signal [AR94, AB98, CH90, KM92a, Lim92, LKL98, PB99a, RB90]. signaling [Leu97a]. signals [Chi94, PL99]. Signature [Ano96b, Dav90, NAH98, Bos90, CPV91, DOR91, FFKI95, Hla92, HMC90, KPI93, KC93a, KLV94, LK91a, MMJ91, MI93, PGK90, Rob91, SFM92, SM97, Wil93, WI95, YL95b, YL95a]. signatures [UR94, Wil93]. Signed [AW93, Par90, PK94, Tho97, CR90a, GM98a, Par93, Pou97, WH99]. Signed-Digit [Par90, PK94, CR90a, Par93, Pou97, WH99]. silent [BES^{+}96]. silicon [RF91]. SIMD [AMS96, CSS90, LPP91, RS96]. similarity [OZL96]. Simple [EL90d, SFM92, CF95b, CCGS98, MC92]. Simplicity [Egg91, ZCT99]. simplified [Ma98]. simulated [ACH^{+}97b, SRHI98]. simulating [GSH^{+}92]. Simulation [CT96a, CCG^{+}91, CRP^{+}96, CAK90, CHH98, FFS92, Gos97, Mah94a, NNHG93, PR95a, PR95c, SS91a, SAMC99, Soh93, TWS92, TWS93, WHL95]. simulation-based [CAK90, Gos97]. simulator [HG96]. Sine [TAY91, Pha98b]. Single [JP98, Lim92, WS95, WI95, AR94, BL92, BBT94, BCLA91, CHH98, ESS98, FFKI95, Fuj99, Joh93, LL96a, OI90a, Rao93b, SW90, SR93, SL90b, Tua92, WG95, XSFR96, vdLS93]. Single- [JP98]. single-bus [OI90a, vdLS93]. single-fault [SR93]. single-input [ESS98]. single-port [Fuj99]. Single-precision [Lim92]. Single-reference [WI95]. single-row [BL92]. singular [EL90e]. site [HJ92b]. situations [KK91c]. size [BOH90, CR93, DAF94, Sav96, SLH97, SRK91, SAMC99, TH99, Weg94b]. sizes [HM94, KM92b]. skew [Fis90, KS90]. Skewed [BS97]. skewing [COK95, Lee92]. skip [CS90a, CSTO92, Kan93a, Kan93b]. skipping [Man90]. slave [GW91, Sah96]. sliding [Sar90, TB99b]. slot [HHS97, JLLR94, KMD97]. Slotted [KH90, BD95, Kam92, MTG93]. slotted-ring [MTG93]. SM [ABK90]. Small [CCD^{+}92, Dro98, Kus90, Lin90, WJR94]. smallest [Yan97]. smartcard [Ara93]. Smith [JBS90]. soft [BS99]. software [AKL90, BZ92, CW99b, CKY^{+}99, DJ99, GPTT98, GOH92, HKC97, IIK^{+}91, KKA95, Kan90, Leu97b, PI93, PZ99, RKS90, ST95a, SZTN91, SE98, SRHI98, WHK97, WK97, WNPM95, XT99]. software-based [KKA95]. software-implemented [SRHI98]. Solomon [HB95, IDI95, KY92, SLN97]. Solution [TLROL93, Ung94, JLC98, Jun95, RCBB96, WL90]. Solutions [Bar92, BHR94]. solvers [LZ94, SZN92]. Solving [WM98, Ber93a, HL90, SF93b, TK96, WM91]. Some [DH92, HNY93, Jun95, LC98a, Man93, Mul94, SI93, TY92b, Ung94, DW90, GY95b, HW95, TSH95, ZPP91]. sorter [BG90]. sorters [LC98a]. Sorting [BG90, FL98, VD92, AK91, Aln93, BSM97, CC94, DDGS97, Fio99, GL91, KL92, LRG91, Pie93, Raj96, SP91, SCG94]. sources [LL97]. Space [Wal97, CMH98, DC93, GL91, ITZ96, TC95, TL99]. space-optimal [TC95]. Space/time [Wal97]. spanning [FA96, FL91, LS92, TS97]. SPAR [TRM95]. spare [CU93a, KF92, TSS95, VRK93a]. spared [LaF99]. spares [BCH93a, CF90, TSH94]. sparse [ABK90, GO98, HL90, WDS^{+}95]. spatial [Tem99, TLH94]. Special [Ano97b, Ano98c, KC93b, FTG92, LLS91]. Special-Purpose [KC93b]. specific [CKY^{+}99]. Specification [JL91, IIK^{+}91, LL91a, PD99, TS99b]. specifications [KS91, TS91]. specified [HHMC91, HM91, ZV95]. spectra [FC97]. Spectral [BC99, BOH90, Cor94b]. spectrum [Cor96, Hei91]. Speculation [CL94c]. speculative [CWM^{+}95, KT99b]. speech [CMD93]. speed [GS98, HG96, KIN^{+}94, MMSS96, OVL96, PKI^{+}96, Pie90, QF92, RBG98, SSK^{+}96, SRHI98, TH99, TKS94, VT97, WS98, XSFR96]. speedup [Hog99, VD92]. split [Vin95]. splitter [LM92b]. spot [DE92]. Square [BNP91, CL94c, Has90, SF96b, CM90, Ell91, EL90b, ITY97, LM92a, LM99a, MH90, MC93, MC95, WS98]. Square-Root [CL94c, ALB98]. squarers [IV94]. Squaring [Zur94]. SREMIS [WI95]. Srikanthan [Hwa97]. SRT [BW95, Man93, MC93, MC95]. stabilized [BB91]. Stabilizing [GM91, CL94b, LZM90]. stack [WHL95]. stage [BCLA91, Dou93, HSL96, HSY97, She95a, She95b, YA90]. staggered [HKL98]. stair [MJ92]. star [BNL93, BDN96, BCGS96, CDP98, CC97a, FA96, Fuj99, GS93, HB94, LY91, LCT97, TS97]. star-type [LY91]. stars [GM98b]. Starting [SF96b]. state [CCC96, Cd93, DOR91, FW93, HB98, IJW92, KWF95, KA94, LY94, McA92, Oik92, PR97b, RFS92, Ros90, SM96a, SK94]. state-table [PR97b]. states [CA91b]. Static [NST96, PM91, SSH^{+}99, SL90a, DRW92, HRP99, KK92, Uht92]. stationarity [Ser99]. Statistical [BS92, DO96, IOP99, PMW92, SAF90]. stealing [BCLR97]. Steiner [SW92, SLW98]. step [KK91a, Pha98b]. sticky [DM97a]. stiff [BT90]. stingy [OI90b]. Stochastic [CDFH93, GS95, Li97, PLM91, CL94b, LG92, LW95a, LW95b, LZM90, NST96, Sch96, WNPM95]. Storage [CG94, DM92, HL91b, MM91b, PMHY97, Vog90, vdBT93]. Storage-efficient [CG94]. store [BBB^{+}94, Gai90, LLS92]. store-and-forward [BBB^{+}94]. Strategic [KSNM91]. strategies [ANR96, BP94, BB92b, BCLR97, BLOS95, DF91, Hog99, MJ91, MY95, MB93, TSS95, VP93]. Strategy [SC90a, CW96, CT92, KK91a, LH91, PR92, PR95b, YW99, vdBT93]. stratified [CPA99]. stream [Uht92]. streams [CL99b, HR95b, HR97, VLPA95]. Street [CSA94]. strengthening [TLE97]. Stress [THZ^{+}99]. Stress-based [THZ^{+}99]. Stretching [BS90b]. strict [WNS96]. String [IJW92, IS90, MS93, MS94]. striping [MY95]. strong [CL98a]. strongly [Hua92, Jha93, WL94]. Structural [TC94]. structure [BP95, IDI95, LS94b, MS93, SLN97, Tze92]. structured [GB90b, O’C95, RSS97, WM91]. structures [BK95, BSS94, CSF90, CLL91, CGTN92, JHK97, KR90a, SA92, SB92b, Tra96, Vij91]. stuck [CH90, LB99, PW95, PR98]. stuck-at [CH90, LB99, PR98]. stuck-open [PW95]. study [Ber92, GB91, HS98c, IYI90, KPB92, Mei91, Soh93, ST96]. subbus [FL98]. Subcube [CB95b, CS90c, CT97b, DH91, BCS92, CB95a]. subcubes [CL92a, TL96]. subgoal [NP90]. Subject [Ano98d, Ano99b, CT93b, RS95]. subnet [AC98]. subset [KBMP99]. subspaces [RT96]. substitution [HT95, O’C95]. substitution-permutation [HT95, O’C95]. subsystem [GP98]. subtraction [ABC92, Lew90]. sufficient [Sch96, WdB92]. Suggestion [JM96]. suitable [CF90, YM92, YL95a]. sum [CY95, HBA98, SB90, Sas91, Wei94]. sum-of-products [Sas91]. summation [CM92]. super [CM92]. Supercube [ARS95]. superpipelined [CLM^{+}95]. superpositions [SWT93]. superscalar [CLM^{+}95, CL96, SIKS95]. supersetting [BDL^{+}99]. superthreaded [THA^{+}99]. Support [SE98, JTY^{+}99, PDBH97, Par93, RPW98, ST95a, XT99]. supporting [Kan90, LH96]. sure [FHA94]. surprise [YF92a]. survey [PS91]. Survival [HW95]. susceptibility [HC94b]. SVD [EL90c]. Svoboda [MPG98]. swap [PLJ94]. sweeping [SSL96]. switch [Wil92b, WC99]. switchable [EAT95]. switchbox [Sar90]. switched [DR96, YB91]. switches [GMP94, TF92]. Switching [CC96a, YLL90, DRS91, HM91, HC94b, JO93, Jun95, Kun92, MY94b, SD96, Szy97, TLROL93, Ung94, VC92, VR94, YM91]. syEC/AUED [LY96]. Symbol [Che92]. Symmetric [SS99, BHSB97, CCD^{+}92, CDFH93, Dre99, KR90b, Lin93a, MPS94, WeBM97]. symmetries [TMS96]. symmetrization [TLE97]. Symposium [Ano98b]. synchronization [CPK90, CL94b, LZM90, OS94, Pal94, RKS90, SRL90]. synchronized [WM92]. synchronizer [WC96]. synchronizers [EANPH93, Kle90]. Synchronizing [HS94a, PR96a]. Synchronous [LD90, AS92a, ACZD94, BJ91, Chi94, KS90, KEA96, Lin93b, LH96, PR93a, PR94a, PR94b, PR95c, PR96a, TW95a, WNPM95]. synchrony [BBD97]. syndromes [KM92a]. Synergistic [SL92]. Synthesis [Cha96, CRC96, CKY^{+}99, FSS99, KM95, LKL97, PS91, AEBVZ91, BB90, BDL^{+}99, BSS94, CCC96, CA92, DJ99, KS91, Kun92, LR97, LS90a, OK96, SJB91, SZTN91, SR96]. Synthesizing [KR90a, RM91, FW93, KT91c]. Synthetic [Mah94a, TWS92, TWS93]. System [DJM99, SJP96, ALPS90, ABC92, AS92b, BB92c, BMP97a, BK90, Bon90, CPRV95, CL99a, CSSP90, CGK93, DKL90, DIP93, DN98, FBL96, GP98, Gos97, JL96, KKA95, KBMP99, KK91c, KS93b, KA96, LW91, Lew90, LS90c, LL91b, LCL95, LP91, LZ94, LC92, LS97, LC98b, OI90a, Pel91, PBND98, RCS92, RDZ95, SKK^{+}90, SN99, SK92a, Som90, TI93, TXD98, TSS95, UP93, VP94, Vog90, WH97b, vdLS93]. system-level [BB92c, DN98, FBL96, GP98, Pel91, RDZ95]. Systematic [KH97, LHF93, MK90, BBT94, GB93, KGB95, KB93, Man90, MM95b, SW90, SW95]. Systems [Par90, PK94, SL90a, Tri98, YLL90, AS92a, AMS96, AV97, AD97, AKL90, ACC^{+}93, ABCW98, AMB96, BBD97, BH94, BR90, BR93, BH97, BL98, BSM92a, BSM92b, BP93b, BB99b, BAA98, Bou90, BNG92, BES^{+}96, BLOS95, CG96a, Cd93, Che92, CL94a, CU94, CS99, Chi94, CPK90, CL99b, CKLY98, DP99, DE90, DA99, DJ99, DD95, DP97, DG91, DM92, DDV90, FP98, FNS95, FR97b, FRK98, GPTT98, GHH93, GE92, GSH^{+}92, GL92, GYY98, HF97, HLH96, HLB^{+}97, HY99, HK95, HL90, HJ92a, HS94d, HS96, HS97, HKC97, HBAB95, IA91, JW91, KRM95, KM97a, Kat97, KK92, KS96c, KK91b, Kos94, KM92b, KS90, KEA96, KA96, KM97b, LHNP94, LY91, LK96, LK98, LS95, LKW97, MTP93, MST98, NHA92, NNHG93, OS94, OK96, OPT92, PBB97, Par93, PS90, PLB93]. systems [Pel98, PMHY97, Pol95, RS94, Rai95, RKS90, Rao93a, RNA93, RR95, RSD^{+}98, RCB96, Sah96, SSH^{+}99, SS93a, SD92, Ser99, SB94, SWG92, SC90b, SES^{+}92, SJ93a, SP96, SI94, SL90b, TB98a, TT90, TI92, TS95, UP93, VT97, VJ93, WRS95, WHL93, XT99, XSFR96, XH95, Yan90, YB91, YF92a, YY97, dG90]. systems-SbEC- [HF97]. Systolic [CM92, CJ90, GW98, Kor94b, MP90, Wal93, CLC92, DOP93, HB92, Hug93, IDI95, JK92, KT91c, KL92, LK91b, LC90, LK90, LHL99, MRB90, OFW92, SM92a, SS93b, ST95b, WL93, Wan94a, Wei94, WS91, WD92]. systolic-array [IDI95]. systolic/pipeline [WS91]. systolic/wavefront [SS93b].

t [RCSSS94, KB96, LY96, THL91]. t-EC [KB96, THL91]. t-EC/AUED [KB96]. t-EC/d-UED [THL91]. t-syEC [LY96]. t-syEC/AUED [LY96]. t-unidirectional [RCSSS94]. T. [PT92]. Table [Tan92, BS90c, MC92, PR97b, SB97, Tak98]. Table-Driven [Tan92]. Tables [DM94, SS99, EAT95, Lin90, PG92, SB92a]. tabular [Jun95, TLE97, TLROL93, Ung94]. tag [RFS92]. tagged [AN90]. tagged-token [AN90]. tags [WSY97, WHL95]. take [TW95c]. tampering [BBRL98]. tape [HMM^{+}90]. target [FM95, Fag97, HC92, LA98, PS93]. Task [KM97a, SWG92, BL98, CS90c, CT97a, HS94d, HS97, IOP99, LLK92, SC90b, TS94, ZB98]. tasks [BLOS95, BS99, DKT96, GPS95, HL92b, KB93, KS96c, KS96b, KM92b, NST96, PM98, SC95b]. TBED [KY92]. TCP [BB97]. TDM [LLJ97]. technique [AE94, Bel90, Ber93a, BM91, BOH90, CS96b, CT93a, DIP93, FL92, HR95b, THL92, Tok94, WJM95, YK98]. techniques [BB92d, Cd93, FNS95, JBA^{+}97, KSNM91, KHW94, LS95, MAH98, NAB96, OCVSV98, OPT92, PHS99, PG91, RFS92, RP99, SAMC99]. Technology [ATF98, DS94]. telescopic [BDL^{+}99]. Temporal [PB99b, Cyr94, Hou99, KKKM99, Tem99]. temporal-based [KKKM99]. term [WS95]. terminal [TPDB98]. ternary [DS94, YM92]. Test [BAA95, CA91b, CCCC96, ORPP93, PR97b, RTZ98, SK92a, TS91, BBP91, BAA98, BK95, CMH98, CG96b, DC93, GPZ99, GRT96, GP96b, HRT^{+}95, HMC90, HRP99, IN96, KTM96, KN90, KLV94, KSL93, LL96a, LL91a, MV95, MBGT91, MAH98, Pie93, PR92, PR94b, PR99, RD96, RT93a, RT93b, Rob91, RP99, Sav98b, SS91a, SS91b, SK96, SSDK95, WG98a]. Test-pattern [ORPP93]. Testability [BH96, Cha96, CA90a, FSS99, KM92a, KMM91, LIO94, PK91, Sav98a, SAF90, Ver92]. testable [CCC96, CRC96, Hon90, LH90, MRT97, RT94, Sas97, SHHJ93, WC90, XK92]. Testing [LY94, PR93b, SMV91, SW94a, BS92, CG96b, CL98b, CCCC96, DOR91, DF90, FHA94, FS94b, HDS91, HJ92a, HG96, KSNM91, LK91a, LS91b, MV95, Maj96, MBGT91, MOM91, NSUR98, Nik98, OPLP93, Pie90, Pie96, PK91, PR94a, PR95b, RFM90, SSL96, SFM92, Tao97, Yan90, ZWL98]. testing-module [PK91]. testing-theory [CG96b]. tests [CT96a, CAK90, CKC96, PR96b, Rag93]. text [LL90]. text-retrieval [LL90]. their [Dam92, Dou93, HNY93, Hla92, LR97, SS95, SISU90, TY92a]. Theorem [NAH98, Sch92, HP95]. theoretic [Bos90, DC93]. theoretical [Bok96, LY96, Nic98]. Theory [BMP^{+}97b, CCC96, DJM99, NKC94, NC97, Nic96, Nik91, NK92, RCSSS94, TB98b, BB92c, BB99b, BCH95, CR90b, CZSM99, Cd93, CG96b, FRU96, KS93b, Pal94, SAF90, Uht91, YK98]. thrashing [FL93, JLC98]. Three [CWM^{+}95, WM91, Dou93, HW97, Maz93, Sar90, SBP90, SBP91, TW95b]. Three-dimensional [WM91, Maz93, TW95b]. three-sided [Sar90]. three-stage [Dou93]. Threshold [GS97, Pie93, RB99, TN93, VCB96]. thresholding [CT93a]. Throughput [LP98, CS92a, CK91, HDS91]. throughput-oriented [CS92a]. Tight [AR94, LH98]. tiling [PNDN99]. Time [GL91, QM93, SF91, THH92, AA92, BOI94, BH97, Ber93a, BLOS95, BS99, CR97, CK94, CC94, CS96b, CK99, CT93b, DKT96, DB91, Fer91, FR97b, GOH92, GPS95, HL91a, HL97, HR95a, HLH96, HHS97, HS98a, HLB^{+}97, HL92b, HS94c, HS94d, HS96, HS97, HKC97, HJ92b, HBAB95, IA91, JL91, JCMH99, JLLR94, Kan90, KMM91, KS96c, KL92, KM97b, LIO94, LO94a, LMT95, MRB90, MC93, MC95, NT92, Pol95, PR92, QSA91, RT94, RHS98, SMC91, SS93a, SRL90, SRSC91, SSS97, SB94, SK94, SC95b, SS94c, SLS95, SC90c, TW95c, VRK93b, Wal97, WRS95, WNS96, WT90, WD92, ZSR90, ZS99, FTG92]. time-analysis [HJ92b]. time-constrained [HHS97, ZSR90]. Time-division [QM93]. time-freeze [CR97]. Time-space [GL91]. time/hardware [Fer91]. Timed [DLW91, LD90, ACZD94, BDL^{+}99, DGY92a, DGY92b]. timeouts [HS94c]. times [FCL98, HKC97, IOP99, KT91b, Li97, vdLS93]. Timing [TS98, ASWW92, CBRS90, Mil91, Pol95, SL95]. TLB [SS95]. TMR [CS99, KS96a, KS96b, SK94]. Token [CK90b, ACZD94, AN90]. Token-Ring [CK90b]. Tolerance [BGM^{+}95, PT92, Als92, BRS^{+}90, CDHT99, CB95b, CB96, DJ99, DA96, DM97b, FRK94, Fit97, FA96, GRR94, HJ92b, HNG94, LNS94, LMJ93, MRB90, NAB96, NG90, PMAC95, RYT95, Red98, RR93, SC94, SL92, Sto93, SCG94, TK97, VC92, WJ94, WG98b, YF92a]. Tolerant [MM95a, AKL90, ACC^{+}93, BR90, BR93, BR94, Bar92, BP98, BC95, BP95, BCH93a, BCH95, Cd93, CF90, CS90b, CU94, CUC97, CW99b, CW96, CC97b, CPK90, CS96c, DG91, DH90, DH92, FP98, Fuj99, GB91, GP97, JHK97, Kan90, KS93a, KLL97, KKS93, KH97, LS96a, LH92, LK96, LK99, LKK^{+}94, LH91, LCT97, LBF92, LKW97, Maz93, Mel92, MS95, NA90, NHA92, OPT92, PM98, PS90, PLB93, PR93b, PV94b, Rai95, RNA93, RB93, SCD^{+}95, ST95a, SJB91, SY91b, SJ93a, ST95b, SR91, Sta93, SS96, SW97, SR96, TT90, TK97, TSH94, TSH95, TK96, TYJ99, TS95, VP93, VJ93, WRS95, WH97b]. Tolerating [BCS92, Pol95, BGJ^{+}96, BBBM95, HD98, PH93]. tool [DA90, GAFN96, TMS96]. toolkit [JTK97]. top [Dal92]. topics [Fal99]. Topological [CDP98, SL91a, BBKA95, LC98a]. Topologies [HSY97, RDZ95, WH94]. topology [DDS99, PS99, SP96, YA90]. topology-based [SP96]. tori [TYJ99, TNS99]. toroidal [Pea96, TP94]. torus [Bae97, CSA94, CG94, Tse97]. torus-based [Bae97]. total [PMW92, WeBM97]. totally [BJ94, CW99a, HPN95, LT90, RFKL93, THL92]. tour [SM96a]. trace [CHH98, KHW94, Mah94a, PFP99, RBS99, TWS92, TWS93]. trace-driven [Mah94a, TWS92, TWS93]. trace-sampling [KHW94]. traces [Mah94a, RM91, TWS92, TWS93]. track [VRK93b, VRK93a]. trade [BOI94, CS96b, CSA94, DAF94, PV97, SAMC99, Wal97]. trade-off [PV97]. trade-offs [BOI94, CS96b, CSA94, DAF94, SAMC99, Wal97]. tradeoff [Fer91]. Tradeoffs [KC93b, MFH91, SRK91, SJ93b, SC90c]. Traffic [KS92, FTL93, HR95a, KBMP99, LL97, RKJZ99, UVM97]. traffic-balanced [UVM97]. TRAM [RK96]. transaction [GL92, LS97]. Transactions [Ano96a, Ano98a, GPTT98]. transcendentals [LMT98]. transfers [BH92]. transform [Man95, ZV95]. transformation [LK90, Pou97]. Transformations [HC94a, ABCW98, FC97, Har92, KRC99]. transforms [CJ90, CJ91, CT91a, Dam92, LM92c, MS99, Mie93]. transient [CRP^{+}96, YS96]. transistor [WM92]. transition [CKLY98, SM96a]. transition-tour [SM96a]. transitive [SM92a]. Translation [BBP91, Dal92, Kat97]. transmission [Kip94, KT91b, ZSR90]. transmission-line [Kip94]. transmitters [CDP98]. transmitting [LS96c]. transparency [SK96]. Transparent [KM94, EZ92, Nic96]. transport [Mei91]. Transposition [LS96a]. traveling [BJK93]. treating [YM92]. tree [BDN96, BHR94, BS90a, BH96, CB99, CU93a, DH90, LK96, LS92, O’C95, RSS97, SM93, SR98, SBP90, SBP91, SY91b, SL91c, SB92b, SCW91, Tra96, TC94, Vij91, Zhe94]. tree-match [SL91c]. tree-structured [O’C95, RSS97]. trees [BB92b, BOH90, FBM93, FA96, FL91, GH91, Jha93, KT91b, KB95, MJ92, MM91b, OZL96, PK90, SW92, SLW98, TS97, WM96, ZS99]. trellises [LV99]. triangular [HL90, LLS92]. triangularization [EL90c]. triangulation [SBP90, SBP91]. tridiagonal [DD95, LZ94, SZN92]. Trigonometric [Fow93, Smi95, Kan96]. trinomials [SK99]. triple [GB93, KGB95]. triple-adjacent [GB93, KGB95]. triplex [Vai96]. true [JLC98]. Truncated [YGI91]. truncation [BW95]. TSC [DSPN95, LF96, PEH90]. tunable [CDP98]. Tung [MPG98]. tuning [CDP98]. TWDM [LLJ97]. twin [LD94]. twin-bus [LD94]. twist [Pea96]. twisted [DMS94, ENS91, Pea96]. Two [MBNSV93, PHK94, RB99, BOH90, CCGS98, CL92a, CG96b, FRK98, Hua92, IO98, IO91, Jha93, Kan93a, LYW94, LKW97, MH90, OFW92, OZL96, Rao93b, SG90, Sas91, SBP90, SBP91, SM96a, SF96a, ST96, TA95, UVM97, Vai98, YF92a]. two-dimensional [IO98, IO91, SF96a, ST96, UVM97]. two-layer [LYW94]. Two-level [MBNSV93, BOH90, FRK98, Kan93a, LKW97, Vai98]. two-pattern [CG96b]. two-phase [TA95]. two-rail [Jha93]. two-state [SM96a]. two-valued [Sas91]. type [DBJ98, HSL96, Hua90, Kan93a, LY91, Man93, She95a, YM99]. Types [Wel94, LaF99, WdB92, WHL93].

UEC [SAR91]. UED [SAR91, THL91]. ULMs [ZV98]. UMFA [GE92]. unanimous [LYS95]. unbounded [ACH^{+}97a]. Underflow-Free [Yok92]. underlying [Tra96]. understanding [CMD93]. Undirected [Pel91]. unequal [FRK98]. unfolding [PM91]. unicasting [Wu97]. Unidirectional [DDV90, Vai95, ABB94a, BAB92, KR90b, LD94, LC90, Lin93a, MK90, Mon91, Nik91, NK92, RCSSS94, SI93, SAR91, SW95]. Unified [ABZ96, LZ94, PK94, ESS98, GSH^{+}92, HTHR94, KKS93, PLB93, WM96]. Uniform [NM94, KT99a, SF91, SF92, WPS96]. uniformization [SHC96]. Unifying [Par90, SWK90]. Unit [Lew94, ALB98, LLS91, LM99a, MC92, SJ95, SCD^{+}95, Soh90]. Unitary [HC94a]. Units [JP98, BH94, BDL^{+}99, Gun97, MC94]. universal [Gre94, Kos94, WC99]. unlimited [HTHR94]. unordered [Bos91, BJ94]. untestable [CKC96]. up/down [STE98]. Update [Leu97a]. upon [BMP97a, LCL95, YLC90]. UPP [GE92]. upper [HM94, KS90, Tem99]. Use [CC97b, KACV94, CF90, CL99a, Har92, KTM96, SST92]. used [CF95b]. user [SZTN91]. Using [GP98, KTM^{+}94, Kat94, MS91, RKJZ99, WF92, WG94, ZV98, ACFH95, ABLZ97, BCSS90, BB92b, BAA95, BT90, BS93a, BCS92, CSTO92, CR97, CT94b, CS96b, CGC95b, CMST90, CGTN92, CN99, DS94, ESS98, FM91, Fio99, GM98a, GLR^{+}99, HL91b, HNR90, Hla92, HBA98, HC96, HLP93, JG99, KS91, KPB92, KM92a, KIN^{+}94, KMD97, Kus90, LPV96, LL96a, Lee92, LS94b, MRB90, Man95, MMM93, Mit92, MY94b, NST96, OVL96, OCVSV98, OdSC96, OM91, PS99, PI93, PKI^{+}96, PB97b, Pie94, PR95a, PLJ94, RS98, RS95, SWT93, SISU90, SN99, SM90b, SK94, SS92, SRHI98, TCP91, TT90, TK97, TS97, VRK93a, VT94, Wal97, Wal98, WM98, Wu97, WHB98, YTR94, ZVDF96, ZPP91]. Utilization [GP96b, Kam92, PM98]. Utilizing [KH90]. UV [LYS95].

Validated [DM97a]. validation

[KS91, MAH98]. value

[EL90e, FFKI95, LA98, Rao93b, Sas91]. valued [AE94, BHSB97, Dam92, GM98a, HNY93, KIN^{+}94, KLPS93, LPV96, PKI^{+}96, Sas91, SB92a, SB97, TCI98, TB91, ZV95]. Valved [LK95]. Variable [Yok92, AC98, BBP^{+}96, BW96b, CS90a, Cd93, DBJ98, FR97a, FS90, IS94, MMSS96]. Variable-Length [Yok92]. variants [HV91]. variation [Efe91]. various [WdB92]. varying [KM92b]. Vector [AK92, DC93, BA93, CL99b, Dek98, Fri95, HL91b, Har92, HC93, PWG94, RH93, SL91b, Soh93, SY99, Yan93]. vector-reduction [SL91b]. vectoring [LA98]. vectorization [CY95]. vectors [DHK93, HRP99, Lin93b, SSDK95]. verifiable [KM95]. verification [Cor94a, LPV96, LY94, Mil91, Pal94, PBND98, Win95, YL95a]. verified [BSM97]. verifier [PD99]. verify [JBA^{+}97]. versa [GP96b]. versatile [IJN94, SP91, SLN97]. versus [CSA94, Egg91]. vertex [SM96b]. Vertical [PI93]. vertices [Bar98, PRS94]. Very [ELM94, LM99a, ALB98, DM97b, HS93, Kan93a]. Very-High [ELM94]. via [DL94, GL92, Ng91, PM91, Red98, SL91a, SJP96, TMS97, Uht92]. vice [GP96b]. victim [SV97]. view [BBD97, SBM^{+}98]. Virtual [BP92b, Wel94, DRS91, Hou99, KS92, WF90, WHL95]. visibility [SS90a]. vision [Sut91, TH91]. visualization [GAFN96, MS93, MS94]. VLSI [SP91, AS92a, AM91, AM93, AB98, AK91, Aln93, BOI94, BB92b, Bry91, CF95a, CJ91, CU94, CUC97, CRS98, CGTN92, CT91b, DM93, FE97, LK91a, LR97, LP90, LH90, Nic98, NV99, NLV91, OK96, RBK90, SR98, SBP90, SK92a, SCW91, SC90c, TB98a, TT90, TF92, Tua92, WP90]. VLSI-efficient [CF95a]. VLSI/ [RBK90]. VLSI/WSI [CU94, CUC97]. VOLTAN [SES^{+}92]. Volume [Ano96a, COK95]. Vote [TK91]. voting [LYS95, TK91]. vs [AB98].

wafer [DLC93, LKL98]. wafer-packing [DLC93]. wafers [RFM90]. Walk [CK99]. Walk-time [CK99]. warp [AA92, FTG92]. warranty [PZ99]. watchdog [SM90b]. waveforms [Cyr94]. wavefront [SS93b, Sto93]. Wavelength [Dow92]. way [IJW92, San93]. WDM [CDP98]. weak [SS90a]. weakly [WHB98]. Web [CL98a]. weight [AW93, Lin93a, TB98a, WH99]. Weighted [Rag96, AM93, Bar98, Cor96, Maj96, NK97, Sav99, TK91, WZXX93]. weights [Maj96]. well [CDFH93, LL91a]. well-defined [LL91a]. well-formed [CDFH93]. while [LRG91]. WICI [PA96]. Wide [CL98a, YW99]. Wide-sense [YW99]. width [MRT97]. Wildcard [BCH95]. window [Ara90, DAF94, MRJ97, SAMC99, TB99b, ZSR90]. window-based [TB99b]. window-consistent [MRJ97]. windows [Sav98b]. winner [TW95c]. winner-take-all [TW95c]. wire [Gre94]. wired [LM92b, VSS91, VS93]. wireless [DRT97, Leu97a, TPDB98]. wires [BL92]. Wiring [SWWW94, BS90b]. within [Joh93]. Without [DM96, CM96b, CL92b, EL90a, EL90b, LP90, MS99, Par94]. word [ABC92, Lew90]. work [AS92b]. workflow [PDBH97]. Workload [Hog99, CS92b]. workloads [SST92]. workstation [SKK^{+}90]. workstations [BCLR97]. World [CL98a]. wormhole [BC95, BD97, FRU96, KD94, LH98, LH91, SW97, Tse97, TYJ99, TNS99, UVM97]. wormhole-routed [Tse97, TYJ99, TNS99]. worst [BHSB97, CK91, Krä98]. worst-case [CK91]. write [BBRL98, Gai90]. write-once [BBRL98, Gai90]. writers [TN93]. writing [ST95a]. WSI [CU94, CUC97, RBK90].

XOR [CRC96, ESS98].

Yield [CU93b, HK90, KKS93, KF92, NV99, Sta93, TSH95, WL90]. yielding [Dro98].

Zero [LG95, CMH98, GL93, MC93, MC95, PG91]. Zero-aliasing [LG95, CMH98].

References

Agrawal:1992:RMP

[AA92] D. Agrawal and J. R. Agre. Recovering from multiple process failures in the time warp mechanism. IEEE Transactions on Computers, 41(12): 1504–1514, December 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                                   ieeexplore.ieee.org/stamp/
                                                   stamp.jsp?tp=&arnumber=214658.

Abu-Amara:1994:EAC

[AAL94] H. Abu-Amara and J. Lokre.

Election in asynchronous complete networks with intermittent link failures. IEEE Transactions on Computers, 43(7):778–788, July 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=293257.

Alippi:1998:AVP

[AB98] C. Alippi and L. Briozzo. Accuracy vs. precision in digital VLSI architectures for signal processing. IEEE Transactions on Computers, 47(4):472–477, April 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=675715.

Al-Bassam:1993:DEE

[ABB93] S. Al-Bassam and B. Bose. Design of efficient error-correcting balanced codes. IEEE Transactions on Computers, 42 (10):1261–1266, October 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=257712.

Al-Bassam:1994:AUE

[ABB94a] S. Al-Bassam and B. Bose.

Asymmetric/unidirectional error correcting and detecting codes. IEEE Transactions on Computers, 43(5):590–597, May 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=280797.

Al-Bassam:1994:DEB

[ABB94b] S. Al-Bassam and B. Bose. Design of efficient balanced codes. IEEE Transactions on Computers, 43(3):362–365, March 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=272436.

Arnold:1992:CAA

[ABC92] M. Arnold, T. Bailey, and J. Cowles. Comments on “An architecture for addition and subtraction of long word length numbers in the logarithmic number system” by D. M. Lewis. IEEE Transactions on Computers, 41(6):786–788, June 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=144632.

Arnold:1990:RLA

[ABCC90] M. G. Arnold, T. A. Bailey, J. R. Cowles, and J. J. Cupal. Redundant logarithmic arithmetic. IEEE Transactions on Computers, 39(8):1077–1086, August 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=57046.

Arnold:1992:AFI

[ABCW92] M. G. Arnold, T. A. Bailey, J. R. Cowles, and M. D. Winkel. Applying features of IEEE 754 to sign/logarithm arithmetic. IEEE Transactions on Computers, 41(8):1040–1050, August 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=156547.

Arnold:1998:ACT

[ABCW98] M. G. Arnold, T. A. Bailey, J. R. Cowles, and M. D. Winkel. Arithmetic co-transformations in the real and complex logarithmic number systems. IEEE Transactions on Computers, 47(7):777– 786, July 1998. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
709377.

Amano:1990:SIL

[ABK90] H. Amano, T. Boku, and T. Kudoh. (SM)^{2}-II: a large-scale multiprocessor for sparse matrix calculations. IEEE Transactions on Computers, 39(7):889–905, July 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=55691.

Antelo:1997:EAR

[ABLZ97] E. Antelo, J. D. Bruguera, T. Lang, and E. L. Zapata. Error analysis and reduction for angle calculation using the CORDIC algorithm. IEEE Transactions on Computers, 46(11):1264–1271, November 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=644300.

Antelo:1996:UMR

[ABZ96] E. Antelo, J. D. Bruguera, and E. L. Zapata. Unified mixed radix 2-4 redundant CORDIC processor. IEEE Transactions on Computers, 45 (9):1068–1073, September 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=537131.

Atallah:1995:OPH

[AC95] M. J. Atallah and D. Z. Chen. Optimal parallel hypercube algorithms for polygon problems. IEEE Transactions on Computers, 44(7):914–922, July 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=392849.

Atallah:1998:AVL

[AC98] M. J. Atallah and D. E. Comer. Algorithms for variable length subnet address assignment. IEEE Transactions on Computers, 47(6):693–699, June 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=689648.

Arlat:1993:FID

[ACC^{+}93] J. Arlat, A. Costes, Y. Crouzet, J. C. Laprie, and D. Powell. Fault injection and dependability evaluation of fault-tolerant systems. IEEE Transactions on Computers, 42(8):913–923, August 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=238482.

Alewine:1995:CAM

[ACFH95] N. J. Alewine, Shyh-Kwei Chen, W. Kent Fuchs, and W.-M. W. Hwu. Compiler-assisted multiple instruction rollback recovery using a read buffer. IEEE Transactions on Computers, 44(9):1096–1107, September 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=464388.

Albrecht:1997:OPFa

[ACH^{+}97a] A. Albrecht, S. K. Cheung, K. C. Hui, K. S. Leung, and C. K. Wong. Optimal placements of flexible objects. I. Analytical results for the unbounded case. IEEE Transactions on Computers, 46(8):890–904, August 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=609278.

Albrecht:1997:OPFb

[ACH^{+}97b] A. Albrecht, S. K. Cheung, K. C. Hui, K. S. Leung, and C. K. Wong. Optimal placements of flexible objects. II. A simulated annealing approach for the bounded case. IEEE Transactions on Computers, 46(8):905–929, August 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=609279.

Agrawal:1994:GSM

[ACZD94] G. Agrawal, B. Chen, W. Zhao, and S. Davari. Guaranteeing synchronous message deadlines with the timed token medium access control protocol. IEEE Transactions on Computers, 43(3):327–339, March 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=272433.

Apon:1997:CPM

[AD97] A. W. Apon and L. W. Dowdy. The circulating processor model of parallel systems. IEEE Transactions on Computers, 46(5):572–587, May 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=589223.

Asari:1994:OTD

[AE94] K. V. Asari and C. Eswaran.

An optimization technique for the design of multiple valued PLA’s. IEEE Transactions on Computers, 43(1):118–122, January 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=250617.

Abd-El-Barr:1990:CRC

[AEBV90] M. H. Abd-El-Barr and Z. G.

Vranesic. Cost reduction in the CCD realization of MVMT functions. IEEE Transactions on Computers, 39(5):702–706, May 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=53584.

Abd-El-Barr:1991:ASM

[AEBVZ91] M. H. Abd-El-Barr, Z. G.

Vranesic, and S. G. Zaky. Algorithmic synthesis of MVL functions for CCD implementation. IEEE Transactions on Computers, 40(8):977–986, August 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=83641.

Arora:1994:DR

[AG94] A. Arora and M. Gouda. Distributed reset. IEEE Transactions on Computers, 43(9): 1026–1038, September 1994.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=312126.

Antonio:1992:FDS

[AHT92] J. K. Antonio, G. M. Huang, and W. K. Tsai. A fast distributed shortest path algorithm for a class of hierarchically clustered data networks. IEEE Transactions on Computers, 41(6):710–724, June 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=144623.

Alnuweiri:1991:OVS

[AK91] H. M. Alnuweiri and V. K. Prasanna Kumar. Optimal VLSI sorting with reduced number of processors. IEEE Transactions on Computers, 40(1):105–110, January 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=67326.

Allen:1992:VRA

[AK92] R. Allen and K. Kennedy. Vector register allocation. IEEE Transactions on Computers, 41 (10):1290–1317, October 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=166606.

Azadmanesh:1996:NHF

[AK96] M. H. Azadmanesh and R. M.

Kieckhafer. New hybrid fault models for asynchronous approximate agreement. IEEE Transactions on Computers, 45(4):439– 449, April 1996. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
494101.

Arlat:1990:DME

[AKL90] J. Arlat, K. Kanoun, and J.-C. Laprie. Dependability modeling and evaluation of software fault-tolerant systems. IEEE Transactions on Computers, 39(4):504–513, April 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=54843.

Antelo:1998:CVH

[ALB98] E. Antelo, T. Lang, and J. D. Bruguera. Computation of \sqrt{(x/d)} in a very high radix combined division/ square-root unit with scaling and selection by rounding. IEEE Transactions on Computers, 47(2):152–161, February 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=663761.

Alnuweiri:1993:NCO

[Aln93]  H. M. Alnuweiri.  A new

class of optimal boundeddegree VLSI sorting networks. IEEE Transactions on Computers, 42(6):746–752, June 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=277299.

Ancilotti:1990:DCP

[ALPS90] P. Ancilotti, B. Lazzerini, C. A. Prete, and M. Sacchi. A distributed commit protocol for a multicomputer system. IEEE Transactions on Computers, 39(5):718–724, May 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=53589.

Alspach:1992:CGO

[Als92] B. Alspach. Cayley graphs with optimal fault tolerance. IEEE Transactions on Computers, 41(10):1337–1339, October 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=166612.

Alia:1991:VMM

[AM91] G. Alia and E. Martinelli. A VLSI modulo m multiplier. IEEE Transactions on Computers, 40(7):873–878, July 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=83626.

Alia:1993:LBV

[AM93] G. Alia and E. Martinelli. On the lower bound to the VLSI complexity of number conversion from weighted to residue representation. IEEE Transactions on Computers, 42(8):962–967, August 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=238486.

Ayoubi:1996:ECC

[AMB96] R. A. Ayoubi, Q. M. Malluhi, and M. A. Bayoumi. The extended cube connected cycles: an efficient interconnection for massively parallel systems. IEEE Transactions on Computers, 45(5):609–614, May 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=509913.

Al-Mouhamed:1996:MMN

[AMS96] M. A. Al-Mouhamed and S. S.

Seiden. Minimization of memory and network contention for accessing arbitrary data patterns in SIMD systems. IEEE Transactions on Computers, 45(6):757–762, June 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=506432.

Arvind:1990:EPT

[AN90] Arvind and R. S. Nikhil. Executing a program on the MIT tagged-token dataflow architecture. IEEE Transactions on Computers, 39(3):300–318, March 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=48862.

Annexstein:1997:GBS

[Ann97] F. S. Annexstein. Generating de Bruijn sequences: an efficient implementation. IEEE Transactions on Computers, 46 (2):198–200, February 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=565596.

Anonymous:1996:IIT

[Ano96a] Anonymous. 1996 index IEEE Transactions on Computers volume 45. IEEE Transactions on Computers, 45(12): INDEX:1–INDEX:15, December 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=545977.

Anonymous:1996:CID

[Ano96b] Anonymous. Correction to “Improved Digital Signature Algorithm”. IEEE Transactions on Computers, 45(7):864, July 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=508326.

See [YL95b].

Anonymous:1997:RL

[Ano97a] Anonymous. 1996 reviewers list. IEEE Transactions on Computers, 46(1):125–128, January 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=559813.

Anonymous:1997:RLS

[Ano97b] Anonymous. Reviewers list for special issue on mobile computing. IEEE Transactions on Computers, 46(3):371, March 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=580434.

Anonymous:1998:IIT

[Ano98a] Anonymous. 1998 index IEEE Transactions On Computers — author index. IEEE Transactions on Computers, 47(12): 1427–1431, December 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=737691.

Anonymous:1998:PIS

[Ano98b] Anonymous. Papers from the 13th IEEE Symposium on Computer Arithmetic. IEEE Transactions on Computers, 47(7):721, July 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=709371.

Anonymous:1998:SSR

[Ano98c] Anonymous. Special section referee list. IEEE Transactions on Computers, 47(7):787, July 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=709378.

Anonymous:1998:SI

[Ano98d] Anonymous. Subject index. IEEE Transactions on Computers, 47(12):1431–1440, December 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=737692.

Anonymous:1999:AI

[Ano99a] Anonymous. Author index. IEEE Transactions on Computers, 48(12):1380–1384, December 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=817404.

Anonymous:1999:SI

[Ano99b] Anonymous. Subject index. IEEE Transactions on Computers, 48(12):1384–1392, December 1999. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=817405.

Atar:1996:RSF

[ANR96] Y. Atar, J. Naor, and R. Rom.

Routing strategies for fast networks. IEEE Transactions on Computers, 45(2):165–173, February 1996. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
485380.

Ali:1994:TLB

[AR94] S. A. Ali and G. R. Redinbo. Tight lower bounds on the detection probabilities of single faults at internal signal lines in combinational circuits. IEEE Transactions on Computers, 43(12):1426–1428, December 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=338103.

Aravena:1990:RMW

[Ara90] J. L. Aravena. Recursive moving window DFT algorithm. IEEE Transactions on Computers, 39(1):145–148, January 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=46292.

Arazi:1992:CBS

[Ara92] B. Arazi. A circular binary search. IEEE Transactions on Computers, 41(1):109–112, January 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=123386.

Arazi:1993:AEA

[Ara93] B. Arazi. Architectures for exponentiation over GD(2^{n}) adopted for smartcard application. IEEE Transactions on Computers, 42(4):494–497, April 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=214694.

Auletta:1995:EGS

[ARS95] L. Auletta, A. A. Rescigno, and V. Scarano. Embedding graphs onto the Supercube. IEEE Transactions on Computers, 44(4):593–597, April 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=376173.

Afghahi:1992:PSA

[AS92a] M. Afghahi and C. Svensson. Performance of synchronous and asynchronous schemes for VLSI systems. IEEE Transactions on Computers, 41(7):858–872, July 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=256454.

Atiquzzaman:1992:MBO

[AS92b] M. Atiquzzaman and W. K.

Shehadah. A microprocessorbased office image processing system — an extension of work. IEEE Transactions on Computers, 41(3):379–381, March 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=127454.

Ahn:1993:NHM

[AS93] K. Ahn and S. Sahni. NPhard module rotation problems. IEEE Transactions on Computers, 42(12):1506–1510, December 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=260642.

Akyurek:1995:MPS

[AS95] S. Akyurek and K. Salem. Management of partially safe buffers. IEEE Transactions on Computers, 44(3):394–407, March 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=372032.

Abonamah:1994:CRF

[ASS94] A. A. Abonamah, F. N. Sibai, and N. K. Sharma. Conflict resolution and fault-free path selection in multicastconnected cube-based networks. IEEE Transactions on Computers, 43(3):374–380, March 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=272439.

Allan:1992:FIS

[ASWW92] V. H. Allan, B. Su, P. Wijaya, and J. Wang. Foresighted instruction scheduling under timing constraints. IEEE Transactions on Computers, 41(9):1169–1172, September 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=165397.

Al-Twaijry:1998:TSE

[ATF98] H. A. Al-Twaijry and M. J. Flynn. Technology scaling effects on multipliers. IEEE Transactions on Computers, 47 (11):1201–1215, November 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=736430.

Alagar:1997:COD

[AV97] S. Alagar and S. Venkatesan. Causal ordering in distributed mobile systems. IEEE Transactions on Computers, 46(3):353–361, March 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=580430.

Antelo:1997:HPR

[AVBZ97] E. Antelo, J. Villalba, J. D.

Bruguera, and E. L. Zapata. High performance rotation architectures based on the radix-4 CORDIC algorithm. IEEE Transactions on Computers, 46(8):855–870, August 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=609275.

Akella:1998:ACB

[AVR98] V. Akella, N. H. Vaidya, and G. R. Redinbo. Asynchronous comparison-based decoders for delay-insensitive codes. IEEE Transactions on Computers, 47 (7):802–811, July 1998. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
709380.

Arno:1993:SDR

[AW93] S. Arno and F. S. Wheeler.

Signed digit representations of minimal Hamming weight. IEEE Transactions on Computers, 42(8):1007–1010, August 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=238495.

Agrawal:1998:FLC

[AZ98] J. Agrawal and Yixin Zhang. A fast and low cost selfrouting permutation network. IEEE Transactions on Computers, 47(9):1033–1036, September 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=713324.

Boeri:1993:OVP

[BA93] F. Boeri and M. Auguin. OP-SILA: a vector and parallel processor. IEEE Transactions on Computers, 42(1):76–82, January 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=192215.

Bhattacharya:1995:TGP

[BAA95] D. Bhattacharya, P. Agrawal, and V. D. Agrawal. Test generation for path delay faults using binary decision diagrams. IEEE Transactions on Computers, 44(3):434–447, March 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=372035.

Bhogavilli:1997:DAH

[BAA97] S. K. Bhogavilli and H. Abu-Amara. Design and analysis of high performance multistage interconnection networks. IEEE Transactions on Computers, 46(1):110–117, January 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=559810.

Bose:1998:DLS

[BAA98] S. Bose, P. Agrawal, and V. D.

Agrawal. Deriving logic systems for path delay test generation. IEEE Transactions on Computers, 47(8):829–846, August 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=707585.

Bose:1992:BUE

[BAB92] B. Bose and S. Al-Bassam.

Byte unidirectional error correcting and detecting codes. IEEE Transactions on Computers, 41(12):1601–1606, December 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=214669.

Bae:1997:RPT

[Bae97] M. M. Bae. Resource placement in torus-based networks. IEEE Transactions on Computers, 46(10):1083–1092, October 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=628393.

Baldoni:1998:PAP

[Bal98] R. Baldoni. A positive acknowledgment protocol for causal broadcasting. IEEE Transactions on Computers, 47(12): 1341–1350, December 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=737681.

Barbour:1992:SMP

[Bar92]  A. E. Barbour.  Solutions

to the minimization problem of fault-tolerant logic circuits. IEEE Transactions on Computers, 41(4):429–443, April 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=135556.

Barbehenn:1998:NCD

[Bar98] M. Barbehenn. A note on the complexity of Dijkstra’s algorithm for graphs with weighted vertices. IEEE Transactions on Computers, 47(2):263, February 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=663776.

Balasubramanian:1990:CAS

[BB90] V. Balasubramanian and P. Banerjee. Compiler-assisted synthesis of algorithm-based checking in multiprocessors. IEEE Transactions on Computers, 39(4):436– 446, April 1990. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
54837.

Brent:1991:SPA

[BB91] R. P. Brent and Z. B. Bing.

A stabilized parallel algorithm for direct-form recursive filters. IEEE Transactions on Computers, 40(3):333–336, March 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=76411.

Belkhale:1992:PAG

[BB92a] K. P. Belkhale and P. Banerjee. Parallel algorithms for geometric connected component labeling on a hypercube multiprocessor. IEEE Transactions on Computers, 41(6):699–709, June 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=144622.

Belkhale:1992:RSV

[BB92b] K. P. Belkhale and P. Banerjee. Reconfiguration strategies for VLSI processor arrays and trees using a modified Diogenes approach. IEEE Transactions on Computers, 41(1):83–96, January 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=123383.

Bianchini:1992:IOD

[BB92c] R. P. Bianchini, Jr. and R. W. Buskens. Implementation of online distributed system-level diagnosis theory. IEEE Transactions on Computers, 41(5):616–626, May 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=142688.

Bruck:1992:NTC

[BB92d] J. Bruck and M. Blaum. New techniques for constructing EC/AUED codes. IEEE Transactions on Computers, 41 (10):1318–1324, October 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=166607.

Blaum:1995:DIP

[BB95] M. Blaum and J. Bruck. Delayinsensitive pipelined communication on parallel buses. IEEE Transactions on Computers, 44(5):660–668, May 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=381951.

Bakre:1997:IPE

[BB97] A. V. Bakre and B. R. Badrinath. Implementation and performance evaluation of Indirect TCP. IEEE Transactions on Computers, 46(3):260–278, March 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=580423.

Battiti:1999:GPR

[BB99a] R. Battiti and A. A. Bertossi.

Greedy, prohibition, and reactive heuristics for graph partitioning. IEEE Transactions on Computers, 48(4):361–385, April 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=762522.

Blough:1999:BCM

[BB99b] D. M. Blough and H. W.

Brown. The broadcast comparison model for on-line fault diagnosis in multicomputer systems: theory and implementation. IEEE Transactions on Computers, 48(5):470–493, May 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=769431.

Blazewicz:1994:OCA

[BBB^{+}94] J. Blazewicz, D. P. Bovet, J. Brzezinski, G. Gambosi, and M. Talamo. Optimal centralized algorithms for storeand-forward deadlock avoidance. IEEE Transactions on Computers, 43(11):1333–1338, November 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=324567.

Blaum:1995:EES

[BBBM95] M. Blaum, J. Brady, J. Bruck, and Jai Menon. EVEN-ODD: an efficient scheme for tolerating double disk failures in RAID architectures. IEEE Transactions on Computers, 44(2):192–202, February 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=364531.

Babaoglu:1997:EVS

[BBD97] O. Babaoglu, A. Bartoli, and G. Dini. Enriched view synchrony: a programming paradigm for partitionable asynchronous distributed systems. IEEE Transactions on Computers, 46(6):642–658, June 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=600823.

Bose:1995:LDT

[BBKA95] B. Bose, B. Broeg, Younggeun Kwon, and Y. Ashir. Lee distance and topological properties of k-ary n-cubes. IEEE Transactions on Computers, 44 (8):1021–1030, August 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=403718.

Bhattacharjee:1991:TPC

[BBP91] P. R. Bhattacharjee, S. K. Basu, and J. C. Paul. Translation of the problem of complete test set generation to pseudo-Boolean programming. IEEE Transactions on Computers, 40(7):864–867, July 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=83624.

Bhatt:1996:BRV

[BBP^{+}96] S. N. Bhatt, G. Bilardi, G. Pucci, A. Ranade, A. L. Rosenberg, and E. J. Schwabe. On bufferless routing of variable length messages in leveled networks. IEEE Transactions on Computers, 45(6):714–729, June 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=506427.

Blaum:1998:CAD

[BBRL98] M. Blaum, J. Bruck, K. Rubin, and W. Lenth. A coding approach for detection of tampering in write-once optical disks. IEEE Transactions on Computers, 47(1):120–125, January 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=656095.

Blaum:1994:NSC

[BBT94] M. Blaum, J. Bruck, and L. Tolhuizen. A note on “A systematic (12, 8) code for correcting single errors and detecting adjacent errors”. IEEE Transactions on Computers, 43(1):125, January 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=250619.

Boppana:1995:FTW

[BC95] R. V. Boppana and S. Chalasani. Fault-tolerant wormhole routing algorithms for mesh networks. IEEE Transactions on Computers, 44(7):848–864, July 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=392844.

Buckles:1997:GEI

[BC97] B. P. Buckles and S. Chanson. Guest Editors’ introduction. IEEE Transactions on Computers, 46(6):641, June 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=600822.

Bernasconi:1999:SAB

[BC99] A. Bernasconi and B. Codenotti. Spectral analysis of Boolean functions as a graph eigenvalue problem. IEEE Transactions on Computers, 48(3):345–351, March 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=755000.

Bettayeb:1996:ESN

[BCGS96] S. Bettayeb, Bin Cong, M. Girou, and I. H. Sudborough. Embedding star networks into hypercubes. IEEE Transactions on Computers, 45 (2):186–194, February 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=485371.

Bruck:1993:FTM

[BCH93a] J. Bruck, R. Cypher, and C. Ho.

Fault-tolerant meshes and hypercubes with minimal numbers of spares. IEEE Transactions on Computers, 42(9): 1089–1104, September 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=241598.

Brunner:1993:CMI

[BCH93b] H. Brunner, A. Curiger, and M. Hofstetter. On computing multiplicative inverses in GF(2^{m}). IEEE Transactions on Computers, 42(8):1010–1015, August 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=238496.

Bruck:1995:WDC

[BCH95] J. Bruck, R. Cypher, and Ching-Tien Ho. Wildcard dimensions, coding theory and faulttolerant meshes and hypercubes. IEEE Transactions on Computers, 44(1):150–155, January 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=367998.

Burke:1991:PAS

[BCLA91] J. R. Burke, C. Chen, T.- Y. Lee, and D. P. Agrawal. Performance analysis of single stage interconnection networks. IEEE Transactions on Computers, 40(3):357–365, March 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=76415.

Bhatt:1997:OSC

[BCLR97] S. N. Bhatt, F. R. K. Chung, F. T. Leighton, and A. L. Rosenberg. On optimal strategies for cycle-stealing in networks of workstations. IEEE Transactions on Computers, 46(5):545– 557, May 1997. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        589220.

Bruck:1992:TFH

[BCS92] J. Bruck, R. Cypher, and D. Soroker. Tolerating faults in hypercubes using subcube partitioning. IEEE Transactions on Computers, 41(5):599–605, May 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=142686.

Bruck:1994:ECC

[BCS94] J. Bruck, R. Cypher, and D. Soroker. Embedding cubeconnected cycles graphs into faulty hypercubes. IEEE Transactions on Computers, 43 (10):1210–1220, October 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=324546.

Barton:1990:FIE

[BCSS90] J. H. Barton, E. W. Czeck, Z. Z. Segall, and D. P. Siewiorek. Fault injection experiments using FIAT. IEEE Transactions on Computers, 39(4):575–582, April 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=54853.

Barroso:1995:PES

[BD95] L. A. Barroso and M. Dubois.

Performance evaluation of the slotted ring multiprocessor. IEEE Transactions on Computers, 44(7):878–890, July 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=392846.

Boura:1997:PAB

[BD97] Y. M. Boura and C. R. Das. Performance analysis of buffering schemes in wormhole routers. IEEE Transactions on Computers, 46(6):687–694, June 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=600826.

Bajard:1998:RMM

[BDK98] J.-C. Bajard, L.-S. Didier, and P. Kornerup. An RNS Montgomery modular multiplication algorithm. IEEE Transactions on Computers, 47(7):766–776, July 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=709376.

Benini:1999:ASL

[BDL^{+}99] L. Benini, G. De Micheli, A. Lioy, E. Macii, G. Odasso, and M. Poncino. Automatic synthesis of large telescopic units based on near-minimum timed supersetting. IEEE Transactions on Computers, 48(8):769–779, August 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=795120.

Bagherzadeh:1996:EAB

[BDN96] N. Bagherzadeh, M. Dowd, and N. Nassif. Embedding an arbitrary binary tree into the star graph. IEEE Transactions on Computers, 45(4):475–481, April 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=494105.

Brzozowski:1992:DSG

[BE92] J. A. Brzozowski and J. C.

Ebergen. On the delaysensitivity of gate networks. IEEE Transactions on Computers, 41(11):1349–1360, November 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=177306.

Belik:1990:EDA

[Bel90] F. Belik. An efficient deadlock avoidance technique. IEEE Transactions on Computers, 39 (7):882–888, July 1990. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
55690.

Berry:1992:CBE

[Ber92] R. Berry. Computer benchmark evaluation and design of experiments: a case study. IEEE Transactions on Computers, 41(10):1279–1289, October 1992. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=166605.

Berkovich:1993:OTS

[Ber93a] S. Y. Berkovich. An overlaying technique for solving linear equations in real-time computing. IEEE Transactions on Computers, 42(5):513–517, May 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=223670.

Bernhard:1993:BPM

[Ber93b]  P. J. Bernhard.  Bounds

on the performance of message routing heuristics. IEEE Transactions on Computers, 42 (10):1253–1256, October 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=257719.

Bernard:1996:EFL

[Ber96] E. G. Bernard. Efficient fault location for globally controlled and comparison-based multistage interconnection networks. IEEE Transactions on Computers, 45(12):1420–1425, December 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=545971.

Brasileiro:1996:IFS

[BES^{+}96] F. V. Brasileiro, P. D. Ezhilchelvan, S. K. Shrivastava, N. A. Speirs, and S. Tao. Implementing fail-silent nodes for distributed systems. IEEE Transactions on Computers, 45 (11):1226–1238, November 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=544479.

Bolding:1997:CCA

[BFS97] K. Bolding, M. Fulgham, and L. Snyder. The case for chaotic adaptive routing. IEEE Transactions on Computers, 46 (12):1281–1292, December 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=641929.

Beigel:1990:SOS

[BG90] R. Beigel and J. Gill. Sorting n objects with a k-sorter. IEEE Transactions on Computers, 39(5):714–716, May 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=53587.

Banatre:1996:ATP

[BGJ^{+}96] M. Banatre, A. Gefflaut, P. Joubert, C. Morin, and P. A. Lee. An architecture for tolerating processor failures in shared-memory multiprocessors. IEEE Transactions on Computers, 45(10):1101–1115, October 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=543705.

Boley:1995:FPF

[BGM^{+}95] D. Boley, G. H. Golub, S. Makar, N. Saxena, and E. J. McCluskey. Floating point fault tolerance with backward error assertions. IEEE Transactions on Computers, 44(2):302–311, February 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=364541.

Bagchi:1992:DTB

[BH92] A. Bagchi and S. L. Hakimi. Data transfers in broadcast networks. IEEE Transactions on Computers, 41(7):842–847, July 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=256452.

Bagchi:1994:IDD

[BH94] A. Bagchi and S. L. Hakimi. Information dissemination in distributed systems with faulty units. IEEE Transactions on Computers, 43(6):698–710, June 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=286303.

Blanton:1996:TCT

[BH96] R. D. Blanton and J. P. Hayes.

Testability of convergent tree circuits. IEEE Transactions on Computers, 45(8):950–963, August 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=536237.

Baruah:1997:SOR

[BH97] S. K. Baruah and J. R. Haritsa. Scheduling for overload in real-time systems. IEEE Transactions on Computers, 46 (9):1034–1039, September 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=620484.

Baruah:1998:CLS

[BH98] S. K. Baruah and M. E. Hickey.

Competitive on-line scheduling of imprecise computations. IEEE Transactions on Computers, 47(9):1027–1032, September 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=713322.

Beivide:1991:ODN

[BHBA91] R. Beivide, E. Herrada, J. L.

Balcazar, and A. Arruabarrena. Optimal distance networks of low degree for parallel computers. IEEE Transactions on Computers, 40(10):1109–1124, October 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=93744.

Bataineh:1994:CFS

[BHR94] S. Bataineh, Te-Yu Hsiung, and T. G. Robertazzi. Closed form solutions for bus and tree networks of processors load sharing a divisible job. IEEE Transactions on Computers, 43(10):1184–1196, October 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=324544.

Boros:1992:PAB

[BHS92] E. Boros, P. L. Hammer, and R. Shamir. A polynomial algorithm for balancing acyclic data flow graphs. IEEE Transactions on Computers, 41(11):1380–1385, November 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=177308.

Bagchi:1993:GDN

[BHS93] A. Bagchi, S. L. Hakimi, and E. F. Schmeichel. Gossiping in a distributed network. IEEE Transactions on Computers, 42(2):253–256, February 1993. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=204799.

Butler:1997:AWC

[BHSB97] J. T. Butler, D. S. Herscovici, T. Sasao, and R. J. Barton III. Average and worst case number of nodes in decision diagrams of symmetric multiple-valued functions. IEEE Transactions on Computers, 46(4):491–494, April 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=588065.

Belfore:1991:AFB

[BJ91] L. A. Belfore II and B. W.

Johnson. The analysis of the faulty behavior of synchronous neural networks. IEEE Transactions on Computers, 40 (12):1424–1429, December 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=106228.

Burns:1994:TSC

[BJ94] S. W. Burns and N. K. Jha.

A totally self-checking checker for a parallel unordered coding scheme. IEEE Transactions on Computers, 43(4):490–495, April 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=278488.

Bhide:1993:BNN

[BJK93] S. Bhide, N. John, and M. R. Kabuka. A Boolean neural network approach for the traveling salesman problem. IEEE Transactions on Computers, 42(10):1271–1278, October 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=257714.

Bonomi:1990:AOL

[BK90] F. Bonomi and A. Kumar. Adaptive optimal load balancing in a nonhomogeneous multiserver system with a central job scheduler. IEEE Transactions on Computers, 39 (10):1232–1250, October 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=59854.

Boubezari:1995:DBS

[BK95] S. Boubezari and B. Kaminska. A deterministic builtin self-test generator based on cellular automata structures. IEEE Transactions on Computers, 44(6):805–816, June 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=391181.

Bajard:1994:BNH

[BKM94] Jean-Claude Bajard, Sylvanus Kla, and Jean-Michel Muller.

BKM: a new hardware algorithm for complex elementary functions. IEEE Transactions on Computers, 43(8):955–963, August 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=295857.

Blair:1992:MEW

[BL92] J. R. S. Blair and E. L.

Lloyd. Minimizing external wires in generalized single-row routing. IEEE Transactions on Computers, 41(6):771–776, June 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=144629.

Baruah:1998:PSG

[BL98] S. K. Baruah and Shun-Shii Lin. Pfair scheduling of generalized pinwheel task systems. IEEE Transactions on Computers, 47(7):812–816, July 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=709381.

Bruguera:1999:LOP

[BL99] J. D. Bruguera and T. Lang.

Leading-one prediction with concurrent position correction. IEEE Transactions on Computers, 48(10):1083–1097, October 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=805157.

Burchard:1995:NSA

[BLOS95] A. Burchard, J. Liebeherr, Yingfeng Oh, and S. H. Son. New strategies for assigning real-time tasks to multiprocessor systems. IEEE Transactions on Computers, 44(12):1429–1442, December 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=477248.

Blough:1990:PAG

[BM90] D. M. Blough and G. M. Masson. Performance analysis of a generalized concurrent error detection procedure. IEEE Transactions on Computers, 39(1):47–62, January 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=46280.

Bollinger:1991:HTP

[BM91] S. W. Bollinger and S. F. Midkiff. Heuristic technique for processor and link assignment in multicomputers. IEEE Transactions on Computers, 40(3):325–333, March 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=76410.

Burns:1994:PSS

[BM94] J. Burns and C. J. Mitchell. Parameter selection for server-aided RSA computation schemes. IEEE Transactions on Computers, 43(2):163–174, February 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=262121.

Bui:1996:GAG

[BM96] Thang Nguyen Bui and Byung Ro Moon. Genetic algorithm and graph partitioning. IEEE Transactions on Computers, 45(7): 841–855, July 1996. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
508322.

Blackburn:1997:CNP

[BMP97a] S. R. Blackburn, S. Murphy, and K. G. Paterson. A comment on “A new public-key cipher system based upon the Diophantine equations”. IEEE Transactions on Computers, 46(4):512, April 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=588077.

See [LCL95].

Blackburn:1997:CTA

[BMP^{+}97b] S. R. Blackburn, S. Murphy, K. G. Paterson, S. Nandi, and P. P. Chaudhuri. Comments on “Theory and applications of cellular automata in cryptography” [and reply]. IEEE Transactions on Computers, 46(5):637–639, May 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=589245.

See [NKC94, NC97].

Bowen:1992:APA

[BNG92] N. S. Bowen, C. N. Nikolaou, and A. Ghafoor. On the assignment problem of arbitrary process systems to heterogeneous distributed computer systems. IEEE Transactions on Computers, 41(3):257–273, March 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=127439.

Bagherzadeh:1993:RBS

[BNL93] N. Bagherzadeh, N. Nassif, and S. Latifi. A routing and broadcasting scheme on faulty star graphs. IEEE Transactions on Computers, 42(11): 1398–1403, November 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=247843.

Bar-Noy:1991:SMA

[BNP91] A. Bar-Noy and D. Peleg. Square meshes are not always optimal. IEEE Transactions on Computers, 40(2):196–204, February 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=73589.

Brandman:1990:SLB

[BOH90] Y. Brandman, A. Orlitsky, and J. Hennessy. A spectral lower bound technique for the size of decision trees and two-level AND/OR circuits. IEEE Transactions on Computers, 39(2):282–287, February 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=45216.

Bajwa:1994:ATT

[BOI94] R. Singh Bajwa, R. M. Owens, and M. J. Irwin. Area time trade-offs in micro-grain VLSI array architectures. IEEE Transactions on Computers, 43 (10):1121–1128, October 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=324538.

Bokhari:1996:MCE

[Bok96]  S. H. Bokhari.  Multiphase

complete exchange: a theoretical analysis. IEEE Transactions on Computers, 45 (2):220–229, February 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=485374.

Bonomi:1990:JAP

[Bon90] F. Bonomi. On job assignment for a parallel system of processor sharing queues. IEEE Transactions on Computers, 39(7):858–869, July 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=55688.

Bose:1990:GTS

[Bos90] B. Bose. Group theoretic signature analysis. IEEE Transactions on Computers, 39(11): 1398–1403, November 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=61051.

Bose:1991:UC

[Bos91] B. Bose. On unordered codes. IEEE Transactions on Computers, 40(2):125–131, February 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=73583.

Boute:1990:CRD

[Bou90]  R. Boute.  Correction to

“Representational and denotational semantics of digital systems”. IEEE Transactions on Computers, 39(2):288, February 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=45218.

Blough:1992:CFD

[BP92a] D. M. Blough and A. Pelc.

Complexity of fault diagnosis in comparison models. IEEE Transactions on Computers, 41(3):318–324, March 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=127443.

Bowen:1992:VCA

[BP92b] N. S. Bowen and D. K. Pradhan. Virtual checkpoints: architecture and performance. IEEE Transactions on Computers, 41(5):516–525, May 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=142677.

Blough:1993:CFM

[BP93a] D. M. Blough and A. Pelc. A clustered failure model for the memory array reconfiguration problem. IEEE Transactions on Computers, 42(5):518–528, May 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=223671.

Blough:1993:DRM

[BP93b] D. M. Blough and A. Pelc.

Diagnosis and repair in multiprocessor systems. IEEE Transactions on Computers, 42 (2):205–217, February 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=204793.

Banerjee:1994:DEH

[BP94] P. Banerjee and M. Peercy. Design and evaluation of hardware strategies for reconfiguring hypercubes and meshes under faults. IEEE Transactions on Computers, 43(7):841–848, July 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=293264.

Bowen:1995:FTH

[BP95] N. S. Bowen and D. K. Pradhan. A fault tolerant hybrid memory structure and memory management algorithms. IEEE Transactions on Computers, 44(3):408–418, March 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=372033.

Bowen:1996:EPB

[BP96] N. S. Bowen and D. K. Pradhan. The effect of program behavior on fault observability. IEEE Transactions on Computers, 45(8):868–880, August 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=536230.

Bettola:1998:HPF

[BP98] S. Bettola and V. Piuri. High performance fault-tolerant digital neural networks. IEEE Transactions on Computers, 47(3):357–363, March 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=660173.

Bhatt:1993:SGM

[BPRR93] S. N. Bhatt, G. Pucci, A. Ranade, and A. L. Rosenberg. Scattering and gathering messages in networks of processors. IEEE Transactions on Computers, 42(8):938–949, August 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=238484.

Balakrishnan:1990:RMC

[BR90] M. Balakrishnan and C. S.

Raghavendra. On reliability modeling of closed faulttolerant computer systems. IEEE Transactions on Computers, 39(4):571–575, April 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=54852.

Balakrishnan:1993:ARM

[BR93] M. Balakrishnan and C. S. Raghavendra. An analysis of a reliability model for repairable fault-tolerant systems. IEEE Transactions on Computers, 42(3):327–339, March 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=210175.

Balakrishnan:1994:RMF

[BR94] M. Balakrishnan and A. Reibman. Reliability models for fault-tolerant private network applications. IEEE Transactions on Computers, 43(9): 1039–1053, September 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=312113.

Banerjee:1990:ABF

[BRS^{+}90] P. Banerjee, J. T. Rahmeh, C. Stunkel, V. S. Nair, K. Roy, V. Balasubramanian, and J. A. Abraham. Algorithm-based fault tolerance on a hypercube multiprocessor. IEEE Transactions on Computers, 39(9): 1132–1145, September 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=57055.

Bryant:1991:CVI

[Bry91] R. E. Bryant. On the complexity of VLSI implementations and graph representations of Boolean functions with application to integer multiplication. IEEE Transactions on Computers, 40(2):205–213, February 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=73590.

Biswas:1990:RTA

[BS90a] N. N. Biswas and S. Srinivas. A reconfigurable tree architecture with multistage interconnection network. IEEE Transactions on Computers, 39 (12):1481–1485, December 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=61069.

Brady:1990:SKK

[BS90b] M. L. Brady and M. Sarrafzadeh. Stretching a knockknee layout for multilayer wiring. IEEE Transactions on Computers, 39(1):148–151, January 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=46293.

Butler:1990:ECF

[BS90c] J. T. Butler and K. A. Schueller.

On the equivalence of cost functions in the design of circuits by cost-table. IEEE Transactions on Computers, 39(6):842–844, June 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=53608.

Boneh:1992:SRD

[BS92] A. Boneh and J. Savir. Statistical resistance to detection [digital circuits testing]. IEEE Transactions on Computers, 41(1):123–126, January 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=123388.

Brand:1993:MEE

[BS93a] D. Brand and T. Sasao. Minimization of AND–EXOR expressions using rewrite rules. IEEE Transactions on Computers, 42(5):568–576, May 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=223676.

Braun:1993:ODA

[BS93b] H. Braun and F. C. Stephan. On optimizing diameter and average distance of directed interconnected networks. IEEE Transactions on Computers, 42(3):353–358, March 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=210177.

Bodin:1997:SAI

[BS97] F. Bodin and A. Seznec. Skewed associativity improves program performance and enhances predictability. IEEE Transactions on Computers, 46(5):530–544, May 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=589219.

Buttazzo:1999:ODA

[BS99] G. C. Buttazzo and F. Sensini.

Optimal deadline assignment for scheduling soft aperiodic tasks in hard real-time environments. IEEE Transactions on Computers, 48(10):1035–1052, October 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=805154.

Blough:1992:EDM

[BSM92a] D. M. Blough, G. F. Sullivan, and G. M. Masson. Efficient diagnosis of multiprocessor systems under probabilistic models. IEEE Transactions on Computers, 41(9):1126–1136, September 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=165394.

Blough:1992:IFD

[BSM92b] D. M. Blough, G. F. Sullivan, and G. M. Masson. Intermittent fault diagnosis in multiprocessor systems. IEEE Transactions on Computers, 41 (11):1430–1441, November 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=177313.

Bright:1997:FVS

[BSM97] J. D. Bright, G. F. Sullivan, and G. M. Masson. A formally verified sorting certifier. IEEE Transactions on Computers, 46(12):1304–1312, December 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=641931.

Banerjia:1998:MMP

[BSMC98] S. Banerjia, S. W. Sathaye, K. N. Menezes, and T. M. Conte. MPS: miss-path scheduling for multiple-issue processors. IEEE Transactions on Computers, 47(12):1382–1397, December 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=737684.

Buonanno:1994:ISC

[BSS94] G. Buonanno, D. Sciuto, and R. Stefanelli. Innovative structures for CMOS combinational gates synthesis. IEEE Transactions on Computers, 43(4):385– 399, April 1994. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
278478.

Bobbio:1990:CCM

[BT90] A. Bobbio and K. Trivedi.

Computing cumulative measures of stiff Markov chains using aggregation. IEEE Transactions on Computers, 39 (10):1291–1298, October 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=59859.

Burgess:1995:COT

[BW95] N. Burgess and T. Williams.

Choices of operand truncation in the SRT division algorithm. IEEE Transactions on Computers, 44(7):933–937, July 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=392852.

Blum:1996:RPD

[BW96a] Manuel Blum and H. Wasserman. Reflections on the Pentium division bug. IEEE Transactions on Computers, 45(4): 385–393, April 1996. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://http.cs.

                                        berkeley.edu/~blum/pentium.
                                        ps;   https://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=494097.

Bollig:1996:IVO

[BW96b] B. Bollig and I. Wegener. Improving the variable ordering of OBDDs is NP-complete. IEEE Transactions on Computers, 45(9):993–1002, September 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=537122.

Bubenik:1992:OMS

[BZ92] R. Bubenik and W. Zwaenepoel. Optimistic make [software design]. IEEE Transactions on Computers, 41(2):207–217, February 1992. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        123396.

Chatterjee:1990:TGC

[CA90a] A. Chatterjee and J. A. Abraham. The testability of generalized counters under multiple faulty cells. IEEE Transactions on Computers, 39(11): 1378–1385, November 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=61053.

Cheng:1990:PSM

[CA90b] K.-T. Cheng and V. D. Agrawal.

A partial scan method for sequential circuits with feedback. IEEE Transactions on Computers, 39(4):544–548, April 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=54847.

Chamberlain:1991:BCE

[CA91a] S. C. Chamberlain and P. D.

Amer. Broadcast channels in Estelle. IEEE Transactions on Computers, 40(4):423–436, April 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=88462.

Chatterjee:1991:TGI

[CA91b] A. Chatterjee and J. A. Abraham. Test generation for iterative logic arrays based on an N-cube of cell states model. IEEE Transactions on Computers, 40(10):1133–1148, October 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=93746.

Cheng:1992:ICS

[CA92] K.-T. Cheng and V. D. Agrawal.

Initializability consideration in sequential machine synthesis. IEEE Transactions on Computers, 41(3):374–379, March 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=127453.

Cheng:1990:SBM

[CAK90] K.-T. Cheng, V. D. Agrawal, and E. S. Kuh. A simulationbased method for generating tests for sequential circuits. IEEE Transactions on Computers, 39(12):1456–1463, December 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=61065.

Chang:1995:CAS

[CB95a] Y. Chang and L. N. Bhuyan. A combinatorial analysis of subcube reliability in hypercubes. IEEE Transactions on Computers, 44(7):952–956, July 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=392856.

Chang:1995:SFT

[CB95b] Y. Chang and L. N. Bhuyan. Subcube fault tolerance in hypercube multiprocessors. IEEE Transactions on Computers, 44 (9):1108–1120, September 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=464389.

Chen:1995:EHB

[CB95c] Tien-Fu Chen and Jean-Loup Baer. Effective hardwarebased data prefetching for high-performance processors. IEEE Transactions on Computers, 44(5):609–623, May 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=381947.

Chowdhury:1996:NEA

[CB96] A.-R. Chowdhury and P. Banerjee. A new error analysis based method for tolerance computation for algorithm-based checks. IEEE Transactions on Computers, 45(2):238–243, February 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=485376.

Chalasani:1997:CMN

[CB97] S. Chalasani and R. V. Boppana. Communication in multicomputers with nonconvex faults. IEEE Transactions on Computers, 46(5):616–622, May 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=589238.

Chang:1999:ETC

[CB99] Yeimkuan Chang and L. N.

Bhuyan. An efficient tree cache coherence protocol for distributed shared memory multiprocessors. IEEE Transactions on Computers, 48(3):352–360, March 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=755001.

Chowdhury:1994:DCC

[CBGC94] D. R. Chowdhury, S. Basu, I. S. Gupta, and P. P. Chaudhuri. Design of CAECC — Cellular Automata Based Error Correcting Code. IEEE Transactions on Computers, 43(6):759–764, June 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=286310.

Champernowne:1990:LLT

[CBRS90] A. F. Champernowne, L. B. Bushard, J. T. Rusterholz, and J. R. Schomburg. Latch-to-latch timing rules. IEEE Transactions on Computers, 39(6):798–808, June 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=53601.

Chan:1992:GSP

[CC92] M. Y. Chan and F. Y. L. Chin. General schedulers for the pinwheel problem based on double-integer reduction. IEEE Transactions on Computers, 41(6):755–768, June 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=144627.

Chen:1994:CTS

[CC94] Y.-C. Chen and W.-T. Chen.

Constant time sorting on reconfigurable meshes. IEEE Transactions on Computers, 43(6):749– 751, June 1994. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
286307.

Chen:1996:SCD

[CC96a] C. L. Chen and B. W. Curran. Switching codes for delta-I noise reduction. IEEE Transactions on Computers, 45(9): 1017–1021, September 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=537124.

Cheng:1996:NSR

[CC96b] Wang-Jiunn Cheng and Wen-Tsuen Chen. A new selfrouting permutation network. IEEE Transactions on Computers, 45(5):630–636, May 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=509917.

Chen:1997:OPR

[CC97a] Chi-Chang Chen and Jianer Chen. Optimal parallel routing in star networks. IEEE Transactions on Computers, 46 (12):1293–1303, December 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=641930.

Chiu:1997:URC

[CC97b] Ge-Ming Chiu and Kai-Shung Chen. Use of routing capability for fault-tolerant routing in hypercube multicomputers. IEEE Transactions on Computers, 46(8):953–958, August 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=609283.

Chattopadhyay:1998:CAA

[CCBC98] S. Chattopadhyay, D. F. Chowdhury, S. Bhattacharjee, and P. P. Chaudhuri. Cellular-automata-array-based diagnosis of board level faults. IEEE Transactions on Computers, 47(8):817–828, August 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=707584.

Chakraborty:1996:TAN

[CCC96] S. Chakraborty, D. Roy Chowdhury, and P. Pal Chaudhuri. Theory and application of nongroup cellular automata for synthesis of easily testable finite state machines. IEEE Transactions on Computers, 45(7):769–781, July 1996. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=508316.

Choy:1996:TGD

[CCCC96] O. C.-S. Choy, Lap-Kong Chan, R. Chan, and C. F. Chan. Test generation with dynamic probe points in high observability testing environment. IEEE Transactions on Computers, 45(1):88–96, January 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=481489.

Chen:1993:ECE

[CCCH93] W. Y. Chen, P. P. Chang, T. M.

Conte, and W. W. Hwu. The effect of code expanding optimizations on instruction cache design. IEEE Transactions on Computers, 42(9):1045–1057, September 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=241594.

Campbell:1992:SDS

[CCD^{+}92] L. Campbell, G. E. Carlsson, M. J. Dinneen, V. Faber, M. R. Fellows, M. A. Langston, J. W. Moore, A. P. Mullhaupt, and H. B. Sexton. Small diameter symmetric networks from linear groups. IEEE Transactions on Computers, 41 (2):218–220, February 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=123397.

Cameron:1991:LEP

[CCG^{+}91] E. J. Cameron, D. M. Cohen, T. M. Guinther, W. M. Keese, Jr., L. A. Ness, C. Norman, and H. N. Srinidhi. The L.0 language and environment for protocol simulation and prototyping. IEEE Transactions on Computers, 40(4):562–571, April 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=88474.

Carpenter:1998:SAA

[CCGS98] T. Carpenter, S. Cosares, J. L. Ganley, and I. Saniee. A simple approximation algorithm for two problems in circuit design. IEEE Transactions on Computers, 47(11):1310–1312, November 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=736445.

Chung:1992:CSA

[CCL92] Kuo-Liang Chung, Wen-Chin Chen, and Ferng-Ching Lin. On the complexity of search algorithms. IEEE Transactions on Computers, 41(9): 1172–1176, September 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=165398.

Chatterjee:1993:DFT

[Cd93] A. Chatterjee and M. A.

d’Abreu. The design of faulttolerant linear digital state variable systems: theory and techniques. IEEE Transactions on Computers, 42(7):794–808, July 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=237720.

Chiola:1993:SWF

[CDFH93] G. Chiola, C. Dutheillet, G. Franceschinis, and S. Haddad. Stochastic well-formed colored nets and symmetric modeling applications. IEEE Transactions on Computers, 42(11): 1343–1360, November 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=247838.

Colbourn:1993:ARM

[CDHK93] C. J. Colbourn, J. S. Devitt, D. D. Harms, and M. Kraetzl. Assessing reliability of multistage interconnection networks. IEEE Transactions on Computers, 42(10):1207–1221, October 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=257707.

Cao:1999:FTP

[CDHT99] Feng Cao, Ding-Zhu Du, F. Hsu, and Shang-Hua Teng. Fault tolerance properties of pyramid networks. IEEE Transactions on Computers, 48(1):88–93, January 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=743415.

Cao:1998:TEW

[CDP98] Feng Cao, D. H. C. Du, and A. Pavan. Topological embedding into WDM optical passive star networks with tunable transmitters of limited tuning range. IEEE Transactions on Computers, 47(12):1404–1413, December 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=737686.

Chean:1990:FUS

[CF90] M. Chean and J. A. B. Fortes. The full-use-of-suitable-spares (FUSS) approach to hardware reconfiguration for fault-tolerant processor arrays. IEEE Transactions on Computers, 39(4):564–571, April 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=54851.

Chlamtac:1994:OCA

[CF94] I. Chlamtac and A. Farago. An optimal channel access protocol with multiple reception capacity. IEEE Transactions on Computers, 43(4):480–484, April 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=278486.

Cam:1995:FVE

[CF95a] H. Cam and J. A. B. Fortes.

A fast VLSI-efficient self-routing permutation network. IEEE Transactions on Computers, 44(3):448–453, March 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=372036.

Cam:1995:FSC

[CF95b] H. Cam and J. A. B. Fortes.

Frames: a simple characterization of permutations realized by frequently used networks. IEEE Transactions on Computers, 44(5):695–697, May 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=381954.

Cherkassky:1991:LAA

[CFV91] V. Cherkassky, K. Fassett, and N. Vassilas. Linear algebra approach to neural associative memories and noise performance of neural classifiers. IEEE Transactions on Computers, 40(12):1429–1435, December 1991. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=106229.

Cypher:1994:SED

[CG94] R. Cypher and L. Gravano. Storage-efficient, deadlock-free packet routing algorithms for torus networks. IEEE Transactions on Computers, 43(12): 1376–1385, December 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=338097.

Chang:1996:HPM

[CG96a] J. M. Chang and E. F. Gehringer. A high performance memory allocator for object-oriented systems. IEEE Transactions on Computers, 45(3):357–366, March 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=485574.

Chen:1996:TPG

[CG96b] Chih-Ang Chen and S. K. Gupta. BIST test pattern generators for two-pattern testingtheory and design algorithms. IEEE Transactions on Computers, 45(3):257–269, March 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=485565.

Cobb:1997:RRF

[CG97] J. A. Cobb and M. G. Gouda.

The request reply family of group routing protocols. IEEE Transactions on Computers, 46(6): 659–672, June 1997. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
600824.

Chowdhury:1995:CBB

[CGC95a] D. R. Chowdhury, I. S. Gupta, and P. P. Chaudhuri. CAbased byte error-correcting code. IEEE Transactions on Computers, 44(3):371–382, March 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=372030.

Chowdhury:1995:LCH

[CGC95b] D. R. Chowdhury, I. S. Gupta, and P. P. Chaudhuri. A low-cost high-capacity associative memory design using cellular automata. IEEE Transactions on Computers, 44 (10):1260–1264, October 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=467703.

Chlamtac:1993:HIS

[CGK93] I. Chlamtac, A. Ganz, and M. G. Kienzle. An HIPPI interconnection system. IEEE Transactions on Computers, 42 (2):138–150, February 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=204790.

Clarkson:1992:LPR

[CGTN92] T. G. Clarkson, D. Gorse, J. G. Taylor, and C. K. Ng. Learning probabilistic RAM nets using VLSI structures. IEEE Transactions on Computers, 41(12):1552–1561, December 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=214663.

Chakravarty:1990:CSP

[CH90] S. Chakravarty and H. B. Hunt III. On computing signal probability and detection probability of stuck-at faults. IEEE Transactions on Computers, 39(11):1369–1377, November 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=61046.

Chakravarty:1993:CBD

[Cha93] S. Chakravarty. A characterization of binary decision diagrams. IEEE Transactions on Computers, 42(2):129–137, February 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/

stamp.jsp?tp=&arnumber=204789.

See comments [Weg94a].

Chakravarty:1996:SDF

[Cha96] S. Chakravarty. Synthesis of delay fault testability circuits. IEEE Transactions on Computers, 45(8):985–991, August 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=536242.

Chakrabarti:1999:PPC

[Cha99] P. P. Chakrabarti. Partial precedence constrained scheduling. IEEE Transactions on Computers, 48(10):1127–1130, October 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=805161.

Chen:1992:SEC

[Che92] C. L. Chen. Symbol errorcorrecting codes for computer memory systems. IEEE Transactions on Computers, 41 (2):252–256, February 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=123405.

Conte:1998:CTS

[CHH98] T. M. Conte, M. A. Hirsch, and W.-M. W. Hwu. Combining trace sampling with single pass methods for efficient cache simulation. IEEE Transactions on Computers, 47(6):714–720, June 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=689650.

Chi:1994:SDC

[Chi94] V. L. Chi. Salphasic distribution of clock signals for synchronous systems. IEEE Transactions on Computers, 43(5):597–602, May 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=280806.

Choi:1992:FEE

[CI92] G. S. Choi and R. K. Iyer. FO-CUS: an experimental environment for fault sensitivity analysis. IEEE Transactions on Computers, 41(12):1515–1526, December 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=214660.

Chakrabarti:1990:SAC

[CJ90] C. Chakrabarti and J. Jaja. Systolic architectures for the computation of the discrete Hartley and the discrete cosine transforms based on prime factor decomposition. IEEE Transactions on Computers, 39 (11):1359–1368, November 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=61045.

Chakrabarti:1991:VAM

[CJ91] C. Chakrabarti and J. JaJa.

VLSI architectures for multidimensional transforms. IEEE Transactions on Computers, 40 (9):1053–1057, September 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=83648.

Casavant:1990:CFA

[CK90a] T. L. Casavant and J. G. Kuhl. A communicating finite automata approach to modeling distributed computation and its application to distributed decisionmaking. IEEE Transactions on Computers, 39(5):628–639, May 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=53576.

Choi:1990:AAE

[CK90b] M. Choi and C. M. Krishna.

An adaptive algorithm to ensure differential service in a tokenring network. IEEE Transactions on Computers, 39(1):19–33, January 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=46278.

Choi:1991:EWC

[CK91] Dug-Kyoo Choi and B. G. Kim. The expected (not worstcase) throughput of the Ethernet protocol. IEEE Transactions on Computers, 40(3):245–252, March 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=76401.

Chen:1994:CAT

[CK94] Chien-In Henry Chen and A. Kumar. Comments on “Areatime optimal adder design”. IEEE Transactions on Computers, 43(4):507–512, April 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=278491.

Chen:1999:WTA

[CK99] Chien-Ming Chen and Chung-Ta King. Walk-time address adjustment for improving the accuracy of dynamic branch prediction. IEEE Transactions on Computers, 48(5):457–469, May 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=769430.

Cheng:1996:GHQ

[CKC96] Kwang-Ting Cheng, A. Krstic, and Hsi-Chuan Chen. Generation of high quality tests for robustly untestable path delay faults. IEEE Transactions on Computers, 45(12): 1379–1392, December 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=545968.

Cortadella:1998:DPN

[CKLY98] J. Cortadella, M. Kishinevsky, L. Lavagno, and A. Yakovlev. Deriving Petri nets from finite transition systems. IEEE Transactions on Computers, 47(8):859–882, August 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=707587.

Choi:1999:SAS

[CKY^{+}99] Hoon Choi, Jong-Sun Kim, Chi-Won Yoon, In-Cheol Park, Sung Ho Hwang, and Chong-Min Kyung. Synthesis of application specific instructions for embedded DSP software. IEEE Transactions on Computers, 48(6):603–614, June 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=773797.

Chang:1990:DAB

[CL90] P. R. Chang and C. S. G. Lee. A decomposition approach for balancing large-scale acyclic data flow graphs. IEEE Transactions on Computers, 39(1):34–46, January 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=46279.

Chen:1992:CPP

[CL92a] G.-I. Chen and T.-H. Lai. Constructing parallel paths between two subcubes. IEEE Transactions on Computers, 41(1):118–123, January 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=123387.

Cortadella:1992:ECC

[CL92b] J. Cortadella and J. M. Llaberia. Evaluation of A + B = K conditions without carry propagation. IEEE Transactions on Computers, 41(11):1484–1488, November 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=177318.

See comments [?].

Chen:1994:DCS

[CL94a] Deng-Jyi Chen and Min-Sheng Lin. On distributed computing systems reliability analysis under program execution constraints. IEEE Transactions on Computers, 43(1):87–97, January 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=250612.

Ciardo:1994:CAS

[CL94b] G. Ciardo and C. Lindemann.

Comments on “Analysis of selfstabilizing clock synchronization by means of stochastic Petri nets”. IEEE Transactions on Computers, 43(12): 1453–1456, December 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=338111.

Cortadella:1994:HRD

[CL94c] J. Cortadella and T. Lang.

High-radix division and squareroot with speculation. IEEE Transactions on Computers, 43(8):919–931, August 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=295854.

Cull:1995:MC

[CL95] P. Cull and S. M. Larson. The Möbius cubes. IEEE Transactions on Computers, 44(5):647–659, May 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=381950.

Chang:1996:EEI

[CL96] Meng-Chou Chang and Feipei Lai. Efficient exploitation of instruction-level parallelism for superscalar processors by the conjugate register file scheme. IEEE Transactions on Computers, 45(3):278–293, March 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=485567.

Chlamtac:1997:MCW

[CL97] I. Chlamtac and Yi-Bing Lin. Mobile computing: When mobility meets computation. IEEE Transactions on Computers, 46(3):257–259, March 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=580421.

Cao:1998:MSC

[CL98a] Pei Cao and Chengjie Liu. Maintaining strong cache consistency in the World Wide Web. IEEE Transactions on Computers, 47(4):445–457, April 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=675713.

Chess:1998:LTB

[CL98b] B. Chess and T. Larrabee. Logic testing of bridging faults in CMOS integrated circuits. IEEE Transactions on Computers, 47(3):338–345, March 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=660170.

Chen:1999:EUA

[CL99a] Chung-Ho Chen and Feng-Fu Lin. An easy-to-use approach for practical bus-based system design. IEEE Transactions on Computers, 48(8):780–793, August 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=795121.

Corral:1999:MCB

[CL99b] A. M. Dal Corral and J. M.

Llaberia. Minimizing conflicts between vector streams in interleaved memory systems. IEEE Transactions on Computers, 48(4):449–456, April 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=762540.

Chen:1992:SAN

[CLC92] Yirng-An Chen, Youn-Long Lin, and Long-Wen Chang. A systolic algorithm for the k-nearest neighbors problem. IEEE Transactions on Computers, 41(1):103–108, January 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=123385.

Chung:1990:SPJ

[CLL90] J.-Y. Chung, J. W. S. Liu, and K.-J. Lin. Scheduling periodic jobs that allow imprecise results. IEEE Transactions on Computers, 39(9):1156–1174, September 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=57057.

Chen:1991:NGH

[CLL91] G. I. Chen, Y. H. Lai, and Y.- N. Lien. A note on “Generalized hypercube and hyperbus structures for a computer network” by L. N. Bhuyan and D. P. Agrawal. IEEE Transactions on Computers, 40(5):667–668, May 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=88492.

Chang:1995:IPC

[CLM^{+}95] P. P. Chang, D. M. Lavery, S. A. Mahlke, W. Y. Chen, and W.- M. W. Hwu. The importance of prepass code scheduling for superscalar and superpipelined processors. IEEE Transactions on Computers, 44(3):353–370, March 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=372029.

Ciminiera:1990:HRS

[CM90] L. Ciminiera and P. Montuschi. Higher radix square rooting.

IEEE Transactions on Computers, 39(10):1220–1231, October 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=59853.

Cappello:1992:SSS

[CM92] P. R. Cappello and W. L. Miranker. Systolic super summation with reduced hardware. IEEE Transactions on Computers, 41(3):339–342, March 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=127445.

Cattell:1996:AOD

[CM96a] K. Cattell and J. C. Muzio.

Analysis of one-dimensional linear hybrid cellular automata over GF(q). IEEE Transactions on Computers, 45(7):782–792, July 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=508317.

Ciminiera:1996:CSM

[CM96b] L. Ciminiera and P. Montuschi. Carry-save multiplication schemes without final addition. IEEE Transactions on Computers, 45(9):1050–1055, September 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=537128.

Chung:1993:PCM

[CMD93] Sang-Hwa Chung, D. I. Moldovan, and R. F. DeMara. A parallel computational model for integrated speech and natural language understanding. IEEE Transactions on Computers, 42(10):1171–1183, October 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=257704.

Chakrabarty:1998:OZA

[CMH98] K. Chakrabarty, B. T. Murray, and J. P. Hayes. Optimal zero-aliasing space compaction of test responses. IEEE Transactions on Computers, 47 (11):1171–1187, November 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=736427.

Chang:1999:PSC

[CMS99] D. Chang and M. Marek-Sadowska. Partitioning sequential circuits on dynamically reconfigurable FPGAs. IEEE Transactions on Computers, 48(6):565–578, June 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=773794.

Ciardo:1990:PAU

[CMST90] G. Ciardo, R. A. Marie, B. Sericola, and K. S. Trivedi. Performability analysis using semi-Markov reward processes. IEEE Transactions on Computers, 39 (10):1251–1264, October 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=59855.

Conway:1999:FCM

[CN99] R. Conway and J. Nelson. Fast converter for 3 moduli RNS using new property of CRT. IEEE Transactions on Computers, 48(8):852–860, August 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=795127.

Cohen:1993:OBD

[Coh93] R. Cohen. One-bit delay in ring networks. IEEE Transactions on Computers, 42(6):735–737, June 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=277292.

See response [KH93].

Cohen-Or:1995:SSS

[COK95] D. Cohen-Or and A. Kaufman. A 3D skewing and deskewing scheme for conflict-free access to rays in volume rendering. IEEE Transactions on Computers, 44(5):707–710, May 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=381958.

Corella:1994:AVB

[Cor94a] F. Corella. Automated verification of behavioral equivalence for microprocessors. IEEE Transactions on Computers, 43(1):115–117, January 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=250616.

Corinthios:1994:OPP

[Cor94b]  M. J. Corinthios.  Optimal

parallel and pipelined processing through a new class of matrices with application to generalized spectral analysis. IEEE Transactions on Computers, 43(4):443–459, April 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=278482.

Corinthios:1996:WSP

[Cor96] M. J. Corinthios. A weighted Z spectrum, parallel algorithm, and processors for mathematical model estimation. IEEE Transactions on Computers, 45(5):513–528, May 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=509904.

Cukier:1999:CEM

[CPA99] M. Cukier, D. Powell, and J. Ariat. Coverage estimation methods for stratified faultinjection. IEEE Transactions on Computers, 48(7):707–723, July 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=780878.

Choi:1990:IHI

[CPK90] B.-R. Choi, K. H. Park, and M. Kim. An improved hardware implementation of the faulttolerant clock synchronization algorithm for large multiprocessor systems. IEEE Transactions on Computers, 39(3):404–407, March 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=48872.

Catania:1995:DPA

[CPRV95] V. Catania, A. Puliafito, S. Riccobene, and L. Vita. Design and performance analysis of a disk array system. IEEE Transactions on Computers, 44(10):1236–1247, October 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=467698.

Caspi:1991:PAE

[CPV91] P. Caspi, J. Piotrowski, and R. Velazco. An a priori approach to the evaluation of signature analysis efficiency. IEEE Transactions on Computers, 40(9):1068–1071, September 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=83653.

Carter:1990:RSD

[CR90a] T. M. Carter and J. E. Robertson. Radix-16 signeddigit division. IEEE Transactions on Computers, 39(12): 1424–1433, December 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=61063.

Carter:1990:STA

[CR90b] T. M. Carter and J. E. Robertson. The set theory of arithmetic decomposition. IEEE Transactions on Computers, 39(8):993–1005, August 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=57037.

Choi:1993:MLS

[CR93] A. Choi and M. Ruschitzka. Managing locality sets: the model and fixed-size buffers. IEEE Transactions on Computers, 42(2):190–204, February 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=204792.

Chatterjee:1997:CED

[CR97] A. Chatterjee and R. K. Roy.

Concurrent error detection in nonlinear digital circuits using time-freeze linearization. IEEE Transactions on Computers, 46(11):1208–1218, November 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=644296.

Chattopadhyay:1996:SHT

[CRC96] S. Chattopadhyay, S. Roy, and P. P. Chaudhuri. Synthesis of highly testable fixed-polarity AND–XOR canonical networks — a genetic algorithm-based approach. IEEE Transactions on Computers, 45(4):487–490, April 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=494107.

Card:1998:CLA

[CRMM98] H. C. Card, G. K. Rosendahl, D. K. McNeill, and R. D. McLeod. Competitive learning algorithms and neurocomputer architecture. IEEE Transactions on Computers, 47(8):847–858, August 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=707586.

Cha:1996:GLS

[CRP^{+}96] Hungse Cha, E. M. Rudnick, J. H. Patel, R. K. Iyer, and G. S. Choi. A gate-level simulation environment for alphaparticle-induced transient faults. IEEE Transactions on Computers, 45(11):1248–1256, November 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=544481.

Cho:1998:FAA

[CRS98] Jun-Dong Cho, S. Raje, and M. Sarrafzadeh. Fast approximation algorithms on maxcut, k-coloring, and k-color ordering for VLSI applications. IEEE Transactions on Computers, 47(11):1253–1266, November 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=736440.

Chan:1990:ADC

[CS90a] P. K. Chan and M. D. F. Schlag. Analysis and design of CMOS Manchester adders with variable carry-skip. IEEE Transactions on Computers, 39(8):983–992, August 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=57038.

Chen:1990:AFT

[CS90b] M.-S. Chen and K. G. Shin. Adaptive fault-tolerant routing in hypercube multicomputers. IEEE Transactions on Computers, 39(12):1406–1416, December 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=61061.

Chen:1990:SAT

[CS90c] M.-S. Chen and K. G. Shin.

Subcube allocation and task migration in hypercube multiprocessors. IEEE Transactions on Computers, 39(9): 1146–1155, September 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=57056.

Chen:1991:PAM

[CS91] Wen Tsuen Chen and Jang-Pin Sheu. Performance analysis of multiple bus interconnection networks with hierarchical requesting model. IEEE Transactions on Computers, 40(7):834–842, July 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=83621.

Chiang:1992:EDC

[CS92a] M.-C. Chiang and G. S. Sohi.

Evaluating design choices for shared bus multiprocessors in a throughput-oriented environment. IEEE Transactions on Computers, 41(3):297–317, March 1992. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=127442.

Czeck:1992:OEF

[CS92b] E. W. Czeck and D. P. Siewiorek. Observations on the effects of fault manifestation as a function of workload. IEEE Transactions on Computers, 41(5):559–566, May 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=142682.

Chang:1996:LSH

[CS96a] Yi-Chieh Chang and K. G. Shin. Load sharing in hypercubeconnected multicomputers in the presence of node failures. IEEE Transactions on Computers, 45(10):1203–1211, October 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=543714.

Chen:1996:ATT

[CS96b] Chung-Ho Chen and A. K. Somani. Architecture technique trade-offs using mean memory delay time. IEEE Transactions on Computers, 45 (10):1089–1100, October 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=543704.

Choi:1996:DPA

[CS96c] Sang Bang Choi and A. K.

Somani. Design and performance analysis of loaddistributing fault-tolerant network. IEEE Transactions on Computers, 45(5):540–551, May 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=509906.

Chen:1999:FCC

[CS99] Chung-Ho Chen and A. K.

Somani. Fault-containment in cache memories for TMR redundant processor systems. IEEE Transactions on Computers, 48(4):386–397, April 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=762529.

Chung:1994:CPT

[CSA94] T. Y. Chung, N. Sharma, and D. P. Agrawal. Cost-performance trade-offs in Manhattan Street Network versus 2-D torus. IEEE Transactions on Computers, 43(2):240–243, February 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=262131.

Chang:1990:ODP

[CSF90] M.-F. Chang, W. Shi, and W. K. Fuchs. Optimal diagnosis procedures for k-out-of-n structures. IEEE Transactions on Computers, 39(4):559–564, April 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=54850.

Chen:1990:ARB

[CSK90] M.-S. Chen, K. G. Shin, and D. D. Kandlur. Addressing, routing, and broadcasting in hexagonal mesh multiprocessors. IEEE Transactions on Computers, 39(1):10–18, January 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=46277.

Cypher:1990:AIC

[CSS90] R. E. Cypher, J. L. C. Sanz, and L. Snyder. Algorithms for image component labeling on SIMD mesh-connected computers. IEEE Transactions on Computers, 39(2):276–281, February 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=45215.

Cheung:1990:DAG

[CSSP90] K. C. Cheung, G. S. Sohi, K. K. Salvia, and D. K. Pradhan. Design and analysis of a gracefully degrading interleaved memory system. IEEE Transactions on Computers, 39(1):63–71, January 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=46281.

Chan:1992:DOC

[CSTO92] P. K. Chan, M. D. F. Schlag, C. D. Thomborson, and V. G. Oklobdzija. Delay optimization of carry-skip adders and block carry-lookahead adders using multidimensional dynamic programming. IEEE Transactions on Computers, 41(8):920–930, August 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=156534.

Choinski:1991:GDR

[CT91a] T. C. Choinski and T. T. Tylaska. Generation of digit reversed address sequences for fast Fourier transforms. IEEE Transactions on Computers, 40(6): 780–784, June 1991. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
90256.

Codenotti:1991:NFA

[CT91b] B. Codenotti and R. Tamassia. A network flow approach to the reconfiguration of VLSI arrays. IEEE Transactions on Computers, 40(1):118–121, January 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=67329.

Chuang:1992:FRC

[CT92] P.-J. Chuang and N.-F. Tzeng. A fast recognition-complete processor allocation strategy for hypercube computers. IEEE Transactions on Computers, 41(4):467– 479, April 1992. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        135559.

Cheng:1993:NNI

[CT93a] S.-C. Cheng and W.-H. Tsai. A neural network implementation of the moment-preserving technique and its application to thresholding. IEEE Transactions on Computers, 42(4):501–507, April 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=214696.

Chimento:1993:CTP

[CT93b] P. F. Chimento and K. S. Trivedi. The completion time of programs on processors subject to failure and repair. IEEE Transactions on Computers, 42(10):1184–1194, October 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=257705.

Chang:1994:FER

[CT94a] Pen-Yuang Chang and Jong-Chuang Tsay. A family of efficient regular arrays for algebraic path problem. IEEE Transactions on Computers, 43(7):769– 777, July 1994. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
293256.

Chen:1994:ERP

[CT94b] Hsing-Lung Chen and Nian-Feng Tzeng. Efficient resource placement in hypercubes using multiple-adjacency codes. IEEE Transactions on Computers, 43(1):23–33, January 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=250606.

Chakravarty:1996:SGI

[CT96a] S. Chakravarty and P. J.

Thadikaran. Simulation and generation of IDDQ tests for bridging faults in combinational circuits. IEEE Transactions on Computers, 45 (10):1131–1140, October 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=543707.

Chen:1996:PER

[CT96b] Shenze Chen and D. Towsley.

A performance evaluation of RAID architectures. IEEE Transactions on Computers, 45 (10):1116–1130, October 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=543706.

Chen:1997:LTM

[CT97a] Hsing-Lung Chen and Nian-Feng Tzeng. On-line task migration in hypercubes through double disjoint paths. IEEE Transactions on Computers, 46(3):379–384, March 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=580437.

Chen:1997:SDF

[CT97b] Hsing-Lung Chen and Nian-Feng Tzeng. Subcube determination in faulty hypercubes. IEEE Transactions on Computers, 46(8):871–879, August 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=609276.

Chang:1998:ADM

[CT98] Pen-Yuang Chang and Jong-Chuang Tsay. An approach to designing modular extensible linear arrays for regular algorithms. IEEE Transactions on Computers, 47(2):212–216, February 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=663767.

Chen:1993:RRS

[CU93a] Y.-Y. Chen and S. J. Upadhyaya. Reliability, reconfiguration, and spare allocation issues in binary-tree architectures based on multiple-level redundancy. IEEE Transactions on Computers, 42(6):713–723, June 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=277283.

Chen:1993:YAR

[CU93b] Y.-Y. Chen and S. J. Upadhyaya. Yield analysis of reconfigurable array processors based on multiple-level redundancy. IEEE Transactions on Computers, 42(9):1136–1141, September 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=241602.

Chen:1994:MRC

[CU94] Y.-Y. Chen and S. J. Upadhyaya.

Modeling the reliability of a class of fault-tolerant VLSI/WSI systems based on multiple-level redundancy. IEEE Transactions on Computers, 43(6):737–748, June 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=286306.

Chen:1997:CRS

[CUC97] Yung-Yuan Chen, S. J. Upadhyaya, and Ching-Hwa Cheng.

A comprehensive reconfiguration scheme for fault-tolerant VLSI/WSI array processors. IEEE Transactions on Computers, 46(12):1363–1371, December 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=641936.

Cosentino:1992:AMJ

[CV92] R. J. Cosentino and J. J. Vaccaro. Adaptation of the Mactaggart and Jack complex multiplication algorithm for floating-point operators. IEEE Transactions on Computers, 41(10):1324–1326, October 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=166608.

Chang:1999:COR

[CVMS99] Shih-Chieh Chang, L. P. P. P. Van Ginneken, and M. Marek-Sadowska. Circuit optimization by rewiring. IEEE Transactions on Computers, 48 (9):962–970, September 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=795224.

Clark:1994:MNM

[CW94] D. W. Clark and Lih-Jyh Weng. Maximal and near-maximal shift register sequences: efficient event counters and easy discrete logarithms. IEEE Transactions on Computers, 43(5):560–568, May 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=280803.

Chiu:1996:FTR

[CW96] Ge-Ming Chiu and Shui-Pao Wu.

A fault-tolerant routing strategy in hypercube multicomputers. IEEE Transactions on Computers, 45(2):143–155, February 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=485379.

Chang:1999:LCM

[CW99a] Wen-Feng Chang and Cheng-Wen Wu. Low-cost modular totally self-checking checker design for m-out-of-n code. IEEE Transactions on Computers, 48(8):815–826, August 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=795123.

Chisholm:1999:AFA

[CW99b] G. H. Chisholm and A. S. Wojcik. An application of formal analysis to software in a fault-tolerant environment. IEEE Transactions on Computers, 48(10):1053–1064, October 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=805155.

Chang:1995:TAM

[CWM^{+}95] P. P. Chang, N. F. Warter, S. A. Mahlke, W. Y. Chen, and W. W. Hwu. Three architectural models for compilercontrolled speculative execution. IEEE Transactions on Computers, 44(4):481–494, April 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=376164.

Chung:1995:FVC

[CY95] Kuo-Liang Chung and Wen-Ming Yan. Fast vectorization for calculating a moving sum. IEEE Transactions on Computers, 44(11):1335–1337, November 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=475130.

Cheng:1991:OMB

[CYH91] C. K. Cheng, S. Z. Yao, and T. C. Hu. The orientation of modules based on graph decomposition. IEEE Transactions on Computers, 40(6):774–780, June 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=90255.

Cyre:1994:CRW

[Cyr94] W. R. Cyre. Conceptual representation of waveforms for temporal reasoning. IEEE Transactions on Computers, 43 (2):186–200, February 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=262123.

Cattell:1999:HCA

[CZSM99] K. Cattell, Shujian Zhang, M. Serra, and J. C. Muzio. 2-by-n hybrid cellular automata with regular configuration: theory and application. IEEE Transactions on Computers, 48(3):285–295, March 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=754995.

Dotan:1990:CLP

[DA90] Y. Dotan and B. Arazi. Concurrent logic programming as a hardware description tool. IEEE Transactions on Computers, 39(1):72–88, January 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=46282.

Dutt:1996:MPO

[DA96] S. Dutt and F. T. Assaad.

Mantissa-preserving operations and robust algorithm based fault tolerance for matrix computations. IEEE Transactions on Computers, 45(4):408–424, April 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=494099.

Dandamudi:1999:PHP

[DA99] S. P. Dandamudi and S. Ayachi. Performance of hierarchical processor scheduling in sharedmemory multiprocessor systems. IEEE Transactions on Computers, 48(11):1202–1213, November 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=811109.

Daehn:1991:LBH

[Dae91] W. Daehn. Load balancing in a hybrid ATPG environment. IEEE Transactions on Computers, 40(7):878–882, July 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=83627.

Dubey:1994:IWS

[DAF94] P. K. Dubey, G. B. Adams III, and M. J. Flynn. Instruction window size trade-offs and characterization of program parallelism. IEEE Transactions on Computers, 43(4):431–442, April 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=278481.

Dally:1990:PAA

[Dal90] W. J. Dally. Performance analysis of k-ary n-cube interconnection networks. IEEE Transactions on Computers, 39(6):775– 785, June 1990. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
53599.

Dally:1991:ECI

[Dal91] W. J. Dally. Express cubes: improving the performance of k-ary n-cube interconnection networks. IEEE Transactions on Computers, 40(9):1016–1023, September 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=83652.

Dally:1992:FTM

[Dal92] W. J. Dally. A fast translation method for paging on top of segmentation. IEEE Transactions on Computers, 41 (2):247–250, February 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=123403.

Damarla:1992:GTM

[Dam92] T. Damarla. Generalized transforms for multiple valued circuits and their fault detection.

IEEE Transactions on Computers, 41(9):1101–1109, September 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=165392.

David:1990:CSA

[Dav90] R. David. Comments on “Signature analysis for multiple output circuits” by R. David. IEEE Transactions on Computers, 39(2):287–288, February 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=45217.

Dubois:1991:RTE

[DB91] M. Dubois and F. A. Briggs. The run-time efficiency of parallel asynchronous algorithms. IEEE Transactions on Computers, 40(11):1260–1266, November 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=102830.

Ding:1994:FBA

[DB94] Jianxun Ding and L. N. Bhuyan. Finite buffer analysis of multistage interconnection networks. IEEE Transactions on Computers, 43(2):243–247, February 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=262132.

Dowd:1993:HSP

[DBAP93] P. W. Dowd, K. Bogineni, K. A.

Aly, and J. A. Perreault. Hierarchical scalable photonic architectures for high-performance processor interconnection. IEEE Transactions on Computers, 42 (9):1105–1120, September 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=241599.

See addendum [DBAP94].

Dowd:1994:AHS

[DBAP94] P. W. Dowd, K. K. Bogineni, K. A. Aly, and J. A. Perreault. Addendum to “Hierarchical scalable photonic architectures for high-performance processor interconnection”. IEEE Transactions on Computers, 43(7):864, July 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=293268.

See [DBAP93].

Das:1993:ICG

[DBD93] N. Das, B. B. Bhattacharya, and J. Dattagupta. Isomorphism of conflict graphs in multistage interconnection networks and its application to optimal routing. IEEE Transactions on Computers, 42(6):665–677, June 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=277286.

Das:1994:HCP

[DBD94] N. Das, B. B. Bhattacharya, and J. Dattagupta. Hierarchical classification of permutation classes in multistage interconnection networks. IEEE Transactions on Computers, 43 (12):1439–1444, December 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=338106.

Drechsler:1998:VOD

[DBJ98] R. Drechsler, B. Becker, and A. Jahnke. On variable ordering and decomposition type choice in OKFDDs. IEEE Transactions on Computers, 47 (12):1398–1403, December 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=737685.

Das:1993:VST

[DC93] A. K. Das and P. P. Chaudhuri. Vector space theoretic analysis of additive cellular automata and its application for pseudoexhaustive test pattern generation. IEEE Transactions on Computers, 42(3):340–352, March 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=210176.

Dekker:1995:PMN

[DD95] E. Dekker and L. Dekker. Parallel minimal norm method for tridiagonal linear systems. IEEE Transactions on Computers, 44(7):942–946, July 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=392854.

De:1997:ESA

[DDGS97] M. De, D. Das, M. Ghosh, and B. P. Sinha. An efficient sorting algorithm on the multi-mesh network. IEEE Transactions on Computers, 46 (10):1132–1137, October 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=628397.

Dahlgren:1998:PEC

[DDS98] F. Dahlgren, M. Dubois, and P. Stenstrom. Performance evaluation and cost analysis of cache protocol extensions for shared-memory multiprocessors. IEEE Transactions on Computers, 47(10):1041–1055, October 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=729785.

Das:1999:NNT

[DDS99] D. Das, M. De, and B. P.

Sinha. A new network topology with multiple meshes.

IEEE Transactions on Computers, 48(5):536–551, May 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=769436.

Dunning:1990:UBE

[DDV90] L. A. Dunning, G. Dial, and M. R. Varanasi. Unidirectional byte error detecting codes for computer memory systems. IEEE Transactions on Computers, 39(4):592–595, April 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=54856.

Dandamudi:1990:HIN

[DE90] S. P. Dandamudi and D. L. Eager. Hierarchical interconnection networks for multicomputer systems. IEEE Transactions on Computers, 39(6):786–797, June 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=53600.

Dandamudi:1992:HSC

[DE92] S. P. Dandamudi and D. L. Eager. Hot-spot contention in binary hypercube networks. IEEE Transactions on Computers, 41(2):239–244, February 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=123401.

Dekker:1998:ASP

[Dek98] E. Dekker. Architecture scalability of parallel vector computers with a shared memory. IEEE Transactions on Computers, 47(5):614–624, May 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=677257.

David:1990:FDR

[DF90] R. David and A. Fuentes.

Fault diagnosis of RAMs from random testing experiments. IEEE Transactions on Computers, 39(2):220–229, February 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=45207.

Dubey:1991:BSM

[DF91] P. K. Dubey and M. J. Flynn.

Branch strategies: Modeling and optimization [pipeline processing]. IEEE Transactions on Computers, 40 (10):1159–1167, October 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=93749.

DeBiase:1995:DAN

[DFM95] G. A. De Biase, C. Ferrone, and A. Massini. An O(log_{2} N) depth asymptotically nonblocking. IEEE Transactions on Computers, 44(8):1047–1051, August 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=403721.

deSouzaeSilva:1990:ASM

[dG90] E. de Souza e Silva and H. R. Gail. Analyzing scheduled maintenance policies for repairable computer systems. IEEE Transactions on Computers, 39(11):1309–1324, November 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=61041.

Donatiello:1991:ECP

[DG91] L. Donatiello and V. Grassi. On evaluating the cumulative performance distribution of fault-tolerant computer systems. IEEE Transactions on Computers, 40(11):1301–1307, November 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=102838.

David:1992:EIB

[DGY92a] I. David, R. Ginosar, and M. Yoeli. An efficient implementation of Boolean functions as self-timed circuits. IEEE Transactions on Computers, 41(1):2–11, January 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=123377.

David:1992:ISM

[DGY92b] I. David, R. Ginosar, and M. Yoeli. Implementing sequential machines as self-timed circuits. IEEE Transactions on Computers, 41(1):12–17, January 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=123378.

Dutt:1990:DRF

[DH90] S. Dutt and J. P. Hayes.

On designing and reconfiguring k-fault-tolerant tree architectures. IEEE Transactions on Computers, 39(4):490–503, April 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=54842.

Dutt:1991:SAH

[DH91] S. Dutt and J. P. Hayes. Subcube allocation in hypercube computers. IEEE Transactions on Computers, 40(3):341–352, March 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=76413.

Dutt:1992:SPI

[DH92] S. Dutt and J. P. Hayes. Some practical issues in the design of fault-tolerant multiprocessors. IEEE Transactions on Computers, 41(5):588–598, May 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=142685.

Duprat:1993:NRR

[DHK93] J. Duprat, Y. Herreros, and S. Kla. New redundant representations of complex numbers and vectors. IEEE Transactions on Computers, 42(7):817–824, July 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=237722.

Du:1996:CLD

[DHL96] Ding-Zhu Du, D. F. Hsu, and Yuh-Dauh Lyuu. Corrigendum to “Line Digraph Iterations and Connectivity Analysis of de Bruijn and Kautz Graphs”. IEEE Transactions on Computers, 45(7):863, July 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=508325.

See [DZLH93, PMF96].

Dimauro:1993:NTF

[DIP93] G. Dimauro, S. Impedovo, and G. Pirlo. A new technique for fast number comparison in the residue number system. IEEE Transactions on Computers, 42(5):608–612, May 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=223680;
 https://ieeexplore.ieee.
org/xpl/tocresult.jsp?isnumber=
5840.

Dave:1999:CHS

[DJ99] B. P. Dave and N. K. Jha.

COFTA: hardware-software cosynthesis of heterogeneous distributed embedded systems for low overhead fault tolerance. IEEE Transactions on Computers, 48(4):417–441, April 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=762534.

Dimitrov:1999:TAD

[DJM99] V. S. Dimitrov, G. A. Jullien, and W. C. Miller. Theory and applications of the double-base number system. IEEE Transactions on Computers, 48(10):1098–1106, October 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=805158.

Daigle:1990:QAO

[DKL90] J. N. Daigle, R. B. Kuehl, and J. D. Langford. Queueing analysis of an optical disk jukebox based office system. IEEE Transactions on Computers, 39(6):819–828, June 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=53603.

Dey:1996:LSP

[DKT96] J. K. Dey, J. Kurose, and D. Towsley. On-line scheduling policies for a class of IRIS (Increasing Reward with Increasing Service) real-time tasks. IEEE Transactions on Computers, 45(7):802–813, July 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=508319.

Demmel:1994:FNA

[DL94] James W. Demmel and Xiaoye Li. Faster numerical algorithms via exception handling. IEEE Transactions on Computers, 43(8):983–992, August 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=295860;
                                         http://www.cs.berkeley.edu/
                                        ~xiaoye/ieee.ps.gz. This is an

expanded version of [?].

Dong:1997:EPC

[DL97] Xuefeng Dong and Ten-Hwang Lai. An efficient protocol for call setup and path migration in IEEE 802.6 based personal communication networks. IEEE Transactions on Computers, 46(3):326–336, March 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=580428.

Du:1993:WPP

[DLC93] D. H. C. Du, Ichiang Lin, and K. C. Chang. On waferpacking problems. IEEE Transactions on Computers, 42(11): 1382–1388, November 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=247840.

Dunne:1995:CBF

[DLN95] P. E. Dunne, P. H. Leng, and G. F. Nwana. On the complexity of Boolean functions computed by lazy oracles. IEEE Transactions on Computers, 44(4):495–502, April 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=376165.

Davidson:1991:TAC

[DLW91] S. B. Davidson, I. Lee, and V. Wolfe. Timed atomic commitment. IEEE Transactions on Computers, 40(5):573–583, May 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=88481.

deSouzaeSilva:1990:NCC

[dM90] E. de Souza e Silva and R. R. Muntz. A note on the computational cost of the Linearizer algorithm for queueing networks. IEEE Transactions on Computers, 39(6):840–842, June 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=53607.

Drakopoulos:1992:PAC

[DM92] E. Drakopoulos and M. J. Merges. Performance analysis of client-server storage systems. IEEE Transactions on Computers, 41(11):1442–1452, November 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=177314.

Duprat:1993:CAN

[DM93] J. Duprat and J.-M. Muller. The CORDIC algorithm: New results for fast VLSI implementation. IEEE Transactions on Computers, 42(2):168–178, February 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=204786.

DasSarma:1994:MAR

[DM94] D. DasSarma and D. W. Matula. Measuring the accuracy of ROM reciprocal tables. IEEE Transactions on Computers, 43(8):932–940, August 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=295855.

Dawid:1996:DCA

[DM96] Herbert Dawid and Heinrich Meyr. The differential CORDIC algorithm: Constant scale factor redundant implementation without correcting iterations. IEEE Transactions on Computers, 45(3):307–318, March 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=485569.

Daumas:1997:VRD

[DM97a] M. Daumas and D. W. Matula. Validated roundings of dot products by sticky accumulation. IEEE Transactions on Computers, 46(5):623–629, May 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=589241.

Dutt:1997:NCE

[DM97b] S. Dutt and N. R. Mahapatra.

Node-covering, error-correcting codes and multiprocessors with very high average fault tolerance. IEEE Transactions on Computers, 46(9):997–1015, September 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=620481.

Das:1994:NFB

[DMS94] R. K. Das, K. Mukhopadhyaya, and B. P. Sinha. A new family of bridged and twisted hypercubes. IEEE Transactions on Computers, 43(10):1240–1247, October 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=324555.

Das:1995:IFC

[DMS95] D. Das, K. Mukhopadhyaya, and B. P. Sinha. Implementation of four common functions on an LNS co-processor. IEEE Transactions on Computers, 44(1):155–161, January 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=367997.

Duarte:1998:HAD

[DN98] E. P. Duarte, Jr. and T. Nanya. A hierarchical adaptive distributed system-level diagnosis algorithm. IEEE Transactions on Computers, 47(1):34–45, January 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=656078.

DeGloria:1996:SCL

[DO96] A. De Gloria and M. Olivieri. Statistical carry lookahead adders.

IEEE Transactions on Computers, 45(3):340–347, March 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=485572.

DiClaudio:1993:SRR

[DOP93] E. D. Di Claudio, G. Orlandi, and F. Piazza. A systolic redundant residue arithmetic error correction circuit. IEEE Transactions on Computers, 42(4):427–432, April 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=214689.

Damiani:1991:ADL

[DOR91] M. Damiani, P. Olivo, and B. Ricco. Analysis and design of linear finite state machines for signature analysis testing. IEEE Transactions on Computers, 40(9):1034–1045, September 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=83659.

Douglass:1993:RTS

[Dou93] B. G. Douglass. Rearrangeable three-stage interconnection networks and their routing properties. IEEE Transactions on Computers, 42(5):559–567, May 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=223675.

Dowd:1992:WDM

[Dow92] P. W. Dowd. Wavelength division multiple access channel hypercube processor interconnection. IEEE Transactions on Computers, 41 (10):1223–1241, October 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=166601.

Dadda:1996:PA

[DP96] L. Dadda and V. Piuri. Pipelined adders. IEEE Transactions on Computers, 45(3):348–356, March 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=485573.

Diks:1997:GOD

[DP97] K. Diks and A. Pelc. Globally optimal diagnosis in systems with random faults. IEEE Transactions on Computers, 46 (2):200–204, February 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=565598.

Dai:1999:EBM

[DP99] Dongai Dai and D. K. Panda. Exploiting the benefits of multiplepath network in DSM systems: architectural alternatives and performance evaluation. IEEE Transactions on Computers, 48(2):236–244, February 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=752665.

DiClaudio:1995:FCR

[DPO95] E. D. Di Claudio, F. Piazza, and G. Orlandi. Fast combinatorial RNS processors for DSP applications. IEEE Transactions on Computers, 44(5):624–633, May 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=381948.

Dierrich:1996:RRB

[DR96] P. Dierrich and R. R. Rao. Request resubmission in a blocking, circuit-switched, interconnection network. IEEE Transactions on Computers, 45(11): 1282–1293, November 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=544484.

Drechsler:1999:PKE

[Dre99] R. Drechsler. Pseudo-Kronecker expressions for symmetric functions. IEEE Transactions on Computers, 48(9):987–990, September 1999. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

                                        stamp/stamp.jsp?tp=&arnumber=
                                        795226.

Drolet:1998:NRE

[Dro98]  G. Drolet.  A new repre-

sentation of elements of finite fields GF(2^{m}) yielding small complexity arithmetic circuits. IEEE Transactions on Computers, 47(9):938–946, September 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=713313.

Dolter:1991:PAV

[DRS91] J. W. Dolter, P. Ramanathan, and K. G. Shin. Performance analysis of virtual cutthrough switching in HARTS: a hexagonal mesh multicomputer. IEEE Transactions on Computers, 40(6):669–680, June 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=90246.

Dube:1997:INP

[DRT97] R. Dube, C. D. Rais, and S. K. Tripathi. Improving NFS performance over wireless links. IEEE Transactions on Computers, 46(3):290–298, March 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=580425.

Davidson:1992:RSD

[DRW92] J. W. Davidson, J. R. Rabung, and D. B. Whalley. Relating static and dynamic machine code measurements. IEEE Transactions on Computers, 41(4):444–454, April 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=135557.

De:1994:FPA

[DS94] M. De and B. P. Sinha. Fast parallel algorithm for ternary multiplication using multivalued I^{2}L technology. IEEE Transactions on Computers, 43(5):603–607, May 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=280807.

Dimakopoulos:1995:TCC

[DSPN95] V. V. Dimakopoulos, G. Sourtziotis, A. Paschalis, and D. Nikolos. On TSC checkers for m-out-of-n codes. IEEE Transactions on Computers, 44(8):1055–1059, August 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=403723.

Day:1993:ECA

[DT93] K. Day and A. Tripathi. Embedding of cycles in arrangement graphs. IEEE Transactions on Computers, 42(8):1002–1006, August 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=238494.

Das:1993:MFD

[DTAL93] A. Das, K. Thulasiraman, V. K. Agarwal, and K. B. Lakshmanan. Multiprocessor fault diagnosis under local constraints. IEEE Transactions on Computers, 42(8):984–988, August 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=238490.

Drechsler:1996:FOB

[DTB96] R. Drechsler, M. Theobald, and B. Becker. Fast OFDDbased minimization of fixed polarity Reed–Muller expressions. IEEE Transactions on Computers, 45(11):1294–1299, November 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=544485.

David:1990:ADP

[DW90] R. David and K. Wagner. Analysis of detection probability and some applications. IEEE Transactions on Computers, 39(10):1284–1291, October 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=59858.

Dubois:1991:SBC

[DW91] M. Dubois and J.-C. Wang.

Shared block contention in a cache coherence protocol. IEEE Transactions on Computers, 40(5):640–644, May 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=88487.

Dolev:1997:CRC

[DW97] S. Dolev and J. L. Welch.

Crash resilient communication in dynamic networks. IEEE Transactions on Computers, 46(1):14–26, January 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=559799.

Dan:1994:PMC

[DYD94] A. Dan, P. S. Yu, and D. M.

Dias. Performance modelling and comparisons of global shared buffer management policies in a cluster environment. IEEE Transactions on Computers, 43(11):1281–1297, November 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=324561.

Ding-Zhu:1993:LDI

[DZLH93] D. Ding-Zhu, Y.-D. Lyuu, and D. F. Hsu. Line digraph iterations and connectivity analysis of de Bruijn and Kautz graphs. IEEE Transactions on Computers, 42(5):612–616, May 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=223681.

See comments [PMF96] and corrigendum [DHL96].

El-Amawy:1998:CDO

[EAK98] A. El-Amawy and P. Kulasinghe. On the complexity of designing optimal branchand-combine clock networks. IEEE Transactions on Computers, 47(2):264–269, February 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=664212.

El-Amawy:1993:NME

[EANPH93] A. El-Amawy, M. Naraghi-Pour, and M. Hegde. Noise modeling effects in redundant synchronizers. IEEE Transactions on Computers, 42(12):1487–1494, December 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=260638.

El-Amawy:1995:MPP

[EAT95] A. El-Amawy and Yi Chang Tseng. Maximum performance pipelines with switchable reservation tables. IEEE Transactions on Computers, 44 (8):1066–1069, August 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=403726.

Efe:1991:VHL

[Efe91] K. Efe. A variation on the hypercube with lower diameter. IEEE Transactions on Computers, 40(11):1312–1316, November 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=102840.

Eggers:1991:SVA

[Egg91]  S. J. Eggers.  Simplicity

versus accuracy in a model of cache coherency overhead. IEEE Transactions on Computers, 40(8):893–906, August 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=83635.

Ercegovac:1990:FMC

[EL90a] M. D. Ercegovac and T. Lang.

Fast multiplication without carry-propagate addition. IEEE Transactions on Computers, 39 (11):1385–1390, November 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=61047.

Ercegovac:1990:RSR

[EL90b] M. D. Ercegovac and T. Lang.

Radix-4 square root without initial PLA. IEEE Transactions on Computers, 39(8):1016–1024, August 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=57040.

Ercegovac:1990:RLC

[EL90c] M. D. Ercegovac and T. Lang. Redundant and on-line CORDIC: application to matrix triangularization and SVD. IEEE Transactions on Computers, 39(6):725– 740, June 1990. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
53594.

Ercegovac:1990:SRD

[EL90d] M. D. Ercegovac and T. Lang. Simple radix-4 division with operands scaling. IEEE Transactions on Computers, 39(9): 1204–1208, September 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=57060.

Ewerbring:1990:CSV

[EL90e] L. M. Ewerbring and F. T. Luk. Computing the singular value decomposition on the Connection Machine. IEEE Transactions on Computers, 39(1):152–155, January 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=46294.

Ercegovac:1992:FRC

[EL92] M. D. Ercegovac and T. Lang.

On-the-fly rounding [computing arithmetic]. IEEE Transactions on Computers, 41(12): 1497–1503, December 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=214659.

Ellis:1991:ERG

[Ell91] J. A. Ellis. Embedding rectangular grids into square grids. IEEE Transactions on Computers, 40(1):46–52, January 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=67319.

Ercegovac:1994:VHR

[ELM94] M. D. Ercegovac, T. Lang, and P. Montuschi. Very-high radix division with prescaling and selection by rounding. IEEE Transactions on Computers, 43(8):909–918, August 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=295853.

Esfahanian:1991:TCA

[ENS91] A.-H. Esfahanian, L. M. Ni, and B. E. Sagan. The twisted Ncube with application to multiprocessing. IEEE Transactions on Computers, 40(1):88–93, January 1991. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=67323.

Eshaghian:1991:PAI

[Esh91] M. M. Eshaghian. Parallel algorithms for image processing on OMC. IEEE Transactions on Computers, 40(7):827–833, July 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=83620.

Elsaholy:1998:UAE

[ESS98] M. S. Elsaholy, S. I. Shaheen, and R. H. Seireg. A unified analytical expression for aliasing error probability using single-input external- and internal-XOR LFSR. IEEE Transactions on Computers, 47 (12):1414–1417, December 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=737687.

Eldridge:1993:HIM

[EW93] S. E. Eldridge and C. D. Walter. Hardware implementation of Montgomery’s modular multiplication algorithm. IEEE Transactions on Computers, 42(6):693–699, June 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=277287.

Elnozahy:1992:MTR

[EZ92] E. N. Elnozahy and W. Zwaenepoel.

Manetho: transparent roll backrecovery with low overhead, limited rollback, and fast output commit. IEEE Transactions on Computers, 41(5):526–531, May 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=142678.

Fragopoulou:1996:EDS

[FA96] P. Fragopoulou and S. G. Akl.

Edge-disjoint spanning trees on the star network with applications to fault tolerance. IEEE Transactions on Computers, 45(2):174–185, February 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=485370.

Fagin:1992:FAL

[Fag92] B. S. Fagin. Fast addition of large integers. IEEE Transactions on Computers, 41(9): 1069–1077, September 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=165389.

Fagin:1997:PRB

[Fag97]  B. Fagin.  Partial resolu-

tion in branch target buffers. IEEE Transactions on Computers, 46(10):1142–1145, October 1997. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=628399.

Falkowski:1999:NPF

[Fal99] B. J. Falkowski. A note on the polynomial form of Boolean functions and related topics. IEEE Transactions on Computers, 48(8):860–864, August 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=795128.

Feng:1996:ASL

[FBL96] Chao Feng, L. N. Bhuyan, and F. Lombardi. Adaptive system-level diagnosis for hypercube multiprocessors. IEEE Transactions on Computers, 45 (10):1157–1170, October 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=543709.

Fernandez-Baca:1993:PMA

[FBM93] D. Fernandez-Baca and A. Medepalli. Parametric module allocation on partial k-trees. IEEE Transactions on Computers, 42(6):738–742, June 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=277293.

Fenn:1996:MDD

[FBT96] S. T. J. Fenn, M. Benaissa, and D. Taylor. GF(2^{m}) multiplication and division over the dual basis. IEEE Transactions on Computers, 45(3):319–327, March 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=485570.

Falkowski:1997:FIT

[FC97] B. J. Falkowski and Chip-Hong Chang. Forward and inverse transformations between Haar spectra and ordered binary decision diagrams of Boolean functions. IEEE Transactions on Computers, 46(11): 1272–1279, November 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=644301.

Fang:1998:COT

[FCL98] Y. Fang, I. Chlamtac, and Yi-Bing Lin. Channel occupancy times and handoff rate for mobile computing and PCS networks. IEEE Transactions on Computers, 47(6):679–692, June 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=689647.

Fagin:1990:PPP

[FD90] B. S. Fagin and A. M. Despain. The performance of parallel Prolog programs. IEEE Transactions on Computers, 39 (12):1435–1445, December 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=61072.

Fernandez:1997:EVL

[FE97] A. Fernandez and K. Efe. Efficient VLSI layouts for homogeneous product networks. IEEE Transactions on Computers, 46(10):1070–1082, October 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=628392.

Ferreira:1991:PTH

[Fer91]  A. G. Ferreira.  A par-

allel time/hardware tradeoff T.H = O(2^{n/2}) for the knapsack problem. IEEE Transactions on Computers, 40 (2):221–225, February 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=73593.

Feng:1995:MVA

[FFKI95] Shou-Ping Feng, T. Fujiwara, T. Kasami, and K. Iwasaki. On the maximum value of aliasing probabilities for single input signature registers. IEEE Transactions on Computers, 44(11):1265–1274, November 1995. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=475122.

Fink:1992:RNP

[FFS92] F. Fink, K. Fuchs, and M. H.

Schulz. Robust and nonrobust path delay fault simulation by parallel processing of patterns. IEEE Transactions on Computers, 41(12):1527–1536, December 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=214661.

Forge:1994:ASD

[FHA94] L. E. La Forge, Kaiyuan Huang, and V. K. Agarwal. Almost sure diagnosis of almost every good element [logic testing]. IEEE Transactions on Computers, 43(3):295–305, March 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=272430.

Fiore:1999:PMU

[Fio99] P. D. Fiore. Parallel multiplication using fast sorting networks. IEEE Transactions on Computers, 48(6):640–645, June 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=773800.

Fishburn:1990:CSO

[Fis90] J. P. Fishburn. Clock skew optimization. IEEE Transactions on Computers, 39(7):945–951, July 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=55696.

Fitzpatrick:1997:EBE

[Fit97]  P. Fitzpatrick.  Extending

backward error assertions to tolerance of large errors in floating point computations. IEEE Transactions on Computers, 46(4):505–510, April 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=588072.

Fu:1991:MST

[FL91] Jyh-Jong Fu and R. C. T. Lee. Minimum spanning trees of moving points in the plane. IEEE Transactions on Computers, 40(1):113–118, January 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=67328.

Fiol:1992:PLD

[FL92] M. A. Fiol and A. S. Llado. The partial line digraph technique in the design of large interconnection networks. IEEE Transactions on Computers, 41(7):848– 857, July 1992. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
256453.

Fang:1993:IPA

[FL93] J. Z. Fang and M. Lu. An iteration partition approach for cache or local memory thrashing on parallel processing. IEEE Transactions on Computers, 42(5):529–546, May 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=223672.

Fix:1998:SPI

[FL98] J. D. Fix and R. E. Ladner.

Sorting by parallel insertion on a one-dimensional subbus array. IEEE Transactions on Computers, 47(11):1267–1281, November 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=736441.

Faloutsos:1991:DAM

[FM91] C. Faloutsos and D. Metaxas.

Disk allocation methods using error correcting codes. IEEE Transactions on Computers, 40(8):907–914, August 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=83636.

Fagin:1995:PCC

[FM95] B. Fagin and A. Mital. The performance of counter- and correlation-based schemes for branch target buffers. IEEE Transactions on Computers, 44 (12):1383–1393, December 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=477244.

Faloutsos:1995:FAB

[FNS95] C. Faloutsos, R. Ng, and T. Sellis. Flexible and adaptable buffer management techniques for database management systems. IEEE Transactions on Computers, 44(4):546–560, April 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=376169.

Fowkes:1993:HEA

[Fow93] R. E. Fowkes. Hardware efficient algorithms for trigonometric functions. IEEE Transactions on Computers, 42 (2):235–239, February 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=204796.

Fabre:1998:MAF

[FP98] J.-C. Fabre and T. Perennou. A metaobject architecture for fault-tolerant distributed systems: the FRIENDS approach.

IEEE Transactions on Computers, 47(1):78–95, January 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=656088.

Fiduccia:1994:PSM

[FR94] C. M. Fiduccia and K. J. Rappoport. Perfect shifters [multiprocessor interconnection networks]. IEEE Transactions on Computers, 43(3):340–349, March 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=272434.

Franaszek:1997:VSP

[FR97a] P. A. Franaszek and J. T. Robinson. On variable scope of parity protection in disk arrays. IEEE Transactions on Computers, 46(2):234–240, February 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=565608.

Fridman:1997:MMT

[FR97b] J. Fridman and S. Rangarajan.

Maximizing mean-time to failure in k-resilient systems with repair. IEEE Transactions on Computers, 46(2):229–234, February 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=565606.

Fraigniaud:1992:AOB

[Fra92] P. Fraigniaud. Asymptotically optimal broadcasting and gossiping in faulty hypercube multicomputers. IEEE Transactions on Computers, 41(11): 1410–1419, November 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=177311.

Friedman:1994:FAE

[Fri94] A. D. Friedman. A functional approach to efficient fault detection in iterative logic arrays. IEEE Transactions on Computers, 43(12):1365–1375, December 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=338096.

Fricker:1995:MCP

[Fri95]  C. Fricker.  On memory

contention problems in vector multiprocessors. IEEE Transactions on Computers, 44(1):92–105, January 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=368007.

Feng:1994:ECC

[FRK94] G. L. Feng, T. R. N. Rao, and M. S. Kolluru. Error correcting codes over Z_{2}(m) for algorithm-based fault tolerance.

IEEE Transactions on Computers, 43(3):370–374, March 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=272438.

Fujiwara:1998:OTL

[FRK98] E. Fujiwara, T. Ritthongpitak, and M. Kitakami. Optimal two-level unequal error control codes for computer systems. IEEE Transactions on Computers, 47(12):1313–1325, December 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=737679.

Felperin:1996:TWR

[FRU96] S. Felperin, P. Raghavan, and E. Upfal. A theory of wormhole routing in parallel computers. IEEE Transactions on Computers, 45(6):704–713, June 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=506426.

Friedman:1990:FOV

[FS90] S. J. Friedman and K. J.

Supowit. Finding the optimal variable ordering for binary decision diagrams. IEEE Transactions on Computers, 39(5):710– 713, May 1990. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

                                        stamp/stamp.jsp?tp=&arnumber=
                                        53586.

Feng:1994:NRA

[FS94a] T.-Y. Feng and S.-W. Seo. A new routing algorithm for a class of rearrangeable networks. IEEE Transactions on Computers, 43(11):1270–1280, November 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=324560.

See comment [MM97a].

Franklin:1994:HCR

[FS94b] M. Franklin and K. K. Saluja. Hypergraph coloring and reconfigured RAM testing. IEEE Transactions on Computers, 43 (6):725–736, June 1994. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL https://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        286305.

Franklin:1996:AHM

[FS96] M. Franklin and G. S. Sohi. ARB: a hardware mechanism for dynamic reordering of memory references. IEEE Transactions on Computers, 45(5):552–571, May 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=509907.

Fummi:1999:STH

[FSS99] F. Fummi, D. Sciuto, and M. Serra. Synthesis for testability of highly complex controllers by functional redundancy removal. IEEE Transactions on Computers, 48(12): 1305–1323, December 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=817386.

Fujimoto:1992:DER

[FTG92] R. M. Fujimoto, J.-J. Tsai, and G. C. Gopalakrishnan. Design and evaluation of the rollback chip: special purpose hardware for Time Warp. IEEE Transactions on Computers, 41(1):68–82, January 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=123382.

Funabiki:1993:CSN

[FTL93] N. Funabiki, Y. Takefuji, and K. C. Lee. Comparisons of seven neural network models on traffic control problems in multistage interconnection networks. IEEE Transactions on Computers, 42(4):497–501, April 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=214695.

Fang:1990:DPS

[FTYZ90] Z. Fang, P. Tang, P.-C. Yew, and C.-Q. Zhu. Dynamic processor self-scheduling for general parallel nested loops. IEEE Transactions on Computers, 39(7):919–929, July 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=55693.

Fujiwara:1990:CCC

[Fuj90] H. Fujiwara. Computational complexity of controllability/ observability problems for combinational circuits. IEEE Transactions on Computers, 39(6): 762–767, June 1990. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        53597.

Fujita:1999:FTB

[Fuj99]  S. Fujita.  A fault-tolerant

broadcast scheme in the star graph under the single-port, halfduplex communication model. IEEE Transactions on Computers, 48(10):1123–1126, October 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=805160.

Fisher:1993:RFS

[FW93] P. D. Fisher and Sheng-Fu Wu. Race-free state assignments for synthesizing largescale asynchronous sequential logic circuits. IEEE Transactions on Computers, 42(9): 1025–1034, September 1993.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=241592.

Fujita:1996:FGM

[FY96] S. Fujita and M. Yamashita.

Fast gossiping on mesh-bus computers. IEEE Transactions on Computers, 45(11): 1326–1330, November 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=544491.

Grimsrud:1996:LVT

[GAFN96] K. Grimsrud, J. Archibald, R. Frost, and B. Nelson. Locality as a visualization tool. IEEE Transactions on Computers, 45(11):1319–1326, November 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=544490.

Gait:1990:CPS

[Gai90]  J. Gait.  A checkpointing

page store for write-once optical disk. IEEE Transactions on Computers, 39(1):2–9, January 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=46276.

Ghosal:1990:PED

[GB90a] D. Ghosal and L. N. Bhuyan. Performance evaluation of a dataflow architecture. IEEE Transactions on Computers, 39 (5):615–627, May 1990. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
53575.

Gupta:1990:BMS

[GB90b] R. Gupta and M. A. Breuer. The Ballast methodology for structured partial scan design. IEEE Transactions on Computers, 39(4):538–544, April 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=54846.

Ghafoor:1991:SOG

[GB91] A. Ghafoor and T. R. Bashkow. A study of odd graphs as faulttolerant interconnection networks. IEEE Transactions on Computers, 40(2):225–232, February 1991. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
73594.

Gulliver:1993:SCC

[GB93] T. A. Gulliver and V. K. Bhargava. A systematic (16, 8) code for correcting double errors and detecting triple-adjacent errors. IEEE Transactions on Computers, 42(1):109–112, January 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=192220.

See comments [KGB95].

Ganley:1997:MCH

[GC97] J. L. Ganley and J. P. Cohoon. Minimum-congestion hypergraph embedding in a cycle. IEEE Transactions on Computers, 46(5):600–602, May 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=589233.

Gale:1990:IAE

[GDY90] W. Gale, S. Das, and C. T.

Yu. Improvements to an algorithm for equipartitioning. IEEE Transactions on Computers, 39(5):706–710, May 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=53585.

Gaughan:1996:DDF

[GDYS96] P. T. Gaughan, B. V. Dao, S. Yalamanchili, and D. E. Schimmel. Distributed, deadlockfree routing in faulty, pipelined, direct interconnection networks. IEEE Transactions on Computers, 45(6):651–665, June 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=506422.

Goldfeld:1992:UGU

[GE92] D. Goldfeld and T. Etzion. UPP graphs and UMFA networksarchitecture for parallel systems. IEEE Transactions on Computers, 41(11):1479–1483, November 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=177317.

Gu:1996:CPO

[GGD96] Jun Gu, Qian-Ping Gu, and Ding-Zhu Du. Convergence properties of optimization algorithms for the SAT problem. IEEE Transactions on Computers, 45(2):209–219, February 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=485373.

Gupta:1991:ECB

[GH91] A. K. Gupta and S. E. Hambrusch. Embedding complete binary trees into butterfly networks. IEEE Transactions on Computers, 40(7):853–863, July 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=83623.

Geihs:1993:POR

[GHH93] K. Geihs, R. Heite, and U. H. Hollberg. Protected object references in heterogeneous distributed systems. IEEE Transactions on Computers, 42(7):809–816, July 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=237721.

Guha:1992:RCB

[GK92] A. Guha and L. L. Kinney. Relating the cyclic behavior of linear and intrainverted feedback shift registers. IEEE Transactions on Computers, 41(9):1088–1100, September 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=165391.

Guan:1991:TSO

[GL91] X. Guan and M. S. Langston.

Time-space optimal parallel merging and sorting. IEEE Transactions on Computers, 40 (5):596–602, May 1991. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
88483.

Groenendijk:1992:PAT

[GL92] W. P. Groenendijk and H. Levy.

Performance analysis of transaction driven computer systems via queueing analysis of polling models. IEEE Transactions on Computers, 41(4):455–466, April 1992. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=135558.

Gonzalez:1993:RBD

[GL93] A. M. Gonzalez and J. M. Llaberia. Reducing branch delay to zero in pipelined processors. IEEE Transactions on Computers, 42(3):363–371, March 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=210179.

Garcia:1999:LSS

[GL99a] A. Garcia and A. Lloris. A look-up scheme for scaling in the RNS. IEEE Transactions on Computers, 48(7):748–751, July 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=780883.

Gaudiot:1999:GEI

[GL99b] J. Gaudiot and F. Lombardi. Guest Editors’ introduction. IEEE Transactions on Computers, 48(6):553–555, June 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=773790.

Goudreau:1999:PEP

[GLR^{+}99] M. W. Goudreau, K. Lang, S. B. Rao, T. Suel, and T. Tsantilas. Portable and efficient parallel computing using the BSP model. IEEE Transactions on Computers, 48(7):670–689, July 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=780876.

Gouda:1991:SCP

[GM91] M. G. Gouda and N. J. Multari. Stabilizing communication protocols. IEEE Transactions on Computers, 40(4):448–458, April 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=88464.

Gergov:1994:EBM

[GM94] J. Gergov and C. Meinel. Efficient Boolean manipulation with OBDD’s can be extended to FBDD’s. IEEE Transactions on Computers, 43 (10):1197–1209, October 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=324545.

Gonzalez:1998:MVS

[GM98a] A. F. Gonzalez and P. Mazumder.

Multiple-valued signed digit adder using negative differential resistance devices. IEEE Transactions on Computers, 47 (9):947–959, September 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=713314.

Gravenstreter:1998:RCC

[GM98b] G. Gravenstreter and R. G. Melhem. Realizing common communication patterns in partitioned optical passive stars (POPS) networks. IEEE Transactions on Computers, 47 (9):998–1013, September 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=713318.

Goldberg:1994:PAR

[GMP94] A. V. Goldberg, B. M. Maggs, and S. A. Plotkin. A parallel algorithm for reconfiguring a multibutterfly network with faulty switches. IEEE Transactions on Computers, 43(3):321–326, March 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=272432.

Guo:1998:RSC

[GO98] W. Guo and A. Y. Oruc. Regular sparse crossbar concentrators. IEEE Transactions on Computers, 47(3):363–368, March 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=660174.

Geist:1992:EER

[GOH92] R. Geist, A. J. Offutt, and F. C. Harris, Jr. Estimation and enhancement of real-time software reliability through mutation analysis. IEEE Transactions on Computers, 41(5):550–558, May 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=142681.

Golic:1996:LMK

[Gol96]  J. D. Golic.  Linear mod-

els for keystream generators. IEEE Transactions on Computers, 45(1):41–49, January 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=481485.

Goswami:1997:DSB

[Gos97] K. K. Goswami. DEPEND:

a simulation-based environment for system level dependability analysis. IEEE Transactions on Computers, 46(1):60–74, January 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=559803.

Golic:1996:CAC

[GP96a] J. D. Golic and S. V. Petrovic. Correlation attacks on clock-controlled shift registers in keystream generators. IEEE Transactions on Computers, 45 (4):482–486, April 1996. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
494106.

Gupta:1996:ULC

[GP96b] S. K. Gupta and D. K. Pradhan. Utilization of on-line (concurrent) checkers during builtin self-test and vice versa. IEEE Transactions on Computers, 45(1):63–73, January 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=481487.

Gu:1997:PCF

[GP97] Qian-Ping Gu and Shietung Peng. K-pairwise cluster fault tolerant routing in hypercubes. IEEE Transactions on Computers, 46(9):1042–1049, September 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=620486.

Ganger:1998:USL

[GP98] G. R. Ganger and Y. N. Patt. Using system-level models to evaluate I/O subsystem designs. IEEE Transactions on Computers, 47(6):667–678, June 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=689646.

Gotze:1993:EJL

[GPS93] J. Gotze, S. Paul, and M. Sauer.

An efficient Jacobi-like algorithm for parallel eigenvalue computation. IEEE Transactions on Computers, 42(9): 1058–1065, September 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=241595.

Gerber:1995:PDH

[GPS95] R. Gerber, W. Pugh, and M. Saksena. Parametric dispatching of hard real-time tasks. IEEE Transactions on Computers, 44(3):471–479, March 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=372041.

Garg:1998:APM

[GPTT98] S. Garg, A. Puliafito, M. Telek, and K. Trivedi. Analysis of preventive maintenance in transactions based software systems. IEEE Transactions on Computers, 47(1):96–107, January 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=656092.

Gizopoulos:1999:EBS

[GPZ99] D. Gizopoulos, A. Paschalis, and Y. Zorian. An effective built-in self-test scheme for parallel multipliers. IEEE Transactions on Computers, 48 (9):936–950, September 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=795222.

Greenberg:1994:FPU

[Gre94] R. I. Greenberg. The fat-pyramid and universal parallel computation independent of wire delay. IEEE Transactions on Computers, 43(12):1358–1364, December 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=338095.

Gu:1994:CCS

[GRR94] D. Gu, D. J. Rosenkrantz, and S. S. Ravi. Construction of check sets for algorithm-based fault tolerance. IEEE Transactions on Computers, 43(6):641–650, June 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=286298.

Gupta:1996:AAG

[GRT96] S. Gupta, J. Rajski, and J. Tyszer. Arithmetic additive generators of pseudo-exhaustive test patterns. IEEE Transactions on Computers, 45(8):939–949, August 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=536236.

Graham:1993:CBS

[GS93] S. W. Graham and S. R. Seidel. The cost of broadcasting on star graphs and k-ary hypercubes. IEEE Transactions on Computers, 42(6):756–759, June 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=277296.

Goes:1995:SMP

[GS95] P. B. Goes and U. Sumita.

Stochastic models for performance analysis of database recovery control. IEEE Transactions on Computers, 44(4):561– 576, April 1995. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
376170.

Gavish:1997:TPP

[GS97] B. Gavish and S. Sridhar.

Threshold priority policy for channel assignment in cellular networks. IEEE Transactions on Computers, 46(3):367–370, March 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=580432.

Gokhale:1998:MOC

[GS98] A. S. Gokhale and D. C.

Schmidt. Measuring and optimizing CORBA latency and scalability over high-speed networks. IEEE Transactions on Computers, 47(4):391–413, April 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=675710.

Goyal:1992:UFS

[GSH^{+}92] A. Goyal, P. Shahabuddin, P. Heidelberger, V. F. Nicola, and P. W. Glynn. A unified framework for simulating Markovian models of highly dependable systems. IEEE Transactions on Computers, 41(1):36–51, January 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=123381.

Gunther:1997:MDF

[Gun97] B. K. Gunther. Multithreading with distributed functional units. IEEE Transactions on Computers, 46(4):399–411, April 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=588034.

Greenberg:1991:DAM

[GW91] A. G. Greenberg and P. E. Wright. Design and analysis of master/slave multiprocessors. IEEE Transactions on Computers, 40(8):963–976, August 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=83640.

Gupta:1992:CIP

[GW92] A. Gupta and W.-D. Weber. Cache invalidation patterns in shared-memory multiprocessors. IEEE Transactions on Computers, 41(7):794–810, July 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=256449.

Guo:1998:SAI

[GW98] Jyh-Huei Guo and Chin-Liang Wang. Systolic array implementation of Euclid’s algorithm for inversion and division in GF(2^{m}). IEEE Transactions on Computers, 47 (10):1161–1167, October 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=729800.

Gaughan:1995:PMP

[GY95a] P. T. Gaughan and S. Yalamanchili. A performance model of pipelined k-ary n-cubes. IEEE Transactions on Computers, 44(8):1059–1063, August 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=403724.

Gong:1995:RAS

[GY95b] Wei-Bo Gong and Hong Yang. Rational approximants for some performance analysis problems. IEEE Transactions on Computers, 44(12):1394–1404, December 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=477245.

Guan:1998:PPN

[GYY98] Sheng-Uei Guan, Hsiao-Yeh Yu, and Jen-Shun Yang. A prioritized Petri net model and its application in distributed multimedia systems. IEEE Transactions on Computers, 47(4):477–481, April 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=675716.

Healy:1999:BPI

[HAM^{+}99] C. A. Healy, R. D. Arnold, F. Mueller, D. B. Whalley, and M. G. Harmon. Bounding pipeline and instruction cache performance. IEEE Transactions on Computers, 48(1):53–70, January 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=743411.

Han:1991:OLL

[Han91] Y. Han. An optimal linked list prefix algorithm on a local memory computer. IEEE Transactions on Computers, 40 (10):1149–1153, October 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=93747.

Harper:1992:IMP

[Har92] D. T. Harper III. Increased memory performance during vector accesses through the use of linear address transformations. IEEE Transactions on Computers, 41(2):227–230, February 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=123399.

Harper:1994:MFB

[Har94] D. T. Harper III. A multiaccess frame buffer architecture. IEEE Transactions on Computers, 43(5):618–622, May 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=280810.

See comments [Wei96].

Hashemian:1990:SRA

[Has90] R. Hashemian. Square rooting algorithms for integer and floating-point numbers. IEEE Transactions on Computers, 39 (8):1025–1029, August 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=57041.

Hasan:1997:DA

[Has97]  M. A. Hasan.  Division-

and-accumulation over GF(2^{m}). IEEE Transactions on Computers, 46(6):705–708, June 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=600829.

Hasan:1998:DBM

[Has98] M. A. Hasan. Double-basis multiplicative inversion over GF(2^{m}). IEEE Transactions on Computers, 47(9):960–970, September 1998. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        713315.

Hasan:1992:BSS

[HB92] M. A. Hasan and V. K. Bhargava. Bit-serial systolic divider and multiplier for finite fields GF(2^{m}). IEEE Transactions on Computers, 41(8):972–980, August 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=156540.

Hoelzeman:1994:GSG

[HB94] D. A. Hoelzeman and S. Bettayeb. On the genus of star graphs. IEEE Transactions on Computers, 43(6):755–759, June 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=286309.

Hasan:1995:ALC

[HB95] M. A. Hasan and V. K. Bhargava. Architecture for a low complexity rate-adaptive Reed– Solomon encoder. IEEE Transactions on Computers, 44(7): 938–942, July 1995. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
392853.

Hasteer:1998:PAS

[HB98] G. Hasteer and P. Banerjee. A parallel algorithm for state assignment of finite state machines. IEEE Transactions on Computers, 47(2):242–246, February 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=663772.

Ho:1998:PSQ

[HBA98] Ching-Tien Ho, J. Bruck, and R. Agrawal. Partialsum queries in OLAP data cubes using covering codes. IEEE Transactions on Computers, 47(12):1326–1340, December 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=737680.

Hulgaard:1995:AEB

[HBAB95] H. Hulgaard, S. M. Burns, T. Amon, and G. Borriello. An algorithm for exact bounds on the time separation of events in concurrent systems. IEEE Transactions on Computers, 44(11):1306–1317, November 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=475126.

Hwu:1992:EIS

[HC92] W. W. Hwu and P. P. Chang. Efficient instruction sequencing with inline target insertion. IEEE Transactions on Computers, 41(12):1537–1551, December 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=214662.

Harper:1993:AEV

[HC93] D. T. Harper III and Y. Costa. Analytical estimation of vector access performance in parallel memory architectures. IEEE Transactions on Computers, 42(5):616–624, May 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=223682.

Hemkumar:1994:RLC

[HC94a] N. D. Hemkumar and J. R. Cavallaro. Redundant and on-line CORDIC for unitary transformations. IEEE Transactions on Computers, 43(8):941–954, August 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=295856.

Hwu:1994:SPC

[HC94b] W.-M. W. Hwu and T. M.

Conte. The susceptibility of programs to context switching. IEEE Transactions on Computers, 43(9):994–1003, September 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=312110.

Hu:1996:NIC

[HC96] Yu Hen Hu and H. H. M.

Chern. A novel implementation of CORDIC algorithm using backward angle recoding (BAR). IEEE Transactions on Computers, 45(12):1370–1378, December 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=545967.

Hsiao:1995:HCA

[HD95] Shen-Fu Hsiao and Jean-Marc Delosme. Householder CORDIC algorithms. IEEE Transactions on Computers, 44(8):990–1001, August 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=403715.

Hanchek:1998:MTC

[HD98] F. Hanchek and S. Dutt. Methodologies for tolerating cell and interconnect faults in FP-GAs. IEEE Transactions on Computers, 47(1):15–33, January 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=656073.

Haddad:1991:ITT

[HDS91] R. W. Haddad, A. T. Dahbura, and A. B. Sharma. Increased throughput for the testing and repair of RAMs with redundancy. IEEE Transactions on Computers, 40 (2):154–166, February 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=73586.

Heidtmann:1991:ASA

[Hei91] K. D. Heidtmann. Arithmetic spectrum applied to fault detection for combinational networks. IEEE Transactions on Computers, 40(3):320–324, March 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=76409.

Hamada:1997:CEC

[HF97] M. Hamada and E. Fujiwara.

A class of error control codes for byte organized memory systems-SbEC-(Sb+S)ED codes. IEEE Transactions on Computers, 46(1):105–109, January 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=559809.

Hsu:1996:SSR

[HG96] Yuan-Chieh Hsu and S. K.

Gupta. A simulator for at-speed robust testing of path delay faults in combinational circuits. IEEE Transactions on Computers, 45(11):1312–1318, November 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=544489.

Holi:1993:FPE

[HH93] J. L. Holi and J.-N. Hwang. Finite precision error analysis of neural network hardware implementations. IEEE Transactions on Computers, 42(3):281–290, March 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=210171.

Hu:1991:ERC

[HHB91] X. Hu, R. G. Harber, and S. C.

Bass. Expanding the range of convergence of the CORDIC algorithm. IEEE Transactions on Computers, 40(1):13–21, January 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=67316.

Heck:1991:HPE

[HHMC91] E. Heck, D. Hogrefe, and B. Muller-Clostermann. Hierarchical performance evaluation based on formally specified communication protocols. IEEE Transactions on Computers, 40(4):500–513, April 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=88469.

Han:1997:SAT

[HHS97] Ching-Chih Han, Chao-Ju Hou, and K. G. Shin. On slot allocation for time-constrained messages in dual-bus networks. IEEE Transactions on Computers, 46(7):756–767, July 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=599896.

Hosseini:1992:EDA

[HJ92a] S. H. Hosseini and N. Jamal. Efficient distributed algorithms for self testing of multiple processor systems. IEEE Transactions on Computers, 41 (11):1397–1409, November 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=177310.

Huang:1992:EFT

[HJ92b] Y. Huang and P. Jalote. Effect of fault tolerance on response time-analysis of the primary site approach. IEEE Transactions on Computers, 41(4):420–428, April 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=135555.

Ha:1990:DHY

[HK90] D. S. Ha and V. P. Kumar. On the design of high-yield reconfigurable PLA’s. IEEE Transactions on Computers, 39(4):470–479, April 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=54840.

Hitz:1995:IDR

[HK95] Markus A. Hitz and Erich Kaltofen. Integer division in residue number systems. IEEE Transactions on Computers, 44(8):983–989, August 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=403714;
 https://ieeexplore.ieee.
org/xpl/tocresult.jsp?isnumber=

Hou:1997:ORT

[HKC97] Rong-Huei Hou, Sy-Yen Kuo, and Yi-Ping Chang. Optimal release times for software systems with scheduled delivery time based on the HGDM. IEEE Transactions on Computers, 46(2):216–221, February 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=565602.

Hurson:1998:CSS

[HKL98] A. R. Hurson, K. Kavi, and J. T. Lim. Cyclic staggered scheme: a loop allocation policy for DOACROSS loops. IEEE Transactions on Computers, 47(2):251–255, February 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=663774.

Harper:1999:AMS

[HKN99] J. S. Harper, D. J. Kerbyson, and G. R. Nudd. Analytical modeling of set-associative cache behavior. IEEE Transactions on Computers, 48(10):1009–1024, October 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=805152.

Ho:1990:PAS

[HL90] C.-W. Ho and R. C. T. Lee. A parallel algorithm for solving sparse triangular systems.

IEEE Transactions on Computers, 39(6):848–852, June 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=53610.

Ha:1991:CTS

[HL91a] S. Ha and E. A. Lee. Compiletime scheduling and assignment of data-flow program graphs with data-dependent iteration. IEEE Transactions on Computers, 40(11):1225–1238, November 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=102826.

Harper:1991:CFV

[HL91b] D. T. Harper III and D. A.

Linebarger. Conflict-free vector access using a dynamic storage scheme. IEEE Transactions on Computers, 40(3):276–283, March 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=76404.

Hansen:1992:IAP

[HL92a] P. Hansen and K.-W. Lih. Improved algorithms for partitioning problems in parallel, pipelined, and distributed computing. IEEE Transactions on Computers, 41(6):769–771, June 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=144628.

Hong:1992:LSR

[HL92b] K. S. Hong and J. Y.-T. Leung. On-line scheduling of real-time tasks. IEEE Transactions on Computers, 41 (10):1326–1331, October 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=166609.

Ha:1997:CTS

[HL97] S. Ha and E. A. Lee. Compiletime scheduling of dynamic constructs in dataflow program graphs. IEEE Transactions on Computers, 46(7):768–778, July 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=599897.

Hlawiczka:1992:PSA

[Hla92] A. Hlawiczka. Parallel signature analyzers using hybrid design of their linear feedbacks. IEEE Transactions on Computers, 41(12):1562–1571, December 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=214664.

Hansson:1997:BAM

[HLB^{+}97] H. Hansson, L. Lawson, O. Bridal, C. Eriksson, S. Larsson, H. Lon, and M. Stromberg. BASE-MENT: an architecture and methodology for distributed automotive real-time systems. IEEE Transactions on Computers, 46(9):1016–1027, September 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=620482.

Han:1996:DCS

[HLH96] Ching-Chih Han, Kwei-Jay Lin, and Chao-Ju Hou. Distanceconstrained scheduling and its applications to real-time systems. IEEE Transactions on Computers, 45(7):814–826, July 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=508320.

Hsu:1991:CLM

[HLHC91] Yu Chin Hsu, Youn Long Lin, Hang Ching Hsieh, and Ting Hai Chao. Combining logic minimization and folding for PLAs. IEEE Transactions on Computers, 40(6):706–713, June 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=90249.

Hua:1993:DHP

[HLP93] K. A. Hua, L. Liu, and J. Peir. Designing highperformance processors using real address prediction. IEEE Transactions on Computers, 42 (9):1146–1151, September 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=241604.

Hong:1991:AMC

[HM91] S. J. Hong and S. Muroga. Absolute minimization of completely specified switching functions. IEEE Transactions on Computers, 40(1):53–65, January 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=67320.

Hao:1993:AGO

[HM93] Hong Hao and E. J. Mc-Cluskey. Analysis of gate oxide shorts in CMOS circuits. IEEE Transactions on Computers, 42(12):1510–1516, December 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=260643.

Heap:1994:LUB

[HM94] M. A. Heap and M. R. Mercer. Least upper bounds on OBDD sizes. IEEE Transactions on Computers, 43(6):764–767, June 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=286311.

Hady:1995:PCB

[HM95] F. T. Hady and B. L. Menezes.

The performance of crossbarbased binary hypercubes. IEEE Transactions on Computers, 44 (10):1208–1215, October 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=467695.

Hortensius:1990:CAB

[HMC90] P. D. Hortensius, R. D. McLeod, and H. C. Card. Cellular automata-based signature analysis for built-in self-test. IEEE Transactions on Computers, 39(10):1273–1283, October 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=59857.

Hornick:1990:ST

[HMM^{+}90] S. W. Hornick, S. R. Maddila, E. P. Mucke, H. Rosenberger, S. S. Skiena, and I. G. Tollis. Searching on a tape. IEEE Transactions on Computers, 39(10):1265–1272, October 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=59856.

Hu:1993:ARM

[HN93] Y. H. Hu and S. Naganathan.

An angle recording method for CORDIC algorithm implementation. IEEE Transactions on Computers, 42(1):99–102, January 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=192217.

Hwang:1994:CNR

[HNG94] F. K. Hwang, W. Najjar, and J. L. Gaudiot. Comments on “Network resilience: a measure of network fault tolerance” [and reply]. IEEE Transactions on Computers, 43 (12):1451–1453, December 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=338109.

Heidelberger:1990:PQU

[HNR90] P. Heidelberger, A. Norton, and J. T. Robinson. Parallel Quicksort using fetch-and-add. IEEE Transactions on Computers, 39(1):133–138, January 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=46289.

Hata:1993:SFP

[HNY93] Y. Hata, K. Nakashima, and K. Yamato. Some fundamental properties of multiple-valued Kleenean functions and determination of their logic formulas. IEEE Transactions on Computers, 42(8):950–961, August 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=238485.

Ho:1992:OBI

[Ho92] C.-T. Ho. An observation on the bisectional interconnection networks. IEEE Transactions on Computers, 41(7):873–877, July 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=256455.

Hoganson:1999:WES

[Hog99] K. E. Hoganson. Workload execution strategies and parallel speedup on clustered computers. IEEE Transactions on Computers, 48(11):1173–1182, November 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=811107.

Hong:1990:DTP

[Hon90] S. J. Hong. The design of a testable parallel multiplier. IEEE Transactions on Computers, 39(3):411–416, March 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=48874.

Hou:1999:RVC

[Hou99] C.-J. J. Hou. Routing virtual circuits with temporal QoS requirements in virtual pathbased ATM networks. IEEE Transactions on Computers, 48 (11):1228–1243, November 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=811111.

Hung:1995:EAA

[HP95] Ching Yu Hung and B. Parhami. Error analysis of approximate Chinese-Remainder-Theorem decoding. IEEE Transactions on Computers, 44(11):1344–1348, November 1995. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
475132.

Haniotakis:1995:ETS

[HPN95] T. Haniotakis, A. Paschalis, and D. Nikolos. Efficient totally self-checking checkers for a class of Borden codes. IEEE Transactions on Computers, 44(11):1318–1322, November 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=475127.

Hamdaoui:1995:DRT

[HR95a] M. Hamdaoui and P. Ramanathan. Deferring real-time traffic for improved non-real-time communication in FDDI networks. IEEE Transactions on Computers, 44(12): 1420–1428, December 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=477247.

Hamdaoui:1995:DPA

[HR95b] M. Hamdaoui and P. Ramanathan. A dynamic priority assignment technique for streams with (m, k)-firm deadlines. IEEE Transactions on Computers, 44 (12):1443–1451, December 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=477249.

Hamsaoui:1997:EDF

[HR97] M. Hamsaoui and P. Ramanathan. Evaluating dynamic failure probability for streams with (m, k)-firm deadlines. IEEE Transactions on Computers, 46 (12):1325–1337, December 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=641933.

Hsiao:1999:FSC

[HRP99] M. S. Hsiao, E. M. Rudnick, and J. H. Patel. Fast static compaction algorithms for sequential circuit test vectors. IEEE Transactions on Computers, 48(3):311–322, March 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=754997.

Hellebrand:1995:BTC

[HRT^{+}95] S. Hellebrand, J. Rajski, S. Tarnick, S. Venkataraman, and B. Courtois. Built-in test for circuits with scan based on reseeding of multiple-polynomial linear feedback shift registers. IEEE Transactions on Computers, 44(2):223–233, February 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=364534.

Hua:1993:DLB

[HS93] K. A. Hua and J. X. W. Su. Dynamic load balancing in very large shared-nothing hypercube database computers. IEEE Transactions on Computers, 42(12):1425–1439, December 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=260633.

Harrington:1994:SHN

[HS94a] M. Harrington and A. K. Somani. Synchronizing hypercube networks in the presence of faults. IEEE Transactions on Computers, 43 (10):1175–1183, October 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=324543.

Holliday:1994:PEH

[HS94b] M. Holliday and M. Stumm. Performance evaluation of hierarchical ring-based shared memory multiprocessors. IEEE Transactions on Computers, 43(1):52–67, January 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=250609.

Hou:1994:IOT

[HS94c] Chao-Ju Hou and K. G. Shin.

Incorporation of optimal timeouts into distributed real-time load sharing. IEEE Transactions on Computers, 43(5):528–547, May 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=280801.

Hou:1994:LSC

[HS94d] Chao-Ju Hou and K. G. Shin.

Load sharing with consideration of future task arrivals in heterogeneous distributed realtime systems. IEEE Transactions on Computers, 43(9): 1076–1090, September 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=312127.

Hou:1996:DOR

[HS96] Chao-Ju Hou and K. G. Shin.

Determination of an optimal retry time in multiplemodule computing systems. IEEE Transactions on Computers, 45(3):374–379, March 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=485576.

Hou:1997:APT

[HS97] Chao-Ju Hou and K. G. Shin. Allocation of periodic task modules with precedence and deadline constraints in distributed real-time systems. IEEE Transactions on Computers, 46(12): 1338–1356, December 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=641934.

Han:1998:PBC

[HS98a] Seungjae Han and K. G. Shin. A primary-backup channel approach to dependable real-time communication in multihop networks. IEEE Transactions on Computers, 47(1):46–61, January 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=656080.

Hiltunen:1998:CMS

[HS98b] M. A. Hiltunen and R. D. Schlichting. A configurable membership service. IEEE Transactions on Computers, 47(5):573– 586, May 1998. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
677244.

Hsu:1998:PSI

[HS98c] W.-C. Hsu and J. E. Smith. A performance study of instruction cache prefetching methods. IEEE Transactions on Computers, 47(5):497–508, May 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=677221.

Heinrich:1999:QAP

[HSHG99] M. Heinrich, V. Soundararajan, J. Hennessy, and A. Gupta. A quantitative analysis of the performance and scalability of distributed shared memory cache coherence protocols. IEEE Transactions on Computers, 48(2):205–217, February 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=752662.

Hu:1996:ORL

[HSL96] Qing Hu, Xiaojun Shen, and Weifa Liang. Optimally routing LC permutations on kextra-stage cube-type networks. IEEE Transactions on Computers, 45(1):97–103, January 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=481490.

Hu:1997:TCS

[HSY97] Qing Hu, Xiaojun Shen, and Jingyu Yang. Topologies of combined (2 log N − 1)- stage interconnection networks. IEEE Transactions on Computers, 46(1):118–124, January 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=559812.

Heys:1995:ACS

[HT95] H. M. Heys and S. E. Tavares. Avalanche characteristics of substitution-permutation encryption networks. IEEE Transactions on Computers, 44(9): 1131–1139, September 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=464391.

Huang:1991:ERA

[HTCT91] Shing-Tsaan Huang, S. K. Tripathi, Nian-Shing Chen, and Yu-Chee Tseng. An efficient routing algorithm for realizing linear permutations on p^{t}-shuffle-exchange networks. IEEE Transactions on Computers, 40(11):1292–1298, November 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=102836.

Hahn:1994:UDF

[HTHR94] H. Hahn, D. Timmermann, B. J. Hosticka, and B. Rix. A unified and division-free CORDIC argument reduction method with unlimited convergence domain including inverse hyperbolic functions. IEEE Transactions on Computers, 43 (11):1339–1344, November 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=324568.

Hamano:1997:DME

[HTYP97] T. Hamano, N. Takagi, S. Yajima, and F. P. Preparata. O(n)- depth modular exponentiation circuit algorithm. IEEE Transactions on Computers, 46(6): 701–704, June 1997. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
600828.

Huang:1990:NSE

[Hua90] S.-T. Huang. Notes on shuffle/exchange type permutation sets. IEEE Transactions on Computers, 39(7):962–965, July 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=55699.

Huang:1992:CTC

[Hua92] Nen-Fu Huang. On the complexity of two circle strongly connecting problems. IEEE Transactions on Computers, 41 (9):1185–1188, September 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=165400.

Hughey:1993:CED

[Hug93] R. Hughey. Concurrent error detection on programmable systolic arrays. IEEE Transactions on Computers, 42(6):752–756, June 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=277295.

Haglin:1991:AIR

[HV91] D. J. Haglin and S. M. Venkatesan. Approximation and intractability results for the maximum cut problem and its variants. IEEE Transactions on Computers, 40(1):110–113, January 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=67327.

Hwang:1995:SRS

[HW95] F. K. Hwang and P. E. Wright. Survival reliability of some double-loop networks and chordal rings. IEEE Transactions on Computers, 44 (12):1468–1471, December 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=477253.

Hwang:1997:CFC

[HW97] F. K. Hwang and P. J. Wan.

Comparing file copies with at most three disagreeing pages. IEEE Transactions on Computers, 46(6):716–718, June 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=600831.

Hwang:1997:MDA

[Hwa97] F. K. Hwang. A modification to a decomposition algorithm of Gordon and Srikanthan. IEEE Transactions on Computers, 46(8):958–960, August 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=609284.

Hasan:1992:MCL

[HWB92] M. A. Hasan, M. Wang, and V. K. Bhargava. Modular construction of low complexity parallel multipliers for a class of finite fields GF(2^{m}). IEEE Transactions on Computers, 41(8):962–971, August 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=156539.

Hasan:1993:MMO

[HWB93] M. A. Hasan, M. Z. Wang, and V. K. Bhargava. A modified Massey–Omura parallel multiplier for a class of finite fields. IEEE Transactions on Computers, 42(10):1278–1280, October 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=257715.

Harada:1999:IAM

[HY99] T. Harada and M. Yamashita. Improving the availability of mutual exclusion systems on incomplete networks. IEEE Transactions on Computers, 48(7):744– 747, July 1999. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
780882.

Islam:1991:PAD

[IA91] S. M. R. Islam and H. H. Ammar. Performability analysis of distributed real-time systems. IEEE Transactions on Computers, 40(11):1239–1251, November 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=102827.

Iwamura:1995:DRS

[IDI95] K. Iwamura, Y. Dohi, and H. Imai. A design of Reed– Solomon decoder with systolicarray structure. IEEE Transactions on Computers, 44 (1):118–122, January 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=368005.

Ichikawa:1991:SIS

[IIK^{+}91] H. Ichikawa, M. Itoh, J. Kato, A. Takura, and M. Shibasaki. SDE: incremental specification and development of communications software. IEEE Transactions on Computers, 40(4):553– 561, April 1991. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
88473.

Iyengar:1994:VAD

[IJN94] S. S. Iyengar, D. N. Jayasimha, and D. Nadig. A versatile architecture for the distributed sensor integration problem. IEEE Transactions on Computers, 43(2):175–185, February 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=262122.

Ibarra:1992:SEO

[IJW92] O. H. Ibarra, T. Jiang, and H. Wang. String editing on a oneway linear array of finite-state machines. IEEE Transactions on Computers, 41(1):112–118, January 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=123389.

Iwasaki:1996:AEM

[IN96] K. Iwasaki and S. Nakamura. Aliasing error for a mask ROM built-in self-test. IEEE Transactions on Computers, 45(3):270–277, March 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=485566.

Irwin:1991:TDD

[IO91] M. J. Irwin and R. M. Owens. A two-dimensional, distributed logic architecture. IEEE Transactions on Computers, 40 (10):1094–1101, October 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=93742.

Ikenaga:1998:CHP

[IO98] T. Ikenaga and T. Ogura. CAM^{2}: a highly-parallel twodimensional cellular automaton architecture. IEEE Transactions on Computers, 47(7):788–801, July 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=709379.

Iverson:1999:SPT

[IOP99] M. A. Iverson, F. Ozguner, and L. Potter. Statistical prediction of task execution times through analytic benchmarking for scheduling in a heterogeneous environment. IEEE Transactions on Computers, 48(12):1374–1379, December 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=817403.

Ibarra:1991:PRP

[IPS91] O. H. Ibarra, T.-C. Pong, and S. M. Sohn. Parallel recognition and parsing on the hypercube. IEEE Transactions on Computers, 40(6):764–770, June 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=90253.

Isenman:1990:PAI

[IS90] M. E. Isenman and D. E.

Shasha. Performance and architectural issues for string matching. IEEE Transactions on Computers, 39(2):238–250, February 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=45209.

Intrater:1994:PED

[IS94] G. D. Intrater and I. Y.

Spillinger. Performance evaluation of a decoded instruction cache for variable instruction length computers. IEEE Transactions on Computers, 43 (10):1140–1150, October 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=324540.

Ishikawa:1995:HMB

[Ish95] T. Ishikawa. Hypercube multiprocessors with bus connections for improving communication performance. IEEE Transactions on Computers, 44 (11):1338–1344, November 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=475131.

Ito:1997:EIA

[ITY97] M. Ito, N. Takagi, and S. Yajima. Efficient initial approximation for multiplicative division and square root by a multiplication with operand modification. IEEE Transactions on Computers, 46(4):495–498, April 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=588066.

Ivanov:1996:PSC

[ITZ96] A. Ivanov, B. K. Tsuji, and Y. Zorian. Programmable BIST space compactors. IEEE Transactions on Computers, 45 (12):1393–1404, December 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=545969.

Ienne:1994:BSM

[IV94] P. Ienne and M. A. Viredaz. Bitserial multipliers and squarers. IEEE Transactions on Computers, 43(12):1445–1450, December 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=338107.

Iyer:1990:ARI

[IYI90] R. K. Iyer, L. T. Young, and P. V. K. Iyer. Automatic recognition of intermittent failures: an experimental study of field data. IEEE Transactions on Computers, 39(4):525–537, April 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=54845.

Jain:1997:IBA

[JBA^{+}97] J. Jain, J. Bitner, M. S. Abadir, J. A. Abraham, and D. S. Fussell. Indexed BDDs: algorithmic advances in techniques to represent and verify Boolean functions. IEEE Transactions on Computers, 46(11):1230–1245, November 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=644298.

Jacob:1990:FCR

[JBS90] J. Jacob, N. N. Biswas, and J. E. Smith. Further comments, with reply, on “Detection of faults in programmable logic arrays” by J. E. Smith. IEEE Transactions on Computers, 39(1):155–157, January 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=46295.

Johnson:1999:RTC

[JCMH99] T. L. Johnson, D. A. Connors, M. C. Merten, and W.-M. W. Hwu. Run-time cache bypassing. IEEE Transactions on Computers, 48(12):1338–1354, December 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=817393.

Jacob:1997:CAM

[JCMS97] B. L. Jacob, P. M. Chen, T. N. Mudge, and S. R. Silverman. A comment on “An analytical model for designing memory hierarchies”. IEEE Transactions on Computers, 46(10):1151, October 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=628401.

See [JCSM96].

Jacob:1996:AMD

[JCSM96] Bruce L. Jacob, Peter M. Chen, Seth R. Silverman, and Trevor N.

Mudge. An analytical model for designing memory hierarchies. IEEE Transactions on Computers, 45(10):1180–1194, October 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=543711.

Simulates a storage hierarchy similar to that of Plan 9 [?, ?]. See comment [JCMS97].

Joseph:1999:PUM

[JG99] D. Joseph and D. Grunwald. Prefetching using Markov predictors. IEEE Transactions on Computers, 48 (2):121–133, February 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=752653.

Johnsson:1995:CBB

[JH95] S. Lennart Johnsson and Ching-Tien Ho. On the conversion between binary code and binaryreflected Gray code on binary cubes. IEEE Transactions on Computers, 44(1):47–53, January 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=368010.

Jha:1993:FDC

[Jha93] N. K. Jha. Fault detection in CVS parity trees with application to strongly self-checking parity and two-rail checkers.

IEEE Transactions on Computers, 42(2):179–189, February 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=204791.

Jiang:1997:CSF

[JHK97] Jehn-Ruey Jiang, Shing-Tsaan Huang, and Yu-Chen Kuo. Cohorts structures for fault-tolerant k entries to a critical section. IEEE Transactions on Computers, 46(2):222–228, February 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=565603.

Jen:1992:DFR

[JK92] Chein-Wei Jen and Ding-Ming Kwai. Data flow representation of iterative algorithms for systolic arrays. IEEE Transactions on Computers, 41(3):351–355, March 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=127448.

Jain:1991:SRT

[JL91] P. Jain and S. S. Lam. Specification of real-time broadcast networks. IEEE Transactions on Computers, 40(4):404–422, April 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=88461.

John:1996:PMP

[JL96] L. K. John and Yu-Cheng Liu.

Performance model for a prioritized multiple-bus multiprocessor system. IEEE Transactions on Computers, 45(5):580–588, May 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=509909.

Jin:1998:ESC

[JLC98] G. Jin, Z. Li, and F. Chen. An efficient solution to the cache thrashing problem caused by true data sharing. IEEE Transactions on Computers, 47(5): 527–543, May 1998. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
677228.

Jordan:1994:SAT

[JLLR94] H. F. Jordan, D. Lee, K. Y.

Lee, and S. V. Ramanan. Serial array time slot interchangers and optical implementations. IEEE Transactions on Computers, 43(11):1309–1318, November 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=324563.

Jiang:1996:SNR

[JM96] Honghai Jiang and J. C. Majithia. Suggestion for a new representation for binary function.

IEEE Transactions on Computers, 45(12):1445–1449, December 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=545975.

Jan:1993:FSR

[JO93] Ching Yuh Jan and A. Y. Oruc. Fast self-routing permutation switching on an asymptotically minimum cost network. IEEE Transactions on Computers, 42(12):1469–1479, December 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=260636.

Johansson:1993:CCC

[Joh93]  R. Johansson.  A class of

(12, 8) codes for correcting single errors and detecting double errors within a nibble. IEEE Transactions on Computers, 42(12):1504–1506, December 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=260641.

Johnson:1995:CPA

[Joh95]  T. Johnson.  Characteriz-

ing the performance of algorithms for lock-free objects. IEEE Transactions on Computers, 44(10):1194–1207, October 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=467694.

Jessani:1998:CSD

[JP98] R. M. Jessani and M. Putrino. Comparison of singleand dual-pass multiply-add fused floating-point units. IEEE Transactions on Computers, 47 (9):927–937, September 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=713312.

Jiang:1992:PPM

[JS92] H. Jiang and K. C. Smith.

PPMB: a partial-multiple-bus multiprocessor architecture with improved cost-effectiveness. IEEE Transactions on Computers, 41(3):361–366, March 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=127450.

Joseph:1997:MCR

[JTK97] A. D. Joseph, J. A. Tauber, and M. F. Kaashoek. Mobile computing with the Rover toolkit. IEEE Transactions on Computers, 46(3):337–352, March 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=580429.

Jean:1999:DRS

[JTY^{+}99] J. S. N. Jean, K. Tomko, V. Yavagal, J. Shah, and R. Cook. Dynamic reconfiguration to support concurrent applications. IEEE Transactions on Computers, 48(6):591–602, June 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=773796.

Jung:1995:CSA

[Jun95]  Gihyan Jung.  Comments

on “Some additions to solution of switching equations based on a tabular algebra”. IEEE Transactions on Computers, 44(11):1357–1358, November 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=475135.

Juang:1991:CBB

[JW91] J.-Y. Juang and B. W. Wah. A contention-based bus-control scheme for multiprocessor systems. IEEE Transactions on Computers, 40(9):1046–1053, September 1991. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
83647.

Jone:1994:MFD

[JW94] Wen-Ben Jone and Cheng-Juei Wu. Multiple fault detection in parity checkers.

IEEE Transactions on Computers, 43(9):1096–1099, September 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=312118.

Kantabutra:1994:SAA

[KA94] V. Kantabutra and A. G. Andreou. A state assignment approach to asynchronous CMOS circuit design. IEEE Transactions on Computers, 43(4):460– 469, April 1994. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
278483.

Kumar:1996:PSE

[KA96] A. Kumar and D. P. Agrawal.

Parameters for system effectiveness evaluation of distributed systems. IEEE Transactions on Computers, 45(6):746–752, June 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=506430.

Karkouri:1994:UFD

[KACV94] Y. Karkouri, E. M. Aboulhamid, E. Cerny, and A. Verreault. Use of fault dropping for multiple fault analysis. IEEE Transactions on Computers, 43(1):98–103, January 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=250613.

Kaliski:1995:MIA

[Kal95] Burton S. Kaliski, Jr. The Montgomery inverse and its applications. IEEE Transactions on Computers, 44(8):1064–1065, August 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=403725.

Kamal:1992:AEU

[Kam92] A. E. Kamal. An algorithm for the efficient utilization of bandwidth in the slotted ring. IEEE Transactions on Computers, 41(12):1620–1627, December 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=214673.

Kant:1990:PAR

[Kan90] K. Kant. Performance analysis of real-time software supporting fault-tolerant operation. IEEE Transactions on Computers, 39(7):906–918, July 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=55692.

Kantabutra:1993:ATL

[Kan93a] V. Kantabutra. Accelerated two-level carry-skip adders — a type of very fast adders.

IEEE Transactions on Computers, 42(11):1389–1393, November 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=247841.

Kantabutra:1993:DOO

[Kan93b] V. Kantabutra. Designing optimum one-level carry-skip adders. IEEE Transactions on Computers, 42(6):759–764, June 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=277297.

Kantabutra:1993:RCL

[Kan93c] V. Kantabutra. A recursive carry-lookahead/carry-select hybrid adder. IEEE Transactions on Computers, 42(12): 1495–1499, December 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=260639.

Kantabutra:1996:HCE

[Kan96]  V. Kantabutra.  On hard-

ware for computing exponential and trigonometric functions. IEEE Transactions on Computers, 45(3):328–339, March 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=485571.

Katti:1994:CDC

[Kat94] R. Katti. Comments on “Decomposition of Complex Multipliers Using Polynomial Encoding”. IEEE Transactions on Computers, 43(3):381–383, March 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=272441.

See [SS92].

Katti:1996:NSA

[Kat96a] R. Katti. A note on SEC/ AUED codes. IEEE Transactions on Computers, 45 (2):244–246, February 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=485377.

Katti:1996:NRA

[Kat96b] Rajendra S. Katti. A new residue arithmetic error correction scheme. IEEE Transactions on Computers, 45(1):13–19, January 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=481482.

Katti:1997:NMS

[Kat97] R. S. Katti. Nonprime memory systems and error correction in address translation. IEEE Transactions on Computers, 46(1):75–79, January 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=559804.

Kavianpour:1993:SAM

[KB93] A. Kavianpour and N. Bagherzadeh.

A systematic approach for mapping application tasks in hypercubes. IEEE Transactions on Computers, 42(6):742–746, June 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=277294.

Kulasinghe:1995:EBT

[KB95] P. Kulasinghe and S. Bettayeb.

Embedding binary trees into crossed cubes. IEEE Transactions on Computers, 44(7):923– 929, July 1995. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
392850.

Katti:1996:ICE

[KB96] R. S. Katti and M. Blaum.

An improvement on constructions of t-EC/AUED codes. IEEE Transactions on Computers, 45(5):607–608, May 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=509912.

Kanoun:1999:ACS

[KBMP99] K. Kanoun, M. Borrel, T. Morteveille, and A. Peytavin. Availability of CAUTRA, a subset of the French air traffic control system. IEEE Transactions on Computers, 48(5):528–535, May 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=769435.

Karpovsky:1993:DSD

[KC93a] M. G. Karpovsky and S. M. Chaudhry. Design of selfdiagnostic boards by multiple signature analysis. IEEE Transactions on Computers, 42 (9):1035–1044, September 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=241593.

Kota:1993:NAH

[KC93b] K. Kota and J. R. Cavallaro. Numerical accuracy and hardware tradeoffs for CORDIC arithmetic for special-purpose processors. IEEE Transactions on Computers, 42(7):769–779, July 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=237718.

Kekckler:1999:CEH

[KCCD99] S. W. Kekckler, A. Chang, W. S. L. S. Chatterjee, and W. J. Dally. Concurrent event handling through multithreading. IEEE Transactions on Computers, 48(9):903–916, September 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=795220.

Kim:1994:HCD

[KD94] Jong Kim and C. R. Das.

Hypercube communication delay with wormhole routing. IEEE Transactions on Computers, 43(7):806–814, July 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=293259.

Kaeli:1997:IAH

[KE97] D. R. Kaeli and P. G. Emma.

Improving the accuracy of history-based branch prediction. IEEE Transactions on Computers, 46(4):469–472, April 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=588060.

Kulasinghe:1995:COB

[KEA95] P. Kulasinghe and A. El-Amawy. On the complexity of optimal bused interconnections. IEEE Transactions on Computers, 44(10):1248–1251, October 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=467700.

Kulasinghe:1996:ORS

[KEA96] P. Kulasinghe and A. El-Amawy. Optimal realization of sets of interconnection functions on synchronous multiple bus systems. IEEE Transactions on Computers, 45(8):964–969, August 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=536238.

Kuo:1992:FDS

[KF92] Sy-Yen Kuo and W. K. Fuchs. Fault diagnosis and spare allocation for yield enhancement in large reconfigurable PLAs. IEEE Transactions on Computers, 41(2):221–226, February 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=123398.

Khanna:1997:GPA

[KF97] S. Khanna and W. K. Fuchs. A graph partitioning approach to sequential diagnosis. IEEE Transactions on Computers, 46(1):39–47, January 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=559801.

Kakugawa:1993:AC

[KFYA93] H. Kakugawa, S. Fujita, M. Yamashita, and T. Ae. Availability of k-coterie. IEEE Transactions on Computers, 42(5):553–558, May 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=223674.

See comments [YC94].

Katti:1995:CSC

[KGB95] R. S. Katti, T. A. Gulliver, and V. K. Bhargava. Comments on “A systematic (16, 8) code for correcting double errors and detecting tripleadjacent errors” [and reply]. IEEE Transactions on Computers, 44(12):1472–1474, December 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=477254.

See [GB93].

Kamal:1990:UBS

[KH90] A. E. Kamal and V. C.

Hamacher. Utilizing bandwidth sharing in the slotted ring. IEEE Transactions on Computers, 39(3):289–299, March 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=48861.

Kamal:1993:ROB

[KH93] A. E. Kamal and V. C.

Hamacher. Response to “Onebit delay in ring networks” by R. Cohen. IEEE Transactions on Computers, 42(6):737–738, June 1993. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=277303.

See [Coh93].

Ku:1997:SDF

[KH97] Hung-Kuei Ku and J. P. Hayes. Systematic design of faulttolerant multiprocessors with shared buses. IEEE Transactions on Computers, 46(4):439–455, April 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=588058.

Kurian:1994:MLE

[KHC94] L. Kurian, P. T. Hulina, and L. D. Coraor. Memory latency effects in decoupled architectures. IEEE Transactions on Computers, 43(10):1129–1139, October 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=324539.

Kessler:1994:CTS

[KHW94] R. E. Kessler, M. D. Hill, and D. A. Wood. A comparison of trace-sampling techniques for multi-megabyte caches. IEEE Transactions on Computers, 43 (6):664–675, June 1994. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL https://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        286300.

King:1991:FPE

[Kin91] P. W. King. Formalization of protocol engineering concepts. IEEE Transactions on Computers, 40(4):387–403, April 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=88460.

Kawahito:1994:HSA

[KIN^{+}94] S. Kawahito, M. Ishida, T. Nakamura, M. Kameyama, and T. Higuchi. High-speed areaefficient multiplier design using multiple-valued current-mode circuits. IEEE Transactions on Computers, 43(1):34–42, January 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=250607.

See comments [PKI^{+}96].

Kawahito:1996:AR

[KIN^{+}96] S. Kawahito, M. Ishida, T. Nakamura, M. Kameyama, and T. Higuchi. Author’s reply. IEEE Transactions on Computers, 45(5):639, May 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=509919.

Kipnis:1994:AAB

[Kip94] S. Kipnis. Analysis of asynchronous binary arbitration on digital transmission-line busses.

IEEE Transactions on Computers, 43(4):484–489, April 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=278487.

Kavianpour:1991:DHU

[KK91a] A. Kavianpour and K. H. Kim. Diagnosabilities of hypercubes under the pessimistic one-step diagnosis strategy. IEEE Transactions on Computers, 40(2):232–237, February 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=73595.

Koren:1991:DCM

[KK91b] I. Koren and Z. Koren. Discrete and continuous models for the performance of reconfigurable multistage systems. IEEE Transactions on Computers, 40(9):1024–1033, September 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=83658.

Kozlowski:1991:CBA

[KK91c] W. E. Kozlowski and H. Krawczyk. A comparison-based approach to multicomputer system diagnosis in hybrid fault situations. IEEE Transactions on Computers, 40(11):1283–1287, November 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=102834.

Kim:1992:AOS

[KK92] Chonggun Kim and H. Kameda.

An algorithm for optimal static load balancing in distributed computer systems. IEEE Transactions on Computers, 41(3):381–384, March 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=127455.

Kanawati:1995:FFS

[KKA95] G. A. Kanawati, N. A. Kanawati, and J. A. Abraham. FER-RARI: a flexible software-based fault and error injection system. IEEE Transactions on Computers, 44(2):248–260, February 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=364536.

Kalamatianos:1999:ATB

[KKKM99] J. Kalamatianos, A. Khalafi, D. R. Kaeli, and W. Meleis. Analysis of temporal-based program behavior for improved instruction cache performance. IEEE Transactions on Computers, 48(2):168–175, February 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=752658.

Kilian:1990:OPA

[KKL90] J. Kilian, S. Kipnis, and C. E. Leiserson. The organization of permutation architectures with bused interconnections. IEEE Transactions on Computers, 39(11):1346–1358, November 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=61044.

Koren:1993:UNB

[KKS93] I. Koren, Z. Koren, and C. H. Stepper. A unified negativebinomial distribution for yield analysis of defect-tolerant circuits. IEEE Transactions on Computers, 42(6):724–734, June 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=277291.

Kuo:1992:CED

[KL92] Sy-Yen Kuo and Sheng-Chiech Liang. Concurrent error detection and correction in realtime systolic sorting arrays. IEEE Transactions on Computers, 41(12):1615–1620, December 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=214672.

Kinoshita:1997:RAE

[KL97] Eisuke Kinoshita and Ki-Ja Lee. A residue arithmetic extension for reliable scientific computation. IEEE Transactions on Computers, 46 (2):129–138, February 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=565587.

Kleeman:1990:JMM

[Kle90] L. Kleeman. The jitter model for metastability and its application to redundant synchronizers. IEEE Transactions on Computers, 39(7):930–942, July 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=55694.

Kwak:1999:EMC

[KLH^{+}99] H. Kwak, B. Lee, A. R. Hurson, Suk-Han Yoon, and Woo-Jong Hahn. Effects of multithreading on cache performance. IEEE Transactions on Computers, 48(2):176–184, February 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=752659.

Kim:1997:RPA

[KLL97] Jong Kim, Heejo Lee, and Sunggu Lee. Replicated process allocation for load distribution in fault-tolerant multicomputers. IEEE Transactions on Computers, 46(4):499–505, April 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=588067.

Kuo:1993:MVC

[KLPS93] T.-H. Kuo, H. C. Lin, R. C. Potter, and D. Schupe. Multiplevalued counter. IEEE Transactions on Computers, 42 (1):106–109, January 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=192219.

Karpovsky:1994:DSA

[KLV94] M. G. Karpovsky, L. B. Levitin, and F. S. Vainstein. Diagnosis by signature analysis of test responses. IEEE Transactions on Computers, 43 (2):141–152, February 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=262119.

Kornerup:1990:ARB

[KM90] Peter Kornerup and David W. Matula. An algorithm for redundant binary bit-pipelined rational arithmetic. IEEE Transactions on Computers, 39 (8):1106–1115, August 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=57048.

Kapur:1992:BSP

[KM92a] R. Kapur and M. R. Mercer. Bounding signal probabilities for testability measurement using conditional syndromes. IEEE Transactions on Computers, 41(12):1580–1588, December 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=214666.

Krishnamurti:1992:AAS

[KM92b] R. Krishnamurti and E. Ma.

An approximation algorithm for scheduling tasks on varying partition sizes in partitionable multiprocessor systems. IEEE Transactions on Computers, 41(12):1572–1579, December 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=214665.

Kwok:1994:TCT

[KM94] C. K. Kwok and B. Mukherjee. Transparent (cut-through) bridging of CSMA/CD networks: performance analysis and implementation. IEEE Transactions on Computers, 43 (2):247–253, February 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=262133.

Ke:1995:SDV

[KM95] W. Ke and P. R. Menon. Synthesis of delay-verifiable combinational circuits. IEEE Transactions on Computers, 44 (2):213–222, February 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=364533.

Kartik:1997:TAA

[KM97a] S. Kartik and C. Siva Ram Murthy. Task allocation algorithms for maximizing reliability of distributed computing systems. IEEE Transactions on Computers, 46(6):719–724, June 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=600888.

Kuo:1997:IRL

[KM97b] Tei-Wei Kuo and A. K. Mok. Incremental reconfiguration and load adjustment in adaptive real time systems. IEEE Transactions on Computers, 46 (12):1313–1324, December 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=641932.

Kazakos:1997:RMA

[KMD97] D. Kazakos, L. F. Merakos, and H. Delic. Random multiple access algorithms using a control mini-slot. IEEE Transactions on Computers, 46(4):473–476, April 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=588061.

Kim:1991:PTA

[KMM91] S. M. Kim, R. McNaughton, and R. McCloskey. A polynomial time algorithm for the local testability problem of deterministic finite automata. IEEE Transactions on Computers, 40(10):1087–1093, October 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=93741.

Karpovsky:1990:ORC

[KN90] M. G. Karpovsky and P. Nagvajara. Optimal robust compression of test responses. IEEE Transactions on Computers, 39(1):138–141, January 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=46290.

Knor:1996:NRM

[Kno96] M. Knor. A note on radially Moore digraphs. IEEE Transactions on Computers, 45(3):381–382, March 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=485578.

Kelliher:1992:EFA

[KOIH92] T. P. Kelliher, R. M. Owens, M. J. Irwin, and T.-T. Hwang. ELM — a fast addition algorithm discovered by a program. IEEE Transactions on Computers, 41(9):1181–1184, September 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=165399.

Kopec:1995:CNC

[Kop95] M. Kopec. Can nonlinear compactors be better than linear ones? IEEE Transactions on Computers, 44(11): 1275–1282, November 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=475123.

Kornerup:1994:DSC

[Kor94a] P. Kornerup. Digit-set conversions: generalizations and applications. IEEE Transactions on Computers, 43(5):622–629, May 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=280811.

Kornerup:1994:SLA

[Kor94b] P. Kornerup. A systolic, linear-array multiplier for a class of right-shift algorithms.

IEEE Transactions on Computers, 43(8):892–898, August 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=295851.

Kostopoulos:1991:ACB

[Kos91] D. K. Kostopoulos. An algorithm for the computation of binary logarithms. IEEE Transactions on Computers, 40 (11):1267–1270, November 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=102831.

Kosko:1994:FSU

[Kos94]  B. Kosko.  Fuzzy systems

as universal approximators. IEEE Transactions on Computers, 43(11):1329–1333, November 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=324566.

Kanoun:1998:GEI

[KP98] K. Kanoun and I. Pomeranz. Guest Editors’ introduction. IEEE Transactions on Computers, 47(1):1, January 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=656586.

Kanopoulos:1992:DSC

[KPB92] N. Kanopoulos, D. Pantzartzis, and F. R. Bartram. Design of self-checking circuits using DCVS logic: a case study. IEEE Transactions on Computers, 41(7):891–896, July 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=256445.

Kameda:1993:NMI

[KPI93] T. Kameda, S. Pilarski, and A. Ivanov. Notes on multiple input signature analysis. IEEE Transactions on Computers, 42(2):228–234, February 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=204795.

Kant:1990:SRD

[KR90a] K. Kant and A. Ravichandran. Synthesizing robust data structures — an introduction. IEEE Transactions on Computers, 39(2):161–173, February 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=45202.

Kundu:1990:SEC

[KR90b] S. Kundu and S. M. Reddy. On symmetric error correcting and all unidirectional error detecting codes. IEEE Transactions on Computers, 39(6):752–761, June 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=53596.

Kim:1993:RBA

[KR93] J. H. Kim and P. K. Rhee. The rule-based approach to reconfiguration of 2-D processor arrays. IEEE Transactions on Computers, 42(11):1403–1408, November 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=247844.

Kramer:1998:PWC

[Krä98] W. Krämer. A priori worst case error bounds for floating-point computations. IEEE Transactions on Computers, 47(7):750– 756, July 1998. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

709374. See [Tan92].

Kandemir:1999:ICL

[KRC99] M. Kandemir, J. Ramanujam, and A. Choudhary. Improving cache locality by a combination of loop and data transformations. IEEE Transactions on Computers, 48 (2):159–167, February 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=752657.

Karpovsky:1995:FDM

[KRM95] M. G. Karpovsky, T. D. Roziner, and C. Moraga. Fault detection in multiprocessor systems and array processors. IEEE Transactions on Computers, 44(3):383–393, March 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=372031.

Kugelmass:1990:UBE

[KS90] S. D. Kugelmass and K. Steighlitz. An upper bound on expected clock skew in synchronous systems. IEEE Transactions on Computers, 39(12): 1475–1477, December 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=61068.

Kakuda:1991:IAD

[KS91] Y. Kakuda and H. Saito. An integrated approach to design of protocol specifications using protocol validation and synthesis. IEEE Transactions on Computers, 40(4):459–467, April 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=88465.

Kandlur:1992:TRM

[KS92] D. D. Kandlur and K. G.

Shin. Traffic routing for multicomputer networks with virtual cut-through capability. IEEE Transactions on Computers, 41(10):1257–1270, October 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=166603.

Kim:1993:DFF

[KS93a] J. Kim and K. G. Shin.

Deadlock-free fault-tolerant routing in injured hypercubes. IEEE Transactions on Computers, 42(9):1078–1088, September 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=241597.

Krishna:1993:TFA

[KS93b] H. Krishna and J.-D. Sun.

On theory and fast algorithms for error correction in residue number system product codes. IEEE Transactions on Computers, 42(7):840–853, July 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=237724;
 https://ieeexplore.ieee.
org/xpl/tocresult.jsp?isnumber=

Konstantinidou:1994:CR

[KS94] S. Konstantinidou and L. Snyder. The Chaos router. IEEE Transactions on Computers, 43(12):1386–1397, December 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=338098.

Kim:1996:DAO

[KS96a] Hagbae Kim and K. G. Shin. Design and analysis of an optimal instruction-retry policy for TMR controller computers. IEEE Transactions on Computers, 45(11):1217–1225, November 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=544478.

Kim:1996:STM

[KS96b] Hagbae Kim and K. G. Shin. Sequencing tasks to minimize the effects of near-coincident faults in TMR controller computers. IEEE Transactions on Computers, 45(11):1331–1337, November 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=544492.

Kim:1996:ETA

[KS96c] Jong Kim and K. G. Shin. Execution time analysis of communicating tasks in distributed systems. IEEE Transactions on Computers, 45(5):572–579, May 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=509908.

Koc:1998:LCB

[KS98] C. K. Koc and B. Sunar. Lowcomplexity bit-parallel canonical and normal basis multipliers for a class of finite fields. IEEE Transactions on Computers, 47(3):353–356, March 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=660172.

Kundu:1996:SCC

[KSGT96] S. Kundu, E. S. Sogomonyan, M. Goessel, and S. Tarnick. Self-checking comparator with one periodic output. IEEE Transactions on Computers, 45(3):379–380, March 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=485577.

Kelsey:1993:EAS

[KSL93] T. P. Kelsey, K. K. Saluja, and S. Y. Lee. An efficient algorithm for sequential circuit test generation. IEEE Transactions on Computers, 42(11):1361–1371, November 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=247839.

Katsuyama:1991:STE

[KSNM91] K. Katsuyama, F. Sato, T. Nakakawaji, and T. Mizuno. Strategic testing environment with formal description techniques. IEEE Transactions on Computers, 40(4):514–525, April 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=88470.

Katcher:1995:FPS

[KSS95] D. I. Katcher, S. S. Sathaye, and J. K. Strosnider. Fixed priority scheduling with limited priority levels. IEEE Transactions on Computers, 44 (9):1140–1144, September 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=464392.

Kim:1991:ADI

[KT91a] M. Y. Kim and A. N. Tantawi. Asynchronous disk interleaving: approximating access delays. IEEE Transactions on Computers, 40(7):801–810, July 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=83618.

Koh:1991:IDT

[KT91b] J. Koh and D. Tcha. Information dissemination in trees with nonuniform edge transmission times. IEEE Transactions on Computers, 40 (10):1174–1177, October 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=93751.

Kumar:1991:SOF

[KT91c] V. K. P. Kumar and Y.-C.

Tsai. On synthesizing optimal family of linear systolic arrays for matrix multiplication. IEEE Transactions on Computers, 40(6):770–774, June 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=90254.

Knessl:1992:AEL

[KT92] C. Knessl and C. Tier. Asymptotic expansions for large closed queueing networks with multiple job classes. IEEE Transactions on Computers, 41(4):480–488, April 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=135560.

Kagaris:1996:RBP

[KT96] D. Kagaris and S. Tragoudas.

Retiming-based partial scan. IEEE Transactions on Computers, 45(1):74–87, January 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=481488.

Ke:1999:ACL

[KT99a] Jenn-Yang Ke and Jong-Chuang Tsay. An approach to checking link conflicts in the mapping of uniform dependence algorithms into lower dimensional processor arrays. IEEE Transactions on Computers, 48(7):732–737, July 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=780880.

Krishnan:1999:CMA

[KT99b] V. Krishnan and J. Torrellas. A chip-multiprocessor architecture with speculative multithreading. IEEE Transactions on Computers, 48(9):866–880, September 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=795218.

Kabuo:1994:ARS

[KTM^{+}94] H. Kabuo, T. Taniguchi, A. Miyoshi, H. Yamashita, M. Urano, H. Edamatsu, and S. Kuninobu. Accurate rounding scheme for the Newton–Raphson method using redundant binary representation. IEEE Transactions on Computers, 43(1):43–51, January 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=250608.

Kagaris:1996:UCR

[KTM96] D. Kagaris, S. Tragoudas, and A. Majumdar. On the use of counters for reproducing deterministic test sets. IEEE Transactions on Computers, 45 (12):1405–1419, December 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=545970.

Krishnamoorthy:1990:IDD

[KTS90] V. Krishnamoorthy, K. Thulasiraman, and M. N. S. Swamy. Incremental distance and diameter sequences of a graph: new measures of network performance. IEEE Transactions on Computers, 39(2):230–237, February 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=45208.

Kuhn:1998:CIS

[Kuh98]  M. G. Kuhn.  Cipher in-

struction search attack on the bus-encryption security microcontroller DS5002FP. IEEE Transactions on Computers, 47 (10):1153–1157, October 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=729797.

Kumar:1991:HQC

[Kum91] A. Kumar. Hierarchical quorum consensus: a new algorithm for managing replicated data. IEEE Transactions on Computers, 40(9):996–1004, September 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=83661.

Kundu:1992:BSS

[Kun92] S. Kundu. Basis sets for synthesis of switching functions. IEEE Transactions on Computers, 41(4):489–493, April 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=135561.

Kuszmaul:1990:FDR

[Kus90] B. C. Kuszmaul. Fast, deterministic routing, on hypercubes, using small buffers. IEEE Transactions on Computers, 39(11):1390–1393, November 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=61048.

Kang:1995:ABG

[KWF95] Jun-Woo Kang, Chin-Long Wey, and P. D. Fisher. Application of bipartite graphs for achieving race-free state assignments. IEEE Transactions on Computers, 44(8):1002–1011, August 1995. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=403716.

Koksal:1992:CDA

[KY92] F. Z. Koksal and M. D. Yucel. Comments on the decoding algorithms of DBEC– TBED Reed–Solomon codes. IEEE Transactions on Computers, 41(2):244–247, February 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=123402.

Kim:1997:CIR

[KYM97] M. K. Kim, H. Yoon, and S. R. Maeng. On the correctness of inside-out routing algorithm. IEEE Transactions on Computers, 46(7):820–823, July 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=599903.

Koren:1990:EEF

[KZ90] I. Koren and O. Zinaty. Evaluating elementary functions in a numerical coprocessor based on rational approximations. IEEE Transactions on Computers, 39(8):1030–1037, August 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=57042.

Lang:1998:CVA

[LA98] T. Lang and E. Antelo. CORDIC vectoring with arbitrary target value. IEEE Transactions on Computers, 47(7):736–749, July 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=709373.

LaForge:1999:CLS

[LaF99]  L. E. LaForge.  Configura-

tion of locally spared arrays in the presence of multiple fault types. IEEE Transactions on Computers, 48(4):398–416, April 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=762532.

Lamport:1997:HMC

[Lam97] L. Lamport. How to make a correct multiprocess program execute correctly on a multiprocessor. IEEE Transactions on Computers, 46(7):779–782, July 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=599898.

Latifi:1993:CAF

[Lat93] S. Latifi. Combinatorial analysis of the fault-diameter of the n-cube. IEEE Transactions on Computers, 42(1):27–33, January 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=192211.

Lalgudi:1996:ABH

[LAY96] H. Lalgudi, L. F. Akyildiz, and S. Yalamanchili. Augmented binary hypercube: a new architecture for processor management. IEEE Transactions on Computers, 45(8):980–984, August 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=536241.

Lee:1990:CDS

[LB90] Joo-Kang Lee and J. T. Butler. A characterization of t/s-diagnosability and sequential t-diagnosability in designs. IEEE Transactions on Computers, 39(10):1298–1304, October 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=59861.

Liebelt:1999:DES

[LB99] M. J. Liebelt and N. Burgess.

Detecting exitory stuck-at faults in semimodular asynchronous circuits. IEEE Transactions on Computers, 48(4):442–448, April 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=762538.

Lopez-Benitez:1992:DMR

[LBF92] N. Lopez-Benitez and J. A. B. Fortes. Detailed modeling and reliability analysis of faulttolerant processor arrays. IEEE Transactions on Computers, 41 (9):1193–1200, September 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=165402.

Lin:1990:DUS

[LC90] F.-C. Lin and K. Chen. On the design of a unidirectional systolic array for key enumeration. IEEE Transactions on Computers, 39(2):266–269, February 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=45213.

Lu:1992:NDA

[LC92] M. Lu and J.-S. Chiang. A novel division algorithm for the residue number system. IEEE Transactions on Computers, 41(8):1026–1032, August 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=156545;
                                         https://ieeexplore.ieee.
                                        org/xpl/tocresult.jsp?isnumber=
                                        4052.

Lee:1998:STP

[LC98a] Tsern-Huei Lee and Jin-Jye Chou. Some topological properties of bitonic sorters. IEEE Transactions on Computers, 47 (9):983–997, September 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=713317.

Lubaszewski:1998:RFS

[LC98b] M. Lubaszewski and B. Courtois. A reliable fail-safe system. IEEE Transactions on Computers, 47(2):236–241, February 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=663771.

Li:1995:CBM

[LCFH95] Chung-Chi Jim Li, Shyh-Kwei Chen, W. K. Fuchs, and W.-M. W. Hwu. Compilerbased multiple instruction retry. IEEE Transactions on Computers, 44(1):35–46, January 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=368011.

Lin:1995:NPK

[LCL95] C. H. Lin, C. C. Chang, and R. C. T. Lee. A new public-key cipher system based upon the Diophantine equations. IEEE Transactions on Computers, 44(1):13–19, January 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=368013.

See comment [BMP97a].

Lim:1993:OJC

[LCS93] A. Lim, S.-W. Cheng, and S. Sahni. Optimal joining of compacted cells. IEEE Transactions on Computers, 42(5):597–607, May 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=223679.

Lo:1997:FTA

[LCT97] N. W. Lo, B. S. Carlson, and D. L. Tao. Fault tolerant algorithms for broadcasting on the star graph network. IEEE Transactions on Computers, 46(12):1357–1362, December 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=641935.

Lee:1990:PAT

[LD90] I. Lee and S. B. Davidson. A performance analysis of timed synchronous communication primitives. IEEE Transactions on Computers, 39 (9):1117–1131, September 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=57054.

Leu:1994:CCP

[LD94] Yean-Shiang Leu and D. H. C.

Du. Cycle compensation protocol: a fair protocol for the unidirectional twin-bus architecture. IEEE Transactions on Computers, 43(1):1–12, January 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=250604.

Lee:1992:AAP

[Lee92] D.-L. Lee. Architecture of an array processor using a nonlinear skewing scheme. IEEE Transactions on Computers, 41(4):499– 505, April 1992. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
135554.

Leung:1997:UAR

[Leu97a]  K. K. Leung.  Update al-

gorithm for replicated signaling databases in wireless and advanced intelligent networks. IEEE Transactions on Computers, 46(3):362–367, March 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=580431.

Leung:1997:PAE

[Leu97b]  Yiu-Wing Leung.  Processor

assignment and execution sequence for multiversion software.

IEEE Transactions on Computers, 46(12):1371–1377, December 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=641937.

Lewis:1990:AAS

[Lew90] D. M. Lewis. An architecture for addition and subtraction of long word length numbers in the logarithmic number system. IEEE Transactions on Computers, 39(11):1325–1336, November 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=61042.

Lewis:1994:IMF

[Lew94] D. M. Lewis. Interleaved memory function interpolators with application to an accurate LNS arithmetic unit. IEEE Transactions on Computers, 43(8):974–982, August 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=295859.

Lo:1996:PAT

[LF96] Jien-Chung Lo and E. Fujiwara. Probability to achieve TSC goal. IEEE Transactions on Computers, 45(4):450–460, April 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=494102.

Lombardi:1992:DLM

[LFH92] F. Lombardi, C. Feng, and W.- K. Huang. Detection and location of multiple faults in baseline interconnection networks. IEEE Transactions on Computers, 41(10):1340–1344, October 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=166613.

Li:1992:EPA

[LG92] M. Li and N. D. Georganas.

Exact parametric analysis of stochastic Petri nets. IEEE Transactions on Computers, 41 (9):1176–1180, September 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=165403.

Lempel:1995:ZAM

[LG95] M. Lempel and S. K. Gupta.

Zero-aliasing for modeled faults. IEEE Transactions on Computers, 44(11):1283–1295, November 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=475124.

Laih:1997:CDE

[LG97] C. S. Laih and M. J. Gau. Cryptanalysis of a Diophantine equation oriented public key cryptosystem. IEEE Transactions on Computers, 46(4):511–512, April 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=588074.

Lombardi:1990:FDD

[LH90] F. Lombardi and W. Huang. Fault detection and design complexity in C-testable VLSI arrays. IEEE Transactions on Computers, 39(12):1477–1481, December 1990. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
61070.

Linder:1991:AFT

[LH91] D. H. Linder and J. C. Harden. An adaptive and fault tolerant wormhole routing strategy for kary n-cubes. IEEE Transactions on Computers, 40(1):2–12, January 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=67315.

Lee:1992:FTC

[LH92] T. C. Lee and J. P. Hayes. A fault-tolerant communication scheme for hypercube computers. IEEE Transactions on Computers, 41(10):1242–1256, October 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=166602.

Linder:1996:PLS

[LH96] D. H. Linder and J. H. Harden.

Phased logic: supporting the synchronous design paradigm with delay-insensitive circuitry. IEEE Transactions on Computers, 45(9):1031–1044, September 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=537126.

Libeskindg-Hadas:1998:TLB

[LH98] R. Libeskindg-Hadas. A tight lower bound on the number of channels required for deadlock-free wormhole routing. IEEE Transactions on Computers, 47(10):1158–1160, October 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=729799.

Lapointe:1993:SDP

[LHF93] M. Lapointe, H. T. Huynh, and P. Fortier. Systematic design of pipelined recursive filters. IEEE Transactions on Computers, 42(4):413–426, April 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=214688.

Liu:1999:SAI

[LHL99] Chih-Wei Liu, Kuo-Tai Huang, and Chung-Chin Lu. A systolic array implementation of the Feng–Rao algorithm. IEEE Transactions on Computers, 48(7):690–706, July 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=780877.

Latifi:1994:CCM

[LHNP94] S. Latifi, M. Hegde, and M. Naraghi-Pour. Conditional connectivity measures for large multiprocessor systems. IEEE Transactions on Computers, 43(2):218–222, February 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=262126.

Lee:1998:ACR

[LHS^{+}98] Chang-Gun Lee, Hoosun Hahn, Yang-Min Seo, Sang Lyul Min, Rhan Ha, Seongsoo Hong, Chang Yun Park, Minsuk Lee, and Chong Sang Kim. Analysis of cache-related preemption delay in fixed-priority preemptive scheduling. IEEE Transactions on Computers, 47(6):700–713, June 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=689649.

Li:1997:SBP

[Li97] Keqin Li. Stochastic bounds for parallel program execution times with processor constraints. IEEE Transactions on Computers, 46(5):630–636, May 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=589243.

Lim:1992:SPM

[Lim92] Y. C. Lim. Single-precision multiplier with reduced circuit complexity for signal processing applications. IEEE Transactions on Computers, 41 (10):1333–1336, October 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=166611.

Ling:1990:AIM

[Lin90] H. Ling. An approach to implementing multiplication with small tables. IEEE Transactions on Computers, 39(5):717–718, May 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=53588.

Lin:1991:CPA

[Lin91] H.-A. Lin. Constructing protocols with alternative functions. IEEE Transactions on Computers, 40(4):376–386, April 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=88459.

Lin:1993:CWC

[Lin93a] Mao-Chao Lin. Constant weight codes for correcting symmetric errors and detecting unidirectional errors. IEEE Transactions on Computers, 42(11): 1294–1302, November 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=247835.

Lin:1993:MGV

[Lin93b] W. Lin. Manipulating general vectors on synchronous binary n-cube. IEEE Transactions on Computers, 42(7):863–871, July 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=237726.

Lioy:1993:EFP

[Lio93]  A. Lioy.  On the equiva-

lence of fanout-point faults. IEEE Transactions on Computers, 42(3):268–271, March 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=210169.

Lee:1994:PTT

[LIO94] Gueesang Lee, M. J. Irwin, and R. M. Owens. Polynomial time testability of circuits generated by input decomposition. IEEE Transactions on Computers, 43(2):201–210, February 1994. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=262124.

Lin:1990:NLT

[LK90] W.-M. Lin and V. K. P. Kumar.

A note on the linear transformation method for systolic array design. IEEE Transactions on Computers, 39(3):393–399, March 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=48870.

Lee:1991:OSS

[LK91a] Y.-H. Lee and C. M. Krishna.

Optimal scheduling of signature analysis for VLSI testing. IEEE Transactions on Computers, 40(3):336–341, March 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=76412.

Lewis:1991:OSA

[LK91b] P. S. Lewis and S.-Y. Kung.

An optimal systolic array for the algebraic path problem. IEEE Transactions on Computers, 40(1):100–105, January 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=67325.

Lee:1993:PPP

[LK93] E. K. Lee and R. H. Katz. The performance of parity placements in disk arrays. IEEE Transactions on Computers, 42(6):651–664, June 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=277289.

Liao:1995:VRE

[LK95] Wei-Kuo Liao and Chung-Ta King. Valved routing: efficient flow control for adaptive nonminimal routing in interconnection networks. IEEE Transactions on Computers, 44 (10):1181–1193, October 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=467693.

Leu:1996:FTT

[LK96] Yuh-Rong Leu and Sy-Yen Kuo. A fault-tolerant tree communication scheme for hypercube systems. IEEE Transactions on Computers, 45(6):641–650, June 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=506421.

Li:1998:LBP

[LK98] Jie Li and H. Kameda. Load balancing problems for multiclass jobs in distributed/parallel computer systems. IEEE Transactions on Computers, 47(3):322–332, March 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=660168.

Leu:1999:DFT

[LK99] Yuh-Rong Leu and Sy-Yen Kuo. Distributed fault-tolerant ring embedding and reconfiguration in hypercubes. IEEE Transactions on Computers, 48(1):81–88, January 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=743414.

Leveugle:1994:HDT

[LKK^{+}94] R. Leveugle, Z. Koren, I. Koren, G. Saucier, and N. Wehn. The Hyeti defect tolerant microprocessor: a practical experiment and its cost-effectiveness analysis. IEEE Transactions on Computers, 43(12):1398–1406, December 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=338099.

Lin:1997:SHF

[LKL97] Kuan-Jen Lin, Chi-Wen Kuo, and Chen-Shang Lin. Synthesis of hazard-free asynchronous circuits based on characteristic graph. IEEE Transactions on Computers, 46(11):

1246–1263, November 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=644299.

Linderman:1998:DHP

[LKL98] R. W. Linderman, R. L. R. Kohler, and M. H. Linderman. A dependable high performance wafer scale architecture for embedded signal processing. IEEE Transactions on Computers, 47(1):125–128, January 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=656096.

Lu:1997:FTI

[LKW97] Shyue-Kung Lu, Sy-Yen Kuo, and Cheng-Wen Wu. Faulttolerant interleaved memory systems with two-level redundancy. IEEE Transactions on Computers, 46(9):1028–1034, September 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=620483.

Lee:1990:HHT

[LL90] D. L. Lee and F. H. Lochovsky. HYTREM — a hybrid text-retrieval machine for large databases. IEEE Transactions on Computers, 39(1):111–123, January 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=46285.

Lee:1991:WDE

[LL91a] D. Y. Lee and J. Y. Lee. A well-defined Estelle specification for the automatic test generation. IEEE Transactions on Computers, 40(4):526–542, April 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=88471.

Lin:1991:NNB

[LL91b] C.-T. Lin and C. S. G. Lee.

Neural-network-based fuzzy logic control and decision system. IEEE Transactions on Computers, 40(12):1320–1336, December 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=106218.

Lee:1992:CFR

[LL92a] J.-A. Lee and T. Lang. Constantfactor redundant CORDIC for angle calculation and rotation. IEEE Transactions on Computers, 41(8):1016–1025, August 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=156544.

Liaw:1992:ORG

[LL92b] Heh-Tyan Liaw and Chen-Shang Lin. On the OBDDrepresentation of general Boolean functions. IEEE Transactions on Computers, 41(6):661–664, June 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=144618.

Lee:1994:NSR

[LL94] S. Lee and M. Lu. New selfrouting permutation networks. IEEE Transactions on Computers, 43(11):1319–1323, November 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=324564.

Lai:1996:MFD

[LL96a] K. Lai and P. K. Lala. Multiple fault detection in fan-out free circuits using minimal single fault test set. IEEE Transactions on Computers, 45(6):763–765, June 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=506433.

Low:1996:NCE

[LL96b] C. P. Low and H. W. Leong. A new class of efficient algorithms for reconfiguration of memory arrays. IEEE Transactions on Computers, 45(5):614–618, May 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=509914.

Lim:1997:OPH

[LL97] Kyungshik Lim and Yann-Hang Lee. Optimal partitioning of heterogeneous traffic sources in mobile communications networks. IEEE Transactions on Computers, 46(3):312–325, March 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=580427.

Liu:1997:TTB

[LLJ97] Guoping Liu, K. Y. Lee, and H. F. Jordan. TDM and TWDM de Bruijn networks and ShuffleNets for optical communications. IEEE Transactions on Computers, 46(6):695–701, June 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=600827.

Lee:1992:OTA

[LLK92] Cheol-Hoon Lee, Dongmyun Lee, and Myunghwan Kim. Optimal task assignment in linear array networks. IEEE Transactions on Computers, 41(7):877–880, July 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=256461.

Lam:1991:SFU

[LLS91] H. Lam, C. Lee, and S. Y. W.

Su. A special function unit for database operations (SFU– DB): design and performance evaluation. IEEE Transactions on Computers, 40(3):263–275, March 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=76403.

LaPaugh:1992:HST

[LLS92] A. S. LaPaugh, R. J. Lipton, and J. S. Sandberg. How to store a triangular matrix. IEEE Transactions on Computers, 41(7):896–899, July 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=256446.

Lang:1992:HRS

[LM92a] T. Lang and P. Montuschi. Higher radix square root with prescaling. IEEE Transactions on Computers, 41(8):996–1009, August 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=156542.

Leighton:1992:FAR

[LM92b] F. T. Leighton and B. M. Maggs. Fast algorithms for routing around faults in multibutterflies and randomly-wired splitter networks. IEEE Transactions on Computers, 41(5):578–587, May 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=142684.

Lui:1992:BMT

[LM92c] P. K. Lui and J. C. Muzio.

Boolean matrix transforms for the minimization of modulo-2 canonical expansions. IEEE Transactions on Computers, 41(3):342–347, March 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=127446.

Lang:1999:VHR

[LM99a] T. Lang and P. Montuschi. Very high radix square root with prescaling and rounding and a combined division/square root unit. IEEE Transactions on Computers, 48(8):827–841, August 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=795124.

Luk:1999:ACI

[LM99b] Chi-Keung Luk and T. C.

Mowry. Automatic compilerinserted prefetching for pointerbased applications. IEEE Transactions on Computers, 48 (2):134–141, February 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=752654.

Laranjeira:1993:NNP

[LMJ93] L. A. Laranjeira, M. Malek, and R. Jenevein. Nest: a nested-predicate scheme for fault tolerance. IEEE Transactions on Computers, 42(11): 1303–1324, November 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=247836.

Lui:1995:BMR

[LMT95] J. C. S. Lui, R. R. Muntz, and D. Towsley. Bounding the mean response time of the minimum expected delay routing policy: an algorithmic approach. IEEE Transactions on Computers, 44(12):1371–1382, December 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=477243.

Lefevre:1998:TCR

[LMT98] Vincent Lefèvre, Jean-Michel Muller, and Arnaud Tisserand. Toward correctly rounded transcendentals. IEEE Transactions on Computers, 47(11): 1235–1243, November 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=736435.

Lagman:1994:AEF

[LNS94] A. Lagman, W. A. Najjar, and P. K. Srimani. An analysis of edge fault tolerance in recursively decomposable regular networks. IEEE Transactions on Computers, 43(4):470–475, April 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=278484.

Lin:1994:CTI

[LO94a] Ming-Bo Lin and A. Yavuz Oruc.

Constant time inner product and matrix computations on permutation network processors. IEEE Transactions on Computers, 43(12):1429–1434, December 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=338104.

Lo:1994:RFP

[Lo94b]  Jien-Chung Lo.  Reliable

floating-point arithmetic algorithms for error-coded operands. IEEE Transactions on Computers, 43(4):400–412, April 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=278479.

Lo:1996:HOE

[Lo96] Jien-Chung Lo. A hyper optimal encoding scheme for selfchecking circuits. IEEE Transactions on Computers, 45(9): 1022–1030, September 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=537125.

Lin:1997:DOA

[LO97a] Ming-Bo Lin and A. Y. Oruc. The design of an optoelectronic arithmetic processor based on permutation networks. IEEE Transactions on Computers, 46 (2):142–153, February 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=565589.

Lo:1997:FBA

[Lo97b] Jien-Chung Lo. A fast binary adder with conditional carry generation. IEEE Transactions on Computers, 46 (2):248–253, February 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=565614.

See correction [Lo98].

Lo:1998:CFB

[Lo98]  J. C. Lo.  Correction to

“A Fast Binary Adder with Conditional Carry Generation”. IEEE Transactions on Computers, 47(12):1425, December 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=737689.

See [Lo97b].

Li:1990:OVD

[LP90] H. F. Li and D. K. Probst. Optimal VLSI dictionary machines without compress instructions. IEEE Transactions on Computers, 39(5):676–693, May 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=53580.

Liu:1991:DPA

[LP91] Y. C. Liu and H. G. Perros. A decomposition procedure for the analysis of a closed fork/join queueing system. IEEE Transactions on Computers, 40(3):365–370, March 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=76416.

Lau:1998:TAB

[LP98] F. C. M. Lau and W. C.

Poon. Throughput analysis of B-networks. IEEE Transactions on Computers, 47(4):482–485, April 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=675717.

Luccio:1999:NBF

[LP99] F. Luccio and L. Pagli. On a new Boolean function with applications. IEEE Transactions on Computers, 48(3):296–310, March 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=754996.

Lin:1991:AMN

[LPP91] W.-M. Lin, V. K. Prasanna, and K. W. Przytula. Algorithmic mapping of neural network models onto parallel SIMD machines. IEEE Transactions on Computers, 40(12):1390–1401, December 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=106224.

Lai:1996:FVU

[LPV96] Yung-Te Lai, M. Pedram, and S. B. K. Vrudhula. Formal verification using edgevalued binary decision diagrams. IEEE Transactions on Computers, 45(2):247–255, February 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=485378.

Lenders:1997:MVA

[LR97] P. Lenders and S. Rajopadhye. Multirate VLSI arrays and their synthesis. IEEE Transactions on Computers, 46(5):515–529, May 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=589218.

Lakshmanan:1991:CEI

[LRG91] K. B. Lakshmanan, B. Ravikumar, and K. Ganesan. Coping with erroneous information while sorting. IEEE Transactions on Computers, 40(9):

1081–1084, September 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=83656.

Leveugle:1990:OSC

[LS90a] R. Leveugle and G. Saucier.

Optimized synthesis of concurrently checked controllers. IEEE Transactions on Computers, 39(4):419–425, April 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=54835.

Lin:1990:LCA

[LS90b] H. X. Lin and H. J. Sips.

On-line CORDIC algorithms. IEEE Transactions on Computers, 39(8):1038–1052, August 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=57043.

Lin:1990:LFM

[LS90c] T.-H. Lin and K. G. Shin.

Location of a faulty module in a computing system. IEEE Transactions on Computers, 39(2):182–194, February 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=45204.

Lai:1991:PPH

[LS91a] T.-H. Lai and A. P. Sprague. Placement of the processors of a hypercube. IEEE Transactions on Computers, 40(6):714–722, June 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=90250.

Liu:1991:RAC

[LS91b] J.-C. Liu and K. G. Shin. A RAM architecture for concurrent access and on chip testing. IEEE Transactions on Computers, 40(10):1153–1159, October 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=93748.

Lynch:1992:STC

[LS92] T. Lynch and E. E. Swartzlander, Jr. A spanning tree carry lookahead adder. IEEE Transactions on Computers, 41(8):931–939, August 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=156535.

Lee:1993:OEP

[LS93] S. Lee and K. G. Shin. Optimal and efficient probabilistic distributed diagnosis schemes. IEEE Transactions on Computers, 42(7):882–886, July 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=237729.

Lin:1994:ORP

[LS94a] Tein-Hsiang Lin and K. G.

Shin. An optimal retry policy based on fault classification. IEEE Transactions on Computers, 43(9):1014–1025, September 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=312112.

Lun:1994:PDR

[LS94b] D. P.-K. Lun and Wan-Chi Siu. A pipeline design for the realization of the prime factor algorithm using the extended diagonal structure. IEEE Transactions on Computers, 43(10):1232–1237, October 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=324549.

Liu:1995:EIT

[LS95] Jyh-Charn Liu and K. G.

Shin. Efficient implementation techniques for gracefully degradable multiprocessor systems. IEEE Transactions on Computers, 44(4):503–517, April 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=376166.

Latifi:1996:TNC

[LS96a] S. Latifi and P. K. Srimani. Transposition networks as a class of fault-tolerant robust networks. IEEE Transactions on Computers, 45(2):230–238, February 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=485375.

Lee:1996:MSG

[LS96b] Chung-Len Lee and Meng-Lieh Sheu. A multiple-sequence generator based on inverted nonlinear autonomous machines. IEEE Transactions on Computers, 45(9):1079–1083, September 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=537133.

Liu:1996:RTP

[LS96c] Zhen Liu and Ting-Yi Sung. Routing and transmitting problems in de Bruijn networks. IEEE Transactions on Computers, 45(9):1056–1062, September 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=537129.

Lu:1997:RCM

[LS97] Qi Lu and M. Satyanarayanan. Resource conservation in a mobile transaction system. IEEE Transactions on Computers, 46(3):299–311, March 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=580426.

Lin:1998:DAO

[LS98] T.-H. Lin and K. G. Shin. Damage assessment for optimal rollback recovery. IEEE Transactions on Computers, 47(5):603– 613, May 1998. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
677255.

Lo:1990:DCT

[LT90] J.-C. Lo and S. Thanawastien.

On the design of combinational totally self-checking 1out-of-3 code checkers. IEEE Transactions on Computers, 39(3):387–393, March 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=48869.

Lin:1993:IMC

[LT93] H.-A. Lin and C.-L. Tarng. An improved method for constructing multiphase communications protocols. IEEE Transactions on Computers, 42(1):15–26, January 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=192210.

Lopez:1995:CCC

[LT95] M. A. Lopez and R. Thurimella. On computing connected components of line segments. IEEE Transactions on Computers, 44 (4):597–601, April 1995. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL https://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        376174.

Lo:1993:BCP

[LTR93] J.-C. Lo, S. Thanawastien, and T. R. N. Rao. Berger check prediction for array multipliers and array dividers. IEEE Transactions on Computers, 42(7):892–896, July 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=237731.

See correction [LTR96].

Lo:1996:CBC

[LTR96] Jien-Chung Lo, S. Thanawastien, and T. R. N. Rao. Correction to “Berger Check Prediction for Array Multipliers and Array Dividers”. IEEE Transactions on Computers, 45(3):383, March 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=485579.

See [LTR93].

Lu:1997:SMK

[Lu97] Chung-Chin Lu. A search of minimal key functions for normal basis multipliers. IEEE Transactions on Computers, 46(5):588– 592, May 1997. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
589230.

Lafferty:1999:OBD

[LV99] J. Lafferty and A. Vardy.

Ordered binary decision diagrams and minimal trellises. IEEE Transactions on Computers, 48(9):971–986, September 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=795225.

Llosa:1998:MSR

[LVAG98] J. Llosa, M. Valero, E. Agyuade, and A. Gonzalez. Module scheduling with reduced register pressure. IEEE Transactions on Computers, 47(6):625–638, June 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=689643.

Lades:1993:DIO

[LVB^{+}93] M. Lades, J. C. Vorbruggen, J. Buhmann, J. Lange, C. von der Malsburg, R. P. Wurtz, and W. Konen. Distortion invariant object recognition in the dynamic link architecture. IEEE Transactions on Computers, 42(3):300–311, March 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=210173.

Lai:1991:HNS

[LW91] F.-S. Lai and C.-F. E. Wu. A hybrid number system processor with geometric and complex arithmetic capabilities. IEEE Transactions on Computers, 40(8):952–962, August 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=83639.

Lakshman:1994:DCR

[LW94] T. V. Lakshman and V. K. Wei. Distributed computing on regular networks with anonymous nodes. IEEE Transactions on Computers, 43 (2):211–218, February 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=262125.

Li:1995:CDSa

[LW95a] Yao Li and C. M. Woodside. Complete decomposition of stochastic Petri nets representing generalized service networks. IEEE Transactions on Computers, 44(4):577–592, April 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=376171.

Li:1995:CDSb

[LW95b] Yao Li and C. M. Woodside. Complete decomposition of stochastic Petri nets representing generalized service networks. IEEE Transactions on Computers, 44(8):1031–1046, August 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=403719.

Lee:1990:BNM

[LY90] K. Y. Lee and H. Yoon. The B-network: a multistage interconnection network with backward links. IEEE Transactions on Computers, 39(7):966–969, July 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=55700.

Lee:1991:IST

[LY91] K. Y. Lee and H. Yoon. Indirect star-type networks for large multiprocessor systems. IEEE Transactions on Computers, 40(11):1277–1282, November 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=102833.

Lee:1994:TFS

[LY94] D. Lee and M. Yannakakis.

Testing finite-state machines: state identification and verification. IEEE Transactions on Computers, 43(3):306–320, March 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=272431.

Laih:1996:ADG

[LY96] Chi-Sung Laih and Ching-Nung Yang. On the analysis and design of group theoretical t-syEC/AUED codes. IEEE Transactions on Computers, 45(1):103–108, January 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=481491.

Lee:1995:AUV

[LYS95] Jae Young Lee, Hee Yong Youn, and A. D. Singh. Adaptive unanimous voting (UV) scheme for distributed selfdiagnosis. IEEE Transactions on Computers, 44(5):730–735, May 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=381964.

Lee:1994:BDP

[LYW94] D. T. Lee, C. D. Yang, and C. K. Wong. On bends and distances of paths among obstacles in two-layer interconnection model. IEEE Transactions on Computers, 43(6):711–724, June 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=286304.

Lopez:1994:UAD

[LZ94] J. Lopez and E. L. Zapata.

Unified architecture for divide and conquer based tridiagonal system solvers. IEEE Transactions on Computers, 43 (12):1413–1425, December 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=338101.

Lu:1990:ASS

[LZM90] M. Lu, D. Zhang, and T. Murata. Analysis of self-stabilizing clock synchronization by means of stochastic Petri nets. IEEE Transactions on Computers, 39 (5):597–604, May 1990. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
53573.

Ma:1998:SAM

[Ma98] Yutai Ma. A simplified architecture for modulo (2^{n} + 1) multiplication. IEEE Transactions on Computers, 47(3):333–337, March 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=660169.

Mahmud:1994:CST

[Mah94a] S. M. Mahmud. Comments on “Synthetic traces for tracedriven simulation of cache memories”. IEEE Transactions on Computers, 43(1):125–126, January 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=250620.

See [TWS92].

Mahmud:1994:PAM

[Mah94b] S. M. Mahmud. Performance analysis of multilevel bus networks for hierarchical multiprocessors. IEEE Transactions on Computers, 43(7):789–805, July 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=293258.

Moundanos:1998:ATV

[MAH98] D. Moundanos, J. A. Abraham, and Y. V. Hoskote. Abstraction techniques for validation coverage analysis and test generation. IEEE Transactions on Computers, 47(1):2–14, January 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=656068.

Majumdar:1996:EOW

[Maj96] A. Majumdar. On evaluating and optimizing weights for weighted random pattern testing. IEEE Transactions on Computers, 45(8):904–916, August 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=536233.

Mandelbaum:1990:SMD

[Man90] D. M. Mandelbaum. A systematic method for division with high average bit skipping. IEEE Transactions on Computers, 39(1):127–130, January 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=46287.

Mandelbaum:1993:SRS

[Man93] D. M. Mandelbaum. Some results on a SRT type division scheme. IEEE Transactions on Computers, 42(1):102–106, January 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=192218.

Mandelbaum:1995:DUL

[Man95] D. M. Mandelbaum. Division using a logarithmic-exponential transform to form a short reciprocal. IEEE Transactions on Computers, 44(11): 1326–1330, November 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=475129.

Matic:1990:EHA

[Mat90] S. Matic. Emulation of hypercube architecture on nearestneighbor mesh-connected processing elements. IEEE Transactions on Computers, 39(5):698– 700, May 1990. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
53582.

Mazumder:1993:DFT

[Maz93] P. Mazumder. Design of a faulttolerant three-dimensional dynamic random-access memory with on-chip error-correcting circuit. IEEE Transactions on Computers, 42(12):1453–1468, December 1993. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
260635.

Mukherjee:1993:ASI

[MB93] B. Mukherjee and S. Banerjee. Alternative strategies for improving the fairness in and an analytical model of the DQDB network. IEEE Transactions on Computers, 42 (2):151–167, February 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=204785.

Menezes:1995:NBR

[MB95] B. L. Menezes and U. Bakhru. New bounds on the reliability of augmented shuffle-exchange networks. IEEE Transactions on Computers, 44(1):123–129, January 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=368004.

Makki:1991:ATP

[MBGT91] R. Z. Makki, S. Bou-Ghazale, and C. Tianshang. Automatic test pattern generation with branch testing. IEEE Transactions on Computers, 40(6):785–791, June 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=90257.

Malik:1993:TLM

[MBNSV93] A. A. Malik, R. K. Brayton, A. R. Newton, and A. L. Sangiovanni-Vincentelli. Twolevel minimization of multivalued functions with large offsets. IEEE Transactions on Computers, 42(11):1325–1342, November 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=247837.

Malluhi:1995:EMA

[MBR95] Q. M. Malluhi, M. A. Bayoumi, and T. R. N. Rao. Efficient mapping of ANNs on hypercube massively parallel machines. IEEE Transactions on Computers, 44(6):769–779, June 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=391184.

See correction [MBR96].

Malluhi:1996:CEM

[MBR96] Q. M. Malluhi, M. A. Bayoumi, and T. R. N. Rao. Correction to “Efficient Mapping of ANNs on Hypercube Massively Parallel Machines”. IEEE Transactions on Computers, 45(4):511, April 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=494113.

See [MBR95].

Montuschi:1992:DRD

[MC92] P. Montuschi and L. Ciminiera. Design of a radix 4 division unit with simple selection table. IEEE Transactions on Computers, 41(12):1606–1611, December 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=214670.

Montuschi:1993:RIT

[MC93] P. Montuschi and L. Ciminiera. Reducing iteration time when result digit is zero for radix 2 SRT division and square root with redundant remainders. IEEE Transactions on Computers, 42(2):239–246, February 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/

stamp.jsp?tp=&arnumber=204797.

See remark [MC95].

Montuschi:1994:RDS

[MC94] P. Montuschi and L. Ciminiera.

Over-redundant digit sets and the design of digit-by-digit division units. IEEE Transactions on Computers, 43(3):269–277, March 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=272428.

Montuschi:1995:RRI

[MC95] P. Montuschi and L. Ciminiera.

A remark on “Reducing iteration time when result digit is zero for radix-2 SRT division and square root with redundant remainders”. IEEE Transactions on Computers, 44(1):144–146, January 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=368000.

See [MC93].

McAuley:1992:FSA

[McA92] A. J. McAuley. Four state asynchronous architectures. IEEE Transactions on Computers, 41 (2):129–142, February 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=123391.

McCrackin:1991:EID

[McC91]  D. C. McCrackin.  Elim-

inating interlocks in deeply pipelined processors by delay enforced multistreaming. IEEE Transactions on Computers, 40(10):1125–1132, October 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=93745.

McCrackin:1995:PDE

[McC95] D. C. McCrackin. Practical delay enforced multistream (DE-MUS) control of deeply pipelined processors. IEEE Transactions on Computers, 44(3):458–462, March 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=372038.

Mohapatra:1995:DEM

[MD95] P. Mohapatra and C. R. Das. On dependability evaluation of mesh-connected processors. IEEE Transactions on Computers, 44(9):1073–1084, September 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=464386.

Mohapatra:1994:PAC

[MDyF94] P. Mohapatra, C. R. Das, and Tse yun Feng. Performance analysis of cluster-based multiprocessors. IEEE Transactions on Computers, 43(1):109–114, January 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=250615.

Megson:1992:FFA

[Meg92] G. M. Megson. A fast Faddeev array. IEEE Transactions on Computers, 41(12): 1594–1600, December 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=214668.

Meister:1991:PSI

[Mei91] B. W. Meister. A performance study of the ISO transport protocol. IEEE Transactions on Computers, 40(3):253–262, March 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=76402.

Melhem:1992:BLR

[Mel92] R. G. Melhem. Bi-level reconfigurations of fault tolerant arrays. IEEE Transactions on Computers, 41(2):231–239, February 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=123400.

Merigot:1997:ANG

[Mer97] A. Merigot. Associative nets:

a graph-based parallel computing model. IEEE Transactions on Computers, 46(5):558–571, May 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=589222.

Metzner:1991:ERR

[Met91] J. J. Metzner. Efficient replicated remote file comparison. IEEE Transactions on Computers, 40(5):651–660, May 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=88489.

Mulder:1992:PAD

[MF92] H. Mulder and M. J. Flynn. Processor architecture and data buffering. IEEE Transactions on Computers, 41(10):1211–1222, October 1992. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
166600.

Milutinovic:1991:PDT

[MFH91] V. M. Milutinovic, D. A. Fura, and W. A. Helbig. Pipeline design tradeoffs in a 32-bit gallium arsenide microprocessor. IEEE Transactions on Computers, 40(11):1214–1224, November 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=102825.

Melhem:1990:ERG

[MH90] R. G. Melhem and G.-Y. Hwang.

Embedding rectangular grids into square grids with dilation two. IEEE Transactions on Computers, 39(12): 1446–1455, December 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=61064.

Morii:1993:NAP

[MI93] M. Morii and K. Iwasaki. A note on aliasing probability for multiple input signature analyzer. IEEE Transactions on Computers, 42(9):1152, September 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=241605.

Miel:1993:CGF

[Mie93] G. Miel. Constant geometry fast Fourier transforms on array processors. IEEE Transactions on Computers, 42(3):371–375, March 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=210180.

Milne:1991:FDV

[Mil91] G. J. Milne. The formal description and verification of hardware timing. IEEE Transactions on Computers, 40(7):811–826, July 1991. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=83619.

Mitchell:1992:AMI

[Mit92]  C. J. Mitchell.  Authenti-

cating multicast Internet electronic mail messages using a bidirectional MAC is insecure. IEEE Transactions on Computers, 41(4):505–507, April 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=135563.

Menezes:1991:KMN

[MJ91] B. L. Menezes and R. Jenevein. The KYKLOS multicomputer network: interconnection strategies, properties, and applications. IEEE Transactions on Computers, 40(6):693–705, June 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=90248.

Mou:1992:OSA

[MJ92] Z.-J. Mou and F. Jutand. ‘overturned-stairs’ adder trees and multiplier design. IEEE Transactions on Computers, 41(8):940–948, August 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=156536.

Montgomery:1990:SRE

[MK90] B. L. Montgomery and B. V.

K. V. Kumar. Systematic random error correcting and all unidirectional error detecting codes. IEEE Transactions on Computers, 39(6):836–840, June 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=53606.

Madabhushi:1993:NOG

[MLD93] S. V. R. Madabhushi, S. Lakshmivarahan, and S. K. Dhall. A note on orthogonal graphs. IEEE Transactions on Computers, 42(5):624–630, May 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=223683.

Moore:1998:MCP

[MLK98] J. S. Moore, T. W. Lynch, and M. Kaufmann. A mechanically checked proof of the AMD5_{K} 86^{T M} floatingpoint division program. IEEE Transactions on Computers, 47 (9):913–926, September 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=713311.

Martel:1991:OPC

[MM91a] C. U. Martel and W.-H. L. M.

Moh. Optimal prioritized conflict resolution on a multiple access channel. IEEE Transactions on Computers, 40 (10):1102–1108, October 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=93743.

Murphy:1991:DSE

[MM91b] O. J. Murphy and R. L. McCraw. Designing storage efficient decision trees. IEEE Transactions on Computers, 40(3):315–320, March 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=76408.

Mazzaferri:1995:CNC

[MM95a] R. Mazzaferri and T. M. Murray. The Connection Network class for Fault Tolerant Meshes. IEEE Transactions on Computers, 44(1):131–138, January 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=368002.

McQuillan:1995:SMD

[MM95b] S. E. McQuillan and J. V. Mc-Canny. A systematic methodology for the design of high performance recursive digital filters. IEEE Transactions on Computers, 44(8):971–982, August 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=403713.

Marom:1997:CNR

[MM97a] D. M. Marom and D. Mendlovic.

Comment on “A new routing algorithm for a class of rearrangeable networks”. IEEE Transactions on Computers, 46(6):734, June 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=600891.

See [FS94a].

Michelucci:1997:LA

[MM97b] D. Michelucci and J.-M. Moreau.

Lazy arithmetic. IEEE Transactions on Computers, 46 (9):961–975, September 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=620478.

Min:1991:ADC

[MMJ91] Y. Min, Y. K. Malaiya, and B. Jin. Analysis of detection capability of parallel signature analyzers. IEEE Transactions on Computers, 40(9): 1075–1081, September 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=83655.

Mazenc:1993:CFU

[MMM93] Christophe Mazenc, Xavier Merrheim, and Jean-Michel Muller. Computing functions cos^{−1} and sin^{−1} using CORDIC. IEEE Transactions on Computers, 42(1):118–122, January 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=192222.

Martel:1996:DPC

[MMM96] C. U. Martel, W. M. Moh, and Teng-Sheng Moh. Dynamic prioritized conflict resolution on multiple access broadcast networks. IEEE Transactions on Computers, 45(9): 1074–1079, September 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=537132.

Merchant:1996:ACM

[MMSS96] A. Merchant, B. Melamed, E. Schenfeld, and B. Sengupta. Analysis of a control mechanism for a variable speed processor. IEEE Transactions on Computers, 45(7):793–801, July 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=508318.

Masuda:1990:CML

[MNKF90] S. Masuda, K. Nakajima, T. Kashiwabara, and T. Fujisawa. Crossing minimization in linear embeddings of graphs. IEEE Transactions on Computers, 39(1):124–127, January 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=46286.

Michael:1999:CCA

[MNL99] M. M. Michael, A. K. Nanda, and Beng-Hong Lim. Coherence controller architectures for scalable shared-memory multiprocessors. IEEE Transactions on Computers, 48 (2):245–255, February 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=752666.

Mourad:1991:CTM

[MOM91] A. Mourad, B. Ozden, and M. Malek. Comprehensive testing of multistage interconnection networks. IEEE Transactions on Computers, 40(8):935–951, August 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=83638.

Montgomery:1991:EUE

[Mon91] B. L. Montgomery. Efficient unidirectional error codes for block memories. IEEE Transactions on Computers, 40(11): 1257–1259, November 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=102829.

Mathias:1990:SEP

[MP90] P. C. Mathias and L. M. Patnaik. Systolic evaluation of polynomial expressions. IEEE Transactions on Computers, 39(5):653–665, May 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=53578.

Montalvo:1998:NST

[MPG98] L. A. Montalvo, K. K. Parhi, and A. Guyot. New Svoboda– Tung division. IEEE Transactions on Computers, 47(9): 1014–1020, September 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=713319.

Miller:1993:PCR

[MPKRS93] R. Miller, V. K. Prasanna-Kumar, D. I. Reisis, and Q. F. Stout. Parallel computations on reconfigurable meshes. IEEE Transactions on Computers, 42(6):678–692, June 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=277290.

Miller:1994:NEH

[MPS94] Z. Miller, D. Pritikin, and I. H. Sudborough. Near embeddings of hypercubes into Cayley graphs on the symmetric group. IEEE Transactions on Computers, 43(1):13–22, January 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=250605.

Majumdar:1990:FTL

[MRB90] A. Majumdar, C. S. Raghavendra, and M. A. Breuer. Fault tolerance in linear systolic arrays using time redundancy. IEEE Transactions on Computers, 39(2):269–276, February 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=45214.

Mehra:1997:DEW

[MRJ97] A. Mehra, J. Rexford, and F. Jahanian. Design and evaluation of a window-consistent replication service. IEEE Transactions on Computers, 46 (9):986–996, September 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=620480.

Mukherjee:1997:DTM

[MRT97] N. Mukherjee, J. Rajski, and J. Tyszer. Design of testable multipliers for fixed-width data paths. IEEE Transactions on Computers, 46(7):795–810, July 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=599900.

McCrackin:1991:UHP

[MS91] D. C. McCrackin and B. Szabados. Using horizontal prefetching to circumvent the jump problem. IEEE Transactions on Computers, 40(11):1287–1291, November 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=102835.

Mehta:1993:DSC

[MS93] D. P. Mehta and S. Sahni. A data structure for circular string analysis and visualization. IEEE Transactions on Computers, 42(8):992–997, August 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=238492.

Mehta:1994:CDC

[MS94] D. P. Mehta and S. Sahni. Computing display conflicts in string visualization. IEEE Transactions on Computers, 43(3):350–361, March 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=272435.

Mukhopadhyaya:1995:FTR

[MS95] K. Mukhopadhyaya and B. P. Sinha. Fault-tolerant routing in distributed loop networks. IEEE Transactions on Computers, 44(12):1452–1456, December 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=477250.

Mans:1998:OEF

[MS98a] B. Mans and N. Santoro.

Optimal elections in faulty loop networks and applications. IEEE Transactions on Computers, 47(3):286–297, March 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=660164.

Mishra:1998:ECR

[MS98b] P. Mishra and M. Srivastava. Effect of connection rerouting on application performance in mobile networks. IEEE Transactions on Computers, 47(4):371–390, April 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=675703.

Marino:1999:PIM

[MS99] F. Marino and E. Swartzlander, Jr. Parallel implementation of multidimensional transforms without interprocessor communication. IEEE Transactions on Computers, 48 (9):951–960, September 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=795223.

Marques-Silva:1999:GSA

[MSS99] J. P. Marques-Silva and K. A. Sakallah. GRASP: a search algorithm for propositional satisfiability. IEEE Transactions on Computers, 48(5):506–521, May 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=769433.

Muller:1998:SLN

[MST98] J.-M. Muller, A. Scherbyna, and A. Tisserand. Semilogarithmic number systems. IEEE Transactions on Computers, 47(2):145–151, February 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=663760.

Manohar:1998:APP

[MT98] R. Manohar and J. A. Tierno. Asynchronous parallel prefix computation. IEEE Transactions on Computers, 47(11): 1244–1252, November 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=736437.

Mukherjee:1993:MPB

[MTG93] S. Mukherjee, S. K. Tripathi, and D. Ghosal. A multiclass prioritybased slotted-ring LAN and its analysis. IEEE Transactions on Computers, 42(8):1015–1020, August 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=238497.

Mendelson:1993:MLD

[MTP93] A. Mendelson, D. Thiebaut, and D. K. Pradhan. Modeling live and dead lines in cache memory systems. IEEE Transactions on Computers, 42(1):1–14, January 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=192209.

Muller:1994:SCF

[Mul94] J.-M. Muller. Some characterizations of functions computable in on-line arithmetic. IEEE Transactions on Computers, 43(6):752–755, June 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=286308.

Majumdar:1995:FCT

[MV95] A. Majumdar and S. B. K.

Vrudhula. Fault coverage and test length estimation for random pattern testing. IEEE Transactions on Computers, 44 (2):234–247, February 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=364535.

Milutinovic:1999:GEI

[MV99] V. Milutinovic and M. Valero. Guest Editors’ introduction: Cache memory and related problems: Enhancing and exploiting the locality. IEEE Transactions on Computers, 48(2):97–99, February 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=752650.

Martin:1990:IND

[MvdS90] A. J. Martin and J. L. A. van de Snepscheut. An interconnection network for distributed recursive computations. IEEE Transactions on Computers, 39(11):1393–1395, November 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=61049.

Mun:1994:PAF

[MY94a] Youngsong Mun and Hee Yong Youn. Performance analysis of finite buffered multistage interconnection networks. IEEE Transactions on Computers, 43(2):153–162, February 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=262120.

Mun:1994:PME

[MY94b] Youngsong Mun and Hee Yong Youn. Performance modeling and evaluation of circuit switching using Clos networks. IEEE Transactions on Computers, 43(7):854–861, July 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=293266.

Merchant:1995:AMC

[MY95] A. Merchant and P. S. Yu.

Analytic modeling and comparisons of striping strategies for replicated disk arrays. IEEE Transactions on Computers, 44(3):419–433, March 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=372034.

Merchant:1996:AMC

[MY96] A. Merchant and P. S. Yu.

Analytic modeling of clustered RAID with mapping based on nearly random permutation. IEEE Transactions on Computers, 45(3):367–373, March 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=485575.

Mohapatra:1996:AMB

[MYD96] P. Mohapatra, Chansu Yu, and C. R. Das. Allocation and mapping based reliability analysis of multistage interconnection networks. IEEE Transactions on Computers, 45(5):600–606, May 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=509911.

Murakawa:1999:GCG

[MYK^{+}99] M. Murakawa, S. Yoshizawa, I. Kajitani, X. Yao, N. Kajihara, M. Iwata, and T. Higuchi. The GRD chip: genetic reconfiguration of DSPs for neural network processing. IEEE Transactions on Computers, 48(6):628–639, June 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=773799.

Nair:1990:RNC

[NA90] V. S. S. Nair and J. A. Abraham. Real-number codes for fault-tolerant matrix operations on processor arrays. IEEE Transactions on Computers, 39(4):426–435, April 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=54836.

Nair:1996:ETA

[NAB96] V. S. S. Nair, J. A. Abraham, and P. Banerjee. Efficient techniques for the analysis of algorithm-based fault tolerance (ABFT) schemes. IEEE Transactions on Computers, 45(4):499–503, April 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=494110.

Nicolici:1998:CPT

[NAH98] N. Nicolici and B. M. Al-Hashimi. Correction to “The Proof of Theorem 2 In ‘Parallel Signature Analysis Design With Bounds On Allasing”’. IEEE Transactions on Computers, 47(12):1426, December 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=737690.

See [SM97].

Nair:1995:OBB

[Nai95] R. Nair. Optimal 2-bit branch predictors. IEEE Transactions on Computers, 44(5):698–702, May 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=381956.

Nanda:1993:DAC

[NB93] A. K. Nanda and L. N.

Bhuyan. Design and analysis of cache coherent multistage interconnection networks. IEEE Transactions on Computers, 42(4):458–470, April 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=214692.

Nandi:1996:API

[NC96] S. Nandi and P. Pal Chaudhuri. Analysis of periodic and intermediate boundary 90/150 cellular automata. IEEE Transactions on Computers, 45(1):1–12, January 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=481481.

Nandi:1997:RCT

[NC97] S. Nandi and P. P. Chaudhuri. Reply to comments on “Theory And Application Of Cellular Automata In Cryptography”. IEEE Transactions on Computers, 46(5):639, May 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=589246.

See [NKC94, BMP^{+}97b].

Najjar:1990:NRM

[NG90] W. Najjar and J.-L. Gaudiot. Network resilience: a measure of network fault tolerance. IEEE Transactions on Computers, 39(2):174–181, February 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=45203.

Ng:1991:IDP

[Ng91] S. W. Ng. Improving disk performance via latency reduction. IEEE Transactions on Computers, 40(1):22–30, January 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=67317.

Narayanan:1993:OCM

[NGB93] S. Narayanan, R. Gupta, and M. A. Breuer. Optimal configuring of multiple scan chains. IEEE Transactions on Computers, 42(9):1121–1131, September 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=241600.

Nair:1992:PEO

[NHA92] V. S. S. Nair, Y. V. Hoskote, and J. A. Abraham. Probabilistic evaluation of online checks in fault-tolerant multiprocessor systems. IEEE Transactions on Computers, 41(5):532–541, May 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=142679.

Nicolaidis:1996:TTR

[Nic96]  M. Nicolaidis.  Theory of

transparent BIST for RAMs. IEEE Transactions on Computers, 45(10):1141–1156, October 1996. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=543708.

Nicolaidis:1998:FSI

[Nic98]  M. Nicolaidis.  Fail-safe in-

terfaces for VLSI: theoretical foundations and implementation. IEEE Transactions on Computers, 47(1):62–77, January 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=656082.

Nikolos:1991:TDE

[Nik91] D. Nikolos. Theory and design of t-error correcting/d-error detecting (d > t) and all unidirectional error detecting codes. IEEE Transactions on Computers, 40(2):132–142, February 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=73584.

Nikolos:1998:OST

[Nik98]  D. Nikolos.  Optimal self-

testing embedded parity checkers. IEEE Transactions on Computers, 47(3):313–321, March 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=660167.

Nikolos:1992:TDE

[NK92] D. Nikolos and A. Krokos. Theory and design of t-error correcting, k-error detecting and d-unidirectional error detecting codes with d > k > t. IEEE Transactions on Computers, 41(4):411–419, April 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=135553.

Neebel:1997:CAW

[NK97] D. J. Neebel and C. R. Kime.

Cellular automata for weighted random pattern generation. IEEE Transactions on Computers, 46(11):1219–1229, November 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=644297.

Nielsen:1999:RRR

[NK99] A. M. Nielsen and P. Kornerup. Redundant radix representations of rings. IEEE Transactions on Computers, 48 (11):1153–1165, November 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=811100.

Nandi:1994:TAC

[NKC94] S. Nandi, B. K. Kar, and P. Pal Chaudhuri. Theory and applications of cellular automata in cryptography.

IEEE Transactions on Computers, 43(12):1346–1357, December 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=338094.

Nuang:1990:CDB

[NL90] W. Nuang and F. Lombardi. On the constant diagnosability of baseline interconnection networks. IEEE Transactions on Computers, 39(12): 1485–1489, December 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=61073.

Nannarelli:1999:LPD

[NL99] A. Nannarelli and T. Lang. Lowpower divider. IEEE Transactions on Computers, 48(1):2– 14, January 1999. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
743407.

Nodine:1991:OPV

[NLV91] M. H. Nodine, D. P. Lopresti, and J. S. Vitter. I/O overhead and parallel VLSI architectures for lattice computations. IEEE Transactions on Computers, 40(7):843–852, July 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=83622.

Ng:1994:UPG

[NM94] S. W. Ng and R. L. Mattson.

Uniform parity group distribution in disk arrays with multiple failures. IEEE Transactions on Computers, 43(4):501–506, April 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=278490.

Nicola:1993:FSH

[NNHG93] V. F. Nicola, M. K. Nakayama, P. Heidelberger, and A. Goyal. Fast simulation of highly dependable systems with general failure and repair processes. IEEE Transactions on Computers, 42(12):1440–1452, December 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=260634.

Nicol:1991:IAM

[NO91] D. M. Nicol and D. R. O’Hallaron.

Improved algorithms for mapping pipelined and parallel computations. IEEE Transactions on Computers, 40(3):295–306, March 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=76406.

Naganuma:1994:HPI

[NO94] J. Naganuma and T. Ogura.

A highly OR-parallel inference machine (multi-ASCA) and its performance evaluation: an architecture and its load balancing algorithms. IEEE Transactions on Computers, 43(9): 1062–1075, September 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=312115.

Nie:1990:ERS

[NP90] X. Nie and D. A. Plaisted. Experimental results on subgoal reordering. IEEE Transactions on Computers, 39(6):845–848, June 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=53609.

Nicol:1990:ODR

[NR90] D. M. Nicol and P. F. Reynolds, Jr. Optimal dynamic remapping of data parallel computations. IEEE Transactions on Computers, 39(2):206–219, February 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=45206.

Nicol:1990:ASD

[NS90] D. M. Nicol and J. H. Saltz. An analysis of scatter decomposition. IEEE Transactions on Computers, 39(11): 1337–1345, November 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=61043.

Nabli:1996:PAN

[NS96] H. Nabli and B. Sericola. Performability analysis: a new algorithm. IEEE Transactions on Computers, 45(4):491–494, April 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=494108.

Nicol:1996:SAS

[NST96] D. M. Nicol, R. Simha, and D. Towsley. Static assignment of stochastic tasks using majorization. IEEE Transactions on Computers, 45(6):730–740, June 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=506428.

Nachman:1998:NAR

[NSUR98] L. Nachman, K. Saluja, S. J.

Upadhyaya, and R. Reuse. A novel approach to random pattern testing of sequential circuits. IEEE Transactions on Computers, 47(1):129–134, January 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=656097.

Nain:1992:CHM

[NT92] P. Nain and D. Towsley. Comparison of hybrid minimum laxity/first-in-first-out scheduling policies for real-time multiprocessors. IEEE Transactions on Computers, 41 (10):1271–1278, October 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=166604.

Nikolos:1999:YVP

[NV99] D. Nikolos and H. T. Vergos. On the yield of VLSI processors with on-chip CPU cache. IEEE Transactions on Computers, 48(10):1138–1144, October 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=805163.

OConnor:1995:DCT

[O’C95] L. O’Connor. A differential cryptanalysis of tree-structured substitution-permutation networks. IEEE Transactions on Computers, 44(9):1150–1152, September 1995. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
464394.

Oliveira:1998:EMB

[OCVSV98] A. L. Oliveira, L. P. Carloni, T. Villa, and A. L. Sangiovanni-Vincentelli. Exact minimization of binary decision diagrams using implicit techniques. IEEE Transactions on Computers, 47(11):1282–1296, November 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=736442.

Oommen:1996:GPU

[OdSC96] B. J. Oommen and E. V.

de St. Croix. Graph partitioning using learning automata. IEEE Transactions on Computers, 45(2):195–208, February 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=485372.

Oberman:1997:DID

[OF97a] S. F. Oberman and M. J.

Flynn. Design issues in division and other floating-point operations. IEEE Transactions on Computers, 46(2):154– 161, February 1997. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL ftp://arith.stanford.
edu/tr/desissues_TOC.pdf;
 https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
565590.

Oberman:1997:DAI

[OF97b] S. F. Oberman and M. J. Flynn.

Division algorithms and implementations. IEEE Transactions on Computers, 46(8): 833–854, August 1997. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL ftp://arith.stanford.
edu/tr/divalgo_TOC.pdf;

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        609274.

OKeefe:1992:RBT

[OFW92] M. T. O’Keefe, J. A. B. Fortes, and B. W. Wah. On the relationship between two systolic array design methodologies. IEEE Transactions on Computers, 41(12):1589–1593, December 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=214667.

Onyuksel:1990:MQN

[OI90a] I. H. Onyuksel and K. B. Irani. Markovian queueing network models for performance analysis of a single-bus multiprocessor system. IEEE Transactions on Computers, 39(7):975–980, July 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=55702.

Owens:1990:BSM

[OI90b] R. M. Owens and M. J. Irwin. Being stingy with multipliers. IEEE Transactions on Computers, 39(6):809–818, June 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=53602.

Oikonomou:1992:AFS

[Oik92] K. N. Oikonomou. Abstractions of finite-state machines and immediately-detectable output faults. IEEE Transactions on Computers, 41(3):325–338, March 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=127444.

Orailoglu:1996:ASS

[OK96] A. Orailoglu and R. Karri.

Automatic synthesis of selfrecovering VLSI systems. IEEE Transactions on Computers, 45 (2):131–142, February 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=485368.

Ould-Khaoua:1999:PMD

[OK99] M. Ould-Khaoua. A performance model for Duato’s fully adaptive routing algorithm in k-ary n-cubes. IEEE Transactions on Computers, 48(12): 1297–1304, December 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=817384.

Oruc:1991:SAC

[OM91] A. Y. Oruc and M. Mittal. Setup algorithms for cubeconnected parallel computers using recursive Karnaugh maps. IEEE Transactions on Computers, 40(2):217–221, February 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=73592.

Olstad:1995:EPS

[OM95] B. Olstad and F. Manne. Efficient partitioning of sequences. IEEE Transactions on Computers, 44(11):1322–1326, November 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=475128.

Olukotun:1997:MOP

[OMB97] K. Olukotun, T. N. Mudge, and R. B. Brown. Multilevel optimization of pipelined caches. IEEE Transactions on Computers, 46(10):1093–1102, October 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=628394.

Ortega:1993:GHN

[OPLP93] J. Ortega, A. Prieto, A. Lloris, and F. J. Pelayo. Generalized Hopfield neural network for concurrent testing. IEEE Transactions on Computers, 42(8):898–912, August 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=238481.

Orton:1992:NFT

[OPT92] G. A. Orton, L. E. Peppard, and S. E. Tavares.

New fault tolerant techniques for residue number systems. IEEE Transactions on Computers, 41(11):1453–1464, November 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=177315;
https://ieeexplore.ieee.
org/xpl/tocresult.jsp?isnumber=
4475.

Ortega:1993:TPG

[ORPP93] J. Ortega, A. L. Ruiz, A. Prieto, and F. J. Pelayo. Testpattern generation based on Reed–Muller coefficients. IEEE Transactions on Computers, 42(8):968–980, August 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=238487.

Olson:1994:PCS

[OS94] A. Olson and K. G. Shin. Probabilistic clock synchronization in large distributed systems. IEEE Transactions on Computers, 43(9):1106–1112, September 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=312120.

Opatrny:1996:DLC

[OSST96] J. Opatrny, D. Sotteau, N. Srinivasan, and K. Thulasiraman. DCC linear congruential graphs: a new class of interconnection networks. IEEE Transactions on Computers, 45 (2):156–164, February 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=485369.

Oklobdzija:1996:MSO

[OVL96] V. G. Oklobdzija, D. Villeger, and S. S. Liu. A method for speed optimized partial product reduction and generation of fast parallel multipliers using an algorithmic approach. IEEE Transactions on Computers, 45(3):294–306, March 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=485568.

OGrady:1991:HOA

[OY91] E. Pearse O’Grady and Baek-Kyu K. Young. A hardwareoriented algorithm for floatingpoint function generation. IEEE Transactions on Computers, 40 (2):237–241, February 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=73596.

Oh:1995:EAB

[OYR95] Choong Gun Oh, Hee Yong Youn, and V. K. Raj. An efficient algorithm-based concurrent error detection for FFT networks. IEEE Transactions on Computers, 44(9):1157–1162, September 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=464396.

Ofek:1997:CAB

[OYY97] Y. Ofek, B. Yener, and M. Yung.

Concurrent asynchronous broadcast on the MetaNet. IEEE Transactions on Computers, 46 (7):737–748, July 1997. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
599894.

Oommen:1996:NSD

[OZL96] B. J. Oommen, K. Zhang, and W. Lee. Numerical similarity and dissimilarity measures between two trees. IEEE Transactions on Computers, 45 (12):1426–1434, December 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=545972.

Park:1996:WEH

[PA96] Hyunmin Park and D. P.

Agrawal. WICI: an efficient hybrid routing scheme for scalable and hierarchical networks. IEEE Transactions on Computers, 45(11):1272–1281, November 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=544483.

Paar:1996:NAP

[Paa96]  C. Paar.  A new architec-

ture for a parallel finite field multiplier with low complexity based on composite fields. IEEE Transactions on Computers, 45(7):856–861, July 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=508323.

Palumbo:1994:DEV

[Pal94] D. L. Palumbo. The derivation and experimental verification of clock synchronization theory. IEEE Transactions on Computers, 43(6):676–686, June 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=286301.

Parhami:1990:GSD

[Par90]  Behrooz Parhami.  General-

ized signed-digit number systems: a unifying framework for redundant number representations. IEEE Transactions on Computers, 39(1):89–98, January 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=46283.

Parhami:1993:IAS

[Par93] B. Parhami. On the implementation of arithmetic support functions for generalized signed-digit number systems. IEEE Transactions on Computers, 42(3):379–384, March 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=210182.

Parhami:1994:CEC

[Par94] B. Parhami. Comments on “Evaluation of A + B = K conditions without carry propagation”. IEEE Transactions on Computers, 43(3):381, March 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=272440.

See [CL92b].

Park:1997:AAB

[PB97a] S. Park and B. Bose. All-to-all broadcasting in faulty hypercubes. IEEE Transactions on Computers, 46(7):749–755, July 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=599895.

Parker:1997:MAU

[PB97b] M. G. Parker and M. Benaissa. Modular arithmetic using low order redundant bases. IEEE Transactions on Computers, 46(5):611–616, May 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=589237.

Prabhakaran:1999:PAF

[PB99a] P. Prabhakaran and P. Banerjee. Parallel algorithms for force directed scheduling of flattened and hierarchical signal flow graphs. IEEE Transactions on Computers, 48(7):762–768, July 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=780886.

Purna:1999:TPS

[PB99b] K. M. G. Purna and D. Bhatia. Temporal partitioning and scheduling data flow graphs for reconfigurable computers. IEEE Transactions on Computers, 48(6):579–590, June 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=773795.

Papadimitriou:1997:SCC

[PBB97] S. Papadimitriou, A. Bezerianos, and T. Bountis. Secure communication with chaotic systems of difference equations. IEEE Transactions on Computers, 46(1):27–38, January 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=559800.

Pong:1998:DVS

[PBND98] Fong Pong, M. Browne, A. Nowatzyk, and M. Dubois. Design verification of the S3.mp cachecoherent shared-memory system. IEEE Transactions on Computers, 47(1):135–140, January 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=656100.

Park:1999:ESV

[PD99] Seungjoon Park and D. L. Dill.

An executable specification and verifier for relaxed memory order. IEEE Transactions on Computers, 48(2):227–235, February 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=752664.

Papazoglou:1997:CLS

[PDBH97] M. Papazoglou, A. Dells, A. Bouguettaya, and M. Haghjoo. Class library support for workflow environments and applications. IEEE Transactions on Computers, 46(6):673–686, June 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=600825.

Pearlmutter:1996:DTD

[Pea96]  B. A. Pearlmutter.  Doing

the twist: diagonal meshes are isomorphic to twisted toroidal meshes. IEEE Transactions on Computers, 45(6):766–767, June 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=506434.

Paschalis:1990:ETC

[PEH90] A. M. Paschalis, C. Efstathiou, and C. Halatsis. An efficient TSC 1-out-of-3 code checker. IEEE Transactions on Computers, 39(3):407–411, March 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=48873.

Pekmestzi:1999:MBA

[Pek99] K. Z. Pekmestzi. Multiplexerbased array multipliers. IEEE Transactions on Computers, 48(1):15–23, January 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=743408.

Pelc:1991:UGM

[Pel91]  A. Pelc.  Undirected graph

models for system-level fault diagnosis. IEEE Transactions on Computers, 40(11): 1271–1276, November 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=102832.

Pelc:1992:OFD

[Pel92] A. Pelc. Optimal fault diagnosis in comparison models. IEEE Transactions on Computers, 41(6):779–786, June 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=144631.

Pelc:1998:ODH

[Pel98] A. Pelc. Optimal diagnosis of heterogeneous systems with random faults. IEEE Transactions on Computers, 47(3):298–304, March 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=660165.

Patel:1999:EDO

[PFP99] S. J. Patel, D. H. Friendly, and Y. N. Patt. Evaluation of design options for the trace cache fetch mechanism. IEEE Transactions on Computers, 48(2):193–204, February 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=752661.

Paar:1998:EMA

[PFR98] C. Paar, P. Fleischmann, and P. Roeise. Efficient multiplier architectures for Galois fields GF(2^{4n}). IEEE Transactions on Computers, 47 (2):162–170, February 1998.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=663762.

Paar:1999:FAP

[PFSR99] C. Paar, P. Fleischmann, and P. Soria-Rodriguez. Fast arithmetic for public-key algorithms in Galois fields with composite exponents. IEEE Transactions on Computers, 48 (10):1025–1034, October 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=805153.

Pradhan:1991:NFD

[PG91] D. K. Pradhan and S. K. Gupta. A new framework for designing and analyzing BIST techniques and zero aliasing compression. IEEE Transactions on Computers, 40(6):743–763, June 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=90252.

Perl:1992:AIS

[PG92] Y. Perl and L. Gabriel. Arithmetic interpolation search for alphabet tables. IEEE Transactions on Computers, 41(4):493– 499, April 1992. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
135562.

Phalke:1997:CBP

[PG97] V. Phalke and B. Gopinath.

Compression-based program characterization for improving cache memory performance. IEEE Transactions on Computers, 46 (11):1174–1186, November 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=644292.

Pradhan:1990:APM

[PGK90] D. K. Pradhan, S. K. Gupta, and M. G. Karpovsky. Aliasing probability for multiple input signature analyzer. IEEE Transactions on Computers, 39(4):586– 591, April 1990. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
54855.

Paul:1995:EAC

[PGS95] S. Paul, J. Gotze, and M. Sauer.

Error analysis of CORDIC-based Jacobi algorithms. IEEE Transactions on Computers, 44(7): 947–951, July 1995. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
392855.

Pour:1993:PIT

[PH93] A. F. Pour and M. D. Hill.

Performance implications of tolerating cache faults. IEEE Transactions on Computers, 42(3):257–267, March 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=210168.

Phatak:1998:CDM

[Pha98a] D. S. Phatak. Comments on Duprat and Muller’s branching CORDIC paper. IEEE Transactions on Computers, 47 (9):1037–1040, September 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=713326.

Phatak:1998:DSB

[Pha98b] D. S. Phatak. Double step branching CORDIC: a new algorithm for fast sine and cosine generation. IEEE Transactions on Computers, 47(5):587–602, May 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=677251.

Park:1994:TCA

[PHK94] In-Cheol Park, Se-Kyoung Hong, and Chong-Min Kyung. Two complementary approaches for microcode bit optimization. IEEE Transactions on Computers, 43(2):234–239, February 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=262130.

Peir:1999:FIT

[PHS99] Jih-Kwon Peir, W. W. Hsu, and A. J. Smith. Functional implementation techniques for CPU cache memories. IEEE Transactions on Computers, 48(2):100–110, February 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=752651.

Papachristou:1993:VMS

[PI93] C. A. Papachristou and V. R.

Immaneni. Vertical migration of software functions and algorithms using enhanced microsequencing. IEEE Transactions on Computers, 42(1):45–61, January 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=192213.

Piestrak:1990:DHS

[Pie90]  S. J. Piestrak.  Design of

high-speed and cost-effective self-testing checkers for lowcost arithmetic codes. IEEE Transactions on Computers, 39(3):360–374, March 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=48866.

Piestrak:1993:MTS

[Pie93] S. J. Piestrak. The minimal test set for multioutput threshold circuits implemented as sorting networks. IEEE Transactions on Computers, 42(6):700–712, June 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=277288.

Piestrak:1994:DRG

[Pie94] S. J. Piestrak. Design of residue generators and multioperand modular adders using carry-save adders. IEEE Transactions on Computers, 43(1):68–77, January 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=250610.

Piestrak:1996:DST

[Pie96]  S. J. Piestrak.  Design of

self-testing checkers for Borden codes. IEEE Transactions on Computers, 45(4):461–469, April 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=494103.

Pinkston:1999:FER

[Pin99]  T. M. Pinkston.  Flexible

and efficient routing based on progressive deadlock recovery. IEEE Transactions on Computers, 48(7):649–669, July 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=780873.

Pramanik:1990:PPL

[PK90] S. Pramanik and M. Kim.

Parallel processing of large node B-trees. IEEE Transactions on Computers, 39(9): 1208–1212, September 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=57061.

Pomeranz:1991:PCA

[PK91] I. Pomeranz and Z. Kohavi.

Polynomial complexity algorithms for increasing the testability of digital circuits by testing-module insertion. IEEE Transactions on Computers, 40 (11):1198–1213, November 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=102824.

Park:1993:GLA

[PK93] S. K. Park and J. H. Kim. Geometrical learning algorithm for multilayer neural networks in a binary field. IEEE Transactions on Computers, 42(8):988–992, August 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=238491.

Phatak:1994:HSD

[PK94] D. S. Phatak and I. Koren. Hybrid signed-digit number systems: a unified framework for redundant number representations with bounded carry propagation chains. IEEE Transactions on Computers, 43(8):880–891, August 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=295850.

Parhami:1996:CHS

[PKI^{+}96] B. Parhami, S. Kawahito, M. Ishida, T. Nakamura, M. Kameyama, and T. Higuchi. Comments on “High-speed area-efficient multiplier design using multiplevalued current-mode circuits”. IEEE Transactions on Computers, 45(5):637–639, May 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=509918.

See [KIN^{+}94].

Pan:1999:PSP

[PL99] P. Pan and C. L. Liu. Partial scan with preselected scan signals. IEEE Transactions on Computers, 48(9):1000–1005, September 1999. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        795228.

Pattipati:1993:UFP

[PLB93] K. R. Pattipati, Y. Li, and H. A. P. Blom. A unified framework for the performability evaluation of fault-tolerant computer systems. IEEE Transactions on Computers, 42(3):312–326, March 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=210174.

Prakash:1994:NAS

[PLJ94] S. Prakash, Yann Hang Lee, and T. Johnson. A nonblocking algorithm for shared queues using compare-and-swap. IEEE Transactions on Computers, 43(5):548–559, May 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=280802.

Perunicic:1991:SMA

[PLM91] B. Perunicic, S. Lakhani, and V. Milutinovic. Stochastic modeling and analysis of propagation delays in GaAs adders. IEEE Transactions on Computers, 40(1):31–45, January 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=67318.

Parhi:1991:SRO

[PM91] K. K. Parhi and D. G. Messerschmitt. Static rate-optimal scheduling of iterative data-flow programs via optimum unfolding. IEEE Transactions on Computers, 40(2):178–195, February 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=73588.

Pandya:1998:MAU

[PM98] M. Pandya and M. Malek. Minimum achievable utilization for fault-tolerant processing of periodic tasks. IEEE Transactions on Computers, 47 (10):1102–1112, October 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=729793.

Powell:1995:EFT

[PMAC95] D. Powell, E. Martins, J. Arlat, and Y. Crouzet. Estimators for fault tolerance coverage evaluation. IEEE Transactions on Computers, 44 (2):261–274, February 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=364537.

Padro:1996:CLD

[PMF96] C. Padro, P. Morillo, and M. A. Fiol. Comments on “Line digraph iterations and connectivity analysis of de Bruijn and Kautz graphs”. IEEE Transactions on Computers, 45(6):768, June 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=506435.

See [DZLH93, DHL96].

Pentakalos:1997:APM

[PMHY97] O. I. Pentakalos, D. A. Menasce, M. Halem, and Y. Yesha. Analytical performance modeling of hierarchical mass storage systems. IEEE Transactions on Computers, 46(10):1103–1118, October 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=628395.

Park:1992:TDF

[PMW92] E. S. Park, M. R. Mercer, and T. W. Williams. The total delay fault model and statistical delay fault coverage. IEEE Transactions on Computers, 41(6):688–698, June 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=144621.

Panda:1999:ALT

[PNDN99] P. R. Panda, H. Nakamura, N. D. Dutt, and A. Nicolau. Augmenting loop tiling with data alignment for improved cache performance. IEEE Transactions on Computers, 48 (2):142–149, February 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=752655.

Poledna:1995:TST

[Pol95] S. Poledna. Tolerating sensor timing faults in highly responsive hard real-time systems. IEEE Transactions on Computers, 44(2):181–191, February 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=364530.

Pourbigharaz:1997:SDA

[Pou97] F. Pourbigharaz. A signeddigit architecture for residue to binary transformation. IEEE Transactions on Computers, 46 (10):1146–1150, October 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=628400.

Pomeranz:1992:MOT

[PR92] I. Pomeranz and S. M. Reddy. The multiple observation time test strategy. IEEE Transactions on Computers, 41(5):627–637, May 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=142689.

Pomeranz:1993:CFS

[PR93a] I. Pomeranz and S. M. Reddy. Classification of faults in synchronous sequential circuits.

IEEE Transactions on Computers, 42(9):1066–1077, September 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=241596.

Pomeranz:1993:TFT

[PR93b] I. Pomeranz and S. M. Reddy.

Testing of fault-tolerant hardware through partial control of inputs. IEEE Transactions on Computers, 42 (10):1267–1271, October 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=257713.

Pomeranz:1994:AHS

[PR94a] I. Pomeranz and S. M. Reddy.

Application of homing sequences to synchronous sequential circuit testing. IEEE Transactions on Computers, 43(5):569–580, May 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=280804.

Pomeranz:1994:RHR

[PR94b] I. Pomeranz and S. M. Reddy.

On the role of hardware reset in synchronous sequential circuit test generation. IEEE Transactions on Computers, 43(9):1100–1105, September 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=312119.

Pomeranz:1995:ACU

[PR95a] I. Pomeranz and S. M. Reddy. Aliasing computation using fault simulation with fault dropping. IEEE Transactions on Computers, 44(1):139–144, January 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=368001.

Pomeranz:1995:INS

[PR95b] I. Pomeranz and S. M. Reddy. INCREDYBLE: a new search strategy for design automation problems with applications to testing. IEEE Transactions on Computers, 44(6):792–804, June 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=391182.

Pomeranz:1995:FSS

[PR95c] I. Pomeranz and S. M. Reddy. On fault simulation for synchronous sequential circuits. IEEE Transactions on Computers, 44(2):335–340, February 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=364543.

Pomeranz:1996:RRS

[PR96a] I. Pomeranz and S. M. Reddy. On removing redundancies from synchronous sequential circuits with synchronizing sequences. IEEE Transactions on Computers, 45(1):20–32, January 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=481483.

Pomeranz:1996:NTD

[PR96b] I. Pomeranz and S. M. Reddy.

On the number of tests to detect all path delay faults in combinational logic circuits. IEEE Transactions on Computers, 45(1):50–62, January 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=481486.

Pomeranz:1997:DBF

[PR97a] I. Pomeranz and S. M. Reddy.

On dictionary-based fault location in digital logic circuits. IEEE Transactions on Computers, 46(1):48–59, January 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=559802.

Pomeranz:1997:TGM

[PR97b] I. Pomeranz and S. M. Reddy.

Test generation for multiple state-table faults in finite-state machines. IEEE Transactions on Computers, 46(7):783–794, July 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=599899.

Pomeranz:1998:LSF

[PR98] I. Pomeranz and S. M. Reddy. Location of stuck-at faults and bridging faults based on circuit partitioning. IEEE Transactions on Computers, 47 (10):1124–1135, October 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=729795.

Pomeranz:1999:CBG

[PR99] I. Pomeranz and S. M. Reddy. A cone-based genetic optimization procedure for test generation and its application to n-detections in combinational circuits. IEEE Transactions on Computers, 48 (10):1145–1152, October 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=805164.

Pai:1999:IEI

[PRASA99] V. S. Pai, P. Ranganathan, H. Abdel-Shafi, and S. Adve. The impact of exploiting instructionlevel parallelism on sharedmemory multiprocessors. IEEE Transactions on Computers, 48 (2):218–226, February 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=752663.

Paik:1994:DVB

[PRS94] D. Paik, S. Reddy, and S. Sahni.

Deleting vertices to bound path length. IEEE Transactions on Computers, 43(9): 1091–1096, September 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=312117.

Pattipati:1990:CAP

[PS90] K. R. Pattipati and S. A.

Shah. On the computational aspects of performability models of fault-tolerant computer systems. IEEE Transactions on Computers, 39(6):832–836, June 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=53605.

Probert:1991:SCP

[PS91] R. L. Probert and K. Saleh.

Synthesis of communication protocols: survey and assessment. IEEE Transactions on Computers, 40(4):468–476, April 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=88466.

Perleberg:1993:BTB

[PS93] C. H. Perleberg and A. J. Smith.

Branch target buffer design and optimization. IEEE Transactions on Computers, 42(4):396– 412, April 1993. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
214687.

Panchapakesan:1999:LNT

[PS99] G. Panchapakesan and A. Sengupta. On a lightwave network topology using Kautz digraphs. IEEE Transactions on Computers, 48(10):1131–1137, October 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=805162.

Peha:1992:CTD

[PT92] J. M. Peha and F. A. Tobagi. Comments on “Tolerance of double-loop computer networks to multimode failures” by H. Masuyama and T. Ichimori. IEEE Transactions on Computers, 41(11):1488–1490, November 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=177319.

Papakostas:1998:IOG

[PT98] A. Papakostas and I. G. Tollis. Interactive orthogonal graph drawing. IEEE Transactions on Computers, 47(11): 1297–1309, November 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=736444.

Purwar:1991:EMC

[Pur91]  S. Purwar.  An efficient

method of computing generalized Reed–Muller expansions from binary decision diagram. IEEE Transactions on Computers, 40(11):1298–1301, November 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=102837.

Preparata:1990:PCD

[PV90] F. P. Preparata and J. E.

Vuillemin. Practical cellular dividers. IEEE Transactions on Computers, 39(5):605–614, May 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=53574.

Phillips:1994:HPI

[PV94a] J. Phillips and S. Vassiliadis. High-performance 3-1 interlock collapsing ALU’s. IEEE Transactions on Computers, 43(3):257–268, March 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=272427.

Pradhan:1994:RFC

[PV94b] D. K. Pradhan and N. H.

Vaidya. Roll-forward checkpointing scheme: a novel faulttolerant architecture. IEEE Transactions on Computers, 43 (10):1163–1174, October 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=324542.

Pradhan:1997:RFR

[PV97] D. K. Pradhan and N. H. Vaidya. Roll-forward and rollback recovery: performance-reliability trade-off. IEEE Transactions on Computers, 46(3):372–378, March 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=580435.

Pilarski:1995:CBC

[PW95] S. Pilarski and K. J. Wiebe. Counter-based compaction: Delay and stuck-open faults. IEEE Transactions on Computers, 44(6):780–791, June 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=391183.

Parhi:1994:SPN

[PWG94] K. K. Parhi, F. H. Wu, and K. Genesan. Sequential and parallel neural network vector quantizers. IEEE Transactions on Computers, 43(1):104–109, January 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=250614.

Porter:1993:NNN

[PZ93] W. A. Porter and Xiaoyan Zheng. A nonbinary neural network design. IEEE Transactions on Computers, 42(9): 1132–1135, September 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=241601.

Pham:1999:SCM

[PZ99] Hoang Pham and Xuemei Zhang. A software cost model with warranty and risk costs. IEEE Transactions on Computers, 48(1):71–75, January 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=743412.

Quach:1992:HSA

[QF92] N. T. Quach and M. J. Flynn.

High-speed addition in CMOS. IEEE Transactions on Computers, 41(12):1612–1615, December 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=214671.

Qiao:1993:TDO

[QM93] C. Qiao and R. G. Melhem.

Time-division optical communications in multiprocessor arrays. IEEE Transactions on Computers, 42(5):577–590, May 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=223677.

Qin:1991:MTC

[QSA91] B. Qin, H. A. Sholl, and R. A. Ammar. Micro time cost analysis of parallel computations. IEEE Transactions on Computers, 40(5):613–628, May 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=88485.

Quinn:1990:AIB

[Qui90] M. J. Quinn. Analysis and implementation of branch-and-bound algorithms on a hypercube multicomputer. IEEE Transactions on Computers, 39(3):384–387, March 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=48868.

Raghavan:1993:AIP

[Rag93] V. Raghavan. On asymmetric invalidation with partial tests. IEEE Transactions on Computers, 42(6):764–768, June 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=277298.

Raghavan:1996:WDA

[Rag96] V. Raghavan. Weighted diagnosis with asymmetric invalidation. IEEE Transactions on Computers, 45(12):1435–1438, December 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=545973.

Rai:1995:EFD

[Rai95] S. Rai. Evaluating FTRE’s for dependability measures in fault tolerant systems. IEEE Transactions on Computers, 44 (2):275–285, February 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=364538.

Rajasekaran:1996:MCC

[Raj96] S. Rajasekaran. Mesh connected computers with fixed and reconfigurable buses: packet routing and sorting. IEEE Transactions on Computers, 45(5):529–539, May 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=509905.

Rao:1993:CCI

[Rao93a] N. S. V. Rao. Computational complexity issues in operative diagnosis of graph-based systems. IEEE Transactions on Computers, 42(4):447–457, April 1993. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=214691.

Rao:1993:EVA

[Rao93b] N. S. V. Rao. Expectedvalue analysis of two single fault diagnosis algorithms. IEEE Transactions on Computers, 42(3):272–280, March 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=210170.

Reddy:1990:ABF

[RB90] A. L. N. Reddy and P. Banerjee. Algorithm-based fault detection for signal processing applications. IEEE Transactions on Computers, 39 (10):1304–1308, October 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=59860.

Raghavendra:1991:SRB

[RB91] C. S. Raghavendra and R. V. Boppana. On self-routing in Benes and shuffle-exchange networks. IEEE Transactions on Computers, 40(9): 1057–1064, September 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=83649.

Rowley:1993:FTR

[RB93] R. A. Rowley and B. Bose.

Fault-tolerant ring embedding in de Bruijn networks. IEEE Transactions on Computers, 42 (12):1480–1486, December 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=260637.

Rowley:1997:DRE

[RB97] R. A. Rowley and B. Bose.

Distributed ring embedding in faulty De Bruijn networks. IEEE Transactions on Computers, 46(2):187–190, February 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=565593.

Ramos:1999:TOB

[RB99] J. F. Ramos and A. G. Bohorquez. Two operand binary adders with threshold logic. IEEE Transactions on Computers, 48(12):1324–1337, December 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=817389.

Rovatti:1998:GAM

[RBG98] R. Rovatti, M. Borgatti, and R. Guerrieri. A geometric approach to maximum-speed ndimensional continuous linear interpolation in rectangular grids.

IEEE Transactions on Computers, 47(8):894–899, August 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=707591.

Roychowdhury:1990:EAR

[RBK90] V. P. Roychowdhury, J. Bruck, and T. Kailath. Efficient algorithms for reconfiguration in VLSI/ WSI arrays. IEEE Transactions on Computers, 39(4): 480–489, April 1990. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
54841.

Rotenberg:1999:TCM

[RBS99] E. Rotenberg, S. Bennett, and J. E. Smith. A trace cache microarchitecture and evaluation. IEEE Transactions on Computers, 48(2):111–120, February 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=752652.

Roy-Chowdhury:1996:ABF

[RCB96] A. Roy-Chowdhury and P. Banerjee. Algorithm-based fault location and recovery for matrix computations on multiprocessor systems. IEEE Transactions on Computers, 45(11): 1239–1247, November 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=544480.

Roy-Chowdhury:1996:ABE

[RCBB96] A. Roy-Chowdhury, N. Bellas, and P. Banerjee. Algorithmbased error-detection schemes for iterative solution of partial differential equations. IEEE Transactions on Computers, 45(4):394– 407, April 1996. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
494098.

Ramani:1992:GMP

[RCS92] A. K. Ramani, P. K. Chande, and P. C. Sharma. A general model for performance investigations of priority based multiprocessor system. IEEE Transactions on Computers, 41(6):747– 754, June 1992. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
144626.

Ray-Chaudhuri:1994:TDU

[RCSSS94] D. K. Ray-Chaudhuri, N. M.

Singhi, S. Sanyal, and P. S. Subramanian. Theory and design of t-unidirectional errorcorrecting and d-unidirectional error-detecting code. IEEE Transactions on Computers, 43 (10):1221–1226, October 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=324547.

Rai:1996:HNE

[RD96] Suresh Rai and Weian Deng. Hyperneural network — an efficient model for test generation in digital circuits. IEEE Transactions on Computers, 45(1):115–121, January 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=481493.

Rangarajan:1995:DSL

[RDZ95] S. Rangarajan, A. T. Dahbura, and E. A. Ziegler. A distributed system-level diagnosis algorithm for arbitrary network topologies. IEEE Transactions on Computers, 44(2):312–334, February 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=364542.

Redinbo:1998:GAB

[Red98] G. R. Redinbo. Generalized algorithm-based fault tolerance: error correction via Kalman estimation. IEEE Transactions on Computers, 47(6):639–655, June 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=689644.

Reyneri:1991:APS

[RF91] L. M. Reyneri and E. Filippi. An analysis on the performance of silicon implementations of backpropagation algorithms for artificial neural networks. IEEE Transactions on Computers, 40(12):1380–1389, December 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=106223.

Rangarajan:1992:DAC

[RF92] S. Rangarajan and D. Fussell.

Diagnosing arbitrarily connected parallel computers with high probability. IEEE Transactions on Computers, 41(5):606–615, May 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=142687.

Rahardja:1999:FLI

[RF99] S. Rahardja and B. J. Falkowski.

Fast linearly independent arithmetic expansions. IEEE Transactions on Computers, 48 (9):991–999, September 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=795227.

Ramakrishna:1997:EHH

[RFB97] M. V. Ramakrishna, E. Fu, and E. Bahcekapili. Efficient hardware hashing functions for high performance computers. IEEE Transactions on Computers, 46(12):1378–1381, December 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=641938.

Rao:1993:NTS

[RFKL93] T. R. N. Rao, G. L. Feng, M. S. Kolluru, and J. C. Lo. Novel totally self-checking Berger code checker designs based on generalized Berger code partitioning. IEEE Transactions on Computers, 42(8):1020–1024, August 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=238498.

Rangarajan:1990:BTI

[RFM90] S. Rangarajan, D. Fussell, and M. Malek. Built-in testing of integrated circuit wafers. IEEE Transactions on Computers, 39(2):195–205, February 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=45205.

Rau:1992:DTR

[RFS92] D. Rau, J. A. B. Fortes, and H. J. Siegel. Destination tag routing techniques based on a state model for the IADM network. IEEE Transactions on Computers, 41(3):274–285, March 1992. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=127440.

Raghavan:1993:RIA

[RH93] R. Raghavan and J. P. Hayes.

Reducing interference among vector accesses in interleaved memories. IEEE Transactions on Computers, 42(4):471–483, April 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=214693.

Rexford:1998:RAR

[RHS98] J. Rexford, J. Hall, and K. G.

Shin. A router architecture for real-time communication in multicomputer networks. IEEE Transactions on Computers, 47(10):1088–1101, October 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=729792.

Rangarajan:1995:CRI

[RHT95] S. Rangarajan, Y. Huang, and S. K. Tripathi. Computing reliability intervals for k-resilient protocols. IEEE Transactions on Computers, 44(3):462–466, March 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=372039.

Rai:1996:MTA

[RK96] S. Rai and V. P. Kirpalani. A modified TRAM architecture. IEEE Transactions on Computers, 45(8):969–974, August 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=536239.

Raha:1999:UTR

[RKJZ99] A. Raha, S. Kamat, Xiaohua Jia, and Wei Zhao. Using traffic regulation to meet end-to-end deadlines in ATM networks. IEEE Transactions on Computers, 48(9):917–935, September 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=795221.

Ramanathan:1990:HAS

[RKS90] P. Ramanathan, D. D. Kandlur, and K. G. Shin. Hardwareassisted software clock synchronization for homogeneous distributed systems. IEEE Transactions on Computers, 39(4):514– 524, April 1990. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        54844.

Rajagopal:1991:SPC

[RM91] M. Rajagopal and R. E. Miller. Synthesizing a protocol converter from executable protocol traces. IEEE Transactions on Computers, 40(4):487–499, April 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=88468.

Ramirez:1994:CAF

[RM94] J. Ramirez and R. Melhem.

Computational arrays with flexible redundancy. IEEE Transactions on Computers, 43(4):413– 430, April 1994. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
278480.

Roy:1995:RFU

[RN95]  K. Roy and S. Nag.  On

routability for FPGAs under faulty conditions. IEEE Transactions on Computers, 44(11): 1296–1305, November 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=475125.

Redinbo:1993:MCD

[RNA93] G. R. Redinbo, L. M. Napolitano, Jr., and D. D. Andaleon. Multibit correcting data interface for fault-tolerant systems. IEEE Transactions on Computers, 42(4):433–446, April 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=214690.

Robinson:1991:APF

[Rob91] J. P. Robinson. Aliasing probabilities for feedback signature compression of test data. IEEE Transactions on Computers, 40(7):867–873, July 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=83625.

Rosenkrantz:1990:HHS

[Ros90] D. J. Rosenkrantz. Half-hot state assignments for finite state machines. IEEE Transactions on Computers, 39(5):700–702, May 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=53583.

Rudnick:1999:ETD

[RP99] E. M. Rudnick and J. H. Patel. Efficient techniques for dynamic test sequence compaction. IEEE Transactions on Computers, 48(3):323–330, March 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=754998.

Reinhardt:1998:HSF

[RPW98] S. K. Reinhardt, R. W. Pfile, and D. A. Wood. Hardware support for flexible distributed shared memory. IEEE Transactions on Computers, 47 (10):1056–1072, October 1998.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=729790.

Rosenkrantz:1993:IBA

[RR93] D. J. Rosenkrantz and S. S.

Ravi. Improved bounds for algorithm-based fault tolerance. IEEE Transactions on Computers, 42(5):630–635, May 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=223684.

Robbins:1995:BBM

[RR95] K. A. Robbins and S. Robbins.

Buffered banks in multiprocessor systems. IEEE Transactions on Computers, 44(4):518–530, April 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=376167.

Radoytchevsky:1994:CFI

[RS94] V. T. Radoytchevsky and A. J.

Shalaev. Comments on “An O(n^{2.5}) fault identification algorithm for diagnosable systems”. IEEE Transactions on Computers, 43(2):254–255, February 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=262134.

Rubino:1995:IAA

[RS95] G. Rubino and B. Sericola. Interval availability analysis using denumerable Markov processes: application to multiprocessor subject to breakdowns and repair. IEEE Transactions on Computers, 44 (2):286–291, February 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=364539.

Raghavendra:1996:GCA

[RS96] C. S. Raghavendra and M. A. Sridhar. Global commutative and associative reduction operations in faulty SIMD hypercubes. IEEE Transactions on Computers, 45(4):495–498, April 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=494109.

Reiter:1998:RAU

[RS98] M. K. Reiter and S. G. Stubblebine. Resilient authentication using path independence. IEEE Transactions on Computers, 47(12):1351–1362, December 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=737682.

Rosti:1998:PSS

[RSD^{+}98] E. Rosti, E. Smirni, L. W. Dowdy, G. Serazzi, and K. C.

Sevcik. Processor saving scheduling policies for multiprocessor systems. IEEE Transactions on Computers, 47 (2):178–189, February 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=663764.

Rosenberg:1997:RRP

[RSS97] A. L. Rosenberg, V. Scarano, and R. K. Sitaraman. The reconfigurable ring of processors: fine-grain tree-structured computations. IEEE Transactions on Computers, 46 (10):1119–1131, October 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=628396.

Raghavan:1991:SDC

[RT91a] V. Raghavan and A. Tripathi. Sequential diagnosability is co-NP complete. IEEE Transactions on Computers, 40(5):584–595, May 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=88482.

Raghavan:1991:IDA

[RT91b] V. Raghavan and A. R. Tripathi. Improved diagnosability algorithms. IEEE Transactions on Computers, 40 (2):143–153, February 1991.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=73585.

Rajski:1993:ABC

[RT93a] J. Rajski and J. Tyszer. Accumulator-based compaction of test responses. IEEE Transactions on Computers, 42(6):643– 650, June 1993. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        277285.

Rajski:1993:RPT

[RT93b] J. Rajski and J. Tyszer. Recursive pseudoexhaustive test pattern generation. IEEE Transactions on Computers, 42 (12):1517–1521, December 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=260644.

Rao:1994:PTT

[RT94] N. S. V. Rao and S. Toida. On polynomial-time testable combinational circuits. IEEE Transactions on Computers, 43 (11):1298–1308, November 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=324562.

Rajski:1996:LDS

[RT96] J. Rajski and J. Tyszer. On linear dependencies in subspaces of LFSR-generated sequences. IEEE Transactions on Computers, 45(10):1212–1216, October 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=543715.

Rajski:1999:DSC

[RT99] J. Rajski and J. Tyszer. Diagnosis of scan cells in BIST environment. IEEE Transactions on Computers, 48(7):724–731, July 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=780879.

Rajski:1998:TDD

[RTZ98] J. Rajski, J. Tyszer, and N. Zacharia. Test data decompression for multiple scan designs with boundary scan. IEEE Transactions on Computers, 47(11):1188–1200, November 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=736428.

Reches:1998:IAP

[RW98] S. Reches and S. Weiss. Implementation and analysis of path history in dynamic branch prediction schemes. IEEE Transactions on Computers, 47(8):907–912, August 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=707596.

Raghavendra:1995:FDE

[RYT95] C. S. Raghavendra, Pei-Ji Yang, and Sing-Ban Tien. Free dimensions — an effective approach to achieving fault tolerance in hypercubes. IEEE Transactions on Computers, 44 (9):1152–1157, September 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=464395.

Somani:1992:DDA

[SA92] A. K. Somani and V. K. Agarwal. Distributed diagnosis algorithms for regular interconnected structures. IEEE Transactions on Computers, 41(7):899–906, July 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=256447.

Sundararajan:1998:IMA

[SA98] D. Sundararajan and M. O. Ahmad. Index mapping approach of deriving the PM DFT algorithms. IEEE Transactions on Computers, 47(12): 1418–1424, December 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=737688.

Saab:1995:FRN

[Saa95] Y. G. Saab. A fast and robust network bisection algorithm. IEEE Transactions on Computers, 44(7):903–913, July 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=392848.

Seth:1990:STD

[SAF90] S. C. Seth, V. D. Agrawal, and H. Farhat. A statistical theory of digital circuit testability. IEEE Transactions on Computers, 39(4):582–586, April 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=54854.

Sahni:1996:SMS

[Sah96] S. Sahni. Scheduling masterslave multiprocessor systems. IEEE Transactions on Computers, 45(10):1195–1199, October 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=543712.

Skadron:1999:BPI

[SAMC99] K. Skadron, P. S. Ahuja, M. Martonosi, and D. W. Clark. Branch prediction, instruction-window size, and cache size: performance tradeoffs and simulation techniques. IEEE Transactions on Computers, 48(11):1260–1281, November 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=811115.

Sanchis:1993:MWN

[San93] L. A. Sanchis. Multiple-way network partitioning with different cost functions. IEEE Transactions on Computers, 42 (12):1500–1504, December 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=260640.

Sarrafzadeh:1990:AMT

[Sar90] M. Sarrafzadeh. Area minimization in a three-sided switchbox by sliding the modules. IEEE Transactions on Computers, 39(11):1395–1398, November 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=61050.

Sakaniwa:1991:NUE

[SAR91] K. Sakaniwa, T. N. Ahn, and T. R. N. Rao. A note on tunidirectional error correcting and d (d ≤ t)-unidirectional error detecting (t-UEC and d-UED) codes. IEEE Transactions on Computers, 40(8):987–988, August 1991. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=83642.

Sasao:1991:BAN

[Sas91] T. Sasao. Bounds on the average number of products in the minimum sum-of-products expressions for multiple-value input two-valued output functions. IEEE Transactions on Computers, 40(5):645–651, May 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=88488.

Sasao:1997:ETR

[Sas97] T. Sasao. Easily testable realizations for generalized Reed– Muller expressions. IEEE Transactions on Computers, 46(6): 709–716, June 1997. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
600830.

Savir:1996:RMS

[Sav96] J. Savir. Reducing the MISR size. IEEE Transactions on Computers, 45(8):930–938, August 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=536235.

Savir:1998:RPT

[Sav98a] J. Savir. Random pattern testability of memory control logic.

IEEE Transactions on Computers, 47(3):305–312, March 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=660166.

Savir:1998:STW

[Sav98b] J. Savir. Salvaging test windows in BIST diagnostics. IEEE Transactions on Computers, 47(4):486–491, April 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=675718.

Savir:1999:DGW

[Sav99] J. Savir. Distributed generation of weighted random patterns. IEEE Transactions on Computers, 48(12):1364–1368, December 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=817399.

Sasao:1990:CMS

[SB90] Tsutomu Sasao and Philipp Besslich. On the complexity of mod-2 sum PLA’s. IEEE Transactions on Computers, 39(2):262–266, February 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=45212.

Schueller:1992:DCT

[SB92a] K. A. Schueller and J. T. Butler. On the design of costtables for realizing multiplevalued circuits. IEEE Transactions on Computers, 41 (2):178–189, February 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=123394.

Srinivas:1992:DAG

[SB92b] S. Srinivas and N. N. Biswas.

Design and analysis of a generalized architecture for reconfigurable m-ary tree structures. IEEE Transactions on Computers, 41(11):1465–1478, November 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=177316.

Shallof:1994:CPA

[SB94] A. Shallof and S. Bennett.

Comments on “Performability analysis of distributed realtime-systems”. IEEE Transactions on Computers, 43 (2):255–256, February 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=262135.

Schueller:1997:CAC

[SB97] K. A. Schueller and J. T. Butler. Complexity analysis of the cost-table approach to the design of multiple-valued logic circuits. IEEE Transactions on Computers, 46(2):205–209, February 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=565599.

Strom:1998:CCV

[SBM^{+}98] R. Strom, G. Banavar, K. Miller, A. Prakash, and M. Ward. Concurrency control and view notification algorithms for collaborative replicated objects. IEEE Transactions on Computers, 47(4):458–471, April 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=675714.

Saxena:1990:EVP

[SBP90] S. Saxena, P. C. P. Bhatt, and V. C. Prasad. Efficient VLSI parallel algorithm for Delaunay triangulation on orthogonal tree network in two and three dimensions. IEEE Transactions on Computers, 39(3):400–404, March 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=48871.

Saxena:1991:CPA

[SBP91] S. Saxena, P. C. P. Bhatt, and V. C. Prasad. Correction to “Parallel algorithm for Delaunay triangulation on orthogonal tree network in two and three dimensions”. IEEE Transactions on Computers, 40(1):122, January 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=67331.

Shin:1990:MOL

[SC90a] K. G. Shin and M.-S. Chen.

Minimal order loop-free routing strategy. IEEE Transactions on Computers, 39(7):870–881, July 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=55689.

Shin:1990:NAT

[SC90b] K. G. Shin and M.-S. Chen.

On the number of acceptable task assignments in distributed computing systems. IEEE Transactions on Computers, 39(1):99–110, January 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=46284.

Sugla:1990:EAT

[SC90c] B. Sugla and D. A. Carlson. Extreme area-time tradeoffs in VLSI. IEEE Transactions on Computers, 39 (2):251–257, February 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=45210.

Segee:1994:CFT

[SC94] B. E. Segee and M. J. Carter. Comparative fault tolerance of parallel distributed processing networks. IEEE Transactions on Computers, 43(11): 1323–1329, November 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=324565.

Shin:1995:CLP

[SC95a] K. G. Shin and Yi-Chieh Chang. A coordinated location policy for load sharing in hypercube-connected multicomputers. IEEE Transactions on Computers, 44(5):669–682, May 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=381952.

Shin:1995:RBA

[SC95b] K. G. Shin and Yi-Chieh Chang. A reservation-based algorithm for scheduling both periodic and aperiodic real-time tasks. IEEE Transactions on Computers, 44(12):1405–1419, December 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=477246.

Sasidhar:1996:CDC

[SCC96] K. Sasidhar, S. Chattopadhyay, and P. P. Chaudhuri. CAA decoder for cellular automata based byte error correcting code. IEEE Transactions on Computers, 45(9):1003–1016, September 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=537123.

Saxena:1995:FTF

[SCD^{+}95] N. R. Saxena, C.-W. D. Chang, K. Dawallu, J. Kohli, and P. Helland. Fault-tolerant features in the HaL memory management unit. IEEE Transactions on Computers, 44 (2):170–180, February 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=364529.

Shi:1996:HRR

[SCF96] Weiping Shi, Ming-Feng Chang, and W. K. Fuchs. Harvest rate of reconfigurable pipelines. IEEE Transactions on Computers, 45(10):1200–1203, October 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=543713.

Sun:1994:FTC

[SCG94] Jianli Sun, E. Cerny, and J. Gecsei. Fault tolerance in a class of sorting networks. IEEE Transactions on Computers, 43(7):827–837, July 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=293261.

Schwabe:1992:BLT

[Sch92]  E. J. Schwabe.  A Benes-

like theorem for the shuffleexchange graph. IEEE Transactions on Computers, 41(12): 1627–1630, December 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=214674.

Schneeweiss:1996:NSC

[Sch96] W. G. Schneeweiss. A necessary and sufficient criterion for the monotonicity of Boolean functions with deterministic and stochastic applications. IEEE Transactions on Computers, 45(11):1300–1302, November 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=544486.

Schneeweiss:1998:PFB

[Sch98] W. G. Schneeweiss. On the polynomial form of Boolean functions: derivations and applications. IEEE Transactions on Computers, 47 (2):217–221, February 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=663768.

Sisto:1991:PML

[SCV91] R. Sisto, L. Ciminiera, and A. Valenzano. A protocol for multirendezvous of LOTOS processes. IEEE Transactions on Computers, 40(4):437–447, April 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=88463.

Su:1991:SDF

[SCW91] S. Y. H. Su, M. Cutler, and M. Wang. Self-diagnosis of failures in VLSI tree array processors. IEEE Transactions on Computers, 40(11): 1252–1257, November 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=102828.

Sengupta:1992:SDM

[SD92] A. Sengupta and A. T. Dahbura. On self-diagnosable multiprocessor systems: diagnosis by the comparison approach. IEEE Transactions on Computers, 41(11):1386–1396, November 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=177309.

Shin:1996:AIH

[SD96] K. G. Shin and S. W. Daniel. Analysis and implementation of hybrid switching. IEEE Transactions on Computers, 45(6):684– 692, June 1996. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        506424.

Slye:1998:SSI

[SE98] J. H. Slye and E. N. Elnozahy. Support for software interrupts in log-based rollback-recovery. IEEE Transactions on Computers, 47(10):1113–1123, October 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=729794.

Semal:1995:RBL

[Sem95] P. Semal. Refinable bounds for large Markov chains. IEEE Transactions on Computers, 44 (10):1216–1222, October 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=467696.

Sericola:1999:AAR

[Ser99] B. Sericola. Availability analysis of repairable computer systems and stationarity detection. IEEE Transactions on Computers, 48(11):1166–1172, November 1999. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=811102.

Shrivastava:1992:PFV

[SES^{+}92] S. K. Shrivastava, P. D.

Ezhilchelvan, N. A. Speirs, S. Tao, and A. Tully. Principal features of the VOLTAN family of reliable node architectures for distributed systems. IEEE Transactions on Computers, 41(5):542–549, May 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=142680.

Seznec:1997:DSC

[Sez97]  A. Seznec.  Decoupled sec-

tored caches. IEEE Transactions on Computers, 46 (2):210–215, February 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=565601.

Shang:1991:TOL

[SF91] W. Shang and J. A. B. Fortes.

Time optimal linear schedules for algorithms with uniform dependencies. IEEE Transactions on Computers, 40(6):723–742, June 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=90251.

Shang:1992:IPA

[SF92] W. Shang and J. A. B. Fortes. Independent partitioning of algorithms with uniform dependencies. IEEE Transactions on Computers, 41 (2):190–206, February 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=123395.

Schwarz:1993:PHR

[SF93a] E. M. Schwarz and M. J. Flynn. Parallel high-radix nonrestoring division. IEEE Transactions on Computers, 42 (10):1234–1246, October 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=257709.

Sun:1993:HNN

[SF93b] K. T. Sun and H. C. Fu. A hybrid neural network model for solving optimization problems. IEEE Transactions on Computers, 42(2):218–227, February 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=204794.

Schlag:1996:DMF

[SF96a] M. Schlag and F. J. Ferguson. Detection of multiple faults in two-dimensional ILAs. IEEE Transactions on Computers, 45(6):741–746, June 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=506429.

Schwarz:1996:HSA

[SF96b] Eric M. Schwarz and Michael J.

Flynn. Hardware starting approximation method and its application to the square root operation. IEEE Transactions on Computers, 45(12): 1356–1369, December 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=545966.

Saxena:1992:SBS

[SFM92] N. R. Saxena, P. Franco, and E. J. McCluskey. Simple bounds on serial signature analysis aliasing for random testing. IEEE Transactions on Computers, 41(5):638–645, May 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=142690.

Sam:1990:GMR

[SG90] H. Sam and A. Gupta. A generalized multibit recoding of two’s complement binary numbers and its proof with application in multiplier implementations. IEEE Transactions on Computers, 39(8):1006–1015, August 1990. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=57039.

Silio:1992:AMP

[SGP92] C. B. Silio, Jr., H. M. Ghafir, and M. R. Parikh. An approximate method for the performance analysis of PLAYTHROUGH rings. IEEE Transactions on Computers, 41 (9):1137–1155, September 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=165395.

Shang:1996:UAD

[SHC96] Weijia Shang, E. Hodzic, and Zhigang Chen. On uniformization of affine dependence algorithms. IEEE Transactions on Computers, 45(7):827–840, July 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=508321.

Shen:1995:OAPb

[She95a]  Xiaojun Shen.  An opti-

mal O(N lg N) algorithm for permutation admissibility to extra-stage cube-type networks. IEEE Transactions on Computers, 44(9):1144–1149, September 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=464393.

Shen:1995:ORB

[She95b] Xiaojun Shen. Optimal realization of any BPC permutation on K-extra-stage Omega networks. IEEE Transactions on Computers, 44(5):714–719, May 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=381960.

Shen:1993:DPT

[SHHJ93] W.-Z. Shen, G.-H. Hwang, W.-J. Hsu, and Y.-J. Jan. Design of pseudoexhaustive testable PLA with low overhead. IEEE Transactions on Computers, 42(7):887–891, July 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=237730.

Saitoh:1993:SCC

[SI93] Y. Saitoh and H. Imai. Some codes for correcting and detecting unidirectional byte errors. IEEE Transactions on Computers, 42(5):547–552, May 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=223673.

Soneoka:1994:LIM

[SI94] T. Soneoka and T. Ibaraki.

Logically instantaneous message passing in asynchronous distributed systems. IEEE Transactions on Computers, 43(5):513– 527, May 1994. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
280800.

Shintani:1995:PCA

[SIKS95] Y. Shintani, K. Inoue, E. Kamada, and T. Shonai. A performance and cost analysis of applying superscalar method to mainframe computers. IEEE Transactions on Computers, 44(7):891–902, July 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=392847.

Sasaki:1990:FMI

[SISU90] M. Sasaki, T. Inoue, Y. Shirai, and F. Ueno. Fuzzy multiple-input maximum and minimum circuits in current mode and their analyses using bounded-difference equations. IEEE Transactions on Computers, 39(6):768–774, June 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=53598.

Sapiecha:1990:MAH

[SJ90] K. Sapiecha and R. Jarocki. Modular architecture for high performance implementation of the FRR algorithm. IEEE Transactions on Computers, 39 (12):1464–1468, December 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=61066.

Sitaraman:1993:ODC

[SJ93a] R. K. Sitaraman and N. K.

Jha. Optimal design of checks for error detection and location in fault-tolerant multiprocessor systems. IEEE Transactions on Computers, 42(7):780–793, July 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=237719.

Spray:1993:PTR

[SJ93b] A. Spray and S. Jones. Performance tradeoffs in rings of data-driven elements. IEEE Transactions on Computers, 42(1):113–118, January 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=192221.

Sammut:1995:AUD

[SJ95] K. M. Sammut and S. R.

Jones. Arithmetic unit design for neural accelerators: cost performance issues. IEEE Transactions on Computers, 44 (10):1256–1260, October 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=467702.

Sengupta:1991:SAD

[SJB91] A. Sengupta, P. D. Joshi, and S. Bandyopadhyay. A synthesis approach to design optimally fault tolerant network architecture. IEEE Transactions on Computers, 40(1):94–100, January 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=67324.

Smith:1996:SDE

[SJP96] D. T. Smith, B. W. Johnson, and J. A. Profeta III. System dependability evaluation via a fault list generation algorithm. IEEE Transactions on Computers, 45(8):974–979, August 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=536240.

Sayah:1992:TSH

[SK92a] J. Y. Sayah and C. R. Kime. Test scheduling in high performance VLSI system implementations. IEEE Transactions on Computers, 41(1):52–67, January 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=123376.

Stevens:1992:DGS

[SK92b] R. S. Stevens and D. J. Kaplan. Determinacy of generalized schema. IEEE Transactions on Computers, 41(6):776–779, June 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=144630.

Shin:1994:TRA

[SK94] K. G. Shin and Hagbae Kim.

A time redundancy approach to TMR failures using faultstate likelihoods. IEEE Transactions on Computers, 43 (10):1151–1162, October 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=324541.

Singh:1996:EDC

[SK96] A. D. Singh and C. M. Krishna. On the effect of defect clustering on test transparency and IC test optimization. IEEE Transactions on Computers, 45(6):753–757, June 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=506431.

Sunar:1999:MMA

[SK99] B. Sunar and Ç. K. Koç. Mastrovito multiplier for all trinomials. IEEE Transactions on Computers, 48(5):522–527, May 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=769434.

Scherson:1992:BPA

[SKA92] I. D. Scherson, D. A. Kramer, and B. D. Alleyne. Bit-parallel arithmetic in a massivelyparallel associative processor. IEEE Transactions on Computers, 41(10):1201–1210, October 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=166599.

Satyanarayanan:1990:CHA

[SKK^{+}90] M. Satyanarayanan, J. J. Kistler, P. Kumar, M. E. Okasaki, E. H. Siegel, and D. C. Steere. Coda: a highly available file system for a distributed workstation environment. IEEE Transactions on Computers, 39(4):447–459, April 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=54838.

Shyu:1990:PAS

[SL90a] S.-C. Shyu and V. O. K. Li. Performance analysis of static locking in distributed database systems. IEEE Transactions on Computers, 39(6):741–751, June 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=53595.

Su:1990:ASS

[SL90b] C.-C. Su and H.-Y. Lo. An algorithm for scaling and single residue error correction in residue number systems. IEEE Transactions on Computers, 39(8):1053–1064, August 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=57044;
https://ieeexplore.ieee.
org/xpl/tocresult.jsp?isnumber=
2065.

Sarrafzadeh:1991:TMR

[SL91a] M. Sarrafzadeh and D. T.

Lee. Topological via minimization revisited. IEEE Transactions on Computers, 40(11): 1307–1312, November 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=102839.

Sips:1991:IVR

[SL91b] H. J. Sips and H. Lin. An improved vector-reduction method. IEEE Transactions on Computers, 40(2):214–217, February 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=73591.

Smith:1991:TMC

[SL91c] D. R. Smith and J. C. Lin. The tree-match chip. IEEE Transactions on Computers, 40(5):629– 639, May 1991. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
88486.

Stapper:1992:SFT

[SL92] C. H. Stapper and H.-S. Lee. Synergistic fault-tolerance for memory chips. IEEE Transactions on Computers, 41(9): 1078–1087, September 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=165390.

Shih:1995:ASI

[SL95] Wei-Kuan Shih and J. W. S. Liu. Algorithms for scheduling imprecise computations with timing constraints to minimize maximum error. IEEE Transactions on Computers, 44(3):466–471, March 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=372040.

Sheu:1995:DFD

[SLD95] Tsang-Ling Sheu, Woei Lin, and C. R. Das. Distributed fault diagnosis in multistage network-based multiprocessors. IEEE Transactions on Computers, 44(9):1085–1095, September 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=464387.

Shen:1997:EBM

[SLH97] Xiaojun Shen, Weifa Liang, and Qing Hu. On embedding between 2D meshes of the same size. IEEE Transactions on Computers, 46(8):880–889, August 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=609277.

Shayan:1997:CSV

[SLN97] Y. R. Shayan and T. Le-Ngoc. A cellular structure for a versatile Reed–Solomon decoder. IEEE Transactions on Computers, 46(1):80–85, January 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=559805.

Strosnider:1995:DSA

[SLS95] J. K. Strosnider, J. P. Lehoczky, and Lui Sha. The deferrable server algorithm for enhanced aperiodic responsiveness in hard real-time environments. IEEE Transactions on Computers, 44(1):73–91, January 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=368008.

Sarrafzadeh:1998:FST

[SLW98] M. Sarrafzadeh, Wei-Liang Lin, and C. K. Wong. Floating Steiner trees. IEEE Transactions on Computers, 47 (2):197–211, February 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=663766.

Saxena:1990:ACE

[SM90a] N. R. Saxena and E. J. Mc-Cluskey. Analysis of checksums, extended-precision checksums, and cyclic redundancy checks. IEEE Transactions on Computers, 39(7):969–975, July 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=55701.

Saxena:1990:CFC

[SM90b] N. R. Saxena and E. J. Mc-Cluskey. Control-flow checking using watchdog assists and extended-precision checksums. IEEE Transactions on Computers, 39(4):554–559, April 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=54849.

Sarkar:1992:DOS

[SM92a] D. Sarkar and A. Mukherjee. Design of optimal systolic algorithms for the transitive closure problem. IEEE Transactions on Computers, 41(4):508–512, April 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=135564.

Savir:1992:MSL

[SM92b] J. Savir and W. H. McAnney. A multiple seed linear feedback shift register. IEEE Transactions on Computers, 41(2):250–252, February 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=123404.

Sakar:1993:GPN

[SM93] A. Sakar and R. J. Mammone.

Growing and pruning neural tree networks. IEEE Transactions on Computers, 42(3):291–299, March 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=210172.

Saxena:1996:CTS

[SM96a] N. R. Saxena and E. J.

McCluskey. Counting twostate transition-tour sequences. IEEE Transactions on Computers, 45(11):1337–1342, November 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=544493.

Smith:1996:GMV

[SM96b] M. D. Smith and P. Mazumder. Generation of minimal vertex covers for row/column allocation in self-repairable arrays. IEEE Transactions on Computers, 45(1):109–115, January 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=481492.

Saxena:1997:PSA

[SM97] N. R. Saxena and E. J. Mc-Cluskey. Parallel signature analysis design with bounds on aliasing. IEEE Transactions on Computers, 46(4):425–438, April 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=588057.

See [NAH98].

Saltz:1991:RTP

[SMC91] J. H. Saltz, R. Mirchandaney, and K. Crowley. Run-time parallelization and scheduling of loops. IEEE Transactions on Computers, 40(5):603–612, May 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=88484.

Smith:1995:CFA

[Smi95] Roger Alan Smith. A continuedfraction analysis of trigonometric argument reduction.

IEEE Transactions on Computers, 44(11):1348–1351, November 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=475133.

Stelling:1998:OCP

[SMOR98] P. F. Stelling, C. U. Martel, V. G. Oklobdzija, and R. Ravi. Optimal circuits for parallel multipliers. IEEE Transactions on Computers, 47(3):273–285, March 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=660163.

Savir:1991:TCC

[SMV91] J. Savir, W. H. McAnney, and S. R. Vecchio. Testing for coupled cells in randomaccess memories. IEEE Transactions on Computers, 40 (10):1177–1180, October 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=93752.

Savva:1999:GDM

[SN99] A. Savva and T. Nanya. A gracefully degrading massively parallel system using the BSP model, and its evaluation. IEEE Transactions on Computers, 48(1):38–52, January 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=743410.

Sohi:1990:IIL

[Soh90] G. S. Sohi. Instruction issue logic for high-performance, interruptible, multiple functional unit, pipelined computers. IEEE Transactions on Computers, 39(3):349–359, March 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=48865.

Sohi:1993:HBI

[Soh93] G. S. Sohi. High-bandwidth interleaved memories for vector processors — a simulation study. IEEE Transactions on Computers, 42(1):34–44, January 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=192212.

Somani:1990:SFO

[Som90]  A. K. Somani.  Sequential

fault occurrence and reconfiguration in system level diagnosis. IEEE Transactions on Computers, 39(12):1472–1475, December 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=61067.

Samatham:1991:CBM

[SP91] M. R. Samatham and D. K. Pradhan. Correction to “The De Bruijn multiprocessor network: a versatile parallel processing and sorting network for VLSI”. IEEE Transactions on Computers, 40(1):122, January 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=67330.

Srinivas:1995:FRD

[SP95] H. R. Srinivas and K. K.

Parhi. A fast radix-4 division algorithm and its architecture. IEEE Transactions on Computers, 44(6):826–831, June 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=391179.

Somani:1996:DLF

[SP96] A. K. Somani and O. Peleg. On diagnosability of large fault sets in regular topology-based computer systems. IEEE Transactions on Computers, 45(8):892–903, August 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=536232.

Srinivas:1997:RDR

[SPM97] H. R. Srinivas, K. K. Parhi, and L. A. Montalvo. Radix 2 division with over-redundant quotient selection. IEEE Transactions on Computers, 46(1):85–92, January 1997. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=559806.

Sridhar:1991:FTN

[SR91] M. A. Sridhar and C. S. Raghavendra. Fault-tolerant networks based on the de Bruijn graph. IEEE Transactions on Computers, 40 (10):1167–1174, October 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=93750.

Skavantzos:1992:NMM

[SR92] A. Skavantzos and P. B. Rao. New multipliers modulo 2^{N} − 1. IEEE Transactions on Computers, 41(8):957–961, August 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=156538.

Stephens:1993:SFS

[SR93] J. Stephens and V. Raghavan. On single-fault set diagnosability in the PMC model. IEEE Transactions on Computers, 42(8):981–983, August 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=238489.

Sung:1996:ABF

[SR96] Jan-Lung Sung and G. R. Redinbo. Algorithm based fault tolerant synthesis for linear operations. IEEE Transactions on Computers, 45(4):425–438, April 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=494100.

Sastry:1998:VAA

[SR98] R. Sastry and N. Ranganathan.

A VLSI architecture for approximate tree matching. IEEE Transactions on Computers, 47(3):346–352, March 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=660171.

Stott:1998:DAH

[SRHI98] D. T. Stott, G. Ries, Mei-Chen Hsueh, and R. K. Iyer. Dependability analysis of a highspeed network using softwareimplemented fault injection and simulated fault injection. IEEE Transactions on Computers, 47(1):108–119, January 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=656094.

Srikant:1990:PPA

[Sri90] Y. N. Srikant. Parallel parsing of arithmetic expressions.

IEEE Transactions on Computers, 39(1):130–132, January 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=46288.

Siu:1991:DST

[SRK91] K. Y. Siu, V. P. Roychowdhury, and T. Kailath. Depth-size tradeoffs for neural computation. IEEE Transactions on Computers, 40(12):1402–1412, December 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=106225.

Sha:1990:PIP

[SRL90] L. Sha, R. Rajkumar, and J. P. Lehoczky. Priority inheritance protocols: an approach to real-time synchronization. IEEE Transactions on Computers, 39(9):1175–1185, September 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=57058.

Sha:1991:RTL

[SRSC91] L. Sha, R. Rajkumar, S. H. Son, and C.-H. Chang. A real-time locking protocol. IEEE Transactions on Computers, 40(7):793– 800, July 1991. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
83617.

Sack:1990:OAD

[SS90a] J.-R. Sack and S. Suri. An optimal algorithm for detecting weak visibility of a polygon. IEEE Transactions on Computers, 39(10):1213–1219, October 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=59852.

Satyanarayanan:1990:PCL

[SS90b] M. Satyanarayanan and E. H.

Siegel. Parallel communication in a large distributed environment. IEEE Transactions on Computers, 39(3):328–348, March 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=48864.

Silberman:1991:FFS

[SS91a] G. M. Silberman and I. Spillinger.

Functional fault simulation as a guide for biased-random test pattern generation. IEEE Transactions on Computers, 40(1):66–79, January 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=67321.

Silberman:1991:RFT

[SS91b] G. M. Silberman and I. Spillinger.

RIDDLE: a foundation for test generation on a high-level design description. IEEE Transactions on Computers, 40(1):80–87, January 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=67322.

Skavantzos:1992:DCM

[SS92] A. Skavantzos and T. Stouraitis. Decomposition of complex multipliers using polynomial encoding. IEEE Transactions on Computers, 41(10):1331–1333, October 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=166610.

See comments [Kat94].

Sasinowski:1993:DPA

[SS93a] J. E. Sasinowski and J. K. Strosnider. A dynamic programming algorithm for cache memory partitioning for real-time systems. IEEE Transactions on Computers, 42(8):997–1001, August 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=238493.

Sha:1993:RRS

[SS93b] E. H.-M. Sha and K. Steiglitz. Reconfigurability and reliability of systolic/wavefront arrays. IEEE Transactions on Computers, 42(7):854–862, July 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=237725.

Schuette:1994:EIL

[SS94a] M. A. Schuette and J. P.

Shen. Exploiting instructionlevel parallelism for integrated control-flow monitoring. IEEE Transactions on Computers, 43 (2):129–140, February 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=262118.

Schulte:1994:HDE

[SS94b] M. J. Schulte and E. E. Swartzlander, Jr. Hardware designs for exactly rounded elementary functions. IEEE Transactions on Computers, 43(8):964–973, August 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=295858.

Spuri:1994:HIP

[SS94c] M. Spuri and J. A. Stankovic.

How to integrate precedence constraints and shared resources in real-time scheduling. IEEE Transactions on Computers, 43(12):1407–1412, December 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=338100.

Saavedra:1995:MCT

[SS95] R. H. Saavedra and A. J. Smith. Measuring cache and TLB performance and their effect on benchmark runtimes. IEEE Transactions on Computers, 44(10):1223–1235, October 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=467697.

Su:1996:AFT

[SS96] Chien-Chun Su and K. G. Shin. Adaptive fault-tolerant deadlock-free routing in meshes and hypercubes. IEEE Transactions on Computers, 45(6):666– 683, June 1996. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        506423.

Schulte:1999:AEF

[SS99] M. J. Schulte and J. E. Stine. Approximating elementary functions with symmetric bipartite tables. IEEE Transactions on Computers, 48(8):842–847, August 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=795125;
                                         http://mesa.ece.wisc.edu/
                                        publications/cp_1999-10.pdf.

Sibai:1994:CRS

[SSA94] F. N. Sibai, N. K. Sharma, and A. A. Abonamah. Comparison of reconfiguration schemes for the C2SC MIN operating in the broadcast mode. IEEE Transactions on Computers, 43(9):1112–1119, September 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=312121.

Soufi:1995:PRI

[SSDK95] M. Soufi, Y. Savaria, F. Darlay, and B. Kaminska. Producing reliable initialization and test of sequential circuits with pseudorandom vectors. IEEE Transactions on Computers, 44(10):1251–1256, October 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=467701.

Sanchez:1999:SDC

[SSH^{+}99] E. Sanchez, M. Sipper, J.-O.

Haenni, J.-L. Beuchat, A. Stauffer, and A. Perez-Uribe. Static and dynamic configurable systems. IEEE Transactions on Computers, 48(6):556–564, June 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=773792.

Shintani:1996:HES

[SSK^{+}96] Y. Shintani, T. Shonai, H. Kurokawa, K. Kuriyama, and A. Yamaoka. Hierarchical execution to speed up pipeline interlock in mainframe computers. IEEE Transactions on Computers, 45(5):589–599, May 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=509910.

Salinas:1996:SLA

[SSL96] J. Salinas, Yinan Shen, and F. Lombardi. A sweeping line approach to interconnect testing. IEEE Transactions on Computers, 45(8):917–929, August 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=536234.

Sha:1997:ADL

[SSS97] L. Sha, S. S. Sathaye, and J. K. Strosnider. Analysis of duallink networks for real-time applications. IEEE Transactions on Computers, 46(1):1–13, January 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=559798.

Singh:1992:MWU

[SST92] J. P. Singh, H. S. Stone, and D. F. Thiebaut. A model of workloads and its use in missrate prediction for fully associative caches. IEEE Transactions on Computers, 41(7):811–825, July 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=256450.

Schlichting:1995:PLS

[ST95a] R. D. Schlichting and V. T.

Thomas. Programming language support for writing faulttolerant distributed software. IEEE Transactions on Computers, 44(2):203–212, February 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=364532.

Smith:1995:FTC

[ST95b] J. C. Smith and F. J. Taylor.

A fault-tolerant CEQRNS processing element for linear systolic array DSP applications. IEEE Transactions on Computers, 44(9):1121–1130, September 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=464390.

Symvonis:1996:ESL

[ST96] A. Symvonis and J. Tidswell.

An empirical study of offline permutation packet routing on two-dimensional meshes based on the multistage routing method. IEEE Transactions on Computers, 45(5):619–625, May 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=509915.

Stapper:1993:IYM

[Sta93] C. H. Stapper. Improved yield models for fault-tolerant memory chips. IEEE Transactions on Computers, 42(7):872–881, July 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=237727.

Stan:1998:LFU

[STE98] M. R. Stan, A. F. Tenca, and M. D. Ercegovac. Long and fast up/down counters. IEEE Transactions on Computers, 47(7):722–735, July 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=709372.

Stouraitis:1993:BFT

[Sto93]  T. Stouraitis.  Borrow: a

fault-tolerance scheme for wavefront array processors. IEEE Transactions on Computers, 42 (10):1257–1261, October 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=257711.

Stone:1992:OPC

[STW92] H. S. Stone, J. Turek, and J. L. Wolf. Optimal partitioning of cache memory. IEEE Transactions on Computers, 41(9):1054–1068, September 1992. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=165388.

Suter:1991:CNV

[Sut91]  D. Suter.  Constraint net-

works in vision. IEEE Transactions on Computers, 40(12): 1359–1367, December 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=106221.

Swaminathan:1994:DM

[SV94] R. Swaminathan and D. Veeramani. Decomposition of {0, 1}- matrices. IEEE Transactions on Computers, 43(5):629–633, May 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=280812.

Sisto:1995:MPN

[SV95] R. Sisto and A. Valenzano.

Mapping Petri nets with inhibitor arcs onto basic LOTOS behavior expressions. IEEE Transactions on Computers, 44 (12):1361–1370, December 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=477242.

Stiliadis:1997:SVC

[SV97] D. Stiliadis and A. Varma.

Selective victim caching: a method to improve the performance of direct-mapped caches. IEEE Transactions on Computers, 46(5):603–610, May 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=589235.

Srbljic:1997:APP

[SVSB97] S. Srbljic, Z. G. Vranesic, M. Stumm, and L. Budin. Analytical prediction of performance for cache coherence protocols. IEEE Transactions on Computers, 46(11):1155–1173, November 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=644291.

Schwartz:1990:SCC

[SW90] J. W. Schwartz and J. K. Wolf. A systematic (12, 8) code for correcting single errors and detecting adjacent errors. IEEE Transactions on Computers, 39(11):1403–1404, November 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=61052.

Sarrafzadeh:1992:BST

[SW92] M. Sarrafzadeh and C. K. Wong. Bottleneck Steiner trees in the plane. IEEE Transactions on Computers, 41(3):370–374, March 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=127452.

Su:1994:TIL

[SW94a] Chih-Yuang Su and Cheng-Wen Wu. Testing iterative logic arrays for sequential faults with a constant number of patterns. IEEE Transactions on Computers, 43(4):495–501, April 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=278489.

Sun:1994:CDA

[SW94b] C.-H. Sun and S.-D. Wang.

Comments on “Distributed algorithms for network recognition problems”. IEEE Transactions on Computers, 43(6):768, June 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=286312.

Skolleborg:1995:CSC

[SW95] A. G. Skolleborg and J. H.

Weber. Construction of systematic codes for unidirectional error control. IEEE Transactions on Computers, 44(11): 1331–1334, November 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=475139.

Sui:1997:IAF

[SW97] Pao-Hwa Sui and Sheng-De Wang. An improved algorithm for fault-tolerant wormhole routing in meshes. IEEE Transactions on Computers, 46 (9):1040–1042, September 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=620485.

Shatz:1992:TAM

[SWG92] S. M. Shatz, J.-P. Wang, and M. Goto. Task allocation for maximizing reliability of distributed computer systems. IEEE Transactions on Computers, 41(9):1156–1168, September 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=165396.

Shih:1990:UMC

[SWK90] Wei-Kuan Shih, Sun Wu, and Y. S. Kuo. Unifying maximum cut and minimum cut of a planar graph. IEEE Transactions on Computers, 39(5):694–697, May 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=53581.

Sullivan:1995:CCR

[SWM95] G. F. Sullivan, D. S. Wilson, and G. M. Masson. Certification of computational results. IEEE Transactions on Computers, 44(7):833–847, July 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=392843.

Saha:1993:ADR

[SWT93] A. Saha, Chuan-Lin Wu, and Dun-Sung Tang. Approximation, dimension reduction, and nonconvex optimization using linear superpositions of Gaussians. IEEE Transactions on Computers, 42(10):1222–1233, October 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=257708.

Sarrafzadeh:1994:WKK

[SWWW94] M. Sarrafzadeh, D. Wagner, F. Wagner, and K. Weihe. Wiring knock-knee layouts: a global approach. IEEE Transactions on Computers, 43(5):581– 589, May 1994. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
280805.

Shen:1995:OAPa

[SXW95] Xiaojun Shen, Mao Xu, and Xiangzu Wang. An optimal algorithm for permutation admissibility to multistage interconnection networks. IEEE Transactions on Computers, 44(4):604–608, April 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=376176.

Salem:1991:NNA

[SY91a] G. J. Salem and T. Y. Young. A neural network approach to the labeling of line drawings. IEEE Transactions on Computers, 40(12):1419–1424, December 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=106227.

Singh:1991:MFT

[SY91b] A. D. Singh and H. Y. Youn. A modular fault-tolerant binary tree architecture with short links. IEEE Transactions on Computers, 40(7):882–890, July 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=83628.

Sun:1999:CAC

[SY99] Tong Sun and Qing Yang. A comparative analysis of cache designs for vector processing. IEEE Transactions on Computers, 48(3):331–344, March 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=754999.

Sun:1992:ETS

[SZN92] X.-H. Sun, H. Zhang, and L. M. Ni. Efficient tridiagonal solvers on multicomputers. IEEE Transactions on Computers, 41(3):286–296, March 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=127441.

Shiratori:1991:UFS

[SZTN91] N. Shiratori, Y.-X. Zhang, K. Takahashi, and S. Noguchi. A user friendly software environment for protocol synthesis. IEEE Transactions on Computers, 40(4):477–486, April 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=88467.

Szymanski:1997:DPP

[Szy97] T. H. Szymanski. Design principles for practical self-routing nonblocking switching networks with O(N ·log N) bit-complexity. IEEE Transactions on Computers, 46(10):1057–1069, October 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=628391.

Tang:1995:CCP

[TA95] K. W. Tang and B. W. Arden. Class-congruence property and two-phase routing of Borel Cayley graphs. IEEE Transactions on Computers, 44 (12):1462–1468, December 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=477252.

Takagi:1992:RMM

[Tak92a] N. Takagi. A radix-4 modular multiplication hardware algorithm for modular exponentiation. IEEE Transactions on Computers, 41(8):949–956, August 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=156537.

Takesue:1992:CMD

[Tak92b] M. Takesue. Cache memories for data flow machines. IEEE Transactions on Computers, 41(6):677–687, June 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=144620.

Takagi:1998:PTL

[Tak98] N. Takagi. Powering by a table look-up and a multiplication with operand modification. IEEE Transactions on Computers, 47(11):1216–1222, November 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=736432.

Tang:1992:TDI

[Tan92] Ping Tak Peter Tang. Tabledriven implementation of the Expm1 function in IEEE floatingpoint arithmetic. ACM Transactions on Mathematical Software, 18(2):211–222, June 1992. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://

www.acm.org/pubs/citations/
journals/toms/1992-18-2/p211-

tang/. See independent analysis and accuracy confirmation of this algorithm in [Krä98].

Tao:1997:STN

[Tao97]  D. L. Tao.  A self-testing

nonincreasing order checker. IEEE Transactions on Computers, 46(7):817–820, July 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=599902.

Tate:1997:BOL

[Tat97] S. R. Tate. Band ordering in lossless compression of multispectral images. IEEE Transactions on Computers, 46(4):477–483, April 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=588062.

Takagi:1991:RCM

[TAY91] N. Takagi, T. Asada, and S. Yajima. Redundant CORDIC methods with a constant scale factor for sine and cosine computation. IEEE Transactions on Computers, 40 (9):989–995, September 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=83660.

Tirumalai:1991:MAM

[TB91] P. P. Tirumalai and J. T. Butler. Minimization algorithms for multiple-valued programmable logic arrays. IEEE Transactions on Computers, 40 (2):167–177, February 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=73587.

Tallini:1998:DBC

[TB98a] L. G. Tallini and B. Bose. Design of balanced and constant weight codes for VLSI systems. IEEE Transactions on Computers, 47(5):556–572, May 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=677239.

Tallini:1998:TDA

[TB98b] L. G. Tallini and B. Bose. Theory and design of adjacent asymmetric error masking codes. IEEE Transactions on Computers, 47(5):544–555, May 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=677235.

Tallini:1999:BCP

[TB99a] L. G. Tallini and B. Bose.

Balanced codes with parallel encoding and decoding. IEEE Transactions on Computers, 48(8):794–814, August 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=795122.

Thibeault:1999:SBC

[TB99b] C. Thibeault and G. Begin.

A scan-based configurable, programmable, and scalable architecture for sliding window-based operations. IEEE Transactions on Computers, 48(6):615–627, June 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=773798.

Tzeng:1994:STE

[TC94] Nian-Feng Tzeng and Hsing-Lung Chen. Structural and tree embedding aspects of incomplete hypercubes. IEEE Transactions on Computers, 43 (12):1434–1439, December 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=338105.

Tsay:1995:DSO

[TC95] Jong-Chuang Tsay and Pen-Yuang Chang. Design of spaceoptimal regular arrays for algorithms with linear schedules. IEEE Transactions on Computers, 44(5):683–694, May 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=381953.

Tang:1998:LMV

[TCI98] Zheng Tang, Qi-Ping Cao, and O. Ishizuka. A learning multiplevalued logic network: algebra, algorithm, and applications. IEEE Transactions on Computers, 47(2):247–251, February 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=663773.

Tagliarini:1991:OUN

[TCP91] G. A. Tagliarini, J. F. Christ, and E. W. Page. Optimization using neural networks. IEEE Transactions on Computers, 40(12):1347–1358, December 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=106220.

Tzeng:1993:CDP

[TCW93] Nian-Feng Tzeng, Po-Jen Chuang, and Chwan-Hwa Wu. Creating disjoint paths in Gamma interconnection networks. IEEE Transactions on Computers, 42 (10):1247–1252, October 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=257710.

Torng:1993:IHO

[TD93] H. C. Torng and M. Day.

Interrupt handling for outof-order execution processors. IEEE Transactions on Computers, 42(1):122–127, January 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=192223.

Temam:1999:AOE

[Tem99]  O. Temam.  An algorithm

for optimally exploiting spatial and temporal locality in upper memory levels. IEEE Transactions on Computers, 48 (2):150–158, February 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=752656.

Tamir:1992:DAM

[TF92] Y. Tamir and G. L. Frazier.

Dynamically-allocated multiqueue buffers for VLSI communication switches. IEEE Transactions on Computers, 41(6):725– 737, June 1992. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

                                        stamp/stamp.jsp?tp=&arnumber=
                                        144624.

Topham:1999:RCP

[TG99] N. Topham and A. Gonzalez. Randomized cache placement for eliminating conflicts. IEEE Transactions on Computers, 48(2):185–192, February 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=752660.

Toborg:1991:CVI

[TH91] S. T. Toborg and K. Hwang. Cooperative vision integration through data-parallel neural computations. IEEE Transactions on Computers, 40(12): 1368–1379, December 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=106222.

Takagi:1999:HSR

[TH99] N. Takagi and T. Horiyama. A high-speed reduced-size adder under left-to-right input arrival. IEEE Transactions on Computers, 48(1):76–80, January 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=743413.

Tsai:1999:SPA

[THA^{+}99] Jenn-Yuan Tsai, Jian Huang, C. Amlo, D. J. Lilja, and Pen-Chung Yew. The superthreaded processor architecture. IEEE Transactions on Computers, 48 (9):881–902, September 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=795219.

Timmermann:1992:LLT

[THH92] D. Timmermann, H. Hahn, and B. J. Hosticka. Low latency time CORDIC algorithms. IEEE Transactions on Computers, 41(8):1010–1015, August 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=156543.

Tao:1991:NED

[THL91] D. L. Tao, C. R. P. Harmann, and P. K. Lala. A note on t-EC/d-UED codes. IEEE Transactions on Computers, 40(5):660–663, May 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=88490.

Tao:1992:GTD

[THL92] D. L. Tao, C. R. P. Hartmann, and P. K. Lala. A general technique for designing totally self-checking checker for 1out-of-N code with minimum gate delay. IEEE Transactions on Computers, 41(7):881–886, July 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=256456.

Thornton:1997:SBA

[Tho97] M. A. Thornton. Signed binary addition circuitry with inherent even parity outputs. IEEE Transactions on Computers, 46(7):811–816, July 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=599901.

Tien:1993:BIH

[THY93] Jenn-Yang Tien, Ching-Tien Ho, and Wei-Pang Yang. Broadcasting on incomplete hypercubes. IEEE Transactions on Computers, 42(11):1393–1398, November 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=247842.

Tsai:1999:SBP

[THZ^{+}99] T. K. Tsai, Mei-Chen Hsueh, Hong Zhao, Z. Kalbarczyk, and R. K. Iyer. Stress-based and path-based fault injection. IEEE Transactions on Computers, 48(11):1183–1201, November 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=811108.

Tang:1992:AMC

[TI92] D. Tang and R. K. Iyer. Analysis and modeling of correlated failures in multicomputer systems. IEEE Transactions on Computers, 41(5):567–577, May 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=142683.

Tang:1993:DMM

[TI93] D. Tang and R. K. Iyer. Dependability measurement and modeling of a multicomputer system. IEEE Transactions on Computers, 42(1):62–75, January 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=192214.

Tong:1991:VAW

[TK91] Z. Tong and R. Y. Kain. Vote assignments in weighted voting mechanisms. IEEE Transactions on Computers, 40(5):664–667, May 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=88491.

Tohma:1996:FTD

[TK96] Y. Tohma and Y. Koyanagi.

Fault-tolerant design of neural networks for solving optimization problems. IEEE Transactions on Computers, 45 (12):1450–1455, December 1996.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=545976.

Tao:1997:ERI

[TK97] D. L. Tao and K. Kantawala. Evaluating reliability improvements of fault tolerant array processors using algorithm-based fault tolerance. IEEE Transactions on Computers, 46(6):725– 730, June 1997. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
600889.

Thitikamol:1998:PNM

[TK98] K. Thitikamol and P. Keleher. Per-node multithreading and remote latency. IEEE Transactions on Computers, 47(4):414–426, April 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=675711.

Tinder:1994:HSM

[TKS94] R. F. Tinder, R. I. Klaus, and J. A. Snodderley. Highspeed microprogrammable asynchronous controller modules. IEEE Transactions on Computers, 43(10):1226–1232, October 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=324548.

Tzeng:1996:EDM

[TL96] Nian-Feng Tzeng and Guanghua Lin. Efficient determination of maximum incomplete subcubes in hypercubes with faults. IEEE Transactions on Computers, 45(11):1303–1308, November 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=544487.

Tse:1999:SRI

[TL99] S. S. H. Tse and F. C. M.

Lau. On the space requirement of interval routing. IEEE Transactions on Computers, 48(7):752–757, July 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=780884.

Tabrizi:1997:TMG

[TLE97] N. Tabrizi, M. J. Liebelt, and K. Eshraghian. A tabular method for guard strengthening, symmetrization, and operator reduction for Martin’s asynchronous design methodology. IEEE Transactions on Computers, 46(9):1050–1054, September 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=620487.

Torrellas:1994:FSS

[TLH94] J. Torrellas, H. S. Lam, and J. L. Hennessy. False sharing and spatial locality in multiprocessor caches. IEEE Transactions on Computers, 43(6):651–663, June 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=286299.

Trabado:1993:SSE

[TLROL93] P. P. Trabado, A. Lloris-Ruiz, and J. Ortega-Lopera. Solution of switching equations based on a tabular algebra. IEEE Transactions on Computers, 42(5):591–596, May 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=223678.

See additions [Ung94].

Tsai:1996:GRM

[TMS96] Chien-Chung Tsai and M. Marek-Sadowska. Generalized Reed– Muller forms as a tool to detect symmetries. IEEE Transactions on Computers, 45(1):33–40, January 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=481484.

Tsai:1997:BFC

[TMS97] Chien-Chung Tsai and M. Marek-Sadowska. Boolean functions classification via fixed polarity Reed–Muller forms. IEEE Transactions on Computers, 46(2):173–186, February 1997. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=565592.

Thomasian:1993:PET

[TN93] A. Thomasian and V. F.

Nicola. Performance evaluation of a threshold policy for scheduling readers and writers. IEEE Transactions on Computers, 42(1):83–98, January 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=192216.

Turner:1995:CFM

[TNC95] S. W. Turner, L. M. Ni, and B. H. C. Cheng. Contention-free 2D-mesh cluster allocation in hypercubes. IEEE Transactions on Computers, 44(8):1051–1055, August 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=403722.

Tseng:1999:TOC

[TNS99] Yu-Chee Tseng, Sze-Yao Ni, and Jang-Ping Sheu. Toward optimal complete exchange on wormhole-routed tori. IEEE Transactions on Computers, 48(10):1065–1082, October 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=805156.

Tokarnia:1994:IMS

[Tok94] A. M. Tokarnia. Identifying minimal shift counters: a search technique. IEEE Transactions on Computers, 43(5):633–639, May 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=280813.

Tang:1994:DTM

[TP94] K. W. Tang and S. A. Padubidri. Diagonal and toroidal mesh networks. IEEE Transactions on Computers, 43(7):815–826, July 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=293260.

Truman:1998:IMT

[TPDB98] T. E. Truman, T. Pering, R. Doering, and R. W. Brodersen. The InfoPad multimedia terminal: a portable device for wireless information access. IEEE Transactions on Computers, 47(10):1073–1087, October 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=729791.

Tragoudas:1996:MCP

[Tra96] S. Tragoudas. Min-cut partitioning on underlying tree and graph structures. IEEE Transactions on Computers, 45(4):470–474, April 1996. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=494104.

Tripathi:1998:GEI

[Tri98] S. K. Tripathi. Guest Editor’s introduction [The 17th IEEE International Conference of Distributed Computing Systems]. IEEE Transactions on Computers, 47(4):369–370, April 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=675678.

Taylor:1995:SNA

[TRM95] V. E. Taylor, A. Ranade, and D. G. Messerschmitt. SPAR: a new architecture for large finite element computations. IEEE Transactions on Computers, 44(4):531–545, April 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=376168.

Tam:1999:AMD

[TRS^{+}99] E. S. Tam, J. A. Rivers, V. Srinivasan, G. S. Tyson, and E. S. Davidson. Active management of data caches by exploiting reuse information. IEEE Transactions on Computers, 48(11):1244–1259, November 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=811113.

Tripathy:1991:TGL

[TS91] P. Tripathy and B. Sarikaya. Test generation from LOTOS specifications. IEEE Transactions on Computers, 40(4):543–552, April 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=88472.

Tsai:1994:ATM

[TS94] Bing-Rung Tsai and K. G. Shin. Assignment of task modules in hypercube multicomputers with component failures for communication efficiency. IEEE Transactions on Computers, 43(5):613–618, May 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=280809.

Tzeng:1995:MOP

[TS95] Hong-Yi Tzeng and Kai-Yeung Siu. Message-optimal protocols for fault-tolerant broadcasts/multicasts in distributed systems with crash failures. IEEE Transactions on Computers, 44(2):346–352, February 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=364545.

Tseng:1997:TOB

[TS97] Yu-Chee Tseng and Jang-Ping Sheu. Toward optimal broadcast in a star graph using multiple spanning trees. IEEE Transactions on Computers, 46(5):593–599, May 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=589231.

Tse:1998:CCP

[TS98] J. Tse and A. J. Smith. CPU cache prefetching: Timing evaluation of hardware implementations. IEEE Transactions on Computers, 47(5):509–526, May 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=677225.

Tang:1999:DPC

[TS99a] Qian-Yu Tang and Xiaoyu Song. Diagnosis of parallel computers arbitrary connectivity. IEEE Transactions on Computers, 48(7):757–761, July 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=780885.

Tomita:1999:IDD

[TS99b] A. W. Tomita and K. Sakamura. Improving design dependability by exploiting an open model-based specification.

IEEE Transactions on Computers, 48(1):24–37, January 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=743409.

Tseng:1997:DDB

[Tse97] Yu-Chee Tseng. A dilateddiagonal-based scheme for broadcast in a wormhole-routed 2D torus. IEEE Transactions on Computers, 46(8):947–952, August 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=609282.

Thibeault:1994:FME

[TSH94] C. Thibeault, Y. Savaria, and J.-L. Houle. A fast method to evaluate the optimum number of spares in defect-tolerant integrated circuits. IEEE Transactions on Computers, 43(6):687–697, June 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=286302.

Thibeault:1995:EPS

[TSH95] C. Thibeault, Y. Savaria, and J. L. Houle. Equivalence proofs of some yield modeling methods for defect-tolerant integrated circuits. IEEE Transactions on Computers, 44(5):724–728, May 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=381962.

Tridandapani:1995:LOM

[TSS95] S. Tridandapani, A. K. Somani, and U. R. Sandadi. Low overhead multiprocessor allocation strategies exploiting system spare capacity for fault detection and location. IEEE Transactions on Computers, 44(7):865–877, July 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=392845.

Thiebaut:1992:IDC

[TSW92] D. Thiebaut, H. S. Stone, and J. L. Wolf. Improving disk cache hit-ratios through cache partitioning. IEEE Transactions on Computers, 41(6):665–676, June 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=144619.

Tamir:1990:HPF

[TT90] Y. Tamir and M. Tremblay.

High-performance fault-tolerant VLSI systems using micro rollback. IEEE Transactions on Computers, 39(4):548–554, April 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=54848.

Tuan:1992:OSJ

[Tua92] T.-C. Tuan. On optimal single jog river routing [VLSI layout]. IEEE Transactions on Computers, 41(3):366–369, March 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=127451.

Tzeng:1991:EH

[TW91] N.-F. Tzeng and S. Wei. Enhanced hypercubes. IEEE Transactions on Computers, 40(3):284–294, March 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=76405.

Takach:1995:AMS

[TW95a] A. Takach and W. Wolf. An automaton model for scheduling constraints in synchronous machines. IEEE Transactions on Computers, 44(1):1–12, January 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=368014.

Tong:1995:RTD

[TW95b] Chao Chi Tong and Chuan-Lin Wu. Routing in a three-dimensional chip. IEEE Transactions on Computers, 44(1):106–117, January 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=368006.

Tseng:1995:CTL

[TW95c] Yuen-Hsien Tseng and Ja-Ling Wu. On a constant-time, lowcomplexity winner-take-all neural network. IEEE Transactions on Computers, 44(4):601–604, April 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=376175.

Thiebaut:1992:STT

[TWS92] D. Thiebaut, J. L. Wolf, and H. S. Stone. Synthetic traces for trace-driven simulation of cache memories. IEEE Transactions on Computers, 41(4):388–410, April 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=135552.

See comments [Mah94a].

Thiebaut:1993:CST

[TWS93] D. Thiebaut, J. Wolf, and H. Stone. Corrigendum to ‘synthetic traces for trace-driven simulation of cache memories’. IEEE Transactions on Computers, 42(5):635–636, May 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=223685.

Torrellas:1998:OIC

[TXD98] J. Torrellas, Chun Xia, and R. L. Daigle. Optimizing the instruction cache performance of the operating system. IEEE Transactions on Computers, 47(12):1363–1381, December 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=737683.

Takagi:1992:MMH

[TY92a] N. Takagi and S. Yajima. Modular multiplication hardware algorithms with a redundant representation and their application to RSA cryptosystem. IEEE Transactions on Computers, 41(7):887–891, July 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=256444.

Tsay:1992:SCA

[TY92b] Jong-Chuang Tsay and Sy Yuan. Some combinatorial aspects of parallel algorithm design for matrix multiplication. IEEE Transactions on Computers, 41(3):355–361, March 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=127449.

Tyagi:1993:RAS

[Tya93]  A. Tyagi.  A reduced-area

scheme for carry-select adders.

IEEE Transactions on Computers, 42(10):1163–1170, October 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=257703.

Tseng:1999:AFT

[TYJ99] Yu-Chee Tseng, Ming-Hour Yang, and Tong-Ying Juang. Achieving fault-tolerant multicast in injured wormholerouted tori and meshes based on Euler path construction. IEEE Transactions on Computers, 48(11):1282–1296, November 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=811118.

Tzeng:1992:CEC

[Tze92] N.-F. Tzeng. A cost-effective combining structure for largescale shared-memory multiprocessors. IEEE Transactions on Computers, 41(11): 1420–1429, November 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=177312.

Tzeng:1993:CCC

[Tze93]  N. F. Tzeng.  A cube-

connected cycles architecture with high reliability and improved performance. IEEE Transactions on Computers, 42 (2):246–253, February 1993.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=204798.

Tzeng:1994:RBD

[Tze94] Nian-Feng Tzeng. Reliable butterfly distributed-memory multiprocessors. IEEE Transactions on Computers, 43(9): 1004–1013, September 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=312111.

Uht:1991:TRM

[Uht91] A. K. Uht. A theory of reduced and minimal procedural dependencies. IEEE Transactions on Computers, 40(6):681–692, June 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=90247.

Uht:1992:CEH

[Uht92] A. K. Uht. Concurrency extraction via hardware methods executing the static instruction stream. IEEE Transactions on Computers, 41(7):826–841, July 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=256451.

Unger:1994:SAS

[Ung94] S. H. Unger. Some additions to “Solution of switching equations based on a tabular algebra”. IEEE Transactions on Computers, 43(3):365–367, March 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=272437.

See [TLROL93].

Unger:1995:HCR

[Ung95] S. H. Unger. Hazards, critical races, and metastability. IEEE Transactions on Computers, 44(6):754–768, June 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=391185.

Upadhyaya:1993:ANS

[UP93] S. J. Upadhyaya and H. Pham.

Analysis of noncoherent systems and an architecture for the computation of the system reliability. IEEE Transactions on Computers, 42(4):484–493, April 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=214699.

Upadhyaya:1994:CPM

[UR94] S. J. Upadhyaya and B. Ramamurthy. Concurrent process monitoring with no reference signatures. IEEE Transactions on Computers, 43(4):475–480, April 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=278485.

Upadhyay:1997:TBA

[UVM97] J. Upadhyay, V. Varavithya, and P. Mohapatra. A trafficbalanced adaptive wormhole routing scheme for two-dimensional meshes. IEEE Transactions on Computers, 46(2):190–197, February 1997. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
565594.

Ural:1997:MLC

[UWZ97] H. Ural, Xiaolin Wu, and Fan Zhang. On minimizing the lengths of checking sequences. IEEE Transactions on Computers, 46(1):93–99, January 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=559807.

Vaidya:1995:UBB

[Vai95]  N. H. Vaidya.  Unidirec-

tional bit/byte error control. IEEE Transactions on Computers, 44(5):710–714, May 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=381959.

Vaidya:1996:CDT

[Vai96] N. H. Vaidya. Comparison of duplex and triplex memory reliability. IEEE Transactions on Computers, 45(4):503–507, April 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=494111.

Vaidya:1997:ICL

[Vai97] N. H. Vaidya. Impact of checkpoint latency on overhead ratio of a checkpointing scheme. IEEE Transactions on Computers, 46(8):942–947, August 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=609281.

Vaidya:1998:CTL

[Vai98] N. H. Vaidya. A case for two-level recovery schemes. IEEE Transactions on Computers, 47(6):656–666, June 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=689645.

Varvarigos:1996:CSR

[VB96] E. A. Varvarigos and D. P. Bertsekas. A conflict sense routing protocol and its performance for hypercubes. IEEE Transactions on Computers, 45(6):693–703, June 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=506425.

Valencia:1995:MAA

[VBH^{+}95] M. Valencia, M. J. Bellido, J. L. Huertas, A. J. Acosta, and S. Sanchez-Solano. Modular asynchronous arbiter insensitive to metastability. IEEE Transactions on Computers, 44 (12):1456–1461, December 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=477251.

Varma:1992:FTA

[VC92] A. Varma and S. Chalasani. Fault-tolerance analysis of onesided crosspoint switching networks. IEEE Transactions on Computers, 41(2):143–158, February 1992. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
123392.

Vassilladis:1996:ARA

[VCB96] S. Vassilladis, S. Contofana, and K. Bertels. 2-1 addition and related arithmetic operations with threshold logic. IEEE Transactions on Computers, 45(9):1062–1067, September 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=537130.

Varman:1992:SLS

[VD92] P. J. Varman and K. Doshi.

Sorting with linear speedup on a pipelined hypercube. IEEE Transactions on Computers, 41(1):97–103, January 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=123384.

vandenBerg:1993:PMR

[vdBT93] J. van den Berg and D. Toswley. Properties of the miss ratio for a 2-level storage model with LRU or FIFO replacement strategy and independent references. IEEE Transactions on Computers, 42(4):508–512, April 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=214697.

vandeLiefvoort:1993:NAP

[vdLS93] A. van de Liefvoort and N. Subramanian. A new approach for the performance analysis of a single-bus multiprocessor system with general service times. IEEE Transactions on Computers, 42(3):358–362, March 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=210178.

Vergis:1992:TOD

[Ver92] A. Vergis. On the testability of one-dimensional ILAs for multiple sequential faults. IEEE Transactions on Computers, 41(7):906–916, July 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=256448.

Vijayan:1991:GMC

[Vij91]  G. Vijayan.  Generalization

of min-cut partitioning to tree structures and its applications. IEEE Transactions on Computers, 40(3):307–314, March 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=76407.

Vinnakota:1995:IMS

[Vin95] B. Vinnakota. Implementing multiplication with split readonly memory. IEEE Transactions on Computers, 44(11): 1352–1356, November 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=475134.

Vinnakota:1993:DDA

[VJ93] B. Vinnakota and N. K. Jha. Diagnosability and diagnosis of algorithm-based fault-tolerant systems. IEEE Transactions on Computers, 42(8):924–937, August 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=238483.

Varma:1998:DAD

[VJ98] A. Varma and Q. Jacobson.

Destage algorithms for disk arrays with nonvolatile caches. IEEE Transactions on Computers, 47(2):228–235, February 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=663770.

Valero:1995:CFA

[VLPA95] M. Valero, T. Lang, M. Peiron, and E. Ayguade. Conflict-free access for streams in multimodule memories. IEEE Transactions on Computers, 44(5):634–646, May 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=381949.

Vogt:1990:BBM

[Vog90]  C. Vogt.  A buffer-based

method for storage allocation in an object-oriented system. IEEE Transactions on Computers, 39(3):375–383, March 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=48867.

Vaidya:1993:FTD

[VP93] N. F. Vaidya and D. K. Pradhan.

Fault-tolerant design strategies for high reliability and safety.

IEEE Transactions on Computers, 42(10):1195–1206, October 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=257706.

Vaidya:1994:SSL

[VP94] N. H. Vaidya and D. K. Pradham. Safe system level diagnosis. IEEE Transactions on Computers, 43(3):367–370, March 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=272443.

Vaidya:1995:DBA

[VP95] N. H. Vaidya and D. K. Pradhan. Degradable Byzantine agreement. IEEE Transactions on Computers, 44(1):146–150, January 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=367999.

Vassiliadis:1993:ICA

[VPB93] S. Vassiliadis, J. Phillips, and B. Blaner. Interlock collapsing ALU’s. IEEE Transactions on Computers, 42(7):825–839, July 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=237723.

Vinnakota:1994:GAR

[VR94] B. Vinnakota and V. V. B. Rao.

Generation of all Reed–Muller expansions of a switching function. IEEE Transactions on Computers, 43(1):122–124, January 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=250618.

Varvarigou:1993:RPA

[VRK93a] T. A. Varvarigou, V. P. Roychowdhury, and T. Kailath. Reconfiguring processor arrays using multiple-track models: the 3-track-1-spare-approach. IEEE Transactions on Computers, 42(11):1281–1293, November 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=247834.

Varvarigou:1993:PTA

[VRK93b] T. A. Varvarigou, V. P. Roychowdhury, and T. Kailth. A polynomial time algorithm for reconfiguring multiple-track models. IEEE Transactions on Computers, 42(4):385–395, April 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=214686.

Vassiliadis:1993:CHW

[VS93] S. Vassiliadis and E. M. Schwarz.

Correction to “Hard-wired multipliers with encoded partial products”. IEEE Transactions on Computers, 42(1):127, January 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=192224.

Vassiliadis:1991:HWM

[VSS91] S. Vassiliadis, E. M. Schwarz, and B. M. Sung. Hard-wired multipliers with encoded partial products. IEEE Transactions on Computers, 40(11): 1181–1197, November 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=102823.

Veeraraghavan:1994:CAP

[VT94] M. Veeraraghavan and K. S. Trivedi. A combinatorial algorithm for performance and reliability analysis using multistate models. IEEE Transactions on Computers, 43 (2):229–234, February 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=262129.

VanScheik:1997:HSE

[VT97] W. S. VanScheik and R. F. Tinderr. High speed externally asynchronous/internally clocked systems. IEEE Transactions on Computers, 46(7):824–829, July 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=599904.

Vuillemin:1990:ERC

[Vui90] J. E. Vuillemin. Exact real computer arithmetic with continued fractions. IEEE Transactions on Computers, 39(8):1087–1105, August 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=57047.

Vuillemin:1994:CN

[Vui94] J. E. Vuillemin. On circuits and numbers. IEEE Transactions on Computers, 43(8):868–879, August 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=295849.

Walter:1993:SMM

[Wal93] C. D. Walter. Systolic modular multiplication. IEEE Transactions on Computers, 42(3):376–378, March 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=210181.

Walter:1997:STT

[Wal97]  C. D. Walter.  Space/time

trade-offs for higher radix modular multiplication using repeated addition. IEEE Transactions on Computers, 46 (2):139–141, February 1997.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=565588.

Walter:1998:EUD

[Wal98]  C. D. Walter.  Exponenti-

ation using division chains. IEEE Transactions on Computers, 47(7):757–765, July 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=709375.

Wang:1994:BLS

[Wan94a] Chin-Liang Wang. Bit-level systolic array for fast exponentiation in GF(2^{m}). IEEE Transactions on Computers, 43(7):838– 841, July 1994. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        293263.

Wang:1994:DEH

[Wan94b] Dajin Wang. Diagnosability of enhanced hypercubes. IEEE Transactions on Computers, 43(9):1054–1061, September 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=312114.

Wang:1996:BAM

[Wan96] Zheng-Ou Wang. A bidirectional associative memory based on optimal linear associative memory. IEEE Transactions on Computers, 45(10):1171–1179, October 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=543710.

Wang:1997:CGC

[Wan97]  Yi-Min Wang.  Consistent

global checkpoints that contain a given set of local checkpoints. IEEE Transactions on Computers, 46(4):456–468, April 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=588059.

Wang:1999:DHE

[Wan99] D. Wang. Diagnosability of hypercubes and enhanced hypercubes under the comparison diagnosis model. IEEE Transactions on Computers, 48 (12):1369–1374, December 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=817401.

Wu:1990:ETI

[WC90] Chen-Wen Wu and P. R. Cappello. Easily testable iterative logic arrays. IEEE Transactions on Computers, 39(5):640–652, May 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=53577.

Walker:1996:NSD

[WC96] J. Walker and A. Cantoni. A new synchronizer design. IEEE Transactions on Computers, 45(11):1308–1311, November 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=544488.

Wu:1999:QUS

[WC99] Guang-Ming Wu and Yao-Wen Chang. Quasi-universal switch matrices for FPD design. IEEE Transactions on Computers, 48(10):1107–1122, October 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=805159.

Wong:1992:OCT

[WD92] Yiwan Wong and J.-M. Delosme. Optimization of computation time for systolic arrays. IEEE Transactions on Computers, 41(2):159–177, February 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=123393.

Weber:1992:NSC

[WdB92] J. H. Weber, C. de Vroedt, and D. E. Boekee. Necessary and sufficient conditions on block codes correcting/ detecting errors of various types. IEEE Transactions on Computers, 41(9):1189–1193, September 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=165401.

Woods:1997:AAA

[WDF^{+}97] J. V. Woods, P. Day, S. B.

Furber, J. D. Garside, N. C. Paver, and S. Temple. AMULET1: an asynchronous ARM microprocessor. IEEE Transactions on Computers, 46(4):385–398, April 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=588033.

Wu:1995:DMC

[WDS^{+}95] J. Wu, R. Das, J. Saltz, H. Berryman, and S. Hiranandan. Distributed memory compiler design for sparse problems. IEEE Transactions on Computers, 44(6):737–753, June 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=391186.

Wang:1997:ATS

[WeBM97] Yuke Wang, M. Abd el Barr, and C. McCrosky. An algorithm for total symmetric OBDD detection. IEEE Transactions on Computers, 46(6):731–733, June 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=600890.

Wegener:1994:CCB

[Weg94a] I. Wegener. Comments on “A characterization of binary decision diagrams”. IEEE Transactions on Computers, 43(3):383–384, March 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=272442.

See [Cha93].

Wegener:1994:SRO

[Weg94b] I. Wegener. The size of reduced OBDD’s and optimal read-once branching programs for almost all Boolean functions. IEEE Transactions on Computers, 43(11):1262–1269, November 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=324559.

Wei:1994:SPS

[Wei94] Shyue-Win Wei. A systolic power-sum circuit for GF(2^{m}). IEEE Transactions on Computers, 43(2):226–229, February 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=262128.

Wei:1996:CMF

[Wei96] Bin Wei. Comments on “A multiaccess frame buffer architecture”. IEEE Transactions on Computers, 45(7):862, July 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=508324.

See [Har94].

Welch:1994:PVM

[Wel94] L. R. Welch. A parallel virtual machine for programs composed of abstract data types. IEEE Transactions on Computers, 43(11):1249–1261, November 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=324558.

Wu:1990:RDS

[WF90] K.-L. Wu and W. K. Fuchs. Recoverable distributed shared virtual memory. IEEE Transactions on Computers, 39(4):460–469, April 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=54839.

Wong:1992:FDU

[WF92] D. Wong and M. Flynn. Fast division using accurate quotient approximations to reduce the number of iterations. IEEE Transactions on Computers, 41(8):981–995, August 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=156541.

Wong:1994:FHB

[WG94] W. F. Wong and E. Goto. Fast hardware-based algorithms for elementary function computations using rectangular multipliers. IEEE Transactions on Computers, 43(3):278–294, March 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=272429.

Wong:1995:FEE

[WG95] W. F. Wong and E. Goto. Fast evaluation of the elementary functions in single precision. IEEE Transactions on Computers, 44(3):453–457, March 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=372037.

Wang:1998:AHD

[WG98a] Seongmoon Wang and S. K. Gupta. ATPG for heat dissipation minimization during test application. IEEE Transactions on Computers, 47 (2):256–262, February 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=663775.

Wu:1998:FTM

[WG98b] Jie Wu and Guanghui Guo.

Fault tolerance measures for mary n-dimensional hypercubes based on forbidden faulty sets. IEEE Transactions on Computers, 47(8):888–893, August 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=707589.

Weller:1994:COS

[WH94] T. Weller and B. Hajek. Comments on “An optimal shortestpath routing policy for network computers with regular mesh-connected topologies”. IEEE Transactions on Computers, 43(7):862–863, July 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=293267.

Wu:1997:EEP

[WH97a] Huapeng Wu and M. A. Hasan.

Efficient exponentiation of a primitive root in GF(2^{m}). IEEE Transactions on Computers, 46(2):162–172, February 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=565591.

Wu:1997:BHC

[WH97b] Jie Wu and Ke Huang. The balanced hypercube: a cubebased system for fault-tolerant applications. IEEE Transactions on Computers, 46(4):484–490, April 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=588063.

Wu:1998:LCB

[WH98] Huapeng Wu and M. A. Hasan. Low complexity bit-parallel multipliers for a class of finite fields. IEEE Transactions on Computers, 47(8):883–887, August 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=707588.

Wu:1999:CFE

[WH99] Huapeng Wu and M. A. Hasan. Closed-form expression for the average weight of signed-digit representations. IEEE Transactions on Computers, 48(8):848–851, August 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=795126.

Wu:1998:NLC

[WHB98] H. Wu, M. A. Hasan, and I. F. Blake. New low-complexity bit-parallel finite field multipliers using weakly dual bases. IEEE Transactions on Computers, 47(11):1223–1234, November 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=736433.

Wang:1997:PRS

[WHK97] Yi-Min Wang, Yennun Huang, and C. Kintala. Progressive retry for software failure recovery in message-passing applications. IEEE Transactions on Computers, 46(10):1137–1141, October 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=628398.

Wu:1993:QEC

[WHL93] C. E. Wu, Y. Hsu, and Y.- H. Liu. A quantitative evaluation of cache types for highperformance computer systems. IEEE Transactions on Computers, 42(10):1154–1162, October 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=257701.

Wu:1995:ESS

[WHL95] C. E. Wu, Yarsun Hsu, and Yew-Huey Liu. Efficient stack simulation for set-associative virtual address caches with real tags. IEEE Transactions on Computers, 44(5):719–723, May 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=381961.

Wu:1995:SRM

[WI95] Yuejian Wu and A. Ivanov. Single-reference multiple intermediate signature (SREMIS) analysis for BIST. IEEE Transactions on Computers, 44(6): 817–825, June 1995. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        391180.

Wilkinson:1992:CDA

[Wil92a] B. Wilkinson. Comments on “Design and analysis of arbitration protocols” by F. El Guibaly. IEEE Transactions on Computers, 41(3):348–351, March 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=127447.

Wilkinson:1992:CSM

[Wil92b] B. Wilkinson. On crossbar switch and multiple bus interconnection networks with overlapping connectivity. IEEE Transactions on Computers, 41(6):738–746, June 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=144625.

Wilken:1993:OGC

[Wil93]  K. D. Wilken.  An op-

timal graph-construction approach to placing program signatures for signature monitoring.

IEEE Transactions on Computers, 42(11):1372–1381, November 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=247847.

Windley:1995:FMV

[Win95] P. J. Windley. Formal modeling and verification of microprocessors. IEEE Transactions on Computers, 44(1):54–72, January 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=368009.

Wigley:1990:MRR

[WJ90] N. M. Wigley and G. A. Jullien. On modulus replication for residue arithmetic computations of complex inner products. IEEE Transactions on Computers, 39(8):1065–1076, August 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=57045.

Wang:1994:ABF

[WJ94] Sying-Jyan Wang and N. K.

Jha. Algorithm-based fault tolerance for FFT networks. IEEE Transactions on Computers, 43(7):849–854, July 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=293265.

Wang:1995:NDT

[WJM95] Zhongde Wang, G. A. Jullien, and W. C. Miller. A new design technique for column compression multipliers. IEEE Transactions on Computers, 44(8):962–970, August 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=403712.

Wigley:1994:LDR

[WJR94] N. M. Wigley, G. A. Jullien, and D. Reaume. Large dynamic range computations over small finite rings. IEEE Transactions on Computers, 43(1):78–86, January 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=250611.

Wilken:1997:CDS

[WK97] K. D. Wilken and T. Kong. Concurrent detection of software and hardware data-access faults. IEEE Transactions on Computers, 46(4):412–424, April 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=588046.

Wang:1990:ECR

[WL90] J.-S. Wang and R. C. T. Lee. An efficient channel routing algorithm to yield an optimal solution. IEEE Transactions on Computers, 39(7):957–962, July 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=55698.

Wu:1992:PAM

[WL92] Chuan-Lin Wu and Manjai Lee.

Performance analysis of multistage interconnection network configurations and operations. IEEE Transactions on Computers, 41(1):18–27, January 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=123379.

Wang:1993:SAC

[WL93] Chin-Liang Wang and Jung-Lung Lin. A systolic architecture for computing inverses and divisions in finite fields GF(2^{m}). IEEE Transactions on Computers, 42(9):1141–1146, September 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=241603.

Wang:1994:PSF

[WL94] J. Q. Wang and P. K. Lala.

Partially strongly fault secure and partially strongly code disjoint 1-out-of-3 code checker. IEEE Transactions on Computers, 43(10):1238–1240, October 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=324550.

Wang:1991:TDS

[WM91] L.-X. Wang and J. M. Mendel. Three-dimensional structured networks for matrix equation solving. IEEE Transactions on Computers, 40(12): 1337–1346, December 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=106219.

Whitaker:1992:SSA

[WM92] S. R. Whitaker and G. K. Maki. Self synchronized asynchronous sequential pass transistor circuits. IEEE Transactions on Computers, 41 (10):1344–1348, October 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=166614.

Wang:1996:NTU

[WM96] Yuke Wang and C. McCrosky. Negation trees: a unified approach to Boolean function complementation. IEEE Transactions on Computers, 45(5):626– 630, May 1996. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        509916.

Wang:1998:SBE

[WM98] Yuke Wang and C. McCrosky.

Solving Boolean equations using ROSOP forms. IEEE Transactions on Computers, 47 (2):171–177, February 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=663763.

Wrzyszcz:1996:NAF

[WMD96] A. Wrzyszcz, D. Milford, and E. L. Dagless. A new approach to fixed-coefficient inner product computation over finite rings. IEEE Transactions on Computers, 45(12):1345–1355, December 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=545965.

Woodside:1995:SRN

[WNPM95] M. Woodside, J. E. Neilson, D. C. Petriu, and S. Majumdar. The stochastic rendezvous network model for performance of synchronous clientserver-like distributed software. IEEE Transactions on Computers, 44(1):20–34, January 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=368012.

Wang:1996:STL

[WNS96] Haigeng Wang, A. Nicolau, and K.-Y. S. Siu. The strict time lower bound and optimal schedules for parallel prefix with resource constraints. IEEE Transactions on Computers, 45(11):1257–1271, November 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=544482.

Wang:1990:VDC

[WP90] C. C. Wang and D. Pei. A VLSI design for computing exponentiations in GF(2^{m}) and its application to generate pseudorandom number sequences. IEEE Transactions on Computers, 39(2):258–262, February 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=45211.

Wang:1996:ODS

[WPS96] Qingyan Wang, N. L. Passos, and Edwin Hsing-Mean Sha. Optimal data scheduling for uniform multidimensional applications. IEEE Transactions on Computers, 45(12):1439–1444, December 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=545974.

Wang:1997:HCA

[WPS97] S. Wang, V. Piuri, and E. E. Swartzlander, Jr. Hybrid CORDIC algorithms. IEEE Transactions on Computers, 46 (11):1202–1207, November 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=644295.

Wu:1996:GPM

[WPZZ96] Haomin Wu, M. A. Perkowski, Xiaoqiang Zeng, and Nan Zhuang. Generalized partiallymixed-polarity Reed–Muller expansion and its fast computation. IEEE Transactions on Computers, 45(9): 1084–1088, September 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=537134.

Wang:1995:DRL

[WRS95] Fuxing Wang, K. Ramamritham, and J. A. Stankovic. Determining redundancy levels for fault tolerant real-time systems. IEEE Transactions on Computers, 44(2):292–301, February 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=364540.

Wong:1991:DRN

[WS91] Kar-Lik Wong and Wan-Chi Siu. Data routing networks for systolic/pipeline realization of prime factor mapping. IEEE Transactions on Computers, 40(9):1072–1074, September 1991. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=83654.

Wang:1995:SRE

[WS95] Shuenn-Shyang Wang and Ming-Yuan Shau. Single residue error correction based on Kterm m_{j}-projection. IEEE Transactions on Computers, 44(1):129–131, January 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=368003.

Wey:1998:DHS

[WS98] Chin-Long Wey and Ming-Der Shieh. Design of a high-speed square generator. IEEE Transactions on Computers, 47(9):1021–1026, September 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=713320.

Wong:1990:PAM

[WSB90] W.-C. Wong, T. Suda, and L. Bic. Performance analysis of a message-oriented knowledgebase. IEEE Transactions on Computers, 39(7):951–957, July 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=55697.

Won:1990:HAM

[WSEZ90] Y. Won, S. Sahni, and Y. El-Ziq.

A hardware accelerator for maze routing. IEEE Transactions on Computers, 39(1):141–145, January 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=46291.

Wang:1997:MAC

[WSY97] H. Wang, T. Sun, and Q. Yang.

Minimizing area cost of onchip cache memories by caching address tags. IEEE Transactions on Computers, 46(11): 1187–1201, November 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=644293.

Wei:1990:ATO

[WT90] B. W. Y. Wei and C. D. Thompson. Area-time optimal adder design. IEEE Transactions on Computers, 39(5):666–675, May 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=53579.

Wu:1995:SLE

[Wu95] Jie Wu. Safety levels — an efficient mechanism for achieving reliable broadcasting in hypercubes. IEEE Transactions on Computers, 44(5):702–706, May 1995. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=381957.

Wu:1997:RUF

[Wu97]  Jie Wu.  Reliable unicast-

ing in faulty hypercubes using safety levels. IEEE Transactions on Computers, 46 (2):241–247, February 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=565613.

Wang:1991:DIS

[WW91] L. Wang and C.-L. Wu. Distributed instruction set computer architecture. IEEE Transactions on Computers, 40(8):915–934, August 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=83637.

Wu:1995:LGI

[WY95] Jie Wu and Kejun Yao. A limited-global-information-based multicasting scheme for faulty hypercubes. IEEE Transactions on Computers, 44(9): 1162–1167, September 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=464397.

Wang:1993:NWL

[WZXX93] Tao Wang, Xinhau Zhuang, XiaoLiang Xing, and Xipeng Xiao. A neuron-weighted learning algorithm and its hardware implementation in associative memories. IEEE Transactions on Computers, 42(5):636–640, May 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=223686.

Xu:1995:SDS

[XH95] Jie Xu and Shi-Ze Huang.

Sequentially t-diagnosable systems: a characterization and its applications. IEEE Transactions on Computers, 44 (2):340–345, February 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=364544.

Xiaoqing:1992:TDL

[XK92] W. Xiaoqing and K. Kinoshita.

A testable design of logic circuits under highly observable condition. IEEE Transactions on Computers, 41(5):654–659, May 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=142692.

Xiao:1996:GSB

[XSFR96] Sihai Xiao, Xiaofa Shi, Guiliang Feng, and T. R. N. Rao.

A generalization of the single b-bit byte error correcting and double bit error detecting codes for high-speed memory systems. IEEE Transactions on Computers, 45(4):508–511, April 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=494112.

Xia:1999:CHS

[XT99] Chun Xia and J. Torrellas. Comprehensive hardware and software support for operating systems to exploit MP memory hierarchies. IEEE Transactions on Computers, 48(5):494–505, May 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=769432.

Youssef:1990:EBF

[YA90] A. Youssef and B. W. Arden. Equivalence between functionality and topology of log N-stage banyan networks. IEEE Transactions on Computers, 39(6):829–832, June 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=53604.

Yang:1990:COP

[Yan90] C.-L. Yang. Constructing optimal procedures for testing series systems. IEEE Transactions on Computers, 39(7):943–945, July 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=55695.

Yang:1993:INC

[Yan93] Quing Yang. Introducing a new cache design into vector computers. IEEE Transactions on Computers, 42(12):1411–1424, December 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=260632.

Yang:1997:SPN

[Yan97] C. D. Yang. The smallest pair of noncrossing paths in a rectilinear polygon. IEEE Transactions on Computers, 46(8):930–941, August 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=609280.

Yang:1998:CIN

[Yan98] Yuanyuan Yang. A class of interconnection networks for multicasting. IEEE Transactions on Computers, 47(8):899–906, August 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=707592.

Yang:1991:APS

[YB91] Q. Yang and L. N. Bhuyan. Analysis of packet-switched multiplebus multiprocessor systems.

IEEE Transactions on Computers, 40(3):352–357, March 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=76414.

Yuan:1994:CAC

[YC94] Shyan-Ming Yuan and Her-Kun Chang. Comments on “Availability of k-Coterie”. IEEE Transactions on Computers, 43(12):1457, December 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=338112.

See [KFYA93].

Yang:1992:RST

[YF92a] Guu-Chang Yang and T. Fuja. The reliability of systems with two levels of fault tolerance: the return of the ‘birthday surprise’. IEEE Transactions on Computers, 41(11):1490–1496, November 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=177320.

Yeh:1992:CRN

[YF92b] Yao-Ming Yeh and T. Feng. On a class of rearrangeable networks. IEEE Transactions on Computers, 41(11):1361–1379, November 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=177307.

Yoshida:1991:PRT

[YGI91] N. Yoshida, E. Goto, and S. Ichikawa. Pseudorandom rounding for truncated multipliers. IEEE Transactions on Computers, 40(9): 1065–1067, September 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=83650.

Yuan:1998:NTO

[YK98] Shih-Yi Yuan and Sy-Yen Kuo.

A new technique for optimization problems in graph theory. IEEE Transactions on Computers, 47(2):190–196, February 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=663765.

Yen:1995:IDSb

[YL95a] S.-M. Yen and C.-S. Laih. Improved digital signature suitable for batch verification. IEEE Transactions on Computers, 44(7):957–959, July 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=392857.

Yen:1995:IDSa

[YL95b] Sung-Ming Yen and Chi-Sung Laih. Improved digital signature algorithm. IEEE Transactions on Computers, 44(5):729–730, May 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=381963.

See correction [Ano96b].

Yang:1998:DCC

[YL98] Ching-Nung Yang and Chi-Sung Laih. DC_{m} codes for constructing t-EC/AUED codes. IEEE Transactions on Computers, 47(4):492–495, April 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=675719.

Yang:1990:PGA

[YLC90] C.-B. Yang, R. C. T. Lee, and W.-T. Chen. Parallel graph algorithms based upon broadcast communications. IEEE Transactions on Computers, 39 (12):1468–1472, December 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=61071.

Yoon:1990:PAM

[YLL90] H. Yoon, K. Y. Lee, and M. T. Liu. Performance analysis of multibuffered packet-switching networks in multiprocessor systems. IEEE Transactions on Computers, 39(3):319–327, March 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=48863.

Yang:1991:NBS

[YM91] Y. Yang and G. M. Masson.

Nonblocking broadcast switching networks. IEEE Transactions on Computers, 40(9): 1005–1015, September 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=83662.

Yamamoto:1992:FTL

[YM92] Y. Yamamoto and M. Mukaidono.

P -functions — ternary logic functions capable of correcting input failures and suitable for treating ambiguities. IEEE Transactions on Computers, 41(1):28–35, January 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=123380.

Yang:1995:BRS

[YM95] Yuanyuan Yang and G. M.

Masson. Broadcast ring sandwich networks. IEEE Transactions on Computers, 44 (10):1169–1180, October 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=467692.

Yang:1999:NCC

[YM99] Yuanyuan Yang and G. M. Masson. The necessary conditions for Clos-type nonblocking multicast networks. IEEE Transactions on Computers, 48 (11):1214–1227, November 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=811110.

Yuan:1999:DPR

[YMG99] X. Yuan, R. Melhem, and R. Gupta. Distributed path reservation algorithms for multiplexed all-optical interconnection networks. IEEE Transactions on Computers, 48(12): 1355–1363, December 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=817397.

Yokoo:1992:OUF

[Yok92] H. Yokoo. Overflow/underflowfree floating-point number representations with self-delimiting variable-length exponent field. IEEE Transactions on Computers, 41(8):1033–1039, August 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=156546.

Yount:1996:MRI

[YS96] C. R. Yount and D. P. Siewiorek. A methodology for the rapid injection of transient hardware errors. IEEE Transactions on Computers, 45(8):881–891, August 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=536231.

Yang:1994:ERM

[YTR94] Pei-Ji Yang, Sing-Ban Tien, and C. S. Raghavendra. Embedding of rings and meshes onto faulty hypercubes using free dimensions. IEEE Transactions on Computers, 43(5):608–613, May 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=280808.

Yang:1999:WSN

[YW99] Yuanyuan Yang and Jianchao Wang. Wide-sense nonblocking Clos networks under packing strategy. IEEE Transactions on Computers, 48(3):265–284, March 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=754994.

Yeung:1997:SBD

[YY97] K. H. Yeung and T. S. Yum.

Selective broadcast data distribution systems. IEEE Transactions on Computers, 46(1):100–104, January 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=559808.

Zarowski:1995:PIS

[Zar95]  C. J. Zarowski.  Parallel

implementation of the Schur Berlekamp–Massey algorithm on a linearly connected processor array. IEEE Transactions on Computers, 44(7):930–932, July 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=392862.

Ziv:1997:LAC

[ZB97] A. Ziv and J. Bruck. An on-line algorithm for checkpoint placement. IEEE Transactions on Computers, 46 (9):976–985, September 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=620479.

Ziv:1998:ACS

[ZB98] A. Ziv and J. Bruck. Analysis of checkpointing schemes with task duplication. IEEE Transactions on Computers, 47 (2):222–227, February 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=663769.

Zarowski:1990:AMH

[ZC90] C. J. Zarowski and H. C. Card. On addition and multiplication with Hensel codes. IEEE Transactions on Computers, 39(12):1417–1423, December 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=61062.

Zaharioudakis:1998:HAC

[ZC98] M. Zaharioudakis and M. J.

Carey. Hierarchical, adaptive cache consistency in a page server OODBMS. IEEE Transactions on Computers, 47(4): 427–444, April 1998. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
675712.

Zhang:1999:ENT

[ZCT99] Zheng Zhang, M. Cintra, and J. Torrellas. Excel-NUMA: toward programmability, simplicity, and high performance. IEEE Transactions on Computers, 48(2):256–264, February 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=752667.

Zheng:1994:CTM

[Zhe94] Si-Qing Zheng. Compressed tree machines. IEEE Transactions on Computers, 43 (2):222–225, February 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=262127.

Zorian:1992:ESR

[ZI92] Y. Zorian and A. Ivanov. An effective BIST scheme for ROM’s. IEEE Transactions on Computers, 41(5):646–653, May 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=142691.

Zissimopoulos:1991:ANC

[ZPP91] V. Zissimopoulos, V. T. Paschos, and F. Pekergin. On the approximation of NP-complete problems by using the Boltzmann machine method: the cases of some covering and packing problems. IEEE Transactions on Computers, 40(12):1413–1418, December 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=106226.

Zorzi:1997:ECE

[ZR97] M. Zorzi and R. R. Rao. Error control and energy consumption in communications for nomadic computing. IEEE Transactions on Computers, 46(3):279–289, March 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=580424.

Zheng:1999:COS

[ZS99] S. Q. Zheng and M. Sun. Constructing optimal search trees in optimal time. IEEE Transactions on Computers, 48(7):738– 743, July 1999. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
780881.

Zhao:1990:WPT

[ZSR90] W. Zhao, J. A. Stankovic, and K. Ramamritham. A window protocol for transmission of time-constrained messages. IEEE Transactions on Computers, 39(9):1186–1203, September 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=57059.

Zuras:1994:MSM

[Zur94] D. Zuras. More on squaring and multiplying large integers. IEEE Transactions on Computers, 43(8):899–908, August 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=295852.

Zilic:1995:MVR

[ZV95] Z. Zilic and Z. G. Vranesic.

A multiple-valued Reed–Muller transform for incompletely specified functions. IEEE Transactions on Computers, 44 (8):1012–1020, August 1995.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=403717.

Zilic:1998:UDD

[ZV98] Z. Zilic and Z. G. Vranesic. Using decision diagrams to design ULMs for FPGAs. IEEE Transactions on Computers, 47 (9):971–982, September 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=713316.

Zhang:1996:SGN

[ZVDF96] Ming Zhang, S. Vassiliadis, and J. G. Delgado-Frias. Sigmoid generators for neural computing using piecewise approximations. IEEE Transactions on Computers, 45(9):1045–1049, September 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=537127.

Zafirovic-Vukotic:1992:PME

[ZVN92] M. Zafirovic-Vukotic and I. G. Niemegeers. A performance modeling and evaluation of the Cambridge Fast Ring. IEEE Transactions on Computers, 41 (9):1110–1125, September 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=165393.

Zhao:1998:TBF

[ZWL98] Lan Zhao, D. M. H. Walker, and F. Lombardi. I_{DDQ} testing of bridging faults in logic resources of reconfigurable field programmable gate arrays. IEEE Transactions on Computers, 47(10):1136–1152, October 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=729796.